A Complete Bibliography of ACM Transactions on Architecture and Code Optimization

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

21 September 2024 Version 1.91

Title word cross-reference

2 [BSL17, LLC22, LD24, SKP^{+}22]. 3 [CAY^{+}18, CWMC16, DVG^{+}23, JHH^{+}23, LGP^{+}16, LLC22, LCW^{+}24, LD24, NRQ16b, SZJK18, SKP^{+}22, ZSLX13]. ^{2} [DZL^{+}24]. ^{3} [CCZ13, DDT^{+}17]. K [Abd20, WZX^{+}24]. QR [BHWN21]. Z [SLM12].

* [SCFD22].

-D [CAY^{+}18]. -means [Abd20]. -mer [WZX^{+}24]. -polytopes [SLM12]. -Tree [WX24].

/channel [LCL^{+}14].

000-core [DAKK19].

2.5D [SKP^{+}22]. 2014 [Aca16, Ano15].

4.0 [KHB^{+}20].

6 [KWM^{+}08]. 64-bit [BWLR06, VED07].

7 [BKM^{+}17]. 754 [LDG^{+}13].

A-DFA [BC13]. Abakus [WZX^{+}24]. Aborts [RLS15]. ABS [AGI^{+}12]. Abstract [LMA^{+}16, PD17]. Abstracting [JSH09]. Abstraction [RLBBN15, ZM15, RCV^{+}12]. Accelerate [CNS^{+}16b, LCW^{+}23]. Accelerated [HS05, SWF16, VZT^{+}20, ZPL^{+}21, ZLM^{+}23, GMZ^{+}21, JED19]. Accelerating [BAZ^{+}19, CGP23, DAKK19, FPK^{+}24, GGYK19, GÁSÁ^{+}13, GR15, HHW^{+}22, JYJ^{+}13, KFJ20, LCP^{+}21, LWF^{+}16, LGH^{+}21, RMA14, SJC^{+}21, TMP16, WPR^{+}22, WZX^{+}24, XSF^{+}23, ZBC^{+}22, HWX^{+}13]. Acceleration [Abd20, BHWN21, GáSÁ^{+}16, HAC13, HBPH24, RVKP19, WFKL10]. Accelerator [CLA^{+}19, DLS22, MCB^{+}12, MMLS21, QAT24, RJK24, SMN22, SNK^{+}23, WWY^{+}24, XLH^{+}24, YXMC23, YZZ^{+}23, YCA18, LHWB12, TWB21, VDSP09]. accelerator-based [LHWB12]. Accelerator-bound [CLA^{+}19]. Accelerators [CKP^{+}22, ICEDR24, JHHM21, KCA^{+}13, KMG14, LWC^{+}22, LSH^{+}23, MTK18, MFSK24, SJD22, USCM16, XLD^{+}24, BKA13, CI13]. Access [CG15b, CSK19, GFD^{+}14, HK14, LGP^{+}16, LHC^{+}17, LWS^{+}19, LTX16, PWE20, SKH^{+}16, WZW^{+}24, XHJY16, XVT20, CLA^{+}19, FTLG11, HLR^{+}13, HCC^{+}14, JSH09, KCKG14, LWH11]. Accesses [CSY20, HEDH21, XHL^{+}24]. Accounting [LMA^{+}16, DEE13, LMCV13]. Accumulate [GG18]. Accumulation [ZBC^{+}22]. Accuracy [AAI^{+}16, ASS17, AMS23]. Accurate [NDP17, PDCS24, SCMU22, SMM^{+}23, WAST16, LMJ^{+}13b]. Achieving [ZNX^{+}24]. ACM [Aca16, Ano13a, Ano15, Bil19]. Across [ELE^{+}23, FDF^{+}14, NDP17, SW17a]. ACTION [MK23]. Activation [DQCK24]. activations [JLCR13]. Active [KHS^{+}14]. Adapt [DGI^{+}14, PGB13]. adaptation [DJB13, LGAZ07, SS04]. Adapting [GHH15, LBJ05]. Adaptive [CG14, CWMC16, FQRG13, GWZ22, GFD^{+}14, HWX^{+}13, JRK16, Lee16, LYH16, LGH24, MK23, Per18, SOAK23, WCI^{+}16, WM11, AGI^{+}12, JML^{+}20, MAN^{+}08, RBM10, SW13, YRGES^{+}19, ZK05]. Adaptively [ZCF18]. Adaptivity [DRHK15]. ADC [XLD^{+}24]. ADC-less [RJK24]. Address [AKBS21, BDB^{+}20, CTY^{+}23, JED19, LHC^{+}24, OAM19, SKAEG16, YSH^{+}22, ZCDD23, CCZ13, VS08, ZPC06]. Address-first [OAM19]. Addressability [YXS^{+}22]. Addressing [WA08, CWCS13]. Admission [LWW^{+}24]. Advancing [FSP^{+}23, TZK18]. Affine [AP17, NCC13, SLM12]. Against [BCHC19, ERAG^{+}16, PHBC17, BVIB12, SDK^{+}22, ZHS^{+}19]. Agent [JPS17, KKL^{+}24]. Aggregate [LY16]. Aggregation [AYC16, JZY^{+}22]. Aggressiveness [PB15]. Aging [DGI^{+}14, KKW^{+}15, LRBG15]. Aging-Aware [LRBG15]. Agnostic [SLJ^{+}18, ZDC^{+}16]. agreement [GMW09]. Ahead [PKPM19]. Ahead-of-Time [PKPM19]. Aho [CW13, PLL10]. AIM [AYC16]. ALEA [MPW^{+}17]. Algebra [APG^{+}23]. Algebraic [WCW^{+}24]. Algorithm [BC13, CRBP24, DGI^{+}14, DTD16, LCW^{+}23, BRSJG12, CW13, CDPD13, HAJ^{+}12, PLL10, XC06, ZGC^{+}12]. Algorithmic [AAI^{+}16, NCC13]. Algorithms [Pro21, OGK^{+}12, VTN13]. All-gather [Pro21]. All-photonic [DLS22]. Alleviate [ZZH^{+}23]. Allocation [DHD^{+}14, JCG^{+}24, KPM21, LDL22, PS12, RTK15, BZS13, CS10, GW09, RB13]. allocator [DHC^{+}13]. ALP [SLA^{+}07]. Alternative [Mic18, SKPD19]. AMX [MFSK24]. Analogue [DSK19]. Analyses [SGS^{+}20]. Analysis [ACG24, AGG22, CLA^{+}19, DZSL20, DSR15, FPK^{+}24, GAM12, GAH22, JK17, KR19, LMZ18, LLS23, MMdS06, SQZK20, SSW^{+}19, VTN13, VGX16, XFS^{+}19, ARS04, AFD12, FER^{+}13, JOA^{+}09b, Nas13, SV05, SMK10, ZCW10]. analytic [XMM04]. Analytical [BEE15, AFD07, CA11]. Analytics [KPP21]. Analyzer [SCMU22]. Analyzing [WLWB19, XYY^{+}24]. Anatomy [LCP^{+}21]. Annotation [MGA^{+}17]. Anomalies [LDC15]. Ant [SGM^{+}22]. Anticipating [LJMG12]. ApHMM [FPK^{+}24]. API [CI13]. Application [GTT^{+}16, LWC^{+}22, OKJ^{+}22, PLT^{+}15, SCFD22, UDLD20, AS13, GÁSÁ^{+}13, RCV^{+}12, SB09, TDP15]. Application-Guided [GTT^{+}16]. Application-Level [PLT^{+}15]. Application-oblivious [LWC^{+}22]. Application-Specific [UDLD20]. Applications [ASS17, AZG17, APS22, CPG21, DMR^{+}16, DTD16, DPBI^{+}19, FWJ^{+}16, GR15, HDW21, JYE^{+}16, KPRK20, LWS^{+}19, LSH^{+}23, MST^{+}21, MPU^{+}23, NKH16, NMPS22, RHLA14, RSU^{+}20, RMA14, RLBBN15, TMSR23, UPR22, WZG^{+}19, XFS^{+}19, CS13, DWDS13, HLR^{+}13, KNBK12, MBKM12, STLM12, SV05, SLA^{+}07, SLM12, YLTL04, ZG05]. Applied [LB10]. applying [ZWHM05]. Approach [AZG17, CNS^{+}16b, CKP^{+}22, EMR14, FDF^{+}14, GGK18, KS16, MRK^{+}22, RKL23, TS15, WAST16, WZG^{+}19, ZX16, FT10, SSR13, WYJL10, YJTF13, ZCS06]. approachable [WHV^{+}13]. Approx [AMS23]. Approx-RM [AMS23]. Approximate [DS12, SPS23, YPT^{+}16]. Approximation [SMM^{+}23, LTG12]. Apps [MPHL22, PCM16]. Arbitrary [PWE20, RHC15, WMGS19, WWGS22]. arbitration [XCC^{+}13]. Architecting [CPB^{+}07, NTV^{+}22]. Architectural [CPS^{+}15, DCP^{+}12, HEMK17, KLA^{+}19, LZZ^{+}22, ME15, MTPK24, QSZ^{+}21, WAST16, WZG^{+}19, YHYBAM20, IMS^{+}08, SB09, ZZQ^{+}05, CWC06]. Architecturally [KBB^{+}14]. Architecture [BOEN23, DNS^{+}24, FBC^{+}22, HK14, HLZ^{+}24, KAC^{+}18, LWS^{+}19, OBdOJ^{+}24, OK21, PVS^{+}17, SLJ^{+}18, SM19, SHY14, SWF16, SAM^{+}23, VC16, VFJ^{+}17, WDX24, XMW^{+}21, XLD^{+}24, XVT20, ZFT^{+}18, ZML^{+}24, ARS04, BVIB12, BWG^{+}12, CPB^{+}07, DJX13, GKP14, GSZI10, JYJ^{+}13, JA14, LNLK13, PM12, STLM12, SNL^{+}04, SRLPV04, SSPL^{+}13, ZK06].

Architecture-Agnostic [SLJ^{+}18]. architecture-independent [BVIB12]. Architectures [ÄJE^{+}16, ASK^{+}16, ASP17, CG15a, CEP^{+}16, CDPN16, GR15, HAM17, HAM19, HHW^{+}22, JLJ^{+}18a, JLL^{+}23, LAS^{+}13, LZM14, MST^{+}21, MK23, PT17, RMA14, SJL^{+}20, ZLYZ16, ZCQ^{+}19, ZMLL24, BBG13, BWLR06, BTS10, CG14, CK11, CDM13, KCP13, LKL^{+}13, LHC^{+}24, OGK^{+}12, RCV^{+}12, SSK11, SD12, SB09, TC07, TDG13, VE13, YXK^{+}12]. Area [LAS^{+}13, MP22, MPU^{+}23, SB09]. Area-Efficient [MPU^{+}23, SB09]. ARI [FQRG13]. Arithmetic [LVR^{+}15, UDLD20, BWG^{+}12]. ARM [GDL16, HZN^{+}22, LHW^{+}19, SHY14, SPH^{+}17]. ARM-to-x86 [LHW^{+}19]. ArmorAll [KPRK20]. Array [DSK19, LY24, WG17, XMW^{+}21, BWLR06, KLMP12]. Arrays [LMSE18, TD16, YSD^{+}23]. Arrival [Pro21]. ARSEC [DDT^{+}17]. Art [MWJ19]. As-Is [SPS23]. ASA [ZBC^{+}22]. ASM [SOAK23]. Assembly [LVR^{+}15]. Assessing [FCVR24]. Assigned [DCL^{+}22]. Assignment [CSW^{+}23, JOB^{+}22]. assistance [JOA^{+}09a]. Assisted [CDPN16, HNKK17, JDZ^{+}13, KKAR16, PHBC17, CST^{+}06, ZZL^{+}21]. associative [HL07, KWCL09]. associativity [YJTF13]. Asymmetric [ZCQ^{+}19, CG14, CCPG13, PCT12, SW13]. Asymmetry [LHW^{+}19]. Asynchronous [WZW^{+}24]. Attack [DHX^{+}22, LFK19]. Attacks [BCHC19, CWK^{+}24, ERAG^{+}16, PHBC17, SKS23, ZHS^{+}19, BVIB12, CCD12, DJL^{+}12]. Attribution [TMSR23]. Augmented [WZY^{+}24a]. AUKE [DSK19]. Auto [APG^{+}23, CG15a, SAT20, WG17]. Auto-Tuning [CG15a, WG17, APG^{+}23]. Auto-Vectorizing [SAT20]. Automata [EZYA23, LZZ^{+}22, VW11]. automatable [AFD07]. Automated [ASS17, BSSS14, BCHC19]. Automatic [AMG16, DSK19, HEDH21, JLER12, LBO14, LT13, MGA^{+}17, NC15, RB13, TPB24, WLZ^{+}13, WGO15, WM10, XZC^{+}20, SPS12, WKCS12]. Automatically [VZT^{+}20]. Automating [TWB21]. Automotive [FWJ^{+}16]. Autonomously [DGI^{+}14]. Autotuning [AMP^{+}16, CCCA20, SYE19, TIP^{+}23, YAG^{+}16, KBR^{+}13, LFC13]. Autovesk [TPB24]. Availability [OK21]. Avionics [DPBI^{+}19]. AVPP [OAM19]. Aware [APG^{+}23, ACA^{+}19, BB21, DGI^{+}14, CG15a, DTD16, DHX^{+}22, DHD^{+}14, GVT^{+}17, JYW22, KFEG18, KMAK22, LK24, LYH16, LHS^{+}24, LRBG15, MBLK24, MN24, PDCS24, PVA^{+}17, PG17, RJK24, RSK^{+}18, SEF^{+}19, SLJ^{+}18, SJD22, SCK^{+}21, SKH^{+}16, SZJK18, SKPD19, SGM^{+}22, UPR22, USCM16, WLZ^{+}13, WJXC17, WZY^{+}24b, ZPL^{+}21, ZNX^{+}24, ZCQ^{+}19, ZZH^{+}23, ZWY17, ZF24, BOEN23, CPB14, CG14, CHD^{+}23, CLA^{+}19, CWCS13, EE09, GGFPRG12, HAM^{+}20, KABS22, NB13, RB24, SLS^{+}21, SSS^{+}04, SAL19, SL20, SEP07, WYJL10, WSC^{+}13, WDXJ14, YZZ^{+}23, ZYCZ10, ZDC^{+}12, ZK06, JZY^{+}22]. Awareness [HLSW17, LKL^{+}13].

B [WX24]. Backup [LYLS24]. Bahurupi [PM12]. Balancing [LLRC17, PGB16, WWH^{+}16]. Band [SPS17]. Band-Pass [SPS17]. Banded [BSL17]. Bandwidth [LGP^{+}16, LDMZ19, ZCCD16, ZCQ^{+}19, DZC^{+}13, WYJL10, XCC^{+}13]. Bandwidth-Asymmetric [ZCQ^{+}19]. Bank [JFK20, LCL^{+}14, XGD^{+}23]. bank- [LCL^{+}14]. bank-/channel-level [LCL^{+}14]. banked [AGI^{+}12]. Banking [MP22]. Banks [ZCF18]. Bare [RKL23]. Bare-wire [RKL23]. Barrier [CDM^{+}22]. Base [AGG21]. Base-2 [AGG21]. Based [ÄJE^{+}16, CNS^{+}16b, CG15a, CTY^{+}23, CCL^{+}24, CG15b, DSR15, DAD16, DAP^{+}15, DWF^{+}23, ELE^{+}23, EZYA23, FDF^{+}14, GAM12, HYYBM16, JPS17, JHQ23, KS16, LCS^{+}19, LLC22, LK24, LTX16, LLLW22, LWW^{+}24, LY16, LYLS24, MNC^{+}16, MTK18, NC15, RAF22, SBS16, TAB^{+}21, WGO15, WDX15, WCI^{+}16, WWC^{+}16, WMGS19, WDW^{+}22, WLLW20, WZY^{+}24a, WZY^{+}24b, XHJY16, XFS^{+}19, XLD^{+}24, YHYBAM20, ZX19, ZLC^{+}15, ZSM^{+}16, ZGX22, AvRF07, AGG22, BCVT13, CPP08, CWK^{+}24, CW13, GK13, HLR^{+}13, HAJ^{+}12, HWM14, HWX^{+}13, JYJ^{+}13, JFK20, JML^{+}20, JLL^{+}23, JRH21, KPRK20, KBR^{+}13, LBO14, LTG12, LCL^{+}14, LHWB12, MRK^{+}22, MPHL22, OLK^{+}23, OK21, PLK^{+}19, PL23, RLS13, SS04, SKKB18, SSP^{+}23, TKJ13, WSC^{+}13, WTFO14, WWGS22, ZHD^{+}04, ZGC^{+}12, ZFT^{+}18, ZXX23, ZMLL24, ZML^{+}24, WGL^{+}24, SNK^{+}23]. Batch [FLW^{+}24, SAG22]. Batched [JYM20]. Batching [SAG22]. Bayesian [AMP^{+}16, MMLS21]. Be [SW17a]. Behavior [HPBS21, AFD07, LS10]. Benchmark [ABB^{+}16, AYL^{+}18, CCM^{+}16, DDT^{+}17, DS16, BE13]. Benchmarking [DAP^{+}15, XZC^{+}20]. benchmarks [JEBJ08]. Benefits [LWWH12]. Benzene [KAC^{+}18]. BestSF [BJWS18]. Better [ZXX23, TBC^{+}12]. Between [EPS17, NMPS22]. Beyond [AGG21, FER^{+}13, LCP^{+}21]. Bias [Lee16]. Big [ZLYW18, ZLC^{+}15]. Big-Memory [ZLC^{+}15]. Bimodal [TD16]. Binary [DGGL16, GDL16, HWL^{+}19, LHW^{+}19, RKC^{+}20, SHY14, XYY^{+}24, CDM13, GHS12, HS06, HLC10, LWH11, PKC12]. bipartite [BZS13]. Bit [TBS06, BWLR06, VED07]. Bit-split [TBS06]. BitSAD [DZSL20]. Bitstream [DZSL20]. bitwidth [NB13]. bitwidth-aware [NB13]. Blaze [PWPD19]. Blaze-Tasks [PWPD19]. Block [GFD^{+}14, HAM^{+}20, KTAE16, LLRC17, LTX16, MPPS18, MK23, TZK18, TAB^{+}21, ZK06]. Block-aware [ZK06]. Blocked [BHWN21]. Blocking [JHQ23, IPSD21]. Blocks [HWJ^{+}15, SYX^{+}15]. body [WPR^{+}22]. Boltzmann [PAVB15]. Bones [NC15]. Boost [KABS22]. Boosting [ASV^{+}16, KH18, RLS13, BTS10]. both [BSWLE13, HP04, MP13]. bottlenecks [MMdS06]. bound [CLA^{+}19, MBKM12]. bounded [HS06]. Bounding [XMM04]. Bounds [ESR^{+}15, BWLR06, JRH21]. BPM [LCL^{+}14]. BPM/BPM [LCL^{+}14]. Brain [vdVSAAS20]. Brain-Simulation [vdVSAAS20]. Branch [EPAG16, LIS20, LWL18, Mic18, CZ07, HWH^{+}11, Jim09, JSM^{+}04, LBJ05, MG12, TS05]. branch-predictor [JSM^{+}04]. branch-target [LBJ05]. Branches [DGGL16]. Breakdown [HYYBM16]. bridging [HCC^{+}14]. Bringing [DDT^{+}17]. buddy [KWCL09, ZJJ^{+}15]. Budget [LWF^{+}16]. Buffer [LDY^{+}21, LSH^{+}23, SLH^{+}20, SL20, LBJ05, RB13]. Buffering [YMM^{+}15, GPL^{+}05]. Bug [WZY^{+}24a, WZY^{+}24b]. Bugs [AAI^{+}16]. build [SSH^{+}13]. Building [KRHK16, SGS^{+}20, WDX15, XZW^{+}22]. BullsEye [SMM^{+}23]. Buri [ZLC^{+}15]. Burst [SLH^{+}20]. Byte [TDS^{+}21]. Byte-Select [TDS^{+}21].

C [CWW^{+}16, NC15, NED^{+}13, ZZB^{+}19]. C-to-CUDA [NC15]. C/C [NED^{+}13]. C1C [LZL^{+}13]. CACF [ZFT^{+}18]. Cache [CKPH19, CS21, CAGS17, DAD16, DVG^{+}23, GFD^{+}14, HK14, HMYZ15, KR19, KAC^{+}18, KAC15, KDMA23, LLRC17, LWS^{+}19, Mic16, MK23, PLK^{+}19, RGK^{+}23, RS21, SSW16, SBS16, SMM^{+}23, SKH^{+}16, SSP^{+}23, SDS^{+}21, SJC^{+}21, SWO21, SLJ^{+}19, TWB21, VPTS19, WSJ^{+}21, WJXC17, YDL^{+}17, ZWY17, ZWL^{+}19, APG13, AGVO05, AGI^{+}12, AFD07, BSWLE13, CA11, CWS06, DJL^{+}12, FTLG11, GGFPRG12, GSZI10, HAJ^{+}12, KS11, KWCL09, LCC11, LZL^{+}13, LHC^{+}24, MMdS06, RFD13, SS04, SBC05, SSH^{+}13, TKJ13, VSP^{+}12, WSC^{+}13, WDXJ14, ZHD^{+}04, ZVYN05, Zha08, NTG13]. cache-coherence [MMdS06]. cache-coherent [APG13]. cache-content-duplication [KS11]. cache-resident [LHC^{+}24]. Cache-Ways [CS21]. CacheInspector [SDS^{+}21]. Caches [CAGS17, CPS^{+}15, GBD^{+}15, JPS17, SBS16, WDX14, AIVL13, DJL^{+}12, HS06, HL07, KS11, KWCL09, LJMG12, MSK05, SSK11, SSC^{+}13, VSP^{+}12, WDXJ14, WLZ^{+}10, WM11, ZDC^{+}12]. Caching [DNT16, LWW^{+}24, SYX^{+}15, DZC^{+}13, JOA^{+}09a, WFKL10]. CACTI [BKM^{+}17]. Caffe [RSK^{+}18]. CAFFEINE [PB15]. CAIRO [HNKK17]. Calculation [SMM^{+}23, YSH^{+}22]. Caliper [KLA^{+}19]. Call [HZN^{+}22, Lee16, MG12]. Call-Site [HZN^{+}22]. Calls [DKK^{+}21]. Camouflage [PDCS24]. Capability [AHA^{+}19, DGI^{+}14]. Capacity [GBD^{+}15, SSK11, WM11]. Captioning [HHW^{+}22]. Captures [MPHL22]. Capturing [XDXL19]. CARL [DCL^{+}22]. CART [CDPD13, CDPD13]. Case [ANS^{+}22, KH18, MMS15, MKC^{+}22, SKAEG16, SSRS15, AFD12, RPS06, WK09, LB10]. CASHT [GAH22]. CATCH [KS11]. Caused [SYX^{+}15]. CAVA [CST^{+}06]. CC [CCZ13]. Cell [SSP^{+}23, YMM^{+}15, STLM12]. Cell-based [SSP^{+}23]. cells [JSM^{+}04]. Center [FXC^{+}15, JYW22]. centers [AVG12]. Centralized [LDY^{+}21]. Centric [CKP^{+}22, JLJ^{+}18a, SJL^{+}20, XGD^{+}23]. Cerberus [HBPH24]. CERE [DAP^{+}15]. CG [MAD17]. CG-OoO [MAD17]. CGRA [HAC13]. CGRAs [CPG21]. Chain [FPMR21]. chains [SSH^{+}13]. Chameleon [WFKL10]. Change [HASA16, JDZ^{+}13, NTV^{+}22, YMM^{+}15, ZDC^{+}12]. Channel [BCHC19, BVIB12, Bis21, DJL^{+}12, JFK20, LFK19]. channel-level [LCL^{+}14]. Channels [DJC16, EPAG16, JHQ23]. chaotic [LTG12]. Characterization [CVB15, HKA^{+}19, HPBS21, DS12, FER^{+}13, VW11]. Characterizing [BCM11, LCW^{+}24, ZNTJE23]. Check [JRH21]. Checking [KK15, BWLR06, MG13, MFSK24]. Checkpoint [GW09, ARS04, CST^{+}06]. checkpoint-assisted [CST^{+}06]. Checkpointing [AEE^{+}19, WZG^{+}19, DXMJ11]. Chief [Kae20]. Chip [BKM^{+}17, CS21, CCL^{+}24, CPS^{+}15, CEP^{+}16, DJC16, EPS18, LBM13, VFW16, ZNTJE23, APG13, BKA13, CK11, EE11, GSZI10, JPS17, LWWH12, LT13, LNLK13, LAS^{+}08, LM05, LPZI12, LDL22, LMMM08, SSH19, SMK10, TDG13, XCC^{+}13]. Chiplet [CWK^{+}24]. Chips [LCS^{+}19, ZM15]. choices [VE13]. Chunking [MG20]. CIB [LDY^{+}21]. CIB-HIER [LDY^{+}21]. CIM [JLL^{+}23]. Circuit [ZFT^{+}18, DJX13]. circuit-architecture [DJX13]. Circuits [KKW^{+}15]. Circuits/Cores [KKW^{+}15]. Citadel [NRQ16a]. Class [AAI^{+}16, PAVB15]. Classes [JHQ23]. Classification [DRHK15, MCB^{+}12, SNN^{+}19, CDPD13, LMJ13a, NCC13]. client [KWM^{+}08]. Climate [GMZ^{+}21]. Clock [CCL^{+}13]. Closer [HZN^{+}22]. Cloud [CSW^{+}23, CKL^{+}24, QYZ^{+}14, RB24, WLX24, XZW^{+}22, XZC^{+}20]. Cloud-scale [CKL^{+}24]. Clouds [SDS^{+}21]. Cluster [SKKB18, YCA18, ZNX^{+}24, TC07]. Cluster-Aware [ZNX^{+}24]. Clustered [LZM14, MMS15, ACGK04, SW13]. Clustering [MNC^{+}16, WMGS19, DS12, JLCR13, SB09]. Clustering-Based [MNC^{+}16, WMGS19]. Clusters [KHS^{+}14, MMS15, ZF24]. CMP [CPB^{+}07, LMCV13, SSK11, SLJ^{+}18, WM11]. CMPs [ABK21, LMJ13a, LY16]. CNN [CSRP22, JHH^{+}23, WWY^{+}24, XLH^{+}24]. CNNs [JML^{+}20, QAT24]. Co [AHA^{+}19, JPS17, KHN^{+}18, LZW23, SOAK23, ZFT^{+}18, ZPH^{+}23, DJX13, YLW08]. Co-iteration [ZPH^{+}23]. Co-located [LZW23, SOAK23]. Co-location [KHN^{+}18, YLW08]. Co-optimization [JPS17, ZFT^{+}18, DJX13]. Co-Processor [AHA^{+}19]. coalescing [SSU^{+}13]. coalescing-lowering [SSU^{+}13]. Coarse [LMSE18, MAD17, TD16, KCP13]. Coarse-Grain [LMSE18, MAD17]. Coarse-Grained [TD16, KCP13]. Coarsening [SF18]. COBAYN [AMP^{+}16]. CODA [KHN^{+}18]. Code [DKK^{+}21, CZ07, DSK19, HZN^{+}22, CDM^{+}22, JCG^{+}24, KL19, PAVB15, PKPM19, SYE19, TMSR23, TPB24, WZY^{+}24a, WZY^{+}24b, ZPH^{+}23, AvRF07, CDM13, GNB08, HLR^{+}13, HS06, JLER12, KBR^{+}13, LKL^{+}13, LBJ05, LZYZ09, LHY^{+}06, PKC12, RCG^{+}10b, VJC^{+}13, ZK05, ZWHM05]. Code-level [TMSR23]. code-positioning [ZWHM05]. coded [ZF24]. Codelet [DAP^{+}15]. Codes [CWMC16, LCW^{+}23, TZK18, AFD07, AFD12]. Codesign [KCA^{+}13, WDX24]. Codesigned [KMG14]. Coding [PM17, ZNX^{+}24]. COER [WDX24]. Coherence [ANS^{+}22, CWK^{+}24, DRHK15, KBB^{+}14, KAC15, PL23, MMdS06, SSH^{+}13, VHKP11]. coherent [APG13]. Collaborative [LLLW22, FT10]. collapse [CWCS13]. Collection [ASV^{+}16]. Collective [FT10]. collector [WK09]. colocated [DWDS13]. Colocation [LSL20, PLLW24]. Colony [SGM^{+}22]. Coloring [YWXW12, LFX09]. Column [ZBC^{+}22]. Column-wise [ZBC^{+}22]. Combination [LDMZ19]. Combinatorial [SKPD19, SSR13]. combined [BWG^{+}12]. Combining [VSP^{+}12, YRGES^{+}19]. CoMeT [SKP^{+}22]. Commodity [GWZ22, WDX15]. common [WK09]. Communication [DSR15, HAM17, TN20, XDXL19, XDW^{+}23, HWX^{+}13, SSPL^{+}13, TC07]. communications [ACGK04]. Comp [DZL^{+}24]. Compact [HEMK17, QAT24, SHC13]. Compaction [DZL^{+}24, WK09]. Comparability [YWXW12]. Comparative [LAS^{+}08]. Comparators [YEI^{+}14]. comparison [FBWS13]. CompEx [PM17]. Compilation [DMR^{+}16, HZN^{+}22, JLL^{+}23, KVH23, LT19, LLLW22, LRBG15, PKPM19, RVKP19, SYE19, SN17, ZC20, CI13, JK13, KHL^{+}13, LBO14, LZYZ09, PC13]. Compile [KTAE16]. Compile-Time [KTAE16]. compiled [NED^{+}13]. Compiler [AMP^{+}16, ABP^{+}17, BKS^{+}22, CGP23, CCD12, DZSL20, DCL^{+}22, DMG13, EAH^{+}20, EPS17, GGK18, GMZ^{+}21, HNKK17, HYBR^{+}15, JRH21, KPRK20, KPP^{+}15, LFX09, MNC^{+}16, MG12, MPHL22, NKH16, NC15, PHBC17, FCVR24, ZSCM08, ZX16, CYXF13, DC07, HWM14, HLC10, JOA^{+}09a, JOA^{+}09b, KBR^{+}13, KWM^{+}08, LZL^{+}13, LCH^{+}04, TR13, YXK^{+}12, ZHD^{+}04]. Compiler-Assisted [HNKK17, PHBC17]. Compiler-based [JRH21, KPRK20, ZHD^{+}04]. Compiler-Directed [HYBR^{+}15, LFX09]. compiler-guided [LZL^{+}13]. Compiler-Oriented [GGK18]. Compiler-support [EAH^{+}20]. Compiler/Runtime [KPP^{+}15]. Compilers [PBCB22, SAT20, CDM13, HEL^{+}09, SD12]. Compiling [LSH^{+}23]. Complex [SSS^{+}24, SHD15, vdVSAAS20, SLA^{+}07]. Complexities [GHH15, ZBH^{+}13]. Complexity [GG18, KAC15, LNFE22, CPP08, DJL^{+}12, RPS06, SRLPV04]. complexity-effective [RPS06]. component [LGAZ07]. Composable [FBC^{+}22]. Comprehensive [CPS^{+}15, HKA^{+}19]. Compressed [SSW16, ZML^{+}24, DZC^{+}13]. Compression [BC13, KPM17, LMSE18, PM17, RS21, SW17a, TDS^{+}21, WYZ^{+}23, KGK10]. Compression-Expansion [PM17]. Compression/Decompression [LMSE18]. Compressive [WCI^{+}16]. Compromising [Bis21]. Computation [CWW^{+}16, DJZ^{+}23, HAM17, JHHM21, JLL^{+}23, KHN^{+}18, VZT^{+}20, WCW^{+}24, DDU12, LFC13]. Computational [CRC^{+}21]. Computationally [DSH^{+}18]. Computations [BKS^{+}22, PAVB15, SQZK20, SHS^{+}20, CYXF13]. Compute [DAKK19]. Computers [LD24]. Computing [BSG23, DZSL20, DSH^{+}18, KHS^{+}14, LCS^{+}19, Lou19, ME17, MPU^{+}23, PKL^{+}24, PWPD19, SW17b, SPS23, TCS16, XZC^{+}20, XGD^{+}23, ZLYW18, ZLC^{+}15, AVG12, LM05]. conceived [APG13]. Concise [WX24]. Concurrency [AAI^{+}16, GMGZP14, ME17, MLB^{+}23]. Concurrent [LDMZ19, MKS22, PCM16, WX24, ZZH^{+}23]. Conditional [Mic18]. conditionals [JSL13]. Configurable [NRQ16b, TGRK21, XMW^{+}21, HVJ06, LZL^{+}13]. Configuration [LL22]. Conflict [JFK20, WZZ^{+}20]. Conflict-based [JFK20]. Conflict-free [WZZ^{+}20]. conflicts [TGAG^{+}12]. Congestion [WDX24, YRGES^{+}19]. connected [BRSJG12]. conscious [LZYZ09]. Conserving [LYYB07]. Considerations [HMYZ15, MTK18, LM05]. considering [AVG12, HP04]. Consistency [HZN^{+}22, LLW^{+}22, NZ15]. Constrained [LZM14, NMPS22, MSF^{+}07, NMKS06, ZK05]. Constraint [RAF22]. Constraints [AEJE16, APS22, AMS23, CSF^{+}20, KCA^{+}13, WYJL10]. Construction [DPBI^{+}19]. Consumer [LYLS24]. Consumption [BNS^{+}21, CS21, FCD^{+}17, GFD^{+}14, LTG12, LYYB07, VED07, ZHD^{+}04]. Contech [RHC15]. Content [LWW^{+}24, KS11]. Contention [DHX^{+}22, GWZ22, GAH22, KMAK22, LLC22, LZW23, CWCS13]. Contention-Free [LLC22]. Context [EPS17, DMG13, LS10]. continual [JA14]. Continuous [TR13]. Continuum [RB24].

Contraction [ZPH^{+}23]. Control [AP17, BRJM15, HAC13, HHC^{+}16, SMKH15, SKH^{+}16, WDX24, YRGES^{+}19, CWC06, FSYA09, IWP^{+}04, MBKM12, TG07]. Control-Flow [SMKH15]. Controlled [ASS17, NTV^{+}22, RCV^{+}05]. controller [AGI^{+}12]. Conventional [NRQ16b]. conversion [CS13]. Converting [HLC10]. Convolution [ADGA20, FSP^{+}23, KDMA23, LCP^{+}21, WLJ^{+}24, FBWS13]. Convolutional [GG18, GLTV23, JHHM21, RAF22, SMN22, TDP15, XMW^{+}21, XSF^{+}23, ZFF^{+}18]. Convolutions [TIP^{+}23]. CoolDC [MBLK24]. Cooled [MBLK24]. cooling [AVG12]. cooling-computing [AVG12]. Cooperation [TZK18]. Cooperative [Abd20, DNT16, JPS17, JDZ^{+}13, LBM13, NMPS22, SHLM14]. Coordinated [LDMZ19, ZDC^{+}16]. Copious [DVG^{+}23]. Coprocessor [MPU^{+}23, LDG^{+}13]. Corasick [CW13, PLL10]. Core [CHE^{+}14, CS21, CC18, FMY^{+}15, JLJ^{+}18a, LNFE22, LK24, LBM13, PVS^{+}17, SPS17, SPH^{+}17, TGRK21, ZLYZ16, DAKK19, FSX^{+}24, JYM20, JHH^{+}23, LNLK13, MLC^{+}23, MRK^{+}22, OGK^{+}12, PM12, QSZ^{+}21, SSS^{+}23, ZGC^{+}12]. Cores [CAY^{+}18, DT17, HYYBM16, JPS17, KKW^{+}15, KABS22, MMS15, TCS20, TDO16b, ZCF18, GB06, NTG13, PCT12, SW13, WYJL10, WFKL10]. CoreUnfolding [APBR16]. Corner [DDT^{+}17]. Correct [DPBI^{+}19]. Correct-by-Construction [DPBI^{+}19]. Correcting [SPM17, TZK18]. Correction [DGI^{+}14, CWMC16, Lee16, LSC^{+}15, LDC15]. Correctness [PD17]. correlating [TKJ13]. Correlation [WPR^{+}22]. coscheduling [PGB13]. Cost [KBB^{+}14, LGP^{+}16, MBLK24, MN24, RB24, SSW16, SKPD19, YEI^{+}14, AGI^{+}12, DC07, FBHN04, MA08, SBC^{+}22]. Cost-aware [RB24]. Cost-Effective [MBLK24]. COTS [RGG^{+}12]. Could [SW17a, ZPR^{+}17]. Counter [WCI^{+}16]. Counter-Based [WCI^{+}16]. Countermeasures [CWK^{+}24]. Counters [LGH24, NDP17, RLS13]. Counting [WZX^{+}24, RBM10]. Coupled [Abd20, PCT12]. covering [PJ13]. Covert [EPAG16]. COWS [MN24]. COX [HLSK22]. CPU [Abd20, BSSS14, LMCV13, PGB16, WLWB19]. CPUs [BHC^{+}16, HLSK22, WLJ^{+}24]. Crash [LLW^{+}22]. Creating [CPG21]. Creation [THA^{+}21]. Critical [EAH^{+}20, LYLS24, RGG^{+}12]. Criticality [FWJ^{+}16]. CRNS [AS13]. Cross [ERAG^{+}16, FSX^{+}24, LGAZ07, LGH24, LVR^{+}15, OTR^{+}18, SWF16, TWB21, VOK^{+}22, WAST16, ZLYZ16]. Cross-accelerator [TWB21]. Cross-Architecture [SWF16]. Cross-component [LGAZ07]. Cross-core [FSX^{+}24]. Cross-Layer [ERAG^{+}16, OTR^{+}18, WAST16, VOK^{+}22]. Cross-Loop [LVR^{+}15]. Cross-Platform [ZLYZ16, LGH24]. Crossbar [XLD^{+}24]. Crowd [MPHL22]. Crowd-sourced [MPHL22]. Crown [MKKE15]. Cryptographic [Bis21]. cryptography [AS13]. CTA [LDMZ19, UPR22]. CUDA [HLSK22, KBR^{+}13, NC15, VJC^{+}13, WG17]. Customized [CPG21]. CXL [WGL^{+}24]. cycle [DEE13, RLS13].

D [LLC22, LD24, BSL17, CAY^{+}18, CWMC16, DZL^{+}24, DVG^{+}23, JHH^{+}23, LGP^{+}16, LCW^{+}24, NRQ16b, SZJK18, SKP^{+}22, ZSLX13]. D-CNN [JHH^{+}23]. d-Packed [BSL17]. D-Stacked [DVG^{+}23, LGP^{+}16, NRQ16b]. D/ [LLC22, LD24]. DAC [RJK24]. DAC/ADC [RJK24]. DAC/ADC-less [RJK24]. DAG [CCL^{+}24]. DAG-Order [CCL^{+}24]. DAPSCO [GGFPRG12]. dark [PCT12]. DarkCache [ZCF18]. DASH [USCM16]. Data [ASH20, ACGH24, AMG16, CKP^{+}22, CDPN16, DZL^{+}24, DAKK19, EPS18, ESR^{+}15, EAH^{+}20, FSX^{+}24, FXC^{+}15, GAM12, GLTV23, HAM17, HAM19, HLSW17, IPSD21, JYW22, JLJ^{+}18a, KPM17, KHN^{+}18, LWL18, LLS23, LYLS24, MST^{+}21, ME15, ME17, MTK18, MNSC16, MGA^{+}17, MGSH16, NKH16, NSF^{+}21, PD17, RMA14, RTK15, SKH^{+}16, SJL^{+}20, SJC^{+}21, SWO21, TMSR23, TDP15, VFJ^{+}17, WGO15, WZG^{+}19, WZY^{+}24b, YXS^{+}22, YMM^{+}15, ZLYW18, ZNTJE23, ZLWS24, AVG12, AGG22, BSWLE13, CS10, CA11, CDPD13, CWC06, FER^{+}13, FLG12, HLR^{+}13, HL07, LWH11, LJMG12, PC13, RB13, RFD13, STLM12, TG07]. Data-Centric [CKP^{+}22]. Data-Driven [ME15, ME17, ASH20]. data-flow [PC13]. Data-Parallel [MGSH16, NKH16]. Data-Race-Free [MNSC16]. Data-Rate [EPS18]. Data-Traversal [RMA14]. Database [BAZ^{+}19]. Datacenter [HZC^{+}23, MBLK24]. Datacenters [CSW^{+}23, XVT20, ZFL18]. Dataflow [CPG21, DT17, FLW^{+}24, KPP^{+}15, MMT^{+}12, ZMLL24, VTN13]. Datapath [IWP^{+}04]. Datasets [WLWB19]. DawnCC [MGA^{+}17]. DCMI [KFJ20]. DCNNs [ESB^{+}20]. DDR4 [TKM14]. DDRNoC [EPS18]. Dead [MPPS18]. Dead-Block [MPPS18]. Deadline [LZM14, USCM16]. Deadline-Aware [USCM16]. Deadline-Constrained [LZM14]. deadlock [BRSJG12]. deadlock-free [BRSJG12]. debugging [VDSP09]. decay [JSM^{+}04, SS04]. Declarative [CZGC20]. decoders [Zha08]. Decoding [CAMJ15, LCW^{+}23]. Decomposition [BHWN21]. Decompression [LMSE18]. Deconstructing [CFH^{+}12]. Decoupled [JHH^{+}23, MPU^{+}23, VPTS19, BZS13, DHC^{+}13, RVOA08]. Decoupling [HAM17]. Decreasing [SWO21]. Dedicated [ICEDR24]. Deep [ASK^{+}16, CKP^{+}22, DLS22, JLJ^{+}18a, MWJ19, QAT24, RSK^{+}18, WWW^{+}21, WZWW23, XDXL19, XDW^{+}23]. Deeply [GKCE17]. DeF [CPG21]. DEFCAM [LCC11]. defect [LCC11]. defect-tolerant [LCC11]. Defined [DMR^{+}16, TGAG^{+}12]. Defragmentation [PVS^{+}17]. DeFT [VHKP11]. Delay [SKS23]. Delay-on-Squash [SKS23]. Delayed [DQCK24]. Delivery [LWW^{+}24, ZZL^{+}21]. Delta [DZC^{+}13]. Delta-compressed [DZC^{+}13]. Demand [ACG24, BRJM15]. Dense [ACGH24, CWW^{+}16]. Dependence [BRJM15, DHD^{+}14, JK17, KABS22, SL09, TG07, VTN13]. Dependence-Aware [DHD^{+}14, KABS22]. dependences [BCVT13]. Dependency [WLZ^{+}13, ZPL^{+}21, ZZH^{+}23]. Dependency-Aware [WLZ^{+}13, ZPL^{+}21, ZZH^{+}23]. dependent [YZL^{+}10]. Deployments [vdVSAAS20]. depth [HP04]. Design [CSF^{+}20, CKPH19, CPS^{+}15, ESB^{+}20, KWM^{+}08, LDY^{+}21, MAY23, RTK15, SZJK18, SPH^{+}17, SL09, UJW15, VHKP11, VKM^{+}21, WLZ^{+}10, WZY^{+}24a, WZY^{+}24b, BE13, CPP08, IMS^{+}08, LB10, LCC11, LHZ13, VE13, ZK05]. Designing [BKA13, BSWLE13, MGSH16]. Destructive [DQCK24]. Details [FMY^{+}15]. Detecting [DSR15, KS11]. Detection [BDB^{+}20, CEP^{+}16, LHC^{+}17, MNSC16, SLH^{+}20, WCI^{+}16, WDW^{+}22, WWY^{+}24, YEI^{+}14, LKL^{+}13, TBS06, TDG13, VHKP11, WTFO14]. Deterministic [CCL^{+}13, MLB^{+}23, VSDL16, VW11]. Detonation [CAY^{+}18]. Devectorization [KMG14]. Development [VCJ^{+}17]. Device [MSFC21, RLBBN15]. Device-Level [RLBBN15]. Devices [LYLS24, TKM14, NMKS06, ZK05]. DFA [BC13]. Diagnosing [JLJ^{+}18b]. diagnosis [BSO07]. DiagSim [JLJ^{+}18b]. Die-Stacked [CWMC16]. die-stacking [ZSLX13]. different [YXK^{+}12]. Dimension [LLC22, RTG^{+}07]. Dimension-Order [LLC22]. dimensional [LT19, LL22]. Direct [DKB^{+}20, FSP^{+}23, LLRC17, YRGES^{+}19]. Direct-Mapped [LLRC17]. Directed [HYBR^{+}15, VZS^{+}18, LFX09, NED^{+}13, SEP07, WM10]. directional [XMW^{+}21]. directives [CXW^{+}12]. Directories [PT17]. Dirty [LLRC17]. Dirty-Block [LLRC17]. Disaggregated [DZL^{+}24, DWF^{+}23, HZC^{+}23, LZS^{+}24]. Disaggregation [MKC^{+}22, WGL^{+}24]. discard [LWWH12]. Discovering [YHYBAM20]. Discrete [ZSM^{+}16]. DisIRer [HLC10]. Disjoint [SJA12]. Disk [LYK^{+}15, WDW^{+}22]. disparate [WLZ^{+}10]. Dispatch [LLRC17]. dispatching [LZ12]. dissemination [LZYZ09]. Distance [DAD16, GGFPRG12, KR19, LK24, SCMU22, FER^{+}13, FTLG11]. Distance-Aware [LK24, GGFPRG12]. Distance-Based [DAD16]. Distilling [JEBJ08]. Distinguished [Aca16, Ano15, Bil19, Ano13a]. distribute [RFD13]. Distributed [JZY^{+}22, KHS^{+}14, KAC^{+}18, LZS^{+}24, MMGS21, MK23, SSS^{+}23, TPN^{+}20, XDXL19, XDW^{+}23, ZPC06]. Distrusting [SOAK23]. Divergence [LWL18, SMKH15]. Divergent [GR15]. Diverse [LP17, SAL19]. diversification [CDM13]. Diversity [TDO16b, KNBK12]. DJ [DDU12]. DJ-graphs [DDU12]. DL [THA^{+}21]. DLP [SNL^{+}04]. DNN [LY24, LWC^{+}22, WYZ^{+}23, XZC^{+}20, XLD^{+}24]. DNNs [ZML^{+}24]. DNNTune [XZC^{+}20]. Do [ZPR^{+}17]. Document [HKA^{+}19]. Doesn’t [LKV12]. Domain [CZGC20, CGP23, FBC^{+}22, GáSÁ^{+}16, GMZ^{+}21, GÁSÁ^{+}13]. Domain-Specific [GMZ^{+}21, CZGC20]. Domains [SW17a]. DPCS [GBD^{+}15]. DPM [GK13]. Dragonfly [CVB15]. DRAM [CKPH19, CAGS17, DQCK24, HCC^{+}14, JLCR13, LLRC17, LCL^{+}14, NCQ14, OLK^{+}23, OBdOJ^{+}24, OTR^{+}18, TKM14, VPTS19, XGD^{+}23, XHJY16]. DRAMCache [PG17]. DRAMs [LSC^{+}15]. Drift [SZJK18]. Drive [MPHL22]. Driven [ME15, ME17, PB15, YXMC23, ZWS^{+}16, ASH20, CDM13, FTLG11, PSL^{+}23, SLP08, WTFO14, XT09, ZCS06]. Droplet [CRBP24]. Dropping [CNAA^{+}22, GFD^{+}14]. DSL [PBY^{+}17]. DSM [JZY^{+}22]. DSM-aware [JZY^{+}22]. DSPs [VCJ^{+}17]. Dual [EPS18, WZG^{+}19]. Dual-Page [WZG^{+}19]. DUCATI [JED19]. duplication [KS11, LKL^{+}13]. DVFS [CS21, EE11, GK13]. DxPU [HZC^{+}23]. DynamAP [LZZ^{+}22]. Dynamic [BHC^{+}16, CCL^{+}24, CCCA20, DGGL16, DD16, DJB13, FER^{+}13, FTLG11, FSYA09, GAM12, GDL16, GBD^{+}15, HWL^{+}19, KE15, KPP^{+}15, KMG14, KKAR16, LKL^{+}13, Lee16, LPZI12, LSL20, LTX16, LHW^{+}19, LZZ^{+}22, MG19, MG20, RHC15, SV05, SGS^{+}20, SHD15, WWH^{+}16, XYY^{+}24, XHJY16, YZZ^{+}23, ZWY17, BBG13, DWDS13, GHS12, HS06, HWH^{+}11, HVJ06, JSH09, LWH11, LJMG12, LCL^{+}14, MG12, NED^{+}13, WSC^{+}13, XMM04, ZZQ^{+}05]. Dynamically [HDW21, LZ12, PGB12, KS11].

E-BATCH [SAG22]. eager [JLCR13].

Early [AKBS21, ZZL^{+}21, JOA^{+}09b, SLP08]. Early-stage [ZZL^{+}21]. Earth [GMZ^{+}21]. Easier [TIP^{+}23]. Easy [TDG13]. ECC [CWMC16]. ECCs [ZWL^{+}19]. Echo [CSF^{+}20]. EchoBay [CSF^{+}20]. ECS [SPM17]. Edge [LK24, RB24, SCK^{+}21, SAM^{+}23]. Edge-Cloud [RB24]. Editor [Kae20]. Editor-in-Chief [Kae20]. Editorial [CT08, Kae20]. EECache [CPS^{+}15]. Effective [GMGZP14, HVJ06, KH18, MBLK24, PGB16, SSW16, SDK^{+}22, SPS17, KHW^{+}05, LWH11, RPS06, SBC05]. Effectiveness [JRK16]. Effects [DRHK15, DVG^{+}23, MGI15, CK11]. Efficiency [AJK^{+}12, CAMJ15, CSK19, DJZ^{+}23, GBD21, HLSW17, JHHM21, LMSE18, LAAMJ15, OTR^{+}18, OAM19, SL20, SCFD22, TCS16, ZJJ^{+}15, BSWLE13, CWS06, RCG^{+}10a, ZSLX13]. Efficient [AYC16, AEE^{+}19, AKBS21, BSG23, BC13, CC13, CMAP22, CPS^{+}15, DDU12, DKB^{+}20, DZL^{+}24, DD16, GáSÁ^{+}16, GNB08, HAM19, HAC13, HEDH21, HEMK17, HFL^{+}23, IMS^{+}08, JYM20, KR19, KAC^{+}18, KH18, KMG14, LLW^{+}22, LWH11, LWS^{+}19, LL22, LDC15, LGH24, MCB^{+}12, MG19, MMLS21, MKKE15, MPU^{+}23, MAD17, MFSK24, NMKS06, NSF^{+}21, OBdOJ^{+}24, PDY^{+}23, PLLW24, PS15, QAT24, SYZZ^{+}14, SN17, SAG22, TDP15, TTS19, WLJ^{+}24, WZG^{+}19, XZW^{+}22, XHL^{+}24, YMM^{+}15, YZZ^{+}23, ZPC06, ZHS^{+}19, ZPL^{+}21, ZZH^{+}23, ZLM^{+}23, ZML^{+}24, ZLJ18, ZZQ^{+}05, APG13, ARS04, CW13, CNAA^{+}22, CWCS13, DLS22, DCP^{+}12, FSX^{+}24, FPK^{+}24, GW08, JSL13, JOA^{+}09a, KHW^{+}05, LZYZ09, LMJ13a, LHZ13, Nas13, PLL10, RFD13, SPGE06, SHC13, SSP^{+}23, SB09, TDG13, WWL^{+}21, WYZ^{+}23, XCC^{+}13, YSH^{+}22, ZGC^{+}12, FSYA09, SLA^{+}07]. Efficiently [NRQ16a, PCT12, RHC15, ZWL^{+}19]. EFGR [TKM14]. Elastic [Per18]. Element [LVR^{+}15]. elementary [LDG^{+}13]. Eliminating [RCG^{+}10b, SWO21]. elimination [JLER12, VED07]. Elision [CDM^{+}22]. Embedded [GTT^{+}16, GKCE17, KE15, KTAE16, CPP08, CDM13, GHS12, MP13, SHC13, SD12, XT09]. embedding [KKM^{+}13]. emergencies [RCG^{+}10b]. emerging [DXMJ11, XCC^{+}13]. empirical [AvRF07]. Emulation [NZ15, TKKM15]. Emulators [HHC^{+}16, TKKM15]. Enable [RAF22]. Enabled [TGRK21, LZS^{+}24]. Enabling [BGG^{+}15, CC18, HNKK17, JYM20, KHN^{+}18, SKAEG16]. Enclaves [PKL^{+}24, YGB21]. Encoding [DNS^{+}24, LCW^{+}23, TDP15, ZX19]. Encoding/Decoding [LCW^{+}23]. Encryption [LIS20]. End [FPMR21, OLK^{+}23, ZJJ^{+}15]. End-to-End [FPMR21, OLK^{+}23]. Endurance [WDXJ14]. Endurance-aware [WDXJ14]. Energy [ABK21, AJK^{+}12, AYC16, ASP17, APS22, AMS23, BSG23, CMAP22, CPS^{+}15, CNAA^{+}22, DLS22, DH16, FSX^{+}24, FPK^{+}24, GKCE17, GFD^{+}14, HMYZ15, JHHM21, JOA^{+}09a, KAC^{+}18, LMSE18, LSC^{+}15, LMA^{+}16, MCB^{+}12, MTK18, MKKE15, MAD17, MPW^{+}17, NMPS22, OBdOJ^{+}24, OTR^{+}18, PM17, RTK15, SW17b, SN17, SAG22, SSP^{+}23, SB09, TCS16, TTS19, YSH^{+}22, ZJJ^{+}15, ZFT^{+}18, ZCF18, AVG12, BSWLE13, CWS06, CWCS13, FBWS13, GWS13, GKP14, LTG12, LGAZ07, LZYZ09, LMJ^{+}13b, LHZ13, SPGE06, SHC13, SAM^{+}23, TDG13, ZHD^{+}04, ZVYN05, ZGC^{+}12, ZSLX13]. Energy- [SB09]. Energy-Efficient [AYC16, CPS^{+}15, KAC^{+}18, MKKE15, MAD17, OBdOJ^{+}24, SN17, SAG22, TTS19, CNAA^{+}22, DLS22, FSX^{+}24, FPK^{+}24, JOA^{+}09a, SSP^{+}23, YSH^{+}22, CWCS13, LZYZ09, LHZ13, SPGE06, SHC13, TDG13, ZGC^{+}12]. Energy-Optimal [SW17b]. Energy-Performance [MTK18, ZCF18]. Energy-Proportional [DH16]. Enforcement [AHA^{+}19, GWM07]. Engine [HKA^{+}19, HLZ^{+}24, LP17, PB15, RMA14, WLZ^{+}13, CW13]. Engineering [SDS^{+}21]. Engines [MGI15, TBS06]. Enhance [GAM12]. Enhanced [GBD21, TKM14, TCR^{+}22, FYI^{+}24]. Entropy [LZW23]. enumeration [SWH09]. Envelope [RWFJ19]. Environment [KMG14]. Environments [KLA^{+}19, SMS23, WLX24, RGG^{+}12, WWWL13]. EOLE [EPS17]. Equality [KS21]. Era [GBD^{+}15, SSS^{+}23, LNLK13, PCT12]. ERASE [CMAP22]. Erasure [LCW^{+}23, ZNX^{+}24, ZF24]. Erasure-coded [ZF24]. Error [BDB^{+}20, DGI^{+}14, CWMC16, DSH^{+}18, LSC^{+}15, OK21, SPM17, TZK18, WWY^{+}24, YEI^{+}14, CCZ13, LKL^{+}13]. Error-Correcting [SPM17]. Error-Tolerant [DSH^{+}18]. Errors [FWJ^{+}16, ZWS^{+}16]. essence [JEBJ08]. Estimation [WAST16, XHJY17, LTG12]. Estimations [Lou19]. Estimator [KLA^{+}19]. Evaluate [MRK^{+}22, TDO16a, VOK^{+}22]. Evaluating [CCM^{+}16, CWS06, HWH^{+}11, SSK11, SAT20, SW17a]. Evaluation [Akr21, BC13, CHE^{+}14, DKB^{+}20, FWJ^{+}16, ZZL^{+}21, AvRF07, KWTD09, LCC11, LAS^{+}08, RGG^{+}12, ZK05]. Evaluator [JSL13]. Evaluator-executor [JSL13]. Even [MP22]. Event [YXMC23, GWM07]. Event-Driven [YXMC23]. Evolving [VGX16]. Examining [ZWS^{+}16]. Example [WDK^{+}24]. exascale [DXMJ11]. ExaStencils [KL19]. exception [HWM14]. Exceptionization [YKM17]. Exclusivity [YDL^{+}17]. Execution [AGG22, ASP17, BNS^{+}21, CC18, DT17, GGYK19, GMGZP14, HAC13, HEMK17, KS16, KABS22, LDMZ19, MG19, MAY23, ME15, MAD17, MKS22, NZ15, PVA^{+}17, PS15, SEF^{+}19, SYE19, SGS^{+}20, VSDL16, WLZ^{+}13, ZX19, ZCCD16, ZLJ18, GB06, LZ12, LHZ13, SJA12, VTN13, XIC12, ZG05]. Executions [NDP17]. executor [JSL13]. exhaustive [KWTD09]. Existing [YEI^{+}14]. Expanding [YBSY19]. Expansion [PM17, ZLC^{+}15]. EXPERTISE [SDK^{+}22]. explicit [STLM12]. Exploit [AAI^{+}16]. Exploiting [AIVL13, ASK^{+}16, HWJ^{+}15, JFK20, KGK10, LHW^{+}19, MA08, NKH16, RSU^{+}20, WWW^{+}21, WZW^{+}24, XSF^{+}23, YEI^{+}14, YZ08, YZL^{+}10, ZX16, LYYB07, PCT12, RLS13, SNL^{+}04, JOA^{+}09b]. Exploration [BKM^{+}17, DNS^{+}24, ESB^{+}20, KL19, MNC^{+}16, QSZ^{+}21, CPP08, IMS^{+}08, KWTD09, VHKP11, WLZ^{+}10].

Explorations [BGG^{+}15]. Exploring [ACGH24, CK11, JK13, JOA^{+}09b, MBKM12, MSK05, SKPD19, WLJ^{+}24, vdVSAAS20, BE13, DJX13]. Exposing [CSK19, HLSK22]. Express [DJC16]. Expression [BC13]. Expressions [VZT^{+}20, JSH09]. Expressiveness [PC13]. Extendable [CXW^{+}12]. extended [SJV08]. Extending [DBH16, DSH^{+}18, JED19, TCS20, VCJ^{+}17]. Extension [ZLWS24, ZC20, DCP^{+}12]. Extensions [FSP^{+}23, KHS^{+}14, KBB^{+}14]. Extractor [DAP^{+}15]. Extreme [CAY^{+}18, JLJ^{+}18a]. Extreme-Scale [CAY^{+}18, JLJ^{+}18a].

Fabric [PBCB22]. Factorization

[MLC^{+}23]. Factorizations [AP17]. Facts [Mic16]. FailAmp [BDB^{+}20]. Failures [NRQ16a]. Fair [LMCV13]. Fairness [GWM07, LY16]. Falcon [CNS16a]. false [BCVT13]. Far [WZW^{+}24]. FASA [DQCK24]. FASA-DRAM [DQCK24]. Fast [ADGA20, BC13, BNS^{+}21, CSSU21, CTY^{+}23, CCPG13, DWF^{+}23, FPK^{+}24, KCP13, KVH23, KHW^{+}05, MKKE15, NRQ16b, NTG13, PRMH13, SCMU22, SZJK18, SNK^{+}23, WLJ^{+}24, XZW^{+}22, LMJ13a, SPGE06, TDG13]. Fast-Drift-Aware [SZJK18]. Fastensor [WZWW23]. Faster [PCM16]. fat [BRSJG12, PRMH13]. fat-trees [BRSJG12]. Fault [CEP^{+}16, PHBC17, RHLA14, TCR^{+}22, WDW^{+}22, RCV^{+}05]. Faults [SDK^{+}22, BSO07, SSC^{+}13]. FaultSim [NRQ16b]. Feature [TKM14, LBO14]. Features [YHYBAM20]. Federation [BTS10]. Feedback [CDM13, NED^{+}13, ZWS^{+}16, WM10]. Feedback-directed [NED^{+}13, WM10]. Feedback-Driven [ZWS^{+}16, CDM13]. Fence [MNSC16]. fetch [EE09, GWS13, JLER12, SRLPV04]. FFT [GS12, ZLM^{+}23]. File [MP22, TS15, VZS^{+}18, YBSY19, GKP14, SJV08]. Files [LZM14, YWXW12]. Filter [SWO21, BSWLE13]. Filtering [LLLW22, ZCCD16]. Filters [WWY^{+}24]. Financial [ABB^{+}16]. Finding [PJ13]. Fine [ACG24, ANS^{+}22, AZG17, BSSS14, CS21, CSRP22, EE11, GWZ22, HYYBM16, MG19, MPW^{+}17, OBdOJ^{+}24, SSS^{+}24, TKM14, WM11, YEI^{+}14, LT13]. Fine-Grain [AZG17, HYYBM16, ACG24, ANS^{+}22]. Fine-Grained [BSSS14, CS21, GWZ22, MG19, MPW^{+}17, OBdOJ^{+}24, YEI^{+}14, CSRP22, EE11, SSS^{+}24, WM11, LT13]. Finite [LVR^{+}15, VW11]. FinPar [ABB^{+}16]. Fire [YXMC23]. First [Lou19, OAM19]. Fixed [DNS^{+}24, CS13]. Fixed-point [DNS^{+}24, CS13]. FLARES [DGI^{+}14]. Flash [DGI^{+}14, LCW^{+}24, LYLS24, SZJK18, ZWL^{+}19]. Flash-Based [LYLS24]. FlexHM [PDY^{+}23]. Flexible [CC13, CSRP22, PDY^{+}23, SNK^{+}23, ZC20, OAB12, SHC13, ZZQ^{+}05]. FlexPointer [CTY^{+}23]. FlexSig [OAB12]. Flextended [ZC20]. flight [SSH^{+}13]. Floating [ASS17, BWG^{+}12, CS13]. floating- [CS13]. Floating-Point [ASS17, BWG^{+}12]. Flow [BRJM15, CWW^{+}16, DMR^{+}16, GAM12, HAC13, LY16, MMT^{+}12, SMKH15, FSYA09, JA14, KHL^{+}13, MBKM12, Nas13, PC13, TG07]. Flow-Based [LY16]. flow-sensitive [Nas13]. FlowPix [CGP23]. FluidCheck [KS16]. fly [VHKP11, WWY^{+}12]. Flynn [TWB21]. Focal [DSK19]. Focal-Plane [DSK19]. Footprint [KDMA23]. Forest [ELE^{+}23]. form [PBCB22]. Format [BJWS18]. Formation [HWL^{+}19, KTAE16, FSYA09]. Formulating [MAN^{+}08]. Forward [LNFE22]. Forwarding [SL20]. Four [TDO16a]. FPDetect [DKB^{+}20]. FPGA [Abd20, CS13, CWW^{+}16, CSRP22, CGP23, CDPD13, MTK18, MRK^{+}22, MTPK24, OLK^{+}23, SNK^{+}23]. FPGA-based [SNK^{+}23, MTK18, MRK^{+}22, OLK^{+}23].

FPGA-processor [CS13]. FPGAs [BHWN21, FBWS13, GNB08, JOB^{+}22, KFJ20, PI12, WZZ^{+}20, XLH^{+}24]. fractal [JYJ^{+}13]. fractal-based [JYJ^{+}13]. Fraction [SPS17]. frame [GK13]. frame-based [GK13]. Framework [ASS17, AMP^{+}16, FYI^{+}24, GTT^{+}16, GáSÁ^{+}16, HDW21, KKL^{+}24, KPP^{+}15, LAS^{+}13, LSC^{+}15, MLC^{+}23, OLK^{+}23, PLLW24, PWPD19, SYE19, SMM^{+}23, SAL19, WMGS19, WPR^{+}22, WWGS22, WYZ^{+}23, XYY^{+}24, ZLYZ16, ZFT^{+}18, ZLYW18, ZLM^{+}23, AS13, BCVN10, CS10, DJX13, HEL^{+}09, KKM^{+}13, LCC11, LCH^{+}04, LFC13, LHWB12, PGB13, YXK^{+}12]. Frameworks [WWW^{+}21]. Franca [MLB^{+}23]. Free [CHD^{+}23, LLC22, MNSC16, YPT^{+}16, BRSJG12, GS12, WZZ^{+}20]. Frequency [BHC^{+}16]. Friendly [LLW^{+}22, CRSP09]. Front [ZJJ^{+}15]. Front-End [ZJJ^{+}15]. FSM [SQZK20]. FTL [HWJ^{+}15]. Full [HHC^{+}16, MMT^{+}12, SWF16, TKKM15]. Full-System [SWF16]. Fully [HWJ^{+}15, BRSJG12]. Function [SKPD19]. Functional [GáSÁ^{+}16, SJD22, GÁSÁ^{+}13, YCCY11]. Functions [HLSK22, SSRS15, HWX^{+}13, LDG^{+}13]. fundamental [VE13]. Fuse [NDP17]. Fused [VPTS19]. Fusing [VPTS19, WM10]. Fusion [PSL^{+}23, SSS^{+}24]. Future [GB06, MMS15, DXMJ11, LMJ13a].

Gadgets [JCG^{+}24]. Gain [SSP^{+}23]. Gaming [QYZ^{+}14, RSU^{+}20, UPR22]. GAN [WDW^{+}22]. gap [HCC^{+}14]. Garbage [ASV^{+}16]. Gated [CS21, LZM14]. gather [Pro21]. Gating [KMG14, ZCF18, WYCC11, YCCY11]. GC [SLS^{+}21, YXS^{+}22]. GC-Triggered [YXS^{+}22]. Gem5 [QSZ^{+}21]. Gem5-X [QSZ^{+}21]. GEMM [SLJ^{+}19]. General [CAMJ15, SW17a, SDZ^{+}21, SMN22, XHL^{+}24, LHY^{+}06]. General-Purpose [CAMJ15, SDZ^{+}21]. Generalized [FDF^{+}14, GGK18, KKL^{+}24, LCW^{+}23, SDH^{+}15]. Generalizing [Jim09]. generate [KBR^{+}13]. Generating [AZG17, RHC15]. Generation [BDB^{+}20, DSK19, CDM^{+}22, HEMK17, JCG^{+}24, TPB24, ZPH^{+}23, GNB08, HLR^{+}13, JLER12, LBO14, LHY^{+}06, VJC^{+}13]. Generator [KL19, PAVB15]. Generic [WMGS19]. GenMatcher [WMGS19]. Genome [FPK^{+}24]. geometry [CNAA^{+}22]. Getting [MWJ19]. GiantVM [JZY^{+}22]. Global [CCL^{+}13, MPPS18, BZS13]. GNN [RJK24]. good [PJ13]. Governors [SW17b]. GP [LRBG15, MYG15, MYKG16]. GP-GPUs [LRBG15]. GP-SIMD [MYKG16]. GPGPU [BGG^{+}15, HLSW17, MBKM12, YXK^{+}12]. GPGPUs [ZJJ^{+}15]. gPPM [LCW^{+}23]. GPU [ADGA20, BJWS18, BNS^{+}21, DS16, FBC^{+}22, GGYK19, GMZ^{+}21, HLR^{+}13, HZC^{+}23, HDW21, JED19, JGSM15, JML^{+}20, KPRK20, KHN^{+}18, LCP^{+}21, LHC^{+}17, LWS^{+}19, LMZ18, LWL18, LDMZ19, LAAMJ15, LFK19, LFC13, QYZ^{+}14, RB13, SEF^{+}19, SNN^{+}19, TBC^{+}12, VC16, VZT^{+}20, VZS^{+}18, WGO15, WWL^{+}21, WPR^{+}22, WDK^{+}24, WZWW23, ZPL^{+}21, ZNTJE23, ZSLX13, ZLM^{+}23, vdVSAAS20]. GPU-Accelerated [ZPL^{+}21, GMZ^{+}21, JED19]. GPU-Based [WGO15, JML^{+}20]. GPUs [ASH20, ACGH24, ASS17, CNAA^{+}22, CSK19, DS16, DHX^{+}22, DJZ^{+}23, DNT16, FSX^{+}24, FBWS13, GWZ22, JAK17, JFK20, KR19, KVH23, LRBG15, LGH^{+}21, MKS22, NSF^{+}21, NC15, PL23, SHLM14, FCVR24, SSB^{+}20, TAB^{+}21, WYCC11, WCW^{+}24, WLLW20, XSF^{+}23, YBSY19, ZSM^{+}16]. gradient [HAJ^{+}12]. gradient-based [HAJ^{+}12]. Gradients [FWJ^{+}16]. Grain [AZG17, HYYBM16, LMSE18, MAD17, ACG24, ANS^{+}22]. Grained [BSSS14, CS21, GWZ22, MG19, MPW^{+}17, OBdOJ^{+}24, TD16, YEI^{+}14, CSRP22, EE11, KCP13, LT13, SSS^{+}24, WM11]. GRAM [HDW21]. Granularity [DRHK15, NRQ16a, TKM14]. Graph [CNS16a, HLZ^{+}24, KPP21, KKAR16, LZZ^{+}22, LGH^{+}21, MST^{+}21, RJK24, TPB24, TAB^{+}21, WZZ^{+}20, WWL^{+}21, YWXW12, YZZ^{+}23, ZPL^{+}21, ZZH^{+}23, ZLJ18, DS12, LFX09]. Graph-Based [TAB^{+}21]. GraphAttack [MST^{+}21]. GraphBLAS [MAY23]. Graphics [ASS17, FSYA09, ZSLX13]. GraphPEG [LGH^{+}21]. Graphs [BRJM15, Lee16, RHC15, VZT^{+}20, VGX16, BZS13, DDU12, MG13]. GraphSER [LK24]. GraphTune [ZZH^{+}23]. Gretch [KPP21]. Grouping [LGH24]. Grus [WWL^{+}21]. gshare [TS05]. Guarded [PS15]. Guidance [OKJ^{+}22]. Guided [GTT^{+}16, HWL^{+}19, YHYBAM20, CS13, LZL^{+}13, RCG^{+}10b, SSU^{+}13].

Hadoop [KHS^{+}14]. HAIR [MP22]. Halide [SSW^{+}19, SSB^{+}20, VCJ^{+}17]. halting [ZVYN05]. Halving [MP22]. Hammer [GBD21]. Hamming [CVB15]. handling [HWM14, HWH^{+}11, LWH11]. HAP [WJXC17]. Hard [DPBI^{+}19, BSO07]. hardened [PKL^{+}24]. Hardening [PHBC17]. Hardware [BGG^{+}15, BOEN23, BAZ^{+}19, CDPN16, DHK18, DPBI^{+}19, DD16, ELE^{+}23, ICEDR24, JDZ^{+}13, KPP21, KAC15, LMJ^{+}13b, LGH24, MMGS21, NDP17, OK21, PVA^{+}17, PLK^{+}19, PKL^{+}24, RHLA14, RAF22, SBC^{+}22, SKAEG16, SDK^{+}22, SWF16, TGAG^{+}12, USCM16, WCI^{+}16, WZY^{+}24a, WZY^{+}24b, ZHS^{+}19, ZLC^{+}15, ZSM^{+}16, Abd20, ATGN^{+}13, CS10, CI13, FSYA09, GNB08, HCC^{+}14, MMdS06, OAB12, RLS13, RPE12, YJTF13, ZSCM08]. Hardware-Accelerated [SWF16]. Hardware-Assisted [CDPN16, JDZ^{+}13]. Hardware-aware [BOEN23]. Hardware-Based [ZLC^{+}15, ZSM^{+}16].

Hardware-hardened [PKL^{+}24]. hardware/software [CS10, HCC^{+}14, MMdS06]. Harvesting [XLD^{+}24]. Hash [SBS16]. Hash-Based [SBS16]. HAShCache [PG17]. Hashing [CHD^{+}23]. HAWS [GGYK19]. HC [CDPD13]. HC-CART [CDPD13]. HDD [CKL^{+}24]. header [VED07]. Healthy [JLJ^{+}18b]. heap [WWY^{+}12]. HeapCheck [SBC^{+}22]. Heterogeneity [PG17, SB09]. Heterogeneity-Aware [PG17]. Heterogeneous [AEJE16, ASV^{+}16, ANS^{+}22, APG^{+}23, ASP17, AMS23, CNS16a, CWW^{+}16, DMR^{+}16, FDF^{+}14, GTT^{+}16, GHH15, GSZY20, HAM17, HAM19, HMYZ15, HHW^{+}22, KRHK16, LP17, MSFC21, OKJ^{+}22, PG17, PDY^{+}23, PBY^{+}17, QAT24, QSZ^{+}21, RKL23, RVKP19, SMS23, SCK^{+}21, SSS^{+}24, SAL19, SL20, TDO16a, TDO16b, TTS19, USCM16, WGO15, ZFL18, ZF24, BBG13, KNBK12, LHC^{+}24, LHZ13, PM12, TDG13, VE13, WFKL10]. Heuristics [MKKE15, TR13]. Hidden [FPK^{+}24]. hide [CST^{+}06]. Hiding [GW08]. HIER [LDY^{+}21]. Hierarchical [ASK^{+}16, CDPN16, JHH^{+}23, LDY^{+}21, ZGP15, SW13]. Hierarchies [GAH22, SKH^{+}16, DJX13]. Hierarchy [AYC16, ELE^{+}23, SSP^{+}23, ZDC^{+}16, ZSM^{+}16]. High [CAY^{+}18, CHE^{+}14, DKK^{+}21, CHD^{+}23, CAMJ15, GGK18, JED19, LNFE22, LDY^{+}21, LL22, ME17, MLB^{+}23, MPU^{+}23, MN24, OBdOJ^{+}24, OK21, SAG22, SWU^{+}15, SSP^{+}23, SLJ^{+}19, TCS16, THA^{+}21, TKM14, UDLD20, USCM16, WZZ^{+}20, WWL^{+}21, YRGES^{+}19, ASK13, BCVN10, CK11, CDM13, GW08, KBR^{+}13, OGK^{+}12, SRLPV04, SD12, ZVYN05]. High-dimensional [LL22]. High-Efficiency [CAMJ15]. High-Level [CHE^{+}14, UDLD20, BCVN10]. High-Order [CAY^{+}18]. High-Performance [DKK^{+}21, GGK18, LNFE22, OBdOJ^{+}24, SLJ^{+}19, TKM14, USCM16, CHD^{+}23, JED19, MLB^{+}23, SSP^{+}23, THA^{+}21, WZZ^{+}20, WWL^{+}21, YRGES^{+}19, CK11, CDM13, GW08, KBR^{+}13, SRLPV04, SD12, ZVYN05]. High-radix [LDY^{+}21, ASK13]. High-Throughput [SAG22, OGK^{+}12]. Higher [SJD22]. Higher-Level [SJD22]. Highly [JYM20, MFSK24, TMP16, TPN^{+}20, ZLM^{+}23]. Histogram [FWJ^{+}16]. hits [CA11]. HMTT [HCC^{+}14]. Holistic [CHD^{+}23, LD24, OLK^{+}23]. Homogeneous [CC18]. Hopping [MSFC21]. hosted [SYZZ^{+}14]. HotSpot^{T M} [KWM^{+}08]. HPar [ZBH^{+}13]. HPC [ACA^{+}19, DVG^{+}23, MP13, MKC^{+}22, PLT^{+}15, SLJ^{+}18, ZPR^{+}17]. HPCG [AYL^{+}18]. HRF [GHH15]. HRF-Relaxed [GHH15]. HTML [ZBH^{+}13]. HTML5 [NKH16]. HW [KMG14, LYK^{+}15, TS15]. HW/SW [KMG14]. Hybrid [BR13, CKPH19, CA11, DXMJ11, HWJ^{+}15, HLZ^{+}24, JYE^{+}16, KAC^{+}18, LYLS24, QAT24, SSP^{+}23, WLL^{+}19, WJXC17, YRGES^{+}19, CS13, CKL^{+}24, DZC^{+}13, HCC^{+}14, MMdS06, RBM10, WLZ^{+}10]. Hybrid-Memory-Aware [WJXC17]. HyGain [SSP^{+}23]. Hypervisor [JZY^{+}22].

I-Cache [ZWY17]. I/O [DCP^{+}12, RHLA14, SLS^{+}21, WLX24, WZWW23]. IATAC [AGVO05]. Identification [WCI^{+}16]. Idiom [KKM^{+}13, TWB21]. Idioms [DKK^{+}21]. Idle [SEF^{+}19, WWY^{+}24, WFKL10]. Idle-Time-Aware [SEF^{+}19]. IEEE [LDG^{+}13]. IEEE-754 [LDG^{+}13]. ILP [SNL^{+}04]. Image [CGP23, LSH^{+}23, PBY^{+}17, CI13]. Imaging [VCJ^{+}17]. Imbalanced [Pro21]. Immersion [MBLK24]. Immersion-Cooled [MBLK24]. Impact [BCVN10, CCM^{+}16, JRK16, FCVR24, SMKH15, ZCDD23, RGG^{+}12, SSC^{+}13]. Impactful [YHYBAM20]. implants [SSPL^{+}13]. Implement [VOK^{+}22]. Implementation [BGG^{+}15, MAY23, CDPD13, LHZ13, PLL10, SSS^{+}04, ZK05, AvRF07]. Implementing [CWW^{+}16, JSM^{+}04, MAN^{+}08, OAB12]. Implications [CVB15, HYYBM16, KAC15, LS10]. Implicit [BWLR06]. Improve [CSK19, CDM^{+}22, LMZ18, OTR^{+}18, VCJ^{+}17, ATGN^{+}13, BSWLE13, KGK10, LBJ05, LZ12, MG12, RWY13, SPS12]. Improved [BCVT13, GMGZP14, NB13, VZS^{+}18, ZJJ^{+}15]. Improvement [SKKB18]. Improvements [LBM13, PM17, SPM17]. Improving [AJK^{+}12, CAGS17, CSW^{+}23, CG15b, DHK18, DJZ^{+}23, FLW^{+}24, HWJ^{+}15, HLSW17, JHHM21, JK17, KLMP12, LGP^{+}16, LMSE18, LYH16, LAAMJ15, OAM19, RJSA18, SL20, YBSY19, ZFT^{+}18, ZWHM05]. In-bounds [JRH21]. In-DRAM [XGD^{+}23]. in-flight [SSH^{+}13]. In-Memory [BAZ^{+}19, WZG^{+}19, YSH^{+}22, ZLYW18]. In-Order [BEE15, MST^{+}21, MAD17, SPH^{+}17, BB04]. in-order/out-of-order [BB04]. in-place [GS12]. In-storage-processing [KKL^{+}24]. inclusive [AIVL13, TKJ13]. Increasing [TZK18]. independent [BVIB12]. Index [WDS^{+}24]. indexing [TS05]. Indirect [CSY20, DGGL16, XVT20, XHL^{+}24, HWH^{+}11, MG12]. indirections [AFD07, AFD12]. Industrial [GHH15]. Ineffective [WLX24]. Infer [UJW15]. Inference [CSRP22, DJZ^{+}23, JHH^{+}23, SCK^{+}21, WYZ^{+}23, LB10]. Inflation [XYY^{+}24]. Influence [ZWS^{+}16]. Information [GAM12, KHL^{+}13, MMT^{+}12, SM19, SAT20, LMJ13a, VSP^{+}12]. Informed [CSY20, SYX^{+}15]. Infrastructures [FCD^{+}17]. Innovative [BKM^{+}17]. Input [LDY^{+}21]. inputs [BE13]. Insights [YHYBAM20]. Instruction [AGG22, CSK19, HNKK17, JHQ23, KBB^{+}14, SPGE06, SKPD19, SGM^{+}22, TCS20, XYY^{+}24, ZCDD23, ACGK04, BVIB12, BR13, CS10, CSVM04, GWS13, HL07, KS11, SSR13, VS11, XL07, ZHD^{+}04, ZK06]. Instruction-Level [HNKK17]. instructions [MG12, RFD13, SHC13]. Instrumentations [JRH21]. Instrumented [SGS^{+}20]. Integer [ÄJE^{+}16, MP22, SLM12, BWG^{+}12]. Integrated [DJC16, LYK^{+}15, PG17, SKP^{+}22, SPH^{+}17, TPN^{+}20, VFJ^{+}17, YJTF13]. Integrating [WSJ^{+}21, WTFO14, XZW^{+}22]. Integration [JDZ^{+}13]. Integrity [KK15]. Intel [Akr21]. Intelligent [SJC^{+}21, TBC^{+}12]. Intensity [LVR^{+}15, SLS^{+}21]. Intensity-aware [SLS^{+}21]. Intensive [RHLA14, ZX19, YLTL04]. Inter [HAM^{+}20, LBM13, TC07]. Inter-cluster [TC07]. Inter-Core [LBM13]. Inter-kernel [HAM^{+}20]. Interaction [FBHN04]. Interactions [EPS17]. Interactive [MPHL22, RSU^{+}20]. Intercepting [SSRS15]. Intercommunication [TMP16, MP13]. Interconnect [BKM^{+}17]. interconnection [SMK10, SEP07]. Interconnects [DH16, YRGES^{+}19, XCC^{+}13]. Interface [WDX24, XVT20, ZSLX13]. Interference [KLA^{+}19]. interferences [LCL^{+}14]. Interleaved [AMG16, ZLWS24]. Interleaving [WWC^{+}16]. Intermediate [JML^{+}20, LHC^{+}24, RJSA18]. Intermediate-oriented [JML^{+}20]. Intermittent [BSG23]. Internal [HWJ^{+}15]. Internet [AVG12]. Interposer [CWK^{+}24]. Interposer-based [CWK^{+}24]. Interpreter [ZXX23, RWY13]. interpreters [SYZZ^{+}14]. interprocedural [SV05]. Intersection [MPHL22]. Interval [SKP^{+}22, SV05]. Intra [MKC^{+}22]. Intra-rack [MKC^{+}22]. Intraprogram [XMM04]. Intrinsic [JRK16]. Introduction [CT04, CT05, CT06, CT07, SD12]. intrusion [TBS06]. Intrusive [FPMR21]. IOV [DCP^{+}12]. IP [Bis21, WYJL10]. IR [GMZ^{+}21, SJD22]. Irregular [KPM21, LWS^{+}19, MN24, RMA14, SN17, AFD12]. Irregularity [ZZH^{+}23]. ISA [CG14, FSP^{+}23, SHC13, VE13]. ISAs [PS15]. Isolating [MTPK24]. Isolation [LDC15, OK21, QYZ^{+}14, SSH19]. Isolation-based [OK21]. ISP [KKL^{+}24]. Issue [DD16, MMS15, BB04, CDM13, GWS13, PI12, SD12]. iSwap [WLX24]. ITAP [SEF^{+}19]. Iteration [WWC^{+}16, ZPH^{+}23]. Iterative [CNS^{+}16b, FXC^{+}15, GGS^{+}17, GGS^{+}19, KFJ20, LLLW22, SYE19, CFH^{+}12]. IVR [ZZL^{+}21]. IVR-assisted [ZZL^{+}21].

Java [HWM14, KWM^{+}08, LBJ05, VED07, WHV^{+}13, YKM17, YLW08]. JavaScript [MGI15, NKH16, PCM16, PKPM19]. JIT [HWM14, JK13, JCG^{+}24, NED^{+}13]. JiuJITsu [JCG^{+}24]. Job [CSW^{+}23, PLLW24, EE12]. Joint [RAF22, TS15, LGAZ07]. Jointly [CSW^{+}23]. jump [MG12]. Just [HZN^{+}22, KHL^{+}13]. Just-In-Time [HZN^{+}22, KHL^{+}13]. JVM [SYZZ^{+}14].

Kernel [CRBP24, DSK19, LP17, LDMZ19, MSFC21, PSL^{+}23, SNN^{+}19, HAM^{+}20]. KernelFaRer [DKK^{+}21]. Kernels [BNS^{+}21, LCP^{+}21, TPB24, VZT^{+}20, WLLW20]. Key [CSSU21]. Key-Value [CSSU21]. kilo [CSVM04]. kilo-instruction [CSVM04]. Knowledge [WZY^{+}24a]. Knowledge-Augmented [WZY^{+}24a].

L1 [HK14, LZL^{+}13]. L2 [AGVO05, CST^{+}06, SLP08, SBC05]. L2-miss-driven [SLP08]. Label [LWW^{+}24]. Lane [WWC^{+}16]. Language [CNS16a, SNK^{+}23]. Languages [DHD^{+}14, YKM17, NED^{+}13]. LAPPS [KFEG18]. Large [HZC^{+}23, NRQ16a, PLLW24, SKH^{+}16, XLH^{+}24, YGB21, ZPL^{+}21, ZLM^{+}23, KWCL09, RCV^{+}12, SMK10]. Large-Scale [PLLW24, SKH^{+}16, ZPL^{+}21, ZLM^{+}23, HZC^{+}23, RCV^{+}12, SMK10]. Large-Tiling [XLH^{+}24]. LargeGraph [ZPL^{+}21]. Last [CPS^{+}15, LBM13, PLK^{+}19, WSJ^{+}21, WDX14, WJXC17, AGI^{+}12, AIVL13, VSP^{+}12, ZDC^{+}12]. Last-Level [CPS^{+}15, LBM13, WSJ^{+}21, WDX14, WJXC17, PLK^{+}19, AGI^{+}12, AIVL13, VSP^{+}12, ZDC^{+}12]. Latency [BAZ^{+}19, DQCK24, HAM17, HK14, KCA^{+}13, PM17, RKL23, SLS^{+}21, LZS^{+}24, MP13, SW13, WYJL10, YLTL04]. Latency-Tolerant [HAM17]. Lattice [CG15b, PAVB15]. Lattice-Based [CG15b]. Lattice-Boltzmann [PAVB15]. Lavender [PLLW24]. Law [DSH^{+}18]. Layer [ERAG^{+}16, JML^{+}20, JLJ^{+}18a, LGP^{+}16, OTR^{+}18, WAST16, VOK^{+}22]. Layer-adaptive [JML^{+}20]. Layer-Centric [JLJ^{+}18a]. Layers [VZT^{+}20]. Layout [ACGH24, CYXF13, RAF22, WG17, ZLWS24]. Layout-oblivious [CYXF13]. Layouts [BSL17]. Layup [JML^{+}20]. LD [LHC^{+}17]. LDAC [SKH^{+}16]. LDPC [LCW^{+}24]. Leakage [Bis21, CS21, JFK20, HL07, MSK05]. Learned [WDS^{+}24]. Learning [ABP^{+}17, CKP^{+}22, DLS22, JPS17, JLJ^{+}18a, LSL20, LLLW22, LSH^{+}23, MCB^{+}12, QAT24, RSK^{+}18, WWW^{+}21, WZWW23, XDXL19, XDW^{+}23, DJB13, LBO14, SPS12, TR13, WO13, WTFO14]. Learning-Based [JPS17]. Leases [RGK^{+}23]. Leasing [DCL^{+}22]. Legacy [MNSC16]. legalization [BR13]. Leighton [WDK^{+}24]. Less [ZPR^{+}17, RJK24]. Level [BGG^{+}15, CHE^{+}14, CPS^{+}15, GMZ^{+}21, HNKK17, HK14, JYE^{+}16, LCS^{+}19, LIS20, LMZ18, LBM13, MGI15, PLT^{+}15, RLBBN15, SJD22, SWU^{+}15, UDLD20, WSJ^{+}21, WDX14, WJXC17, AGI^{+}12, AIVL13, BCVN10, EE09, GMW09, GPL^{+}05, HLSK22, LZW23, LCL^{+}14, Lou19, PLK^{+}19, PCT12, SDK^{+}22, TMSR23, VSP^{+}12, YBSY19, ZDC^{+}12, ZZL^{+}21]. Level-1 [HK14]. Leveling [JDZ^{+}13]. Levels [RJSA18, RCV^{+}12, SLA^{+}07]. Leveraging [AGG22, GAM12, KS21, LMJ13a, NZ15, SHLM14, SMN22]. Liberalization [MY16]. libraries [BCM11]. Library [DKK^{+}21, FDF^{+}14]. Library-Based [FDF^{+}14]. Lifetime [PM17, SPM17, TZK18, XC06]. Lift [SHS^{+}20]. LIGERO [APG13]. Light [CBD15, APG13]. Lightweight [DT17, SLJ^{+}18, WLL^{+}19, ZMLL24, BWG^{+}12, DMG13, LNLK13]. like [Mic18]. limitation [DZC^{+}13]. Limitations [JRK16]. limited [CZ07]. limits [JOA^{+}09b, MBKM12, MSK05]. line [WDXJ14]. Linear [ÄJE^{+}16, APG^{+}23, MG19, MG20]. lines [AGVO05]. Lingua [MLB^{+}23]. Linked [IPSD21, FLG12]. Links [ACA^{+}19]. List [Aca16, Ano13a, Ano15, Bil19]. LiteCON [DLS22]. Live [ZPR^{+}17]. liveness [BZS13, DDU12]. LLC [FQRG13, VPTS19, ZCF18]. LLC-memory [FQRG13]. LLVM [DAP^{+}15]. LLVM-Based [DAP^{+}15]. Load [OAM19, PGB16]. Load-Balancing [PGB16]. Loading [PCM16]. Loads [YPT^{+}16]. Local [LVR^{+}15, XZW^{+}22, DHC^{+}13]. Locality [ASK^{+}16, CG15a, KFEG18, SKH^{+}16, SL20, TAB^{+}21, UPR22, YDS^{+}19, ZCQ^{+}19, AIVL13, FER^{+}13]. Locality-Aware [CG15a, KFEG18, SKH^{+}16, UPR22, SL20]. Localization [CEP^{+}16, WZY^{+}24a, WZY^{+}24b]. located [LZW23, SOAK23]. location [KHN^{+}18, YLW08]. Lock [CHD^{+}23, CWCS13]. Lock-contention-aware [CWCS13].

Lock-Free [CHD^{+}23]. Locking [ZWY17]. Locus [DVG^{+}23]. Logarithmic [AGG21]. Long [SLS^{+}21]. Long-tail [SLS^{+}21]. Look [HZN^{+}22]. Lookups [CSSU21]. Loop [ASP17, CZGC20, JK17, LVR^{+}15, MN24, PHBC17, BCVT13, NCC13, SHLM14, SLM12, YZL^{+}10]. loop-dependent [YZL^{+}10]. Loops [CNS^{+}16b, CLA^{+}19, KFJ20, RGK^{+}23, SN17, SRC16, JSL13, KLMP12, RTG^{+}07]. Low [AGG21, AGG22, BGG^{+}15, CAMJ15, DJL^{+}12, ESB^{+}20, GG18, GáSÁ^{+}16, GDL16, KBB^{+}14, KDMA23, LNFE22, LGP^{+}16, LHC^{+}17, Lou19, LZS^{+}24, OK21, PLK^{+}19, RTK15, SBC^{+}22, SSW16, SLS^{+}21, SW13, SWU^{+}15, WLJ^{+}24, YEI^{+}14, AGI^{+}12, BB04, CCZ13, GKP14, MA08, SRLPV04, SAM^{+}23, ZVYN05]. Low-Complexity [LNFE22, DJL^{+}12, SRLPV04]. Low-Cost [KBB^{+}14, SSW16, YEI^{+}14, SBC^{+}22, AGI^{+}12, MA08]. low-energy [GKP14, ZVYN05]. Low-latency [LZS^{+}24, SW13]. Low-Level [BGG^{+}15, Lou19]. Low-Overhead [GDL16, LHC^{+}17]. Low-Power [CAMJ15, GáSÁ^{+}16, AGG22, PLK^{+}19, BB04, CCZ13]. Low-precision [AGG21]. Lower [ESR^{+}15]. lowering [SSU^{+}13]. LP [GFD^{+}14]. LSM [DZL^{+}24, HFL^{+}23, LHS^{+}24, XZW^{+}22]. LSM-Tree [HFL^{+}23, LHS^{+}24, DZL^{+}24, XZW^{+}22]. LSTM [WDW^{+}22]. LSTM-GAN [WDW^{+}22].

Machine [ABP^{+}17, DWF^{+}23, LSH^{+}23, DJB13, LBO14, SCEG08, SPS12, WO13, WTFO14, WHV^{+}13]. machine-learning-based [WTFO14]. Machines [BSSS14, JK13, RB13, VED07]. MAGIC [KKW^{+}15]. Main [AEE^{+}19, WSJ^{+}21, ZFT^{+}18, ZPR^{+}17, DZC^{+}13, WSC^{+}13, ZDC^{+}12]. Maintaining [YCCY11]. Maintenance [CKL^{+}24]. makespan [CPB14]. makespan-preserving [CPB14]. Making [CRSP09, PLT^{+}15, PI12, SGS^{+}20]. Malicious [KKW^{+}15]. Malware [WCI^{+}16]. MAMBO [GDL16]. Managed [Akr21, YWXW12]. Management [CMAP22, GWZ22, GTT^{+}16, GMGZP14, HYBR^{+}15, HMYZ15, LHS^{+}24, MPPS18, MRK^{+}22, NMPS22, OTR^{+}18, SEF^{+}19, SSS^{+}23, SAL19, SPS17, SJC^{+}21, TTS19, VOK^{+}22, WPR^{+}22, ZDC^{+}16, AVG12, FQRG13, GSZI10, HVJ06, KCKG14, LGAZ07, LFX09, LPZI12, RCG^{+}10a, RB13, SW13, VS08, WWWL13, WSC^{+}13, WDXJ14, WM11, ZYCZ10]. Manager [APS22, ELE^{+}23, Per18]. Managing [APBR16, HS06, KNBK12, SCFD22, VS11, ZFL18, SSK11]. Manipulation [CNS16a, ZHB18]. Many [DT17, FMY^{+}15, JYM20, JHH^{+}23, JLJ^{+}18a, LK24, MLC^{+}23, MRK^{+}22, PVS^{+}17, QSZ^{+}21, SSS^{+}23, WPR^{+}22, ZLYZ16, LNLK13, OGK^{+}12]. Many-body [WPR^{+}22]. Many-Core [FMY^{+}15, JLJ^{+}18a, LK24, PVS^{+}17, ZLYZ16, JYM20, JHH^{+}23, MLC^{+}23, MRK^{+}22, QSZ^{+}21, SSS^{+}23, LNLK13, OGK^{+}12]. Many-Cores [DT17]. Manycore [KS16, KAC^{+}18, LAS^{+}13, MKKE15, ZCQ^{+}19, BTS10]. map [WYJL10]. Mapi [BSG23]. Mapi-Pro [BSG23]. Mapped [LLRC17]. MAPPER [SCFD22]. Mapping [BSG23, CKP^{+}22, CMAP22, CDPN16, DWDS13, DJC16, ESB^{+}20, LLC22, LD24, MKKE15, SSH19, SKAEG16, WGO15, YMM^{+}15, CCZ13, WYJL10, WTFO14]. MapReduce [CC13]. MAPS [RLBBN15]. Markov [FPK^{+}24]. Marvel [CKP^{+}22]. Masking [BAZ^{+}19, WPJ19]. Masses [BCHC19]. Massive [WZW^{+}24]. Massively [MCB^{+}12, RLBBN15]. Matcher [WWGS22]. Matching [EZYA23, UJW15, WMGS19, WWGS22, CW13, PLL10, TBS06, VW11]. Mathematical [Mic16, VZT^{+}20]. MATOG [WG17]. Matrices [BHWN21, ICEDR24].

Matrix

[ASH20, ACGH24, BSL17, HBPH24, JYM20, LCW^{+}23, MLC^{+}23, MFSK24, SMN22, WCW^{+}24, YAG^{+}16, CYXF13, SJV08]. Matrix-Vector [YAG^{+}16]. maximize [RCG^{+}10a]. Maximizing [AEJE16, LWF^{+}16, LDL22]. Maxine [WHV^{+}13]. MaxPB [LWF^{+}16]. MBZip [KPM17]. MC [CPG21]. MC-DeF [CPG21]. McPAT [LAS^{+}13]. means [Abd20]. Measuring [FMY^{+}15]. Mechanism [CEP^{+}16, SPS17, WLX24, ZHS^{+}19, ZCCD16, GB06, HWX^{+}13, KS11, RFD13, SBC05]. mechanisms [HWH^{+}11, LCL^{+}14, LMMM08]. Mechanistic [BEE15, CHE^{+}14]. media [SLA^{+}07]. Meets [WLJ^{+}24, KHL^{+}13]. MemHC [WPR^{+}22]. Memoization [SSRS15]. Memories [BKM^{+}17, DGI^{+}14, KRHK16, LCW^{+}24, SPM17, TZK18, WDX15, YMM^{+}15, CCZ13, DXMJ11, LCC11]. Memory [ADGA20, AJK^{+}12, AYC16, Akr21, ACG24, AEE^{+}19, AHA^{+}19, BSG23, BAZ^{+}19, CSY20, CSF^{+}20, CKPH19, CWMC16, CHD^{+}23, CLA^{+}19, CG15b, CSK19, DHK18, DD16, DHD^{+}14, DWF^{+}23, DJZ^{+}23, ERAG^{+}16, ELE^{+}23, EE09, FMY^{+}15, GHH15, GMGZP14, GHS12, HNKK17, HPBS21, HEDH21, HHC^{+}16, HASA16, JDZ^{+}13, JML^{+}20, JLJ^{+}18a, JRH21, KHB^{+}20, LYK^{+}15, LGP^{+}16, LWS^{+}19, LWC^{+}22, LP17, LSH^{+}23, LZS^{+}24, MYG15, MYKG16, NRQ16a, NRQ16b, NTV^{+}22, NZ15, OTR^{+}18, OKJ^{+}22, PWE20, PDY^{+}23, RJK24, RKC^{+}20, RLBBN15, SBC^{+}22, SW17a, SMKH15, SJD22, SKP^{+}22, SDZ^{+}21, SL20, SJL^{+}20, TKKM15, USCM16, WSJ^{+}21, WWH^{+}16, WLL^{+}19, WPR^{+}22, WZW^{+}24, WLX24, WGL^{+}24, WDS^{+}24, WJXC17, WX24, WZG^{+}19, WYZ^{+}23, XGD^{+}23, XHJY16, XVT20, XHL^{+}24, YXS^{+}22, YSH^{+}22, YBSY19, ZZB^{+}19, ZFT^{+}18, ZLYW18, ZLWS24, ZLC^{+}15, ZCQ^{+}19, ZDC^{+}16, ZWL^{+}19, ZSM^{+}16, ZPR^{+}17, AFD12, ATGN^{+}13, CS10, CCZ13, DHC^{+}13, DJX13, DZC^{+}13, FQRG13]. memory [GPL^{+}05, JSH09, JSM^{+}04, KGK10, KCKG14, LAS^{+}08, LGAZ07, LFX09, LCL^{+}14, LHC^{+}24, LHWB12, MA08, NCQ14, PLL10, PCT12, RLS13, SV05, SL09, TBC^{+}12, TGAG^{+}12, VDSP09, VED07, WKCS12, WWWL13, WSC^{+}13, WWL^{+}21, WLZ^{+}10, YJTF13, YLTL04, YLW08, ZPC06, ZSLX13, ZDC^{+}12]. Memory-access-aware [CLA^{+}19]. Memory-Aware [SJD22]. Memory-centric [SJL^{+}20, XGD^{+}23]. Memory-Disk [LYK^{+}15]. Memory-efficient [WYZ^{+}23, PLL10]. Memory-level [EE09]. Memory-Reliability [NRQ16b]. Memory-Side [AHA^{+}19]. MemTracker [VDSP09]. mer [WZX^{+}24]. merge [DDU12]. Merging [TS05, SSU^{+}13]. Merkle [WDK^{+}24]. Message [Kae20, ZM15]. Message-Passing [ZM15]. Meta [BJWS18]. Meta-Format [BJWS18]. Metadata [CKL^{+}24, LLW^{+}22, VOK^{+}22]. metafunctions [LT13]. MetaStrider [SJL^{+}20]. MetaSys [VOK^{+}22]. Metering [LMA^{+}16, LMJ^{+}13b]. Method [ADGA20, KTAE16, YHYBAM20, CWCS13, SHC13]. Methodology [TCS16]. Metric [LLLW22, SNN^{+}19, SPS17, YHYBAM20]. Metric-Guided [YHYBAM20]. Metrics [EMR14, TDO16a]. MFFT [ZLM^{+}23]. MH [PLK^{+}19]. MIAOW [BGG^{+}15]. Micali [WDK^{+}24]. MiCOMP [ABP^{+}17]. Micro [CAGS17]. Micro-Sector [CAGS17]. Microarchitectural [FMY^{+}15, SKS23, DJB13, LB10]. Microarchitecture [DHX^{+}22, LNFE22, MMS15, ZCDD23, ASK13, HS05, RPS06, SSS^{+}04]. Microarchitecture-Aware [DHX^{+}22]. microarchitectures [ACGK04].

Microbenchmarking [FMY^{+}15].

Microprocessor

[KCA^{+}13, BE13, YCCY11].

Microprocessors

[GSZY20, SDZ^{+}21, BSO07, RCG^{+}10a]. MicroProf [TMSR23]. Microservice [TMSR23]. Mid [MSFC21]. Mid-Kernel [MSFC21]. Migration [JLJ^{+}18a, LTX16, MK23, WLL^{+}19, LJMG12, MSF^{+}07]. Million [CAY^{+}18]. MIMD [FSYA09, GSZY20]. MInGLE [GáSÁ^{+}16]. miniature [JEBJ08]. minimal [XL07]. MINIME [DS16]. MINIME-GPU [DS16]. Minimization [KMAK22, CH06, SSR13]. Minimizing [KHB^{+}20]. mining [CDPD13]. Minor [TCR^{+}22]. Minos [CWC06]. MIPS [SHD15]. misaligned [LWH11]. Mismatches [APBR16]. misprediction [GW08]. Miss [SMM^{+}23, SWO21, SLP08]. misses [CST^{+}06, LS10, VHKP11, Zha08]. Mitigating [ABP^{+}17, DHX^{+}22, EPAG16, GBD21, SYX^{+}15, LCL^{+}14]. mitigation [DJL^{+}12]. mitigations [CCD12]. Mixed [ASH20, LLC22, XIC12, ZLM^{+}23]. Mixed-Precision [ZLM^{+}23]. Mixing [HDW21]. MLC [PM17, RJSA18]. MLC/TLC [PM17]. MLIR [BKS^{+}22, EZYA23]. MLP [KABS22]. Mobile [CNAA^{+}22, PLK^{+}19, XZC^{+}20, AvRF07, TBC^{+}12]. Mobile-cloud [XZC^{+}20]. Mode [HBPH24, SW13]. Model [BNS^{+}21, CC18, DAKK19, ESR^{+}15, GGS^{+}17, JHH^{+}23, NZ15, SRC16, WDW^{+}22, WLLW20, WYZ^{+}23, XHJY17, YCA18, ZHB18, DC07, MG13]. Model-Based [WLLW20]. Modeling [BEE15, KR19, LAS^{+}13, LL22, SSC^{+}13, ZZL^{+}21, AFD07, CA11, EE12, IMS^{+}08, XMM04, SSS^{+}04]. Models [BOEN23, CHE^{+}14, FCD^{+}17, FPK^{+}24, GGS^{+}19, GHH15, LCP^{+}21, VFW16, XZC^{+}20, LAS^{+}08, XIC12]. Modern [ABK21, HYYBM16, WLJ^{+}24, CCD12, JK13, KNBK12]. Modification [GDL16]. Modify [RLS15]. Modular [RKL23]. Modulo [LMSE18, KCP13]. Moldable [MKKE15]. Monitoring [LHC^{+}17, MMGS21, LMMM08, VDSP09, ZZQ^{+}05]. Monolithically [WSJ^{+}21]. monopolizable [DJL^{+}12]. Moore [DSH^{+}18]. Morphable [CKPH19]. Most [PLT^{+}15]. Movement [ESR^{+}15]. Movements [YXS^{+}22]. Moving [DAKK19]. MP [WLZ^{+}13]. MP-Tomasulo [WLZ^{+}13]. MPI [HWX^{+}13, MP13]. MPSoC [FPMR21]. MPSoCs [DMR^{+}16, MMGS21, SL20]. MPU [XGD^{+}23]. MRAM [WDX15]. MRAM-Based [WDX15]. MSHRs [CA11]. MUA [LDL22]. MUA-Router [LDL22]. Multi [BOEN23, CC18, FLW^{+}24, FMY^{+}15, FCD^{+}17, GVT^{+}17, GMZ^{+}21, JPS17, JML^{+}20, KLA^{+}19, LT19, LGP^{+}16, MMGS21, PLK^{+}19, PGB16, SPS17, TCS20, WZZ^{+}20, XMW^{+}21, ZNTJE23, ZCF18, vdVSAAS20, CDPD13, GWS13, LFC13, PM12, RB13, RPE12, ZGC^{+}12]. Multi- [FMY^{+}15]. Multi-Agent [JPS17]. Multi-Batch [FLW^{+}24]. Multi-Chip [ZNTJE23]. Multi-Core [CC18, SPS17, PM12, ZGC^{+}12]. Multi-Cores [ZCF18]. Multi-CPU [PGB16]. Multi-dimensional [LT19]. Multi-directional [XMW^{+}21]. multi-FPGA [CDPD13]. Multi-GPU [vdVSAAS20, LFC13, RB13]. multi-issue [GWS13]. Multi-Layer [LGP^{+}16]. Multi-Level [GMZ^{+}21]. Multi-objective [BOEN23]. Multi-pipeline [WZZ^{+}20]. Multi-programming [LD24]. Multi-retention [PLK^{+}19]. multi-server [RPE12]. Multi-Tenant [FCD^{+}17, KLA^{+}19]. Multi-Threaded [GVT^{+}17]. Multi-Threading [TCS20]. Multi-Tile [MMGS21]. Multi-type [JML^{+}20]. Multibank [CG15b]. Multiblock [KPM17]. multicharacter [CW13]. Multicore [ASV^{+}16, AMS23, BHC^{+}16, CC13, CG15a, CDPN16, DS16, DAKK19, HMYZ15, HEMK17, KE15, KK15, KMAK22, LAS^{+}13, LMA^{+}16, LYH16, MST^{+}21, NMPS22, OK21, PT17, PGB16, SOAK23, SLJ^{+}18, SCMU22, SKH^{+}16, SAL19, ZDC^{+}16, CG14, CK11, CWCS13, DEE13, FBWS13, HWX^{+}13, LMJ^{+}13b, LCL^{+}14, LHZ13, RCG^{+}10a, VE13, WFKL10, ZCW10]. Multicores [HK14, PB15, TDO16a, TTS19, MSF^{+}07]. multidimensional [RTG^{+}07]. Multigrain [AZG17]. Multigrid [WCW^{+}24]. Multilevel [XHJY16, YMM^{+}15, JK13, TKJ13]. multimedia [SV05]. multiobjective [CPP08]. multiplatform [HLC10]. Multiple [KKL^{+}24, KHN^{+}18, LY24, WLLW20, ZSM^{+}16, GB06, HVJ06, RCV^{+}12]. Multiplexing [LGH24, NDP17]. Multiplication [ASH20, HBPH24, MFSK24, SMN22, YAG^{+}16]. Multiplications [JYM20]. Multiply [GG18, YXMC23]. Multiply-Accumulate [GG18]. Multiply-and-Fire [YXMC23]. multiprocessor [BBG13, GSZI10, LT13]. Multiprocessors [CPS^{+}15, LBM13, APG13, GPL^{+}05, LAS^{+}08, LM05, LPZI12, LMMM08, SMK10]. Multiprogram [EMR14]. Multisocket [CG15a]. Multithreaded [AZG17, JYE^{+}16, LYH16, DWDS13, GMW09, NTG13, PGB13, RGG^{+}12, RCG^{+}10a, XIC12]. Multithreading [SDK^{+}22, EE09, GWM07]. Mutation [WZY^{+}24a]. Mutation-Based [WZY^{+}24a]. Mutually [SOAK23].

Namespace [HFL^{+}23]. NAND

[DGI^{+}14, LCW^{+}24, SZJK18, ZWL^{+}19].

Nanoscale [GBD^{+}15]. Native [DKK^{+}21, RPE12]. Native-Code [DKK^{+}21]. Natural [SNK^{+}23]. Near [AGG22, HK14, KCA^{+}13, LP17, MAD17, RJK24, VFJ^{+}17, XGD^{+}23, KCKG14, RPE12]. Near-bank [XGD^{+}23].

Near-Data [VFJ^{+}17, AGG22]. Near-Memory [LP17, RJK24]. Near-Optimal [KCA^{+}13, KCKG14]. Near-Threshold [HK14]. Nearest [NSF^{+}21]. Nearest-Neighbor [NSF^{+}21]. Need [ZPR^{+}17]. Neighbor [NSF^{+}21]. NEM [RJK24]. NEM-GNN [RJK24]. nest [SLM12]. Nested [GSZY20, MGSH16, KLMP12]. nests [NCC13]. Network [CEP^{+}16, DJC16, EPS18, JHHM21, JPS17, LWW^{+}24, PWE20, SSH19, TDP15, VFW16, VZT^{+}20, WDX24, XSF^{+}23, XVT20, YXMC23, YZZ^{+}23, ZCCD16, ZM15, ASK13, LNLK13, LYYB07]. Network-on-Chip [CEP^{+}16, DJC16, EPS18]. Network-on-Chips [ZM15]. Networks [ACA^{+}19, AMP^{+}16, CVB15, CSF^{+}20, CCL^{+}24, CRC^{+}21, GG18, GLTV23, GR15, MWJ19, RJK24, RKC^{+}20, RSK^{+}18, SMS23, SMN22, SAM^{+}23, SNK^{+}23, WYZ^{+}23, XMW^{+}21, ZFF^{+}18, ZMLL24, ZF24, BKA13, LWWH12, PRMH13, SMK10, SEP07]. Networks-on-Chip [CCL^{+}24, LWWH12]. Neural [BOEN23, CRC^{+}21, GG18, GLTV23, GR15, JHHM21, MWJ19, PWE20, RJK24, RKC^{+}20, RSK^{+}18, SMS23, SMN22, SAM^{+}23, SNK^{+}23, TDP15, WYZ^{+}23, XMW^{+}21, XSF^{+}23, YXMC23, YZZ^{+}23, ZFF^{+}18, ZMLL24, Jim09]. Neuromorphic [DLS22, LCS^{+}19]. Next [VZT^{+}20, OAM19]. NISQ [LD24]. no [HL07]. NoC [HWX^{+}13, LLC22]. NoC-based [HWX^{+}13]. NoCMsg [ZM15]. NOCs [RKL23, WYJL10]. Node [CSSU21]. Noise [AAI^{+}16]. Non [AEE^{+}19, DJL^{+}12, FPMR21, HK14, IPSD21, WSJ^{+}21, YKM17, YXS^{+}22, BZS13, WDXJ14]. Non-blocking [IPSD21]. Non-Intrusive [FPMR21]. Non-Java [YKM17]. Non-monopolizable [DJL^{+}12]. non-SSA [BZS13]. Non-Uniform [HK14]. Non-Volatile [WSJ^{+}21, YXS^{+}22, AEE^{+}19, WDXJ14]. Nonaffine [SGS^{+}20]. Nonblocking [MAY23]. Nonlinear [SRC16]. nonuniformity [WA08]. Nonvolatile [SPM17, DXMJ11, DJX13]. Not-taken [PS12]. Novel [HLZ^{+}24, JZY^{+}22, LMZ18, SAM^{+}23, TPN^{+}20, ZFT^{+}18, ZWL^{+}19, CCZ13]. NUCA [GFD^{+}14, HK14, LJMG12]. NUCA-L1 [HK14]. NUMA [RSK^{+}18]. NUMA-Aware [RSK^{+}18]. NUMA-Caffe [RSK^{+}18]. Number [AGG21, DNS^{+}24]. NVIDIA [KVH23]. NVM [EAH^{+}20, LLW^{+}22, WSC^{+}13]. NVM-based [WSC^{+}13]. NVMs [PM17]. NVRAM [ZLYW18].

O [DCP^{+}12, RHLA14, SLS^{+}21, WLX24, WZWW23]. Obfuscation [PDCS24]. Object [CKL^{+}24, MPHL22, YLW08, ZLYW18, TDG13, VED07, WM10]. Objective [SAT20, BOEN23]. objects [WWY^{+}12]. Oblivious [YRGES^{+}19, CYXF13, LWC^{+}22]. Obstruction [WDX14]. Occam [GLTV23]. Occluded [CNAA^{+}22]. Occluded-geometry [CNAA^{+}22]. Occurring [LTX16]. Odd [MP22]. Odd/Even [MP22]. ODE [HLR^{+}13]. ODE-based [HLR^{+}13]. Off [ACA^{+}19, BKM^{+}17, DPBI^{+}19, AVG12, AGVO05, ABK21]. Off-Chip [BKM^{+}17]. Off-the-Shelf [DPBI^{+}19]. Offload [DZL^{+}24]. Offloading [HNKK17, JLL^{+}23, KKL^{+}24, MTK18, MGA^{+}17, WDX24]. offset [CZ07]. On-Chip [CS21, VFW16, JPS17, LDL22, SSH19, BKA13, CK11, EE11, LNLK13, SMK10, TDG13, XCC^{+}13]. On-GPU [LWL18]. On-Package [FBC^{+}22]. On-the-fly [WWY^{+}12, VHKP11]. On/Off [ACA^{+}19]. One [DWF^{+}23]. One-Sided [DWF^{+}23]. Online [BSO07, CG15a, CEP^{+}16, OKJ^{+}22, PSL^{+}23, TTS19, WAST16]. onto [WYJL10]. OoO [MAD17]. Opcode [AGG22]. Opcode-based [AGG22]. Open [BGG^{+}15, HKA^{+}19, VOK^{+}22, GMZ^{+}21]. Open-Source [BGG^{+}15, VOK^{+}22]. OpenCL [RVKP19, WGO15]. OpenMP [CLA^{+}19, PC13, YCA18]. OpenStream [PC13]. Operands [RSU^{+}20]. Operating [HK14]. Operation [LCW^{+}23]. Operations [BSL17, GGK18, LP17, WLX24]. Operators [CKP^{+}22, RAF22]. Opportunities [KKL^{+}24, WWW^{+}21, KGK10, XMM04]. Optane [Akr21]. Optical [CWW^{+}16]. Optically [NTV^{+}22]. Optimal [CH06, CBD15, GK13, GLTV23, KCA^{+}13, Mic16, SMS23, SW17b, SWH09, ZGP15, KCKG14, XC06]. Optimise [WZWW23]. optimised [RWFJ19]. optimising [LBO14]. Optimization [AYL^{+}18, ABP^{+}17, BSL17, CSF^{+}20, CZGC20, CHD^{+}23, CKL^{+}24, DZSL20, DAP^{+}15, FXC^{+}15, FSP^{+}23, GGS^{+}17, GGS^{+}19, JML^{+}20, JRH21, KTAE16, KKL^{+}24, LLLW22, LHS^{+}24, LVR^{+}15, MNC^{+}16, MMLS21, MPHL22, QSZ^{+}21, RMA14, SGM^{+}22, VFW16, WCW^{+}24, YKM17, YDL^{+}17, ZCF18, CFH^{+}12, CXW^{+}12, CYXF13, DJX13, FT10, GHS12, HS06, HEL^{+}09, HVJ06, JPS17, KHW^{+}05, KWTD09, LHC^{+}24, PJ13, SLM12, SSR13, SL09, VW11, ZFT^{+}18, ZWHM05, ZCS06]. optimization-phase [KHW^{+}05]. Optimizations [EPS17, JRK16, JZY^{+}22, PDY^{+}23, FCVR24, SHS^{+}20, THA^{+}21, VOK^{+}22, ZWS^{+}16, LCH^{+}04, LHY^{+}06]. Optimize [DBH16, FPMR21]. Optimized [MLC^{+}23, PKPM19, WPR^{+}22, WDS^{+}24, GS12]. Optimizer [LYK^{+}15]. Optimizing [AP17, BJWS18, CSW^{+}23, DGGL16, HHC^{+}16, JHH^{+}23, LCW^{+}24, MST^{+}21, PAVB15, RLBBN15, STLM12, SLH^{+}20, TN20, TKKM15, WDX15, WDW^{+}22, YWXW12, YRHBL13, ZSLX13, ZFF^{+}18, ZMLL24, YXK^{+}12, WK09]. optimum [HP04]. Orchard [SSS^{+}24]. Orchestrating [MG13]. Orchestration [GVT^{+}17]. Order [BEE15, CAY^{+}18, CCL^{+}24, HYYBM16, LLC22, MST^{+}21, MAD17, PS15, SPH^{+}17, TCS20, BB04, GGYK19, KWTD09, KABS22, SJA12, YJTF13, CCL^{+}24]. Order-Based [CCL^{+}24]. order/out [BB04]. Ordering [ABP^{+}17]. organization [ASK13, GGFPRG12]. Oriented [FWJ^{+}16, GGK18, BTS10, CXW^{+}12, JML^{+}20]. Orlando [ESB^{+}20]. OS- [CRSP09]. Out-of-Order [HYYBM16, MAD17, PS15, TCS20, GGYK19, BB04, SJA12]. overcoming [DZC^{+}13]. overflow [CH06]. Overhead [DSR15, GDL16, KRHK16, LHC^{+}17, ZCDD23, MP13]. Overheads [OK21, TCR^{+}22, BCM11, SSU^{+}13]. Overlap [ADGA20]. Overlap-and-Save [ADGA20]. Overlapped [ZC20]. Overlay [CSRP22, CGP23, JLER12]. Overlong [ZWL^{+}19].

P [DDT^{+}17]. Package [FBC^{+}22]. Packed [BSL17]. packet [LWWH12]. packing [NB13, SPGE06]. Page [TCR^{+}22, WLL^{+}19, WLX24, WZG^{+}19, ZCDD23, LMJ13a]. Pages [YGB21]. Paging [ACG24]. PARALiA [APG^{+}23]. Parallel [ASK^{+}16, ABB^{+}16, APS22, BHWN21, DTD16, DDT^{+}17, DHD^{+}14, HAM19, LCW^{+}23, MCB^{+}12, MPPS18, MN24, MGSH16, NKH16, PWPD19, RHC15, RLBBN15, SN17, SCFD22, TMP16, TPN^{+}20, UJW15, WLZ^{+}13, WGO15, WDK^{+}24, ZLJ18, CDPD13, JYJ^{+}13, LM05, NCC13, STLM12, VJC^{+}13, ZBH^{+}13]. Parallel-Blocked [BHWN21]. Parallelism [CCM^{+}16, CSRP22, CG15b, DHK18, GVT^{+}17, HWJ^{+}15, JHH^{+}23, LMZ18, MGA^{+}17, NKH16, SSS^{+}24, SDH^{+}15, WWW^{+}21, WZW^{+}24, YBSY19, ZX16, EE09, FLG12, PCT12, SLA^{+}07, WTFO14]. Parallelization [BCM11, DPBI^{+}19, GGS^{+}17, GSZY20, KPP^{+}15, DC07, LT13, PKC12, YRHBL13]. Parallelizing [NKH16]. Parallelogram [ZGP15]. Parameter [MGI15]. parametric [SLM12]. Pareto [BOEN23, SW17b]. PARSEC [CCM^{+}16]. PARSECSs [CCM^{+}16]. parser [ZBH^{+}13]. Parsing [PCM16, ZBH^{+}13]. PARTANS [LFC13]. Partial [SLS^{+}21, ZX16]. partially [GGFPRG12, JLER12]. Partition [WWC^{+}16, WJXC17, WO13]. partitioned [RPS06]. Partitioning [CSW^{+}23, CG15b, FLG12, LDMZ19, PLLW24, SBS16, SLJ^{+}19, HAJ^{+}12, LCL^{+}14, ZDC^{+}12]. Pass [SPS17]. Passing [ZM15]. PATCH [RBM10]. Path [WZWW23, ZX19, TS05]. paths [PS12]. pattern [CXW^{+}12, KPM21, PRMH13, VW11]. pattern-oriented [CXW^{+}12]. pattern-specific [PRMH13]. patternized [KCP13]. Patterns [CSK19, DDT^{+}17, LWS^{+}19, LTX16, PWE20, Pro21, UJW15, HLR^{+}13, JSH09]. pausing [NCQ14]. PAVER [TAB^{+}21]. PCantorSim [JYJ^{+}13]. PCIe [MTK18]. PCM [LWF^{+}16, RJSA18]. penalties [HL07]. Penalty [SWO21, GW08]. pending [CA11]. Per-Core [CS21]. per-task [LMJ^{+}13b]. Per-thread [DEE13, BTS10]. perceptron [TS05]. Perfect [BRJM15]. Performance [ABK21, AEJE16, Akr21, APG^{+}23, AYL^{+}18, Bis21, BEE15, DKK^{+}21, DVG^{+}23, FDF^{+}14, GGS^{+}19, GGK18, HMYZ15, JGSM15, KR19, LNFE22, LL22, LCW^{+}24, LMZ18, LYH16, LGH24, LY16, MAY23, ME17, MTK18, MPU^{+}23, MN24, MAD17, MKS22, NDP17, NMPS22, OBdOJ^{+}24, OK21, PDY^{+}23, Per18, RVOA08, RJSA18, SCFD22, SWO21, SLJ^{+}19, TCS16, TKM14, USCM16, WCI^{+}16, WLWB19, XHJY17, XFS^{+}19, YGB21, ZFT^{+}18, ZXX23, ZYCZ10, ZCF18, AFD12, ATGN^{+}13, BSWLE13, BTS10, CHD^{+}23, CK11, CRSP09, CDM13, FBWS13, GW08, HP04, HL07, JED19, KBR^{+}13, KLMP12, KGK10, LM05, MLB^{+}23, PGB12, RWY13, SRLPV04, SSP^{+}23, SD12, THA^{+}21, WKCS12, WZZ^{+}20, WWL^{+}21, XT09, YRGES^{+}19, YCCY11, ZVYN05].

Performance-aware [ZYCZ10]. performance-driven [XT09]. Performance-Energy [ABK21, HMYZ15]. performance-friendly [CRSP09]. PERI [TGRK21]. permanent [SSC^{+}13]. Permissions [ERAG^{+}16]. Permutation [ZX19]. Permutation-Based [ZX19]. Persistence [EAH^{+}20, WZG^{+}19]. Persistent [CHD^{+}23, IPSD21, WDS^{+}24, WX24, ZLYW18]. Perspectives [PLT^{+}15]. PEs [WWY^{+}24]. PETRA [IPSD21]. PGAS [KFEG18, SKAEG16]. Phase [ABP^{+}17, HASA16, JDZ^{+}13, NTV^{+}22, YMM^{+}15, KHW^{+}05, KWTD09, ZDC^{+}12]. Phase-Change [YMM^{+}15]. Phase-Ordering [ABP^{+}17]. phased [HLR^{+}13]. Photonic [DH16, DLS22]. Phronesis [LL22]. Physically [RKL23]. PICO [JRH21]. PiDRAM [OLK^{+}23]. Piecewise [DAP^{+}15]. pilot [DHX^{+}22]. PIMBALL [RKC^{+}20]. PiPA [ZCW10]. Pipeline [AKBS21, ZJJ^{+}15, HP04, JA14, WZZ^{+}20]. pipelined [PLL10, ZCW10]. Pipelines [CGP23, MG19, MG20, SSW^{+}19, SSB^{+}20]. Pipelining [LDL22, CPB14, JSL13, RVOA08, RTG^{+}07]. place [GS12]. Placement [MNSC16, RB24, MA08, SSK11]. Places [Per18]. Plane [DSK19, LLS23, ZGC^{+}12]. Platform [MLC^{+}23, QSZ^{+}21, ZLYZ16, LGH24]. Platforms [RVKP19, SCK^{+}21]. PLDS [FLG12]. PM [CHD^{+}23]. PM-aware [CHD^{+}23]. Point [ASS17, BWG^{+}12, CS13, DNS^{+}24]. pointer [SV05, YLTL04]. pointer-intensive [YLTL04]. Pointers [CTY^{+}23]. points [Nas13]. points-to [Nas13]. Poker [ZX19]. Policies [GFD^{+}14, LSL20, SYX^{+}15, EE09, SSK11]. Policy [KMAK22, LWW^{+}24, JK13]. Pollution [SYX^{+}15]. PolyDL [THA^{+}21]. Polyhedral [GGS^{+}19, KL19, LT19, PKC12, SYE19, SGS^{+}20, SRC16, THA^{+}21, VJC^{+}13, ZC20, ZPH^{+}23, ZHB18]. Polyhedron [GGS^{+}17]. polymorphic [PM12]. polymorphous [SNL^{+}04]. polytopes [SLM12]. Pools [HZC^{+}23]. Port [WDX14, GKP14]. Portability [FDF^{+}14]. Portable [BNS^{+}21, Per18, RMA14, WGO15, KNBK12]. Posit [TGRK21]. positioning [ZWHM05]. Post [WDK^{+}24]. Post-Quantum [WDK^{+}24]. Pot [VSDL16]. potential [FER^{+}13]. POWAR [ACA^{+}19]. Power [AEJE16, ACA^{+}19, BNS^{+}21, CAMJ15, DTD16, DD16, ESB^{+}20, FCD^{+}17, GáSÁ^{+}16, GBD^{+}15, HYBR^{+}15, HYYBM16, HAC13, JYW22, JGSM15, KH18, KMG14, LM05, LAS^{+}13, LWF^{+}16, LZM14, MKS22, RWFJ19, SEF^{+}19, SSS^{+}23, WYCC11, ZCF18, ZZL^{+}21, AVG12, AGG22, BB04, CCZ13, HP04, HL07, LYYB07, MP13, MSK05, PLK^{+}19, SW13, SEP07, WYJL10, XL07, YCCY11]. Power-Aware [ACA^{+}19, DTD16, SEP07, WYJL10]. Power-Efficient [HAC13, KH18]. Power-Gated [LZM14]. Power-Gating [ZCF18]. Power-optimised [RWFJ19]. Power-performance [LM05]. Power/Capacity [GBD^{+}15]. POWER8 [XFS^{+}19]. PowerMorph [JYW22]. Practical [FXC^{+}15, KWTD09, OBdOJ^{+}24, PDY^{+}23, VOK^{+}22, ZGX22, BSWLE13, FT10, ZBH^{+}13]. pre [YCCY11, XC06]. pre-wakeup [YCCY11]. Preallocation [SSR13]. Precise [AFD07]. Precision [ASH20, CCCA20, WLJ^{+}24, ZLM^{+}23, AGG21, LDG^{+}13]. Precisions [HDW21]. Predicate [CPB14]. Predicate-aware [CPB14]. Predication [HAC13]. Predictability [BB21, LBJ05]. Predictable [DPBI^{+}19, SF18, VKM^{+}21, XHJY17]. Predicting [WLWB19]. Prediction [AKBS21, BNS^{+}21, EPS17, GAM12, KS21, MKS22, OAM19, PLG19, YPT^{+}16, CST^{+}06, Jim09, MG12, TS05]. Predictive [LCP^{+}21, IMS^{+}08, RBM10, YCCY11]. predictive/adaptive [RBM10]. Predictor [CNAA^{+}22, Mic18, OAM19, AGVO05, JSM^{+}04, SL09]. Predictors [EPAG16, LIS20]. Prefetch [AKBS21, SPS17]. Prefetch-Fraction [SPS17]. Prefetched [SYX^{+}15]. Prefetcher [KPP21, LYH16, PB15, PWE20, SYX^{+}15, XHL^{+}24, LJMG12, SBC05]. Prefetcher-Caused [SYX^{+}15]. Prefetchers [ELE^{+}23, LBM13]. Prefetching [CSY20, KFEG18, LKV12, OAM19, SPS17, WPJ19, AGI^{+}12, CA11, GB06, SBC05, WFKL10, YLTL04]. Presburger [JRH21]. Preserving [YXS^{+}22, BOEN23, CPB14]. Pressure [KMAK22, SKPD19, SGM^{+}22, SLP08, SSR13, YZ08]. Pressure-Aware [KMAK22]. Preventing [WDX14]. prevention [TBS06]. Primitives [THA^{+}21]. Priority [ASV^{+}16, XHJY16]. PRISM [OK21]. Private [DRHK15, SSK11]. Private/Shared [DRHK15]. Pro [BSG23, FYI^{+}24]. Probabilistic [DAD16, EE12]. Problem [ABP^{+}17, DBH16]. Problems [JOB^{+}22, VFW16]. Process [LCW^{+}23, LTX16, Pro21, KWCL09]. Processes [SOAK23]. Processing [CC13, CGP23, DZL^{+}24, FLW^{+}24, FYI^{+}24, HNKK17, ICEDR24, LT19, LSH^{+}23, LGH^{+}21, MYG15, MYKG16, OLK^{+}23, PBY^{+}17, SNK^{+}23, WZZ^{+}20, WWL^{+}21, ZPL^{+}21, ZZH^{+}23, ZLJ18, KKL^{+}24]. Processing-in-DRAM [OLK^{+}23]. Processing-In-Memory [HNKK17, MYKG16, MYG15]. Processor [AEJE16, AHA^{+}19, BEE15, DSK19, HMYZ15, HWL^{+}19, JYM20, JHH^{+}23, LP17, LZZ^{+}22, SKP^{+}22, XGD^{+}23, XFS^{+}19, ZZL^{+}21, CS13, GW08, LGAZ07, LYYB07, SJA12, SHC13, SSPL^{+}13, WFKL10]. Processor-Memory [SKP^{+}22].

Processor-Tracing [HWL^{+}19]. Processors [ASV^{+}16, AMS23, CAMJ15, DBH16, KS16, KK15, MRK^{+}22, NMPS22, SM19, SCK^{+}21, SHD15, VFJ^{+}17, YWXW12, YHYBAM20, CRSP09, CCD12, CSVM04, DEE13, EE09, EE12, FBWS13, GMW09, GWS13, GKP14, HWX^{+}13, KLMP12, LMCV13, PI12, RGG^{+}12, SRLPV04, SLP08, XT09, YZL^{+}10]. Productive [KFEG18]. Productivity [SKAEG16]. Profile [CS13, FPK^{+}24, SS04, SKKB18, SSU^{+}13, WTFO14]. Profile-based [SS04, SKKB18]. profile-driven [WTFO14]. Profile-guided [CS13, SSU^{+}13]. Profiling [CG15a, JRK16, MPW^{+}17, FBHN04, MAN^{+}08, NMKS06, ZCW10]. profit [ZCS06]. profit-driven [ZCS06]. Profitability [CLA^{+}19]. Program [BB21, DSR15, PDCS24, PVA^{+}17, RAF22, ZHB18, DS12, PJ13]. Programmable [MCB^{+}12, AS13, Zha08]. Programming [ÄJE^{+}16, MGSH16, PBY^{+}17, RGK^{+}23, RAF22, TWB21, YCA18, LD24, NCC13]. Programming-Based [ÄJE^{+}16]. Programs [DKB^{+}20, GKCE17, KPM21, KPP^{+}15, LLS23, MPPS18, MNSC16, RHC15, SGS^{+}20, WLZ^{+}13, WGO15, PC13, PGB13, WO13, YLW08]. Projection [TTS19]. promotion [LJMG12]. Proportional [DH16]. proportionality [AVG12]. proprietary [JEBJ08]. protect [BVIB12]. Protecting [NRQ16a, CWC06]. Protection [AHA^{+}19, BCHC19, Bis21, ERAG^{+}16, CCZ13, MA08]. Protocol [WDX24, SSPL^{+}13, SSH^{+}13]. Providing [KKL^{+}24, XHJY17]. Provisioning [BSSS14]. PS [LMJ13a]. PS-TLB [LMJ13a]. pseudo [YJTF13]. pseudo-associativity [YJTF13]. Public [SDS^{+}21, WLWB19]. Puppeteer [ELE^{+}23]. Purification [WZY^{+}24b]. Purpose [CAMJ15, SDZ^{+}21]. Push [LSH^{+}23, YLTL04]. Push-Memory [LSH^{+}23]. Python [ZXX23].

QoS [ASP17, FYI^{+}24, JYW22, LPZI12, NMPS22, SAL19]. QoS-Aware [JYW22]. QoS-Constrained [NMPS22]. QoS-enhanced [FYI^{+}24]. QoS-pro [FYI^{+}24]. QoS-Supervised [ASP17]. quadruple [LDG^{+}13]. quadruple-precision [LDG^{+}13]. Quality [APS22, GSZI10]. Quantifying [DVG^{+}23, LZW23]. Quantitative [ACG24, TCS16]. Quantized [WLJ^{+}24]. Quantum [LD24, Lou19, SM19, WDK^{+}24, IWP^{+}04]. quasi [JSM^{+}04]. quasi-static [JSM^{+}04]. Qubit [LD24]. QuCloud [LD24]. Query [KVH23]. Queue [HLSW17, BB04]. QuMan [SKKB18].

R [VC16]. R-GPU [VC16]. Race

[LHC^{+}17, MNSC16, YZZ^{+}23]. Racetrack [KHB^{+}20]. rack [MKC^{+}22]. Radio [DMR^{+}16]. radix [ASK13, LDY^{+}21]. RAGuard [ZHS^{+}19]. RAM [CRC^{+}21, LZL^{+}13, PLK^{+}19, RTK15, WDX14]. Random [ELE^{+}23, HLZ^{+}24, VSP^{+}12]. Range [CTY^{+}23]. ranges [MAN^{+}08]. Rank [AJK^{+}12, BOEN23]. Rank-preserving [BOEN23]. Rate [CWMC16, EPS18, SWO21, SHD15]. RATT [CWMC16]. RATT-ECC [CWMC16]. Rcmp [WGL^{+}24]. RDMA [DWF^{+}23, LZS^{+}24, WGL^{+}24, WDX24]. RDMA-Based [WGL^{+}24, DWF^{+}23]. RDMA-enabled [LZS^{+}24]. Reach [JED19]. Reactions [PBCB22]. Read [MNSC16, RJSA18, RLS15, JLCR13]. Read-Modify-Write [RLS15]. read/write [JLCR13]. Real [CCL^{+}24, CEP^{+}16, DPBI^{+}19, DJZ^{+}23, KE15, KTAE16, GK13, YZ08, ZGC^{+}12]. Real-Time [CCL^{+}24, CEP^{+}16, DPBI^{+}19, KE15, KTAE16, GK13, ZGC^{+}12]. Real-world [DJZ^{+}23]. Reasoning [DKB^{+}20]. reassignment [CH06]. recency [VSP^{+}12]. recognition [KKM^{+}13]. recompilation [NED^{+}13]. Recompute [AEE^{+}19]. Reconciliation [TWB21]. Reconfigurable [DBH16, KHS^{+}14, LMSE18, LY24, PT17, RJK24, TD16, VC16, VKM^{+}21, AS13, KLMP12, KCP13, ZSLX13]. Reconfiguration [DTD16]. Reconstructability [BRJM15]. Reconstructing [WGL^{+}24]. Recovery [LHY^{+}06, RHLA14]. Recycling [KKAR16, WWY^{+}24]. ReDirect [PT17]. Reduce [ASP17, DSR15, SLS^{+}21, ZCCD16, YZ08]. Reduced [CS21, VED07]. Reducing [AMS23, CPP08, DQCK24, GWS13, HL07, JLCR13, SLP08, TS15, TCR^{+}22, WLX24, ZHD^{+}04, Zha08, ZWS^{+}16, BCM11, MP13, PGB12, ZSCM08]. Reduction [ASS17, APS22, KTAE16, LSC^{+}15, LWL18, SJL^{+}20, MSK05, XT09]. Reductions [PWPD19]. Redundancy [YZZ^{+}23, ZNX^{+}24]. Redundancy-aware [YZZ^{+}23]. Redundant [KS16, SDK^{+}22, JLER12]. Reference [DCL^{+}22]. references [YZL^{+}10]. referent [WK09]. Refresh [JHHM21, LSC^{+}15, NCQ14, TKM14]. RegCPython [ZXX23]. Region [HWL^{+}19]. Register [JCG^{+}24, KPM21, LZM14, MP22, SKPD19, SGM^{+}22, TS15, TWB21, VZS^{+}18, YWXW12, YBSY19, ZXX23, BZS13, CH06, GKP14, JOA^{+}09a, JOA^{+}09b, JA14, SJV08, SLP08, SSR13]. Register-based [ZXX23]. Register-Pressure-Aware [SKPD19, SGM^{+}22]. registers [SCEG08, YZ08]. Regression [JGSM15, CDPD13]. Regular [BC13, JSH09]. Regulation [JYW22, SCFD22]. regulators [EE11]. ReHarvest [XLD^{+}24]. Reinforcement [JPS17, LSL20]. ReIPE [WWY^{+}24]. Relational [YDS^{+}19]. Relativization [BDB^{+}20]. Relaxed [GHH15, RJSA18, YJTF13]. relaxed-order [YJTF13]. release [GW09, JOA^{+}09b, SLP08]. Reliability [NRQ16b, SQZK20, FCVR24, ZFT^{+}18]. Reliable [CWMC16, KS16, KK15, ZLYW18, CPB^{+}07]. Remapping [LWL18, ZPC06]. Rematerialization [SMS23]. Remote [TN20, XVT20, NMKS06]. removal [BCVT13]. Removing [ACGK04, JCG^{+}24]. renaming [JA14]. Rendering [PLK^{+}19]. ReNIC [DCP^{+}12]. reordering [CZ07]. Repair [ZF24]. Repartition [LK24]. Replacement [DAD16, Mic16, FTLG11, TKJ13, WM11, ZDC^{+}12]. Replacing [DKK^{+}21]. Replay [CCL^{+}13, MPHL22, SKS23]. Replay-based [MPHL22]. REplayer [DAP^{+}15]. Replication [DWF^{+}23, ACGK04, DCP^{+}12]. Representation [SGS^{+}20, KCKG14]. representative [BE13]. requester [ATGN^{+}13]. requester-wins [ATGN^{+}13]. ReRAM [JLL^{+}23, XLD^{+}24, ZFT^{+}18]. ReRAM-Based [XLD^{+}24, JLL^{+}23, ZFT^{+}18]. ReSA [LY24]. ReSense [DWDS13]. Reshaping [JYW22]. resident [LHC^{+}24]. Residue [DNS^{+}24]. Resilience [KPRK20, TCS16]. Resilient [OK21, Pro21, SZJK18]. Resistance [RJSA18]. Resistive [MYKG16, TZK18]. Resource [APS22, CMAP22, CSW^{+}23, JZY^{+}22, LZW23, MRK^{+}22, MKC^{+}22, PLLW24, Per18, PS12, QYZ^{+}14, SAL19, XLD^{+}24, ARS04, DWDS13, GW08, NMKS06, VS11, ZK05]. resource-constrained [NMKS06, ZK05]. resource-efficient [GW08]. Resource-Harvesting [XLD^{+}24]. Resources [KLA^{+}19, MTPK24, SDS^{+}21, RGG^{+}12]. Restoration [DQCK24]. Retargetable [SHY14, HEL^{+}09, HLC10]. retention [PLK^{+}19]. Rethinking [ERAG^{+}16]. return [VS08]. Reusable [HPBS21]. Reuse [AKBS21, DAD16, GLTV23, HAM^{+}20, JLJ^{+}18a, KE15, KR19, LWW^{+}24, SCMU22, SSW^{+}19, AIVL13, FER^{+}13, YZL^{+}10, YLW08]. Reuse-aware [HAM^{+}20]. Reuse-Time-Label [LWW^{+}24]. ReuseTracker [SCMU22]. Reusing [PKPM19]. ReveNAND [SZJK18]. Reverse [SDS^{+}21]. Reviewers [Aca16, Ano13b, Ano15, Bil19, Ano13a]. Revisited [AMG16, MBY13, VS08]. Revisiting [GFD^{+}14, KAC15, MMS15, WWWL13]. Rewrite [SHS^{+}20]. Rewriting [EZYA23, GMZ^{+}21]. RF [TBC^{+}12]. RF-I [TBC^{+}12]. RFVP [YPT^{+}16]. RIMA [XVT20]. RISC [JHQ23, MPU^{+}23, TGRK21]. RISC-V [JHQ23, MPU^{+}23, TGRK21]. rNdN [KVH23]. RNN [SAG22]. Road [SWU^{+}15]. ROCCC [BCVN10]. Rollback [YPT^{+}16]. Rollback-Free [YPT^{+}16]. Roofline [ESR^{+}15]. ROP [ZHS^{+}19]. router [APG13, ASK13, LDL22]. Routers [LDY^{+}21, LDL22]. routes [KCP13]. Routing [ACA^{+}19, CVB15, LLC22, YRGES^{+}19, BRSJG12, PRMH13]. Row [GBD21, SL20, JLCR13]. Row-Buffer [SL20]. RRAM [LCS^{+}19]. RRAM-Based [LCS^{+}19]. RSA [LFK19]. RTL [BGG^{+}15]. Rules [SHS^{+}20]. Runtime [APG^{+}23, DBH16, DT17, ESB^{+}20, KPP^{+}15, LTG12, MMGS21, MSFC21, SSH19, TTS19, YAG^{+}16, YRHBL13]. Runtime-Reconfigurable [DBH16]. Runtimes [CMAP22].

Sabrewing [BWG^{+}12]. SAC [ZML^{+}24]. Safe [JCG^{+}24, YPT^{+}16]. Safe-to-Approximate [YPT^{+}16]. Safety [CLA^{+}19, JRH21, SBC^{+}22]. SAL [ZMLL24]. Salvaging [JDZ^{+}13]. Sample [WDW^{+}22]. Sampled [JYE^{+}16, HS05]. Sampling [Lee16, ZWS^{+}16, JYJ^{+}13]. Sandbox [PKL^{+}24]. SAQIP [SM19]. Save [ADGA20]. Saving [NMPS22]. Scalability [GVT^{+}17, LMZ18, CWCS13, RVOA08].

Scalability-Aware [GVT^{+}17]. Scalable [ASK13, CNS^{+}16b, KFJ20, MG19, Per18, RJK24, SM19, SYE19, SGS^{+}20, SMM^{+}23, SSS^{+}23, SJL^{+}20, TCS16, ZLYW18, ZLJ18, ZM15, CWCS13, KCKG14, LNLK13, LMJ13a, SSH^{+}13, VW11]. Scalar [SPH^{+}17]. Scalarization [LAAMJ15]. Scale [CAY^{+}18, DAKK19, JLJ^{+}18a, PLLW24, SKH^{+}16, YSD^{+}23, ZPL^{+}21, ZLM^{+}23, CKL^{+}24, HZC^{+}23, RCV^{+}12, SMK10]. Scale-out [YSD^{+}23]. Scaling [BHC^{+}16, GBD^{+}15, MKKE15, MBLK24, ZLC^{+}15, XMM04]. SCALO [GVT^{+}17]. Scenario [BB21]. Scenario-Aware [BB21]. Schedule [GGS^{+}17, GGS^{+}19, LMSE18, SSW^{+}19, SSB^{+}20, ZF24]. Scheduler [SCK^{+}21, TD16, USCM16, WZZ^{+}20, CWCS13, KCP13]. Schedulers [KKAR16]. Scheduling [ÄJE^{+}16, ASV^{+}16, CRBP24, CCL^{+}24, DHX^{+}22, DHD^{+}14, HAM^{+}20, LLC22, LWC^{+}22, LZM14, MKKE15, QYZ^{+}14, RB24, SLS^{+}21, SKPD19, SGM^{+}22, TAB^{+}21, UPR22, XHJY16, BBG13, CPB14, CG14, EE12, MBKM12, SPGE06, SWH09, SSR13, TBC^{+}12, XL07, ZGC^{+}12, ZYCZ10]. Scheme [AEE^{+}19, LLW^{+}22, LD24, SDK^{+}22, WPJ19, ZWL^{+}19, BBG13, CCZ13]. schemes [KCKG14]. Scientific [RGK^{+}23]. SCIN [NTG13]. SCIN-cache [NTG13]. SCORE [ZWL^{+}19]. SCP [SLJ^{+}19]. Scratchpad [JAK17, RTK15, SDZ^{+}21, YBSY19, CS10, LFX09]. script [KBR^{+}13]. script-based [KBR^{+}13]. Scythe [LZS^{+}24]. Seamlessly [KNBK12]. Search [BOEN23, CRBP24, HKA^{+}19, KL19, ZX19]. searches [KHW^{+}05]. SecNVM [LLW^{+}22]. SECRET [LSC^{+}15]. Section [DSR15]. Section-Based [DSR15]. Sector [CAGS17]. Sectored [CAGS17, OBdOJ^{+}24]. Secure [LLW^{+}22, SOAK23, CRSP09, SSPL^{+}13]. Securing [LIS20]. Security [SSH19, YGB21]. Segmented [LWW^{+}24]. Select [TDS^{+}21]. Selecting [BE13, TDO16b]. Selection [MNC^{+}16, SNN^{+}19, ZGP15, MBY13]. Selections [BAZ^{+}19]. Selective [DKB^{+}20, GGYK19, KMG14, LSC^{+}15, WPJ19, LWWH12, MA08, VSP^{+}12]. Self [LLRC17, MFSK24, SAL19, BBG13]. Self-aware [SAL19]. Self-Balancing [LLRC17]. Self-checking [MFSK24]. self-scheduling [BBG13]. SelSMaP [WPJ19]. Semantic [AP17, HCC^{+}14]. Sensible [LMA^{+}16]. Sensing [WCI^{+}16]. Sensitive [PDCS24, Nas13]. sensitivity [DWDS13]. Sensor [DSK19]. Sensor-Processor [DSK19]. Sequences [ABP^{+}17, MNC^{+}16, KHW^{+}05, PJ13]. Sequential [WLZ^{+}13, LZ12]. series [LTG12]. Server [AVG12, FCD^{+}17, JYW22, LTG12, RPE12]. Serverless [PKL^{+}24]. Servers [LTX16]. Service [APS22, GMW09, JYW22, RB24, GSZI10]. Set [KBB^{+}14, BR13, HL07, KWCL09, ZK06]. set-associative [HL07, KWCL09]. sets [DDU12]. setups [RPE12]. sFtree [BRSJG12]. SG [YGB21]. SGD [XDW^{+}23]. Shadow [ZGX22]. Shape [MWJ19]. Shared [ADGA20, DRHK15, FYI^{+}24, GKP14, GAH22, HMYZ15, KE15, LBM13, PG17, SKAEG16, SLJ^{+}19, WJXC17, XHJY16, AGI^{+}12, AIVL13, GGFPRG12, GSZI10, HLR^{+}13, KGK10, LHWB12, RGG^{+}12, WM11, ZPC06]. shared-data [HLR^{+}13]. shared-memory [ZPC06]. Shared-port [GKP14]. Sharing [FSX^{+}24, GG18, JAK17, KLA^{+}19, MTPK24, NSF^{+}21, YDL^{+}17, ZJJ^{+}15, ZNTJE23, SSK11]. Shelf [DPBI^{+}19]. Shifts [KHB^{+}20]. ShiftsReduce [KHB^{+}20]. shotgun [FBHN04]. showdown [SCEG08]. shuffler [BVIB12]. Side [AHA^{+}19, BCHC19, Bis21, JFK20, JHQ23, LFK19, BVIB12, DJL^{+}12]. Side-Channel [BCHC19, Bis21, JFK20, LFK19, BVIB12].

Sided [DWF^{+}23]. Signature [WDK^{+}24]. signatures [OAB12]. Significance [PVA^{+}17]. Significance-Aware [PVA^{+}17]. Significantly [MP13]. Silent [PLG19]. silicon [PCT12]. SIMD [BR13, DSK19, FSYA09, GS12, GSZY20, GR15, HEL^{+}09, KMG14, LHW^{+}19, MYG15, MYKG16, RMA14, SMKH15, WWC^{+}16, WWGS22, ZX19, ZX16]. SIMD-based [WWGS22]. SIMD-Matcher [WWGS22]. Simple [BNS^{+}21]. Simplifying [ZZB^{+}19]. SIMPO [ZLYW18]. SIMT [CC18, LAAMJ15, TCS20, XGD^{+}23]. SIMT-X [TCS20]. Simulating [RPE12]. Simulation [GMZ^{+}21, JYE^{+}16, PDCS24, QSZ^{+}21, SLJ^{+}18, SKP^{+}22, vdVSAAS20, HS05, JYJ^{+}13, RCV^{+}12]. Simulations [CAY^{+}18, HEMK17, JLJ^{+}18b]. Simulator [LCS^{+}19, NRQ16b, TPN^{+}20]. Simulators [JLJ^{+}18b]. Simultaneous [LGP^{+}16, WLLW20, EE09, RCG^{+}10a]. Simultaneously [LAS^{+}13]. Single [LD24, RTG^{+}07, TCS20, ZWY17, CG14, GB06, JK13, VE13, WK09]. Single-dimension [RTG^{+}07]. Single-Instruction [TCS20]. single-ISA [CG14, VE13]. single-referent [WK09]. Single/Multi [LD24]. Single/Multi-programming [LD24]. Site [HZN^{+}22]. Size [ZCDD23, MBY13]. Skeleton [NC15]. Skeleton-Based [NC15]. Sketch [XDXL19]. SketchDLC [XDXL19]. Skinny [BHWN21]. Skylake [HYYBM16, YHYBAM20]. Skylake-Based [HYYBM16, YHYBAM20]. Slack [NMPS22]. SLAP [LWW^{+}24]. Slice [KABS22, LNFE22]. Slice-out-of-order [KABS22]. Slicing [FSP^{+}23]. SLO [SCK^{+}21]. SLO-Aware [SCK^{+}21]. SLOOP [ASP17]. Slowdown [XHJY17]. SM [ZJJ^{+}15]. Small [WDW^{+}22]. Small-Sample [WDW^{+}22]. SMART [LLC22, AGVO05, WYZ^{+}23]. Smart-DNN [WYZ^{+}23]. SMT [EE12, LLC22, LMCV13, PLT^{+}15, SLP08, VS11, WA08]. SMT-Based [LLC22]. Snapshot [LDC15]. Snippets [SWU^{+}15]. Snug [HL07]. SoC [CWW^{+}16, ESB^{+}20]. SoCs [FDF^{+}14, SAL19]. Soft [BDB^{+}20, FWJ^{+}16, OK21, WWY^{+}24, LKL^{+}13]. Soft-Error [OK21, WWY^{+}24]. Software [Abd20, BCHC19, Bis21, DMR^{+}16, GSC17, LCL^{+}14, MGI15, RCV^{+}05, RWFJ19, SBS16, SDK^{+}22, SEP07, VCJ^{+}17, VZS^{+}18, WLLW20, YWXW12, ZGX22, CPB14, CS10, HWH^{+}11, HCC^{+}14, MMdS06, RVOA08, RCG^{+}10b, RTG^{+}07, TGAG^{+}12, YRHBL13]. Software-Based [ZGX22, LCL^{+}14]. Software-controlled [RCV^{+}05]. Software-Defined [DMR^{+}16, TGAG^{+}12]. Software-Directed [VZS^{+}18, SEP07]. software-guided [RCG^{+}10b]. Software-hardware [Abd20]. Software-level [SDK^{+}22]. Software-Managed [YWXW12]. Solution [WLLW20]. Solving [JOB^{+}22]. Some [KAC15, Mic16]. SortCache [SJC^{+}21]. Source [BGG^{+}15, EZYA23, HKA^{+}19, YRGES^{+}19, VOK^{+}22]. Source-adaptive [YRGES^{+}19]. sourced [MPHL22]. Space [BC13, CAGS17, ESB^{+}20, KL19, LHC^{+}24, CPP08, IMS^{+}08, Nas13, PJ13, VHKP11]. Space-Efficient [BC13, Nas13]. spaces [BE13]. Sparse [ASH20, ACGH24, BJWS18, BKS^{+}22, HEDH21, HBPH24, ICEDR24, JOB^{+}22, SMN22, SJL^{+}20, SJC^{+}21, WCW^{+}24, XLH^{+}24, YAG^{+}16, YXMC23, ZBC^{+}22, ZPH^{+}23, BR13]. Sparsification [XDW^{+}23]. Sparsity [RJK24, XSF^{+}23]. Sparsity-Aware [RJK24]. Spatial [CKP^{+}22]. Spatiotemporal [LAAMJ15, PL23]. SPCM [HASA16]. SPEC2006 [HPBS21]. SPEC2017 [HPBS21]. special [CDM13, SHC13, SD12]. Specialization [ANS^{+}22, BB21, FBC^{+}22, LLS23, YAG^{+}16]. Specialized [GáSÁ^{+}16, RAF22, GÁSÁ^{+}13]. species [NCC13]. Specific [CGP23, GMZ^{+}21, UDLD20, CZGC20, PRMH13]. Specification [ZPH^{+}23]. SpecTerminator [JHQ23]. Spectral [SBC05]. Spectrum [WZY^{+}24b]. Spectrum-Based [WZY^{+}24b]. Speculation [KS21, MGI15, GPL^{+}05, SHLM14]. Speculative [JHQ23, VS08, DC07, GPL^{+}05, LCH^{+}04, LHY^{+}06, LZ12, LHZ13, NTG13, VS11, XIC12, XC06, YRHBL13, ZSCM08]. speed [GB06, RPE12]. Speeding [GGS^{+}19]. SpGEMM [ZBC^{+}22]. Spiking [CRC^{+}21]. spill [XT09]. Spilling [CBD15]. Spin [ZMLL24, ZML^{+}24]. Spin-based [ZMLL24, ZML^{+}24]. Spintronic [CRC^{+}21, RKC^{+}20]. split [RFD13, TBS06]. splitting [WWY^{+}12]. SplitZNS [HFL^{+}23]. SPM [KE15]. SpMV [BJWS18, ZLYZ16]. SpMxV [KGK10]. sporadic [ZGC^{+}12]. spurious [BCVT13]. SPX64 [SDZ^{+}21]. Squash [SKS23]. SR [DCP^{+}12]. SR-IOV [DCP^{+}12]. SRAM [GBD^{+}15]. SSA [AvRF07, BZS13, CBD15, PBCB22]. SSA-based [AvRF07]. SSA-form [PBCB22]. SSD [CKL^{+}24, HWJ^{+}15, KHS^{+}14, SLH^{+}20, WZWW23, XDW^{+}23]. SSD-HDD-hybrid [CKL^{+}24]. SSD-SGD [XDW^{+}23]. SSDs [FYI^{+}24, HFL^{+}23, LHS^{+}24, SLS^{+}21]. Stabilization [SHD15]. stack [CH06, VS08, SCEG08]. Stacked [CWMC16, DVG^{+}23, LGP^{+}16, NRQ16a, NRQ16b]. Stacking [APBR16, ZSLX13]. Stacks [ZGX22]. stage [ZZL^{+}21]. Star [ZF24]. State [CSF^{+}20, DWF^{+}23, GPL^{+}05]. Static [AFD12, BHC^{+}16, PLG19, SHY14, TPB24, JSM^{+}04]. statically [NED^{+}13]. Stealing [CG15a, CMAP22, MN24, ZCQ^{+}19]. Stencil [CNS^{+}16b, DKB^{+}20, KFJ20, SHS^{+}20, XFS^{+}19, LFC13]. Stencil-Based [XFS^{+}19]. Step [Lou19]. Stopping [SKS23]. Storage [CKL^{+}24, DZL^{+}24, HLZ^{+}24, KBB^{+}14, LTX16, WZX^{+}24, XZW^{+}22, KKL^{+}24]. Store [KKAR16, XZW^{+}22, LHWB12, SL09].

Stores [PLG19]. Strategies [MRK^{+}22, WYCC11]. Strategy [KFJ20, YCCY11, ZHD^{+}04]. Stream [LT19, LK24, XCC^{+}13, YWXW12, MG13, YZL^{+}10]. Stream-Based [LK24]. Streaming [CNS^{+}16b, MKKE15, PC13, WO13]. Streaming-Based [CNS^{+}16b]. Streams [SJL^{+}20]. Strength [GAM12]. Strength-Based [GAM12]. Stride [WPJ19]. String [EZYA23, CW13, PLL10, TBS06]. String-Based [EZYA23]. string-matching [CW13, PLL10, TBS06]. Strings [SPM17]. Stripe [ZF24]. Stripe-schedule [ZF24]. Striped [HASA16]. Strong [OK21]. Structure [JHH^{+}23, WWY^{+}12]. Structured [BDB^{+}20]. Structures [IPSD21, FLG12]. STT [LZL^{+}13, PLK^{+}19, RTK15, WDX14]. STT-RAM [LZL^{+}13, WDX14]. STT-RAM-based [PLK^{+}19]. studies [LB10]. Study [CPS^{+}15, SKAEG16, SSRS15, MSF^{+}07]. Studying [CBD15]. Sub [ABP^{+}17]. Sub-Sequences [ABP^{+}17]. Sublining [HEDH21]. subranked [CCZ13]. Subsetting [AJK^{+}12]. Substrate [ZZH^{+}23]. subwords [SJV08]. Suite [CCM^{+}16, DDT^{+}17]. Sunway [AYL^{+}18, MLC^{+}23, ZFF^{+}18]. Supercomputer [AYL^{+}18, ZFF^{+}18]. Superpages [WLL^{+}19]. Superscalar [BEE15, MMS15, SRLPV04]. Superscalars [HYBR^{+}15]. Supervised [ASP17]. supplied [YZL^{+}10]. Supply [HAM17, HAM19, MST^{+}21]. Support [BKS^{+}22, KFEG18, LZZ^{+}22, ME15, MTPK24, SBC^{+}22, SKAEG16, ZZB^{+}19, CWC06, DMG13, EAH^{+}20, LMJ^{+}13b, SLA^{+}07, ZSCM08, ZZQ^{+}05]. Supporting [WLL^{+}19, SHC13]. Surrogate [BOEN23, GGS^{+}19]. Survey [ICEDR24]. SW [KMG14, TS15]. SW/HW [TS15]. SW26010 [JYM20]. Swap [WLX24]. switch [ASK13, BRSJG12, CPB^{+}07, GWM07, LS10]. switch-to-switch [BRSJG12]. SWITCHES [DT17]. Switching [MSFC21, DMG13]. SYCL [PSL^{+}23]. symbiosis [EE12]. Symbolic [LLS23, ZLJ18]. SYmmetric [PS12]. Symmetry [GSC17, ZDC^{+}16]. Symmetry-Agnostic [ZDC^{+}16]. Synchronization [DAKK19, GWZ22, MNSC16, SLJ^{+}18, CCPG13, ZSCM08]. Synchronization-Aware [SLJ^{+}18]. Synchronous [PBCB22]. SynchroTrace [SLJ^{+}18]. Synergistic [VGX16]. Synthesis [DJC16, GSC17, MMLS21, SJD22, SSW^{+}19, SSB^{+}20, UDLD20]. Synthesizer [DS16]. SYRANT [PS12]. System [Abd20, AJK^{+}12, CC18, CSW^{+}23, GBD21, HPBS21, HHC^{+}16, HHW^{+}22, LYK^{+}15, LCS^{+}19, LWC^{+}22, LZW23, LZS^{+}24, MMGS21, MGSH16, PDY^{+}23, PLT^{+}15, SBS16, SWF16, TKKM15, VOK^{+}22, VKM^{+}21, ZFT^{+}18, ZZL^{+}21, CDPD13, HCC^{+}14, KBR^{+}13, LWH11, SSPL^{+}13, TBC^{+}12, WSC^{+}13]. System- [PLT^{+}15]. System-Level [LCS^{+}19, LZW23, ZZL^{+}21]. Systematic [EMR14]. Systematically [JLJ^{+}18b]. Systems [AGG21, ANS^{+}22, APG^{+}23, CNS16a, CKPH19, CWK^{+}24, DNS^{+}24, FMY^{+}15, FPMR21, GTT^{+}16, HYYBM16, JED19, KE15, KTAE16, KAC^{+}18, KHN^{+}18, KMAK22, LK24, LMA^{+}16, LYH16, MMT^{+}12, MKKE15, MSFC21, NRQ16b, OKJ^{+}22, PLK^{+}19, PG17, PBY^{+}17, PGB16, SKP^{+}22, SPS17, TMP16, TPN^{+}20, TCS16, USCM16, WGO15, WLL^{+}19, XHJY16, ZDC^{+}16, ZSM^{+}16, CPP08, CWCS13, DXMJ11, GK13, GHS12, HS06, HWH^{+}11, KNBK12, KGK10, LMJ^{+}13b, LCL^{+}14, LHZ13, LFC13, LHWB12, MP13, NCQ14, YRHBL13, ZVYN05, ZPC06, ZCW10, ZDC^{+}12]. Systolic [LY24, SMN22, XMW^{+}21, YSD^{+}23]. tables [CPB14]. TACO [Aca16, Ano15, Ano13a, Ano13b, Bil19]. TACOMA [AVG12]. Tactics [CZGC20]. TAFFO [CCCA20]. TAGE [Mic18]. TAGE-like [Mic18]. Tagged [CTY^{+}23]. TaihuLight [AYL^{+}18, ZFF^{+}18]. tail [SLS^{+}21]. taken [PS12, PS12]. Taking [SWU^{+}15]. Tall [BHWN21]. Tall-and-Skinny [BHWN21]. taming [ZBH^{+}13]. target [LBJ05]. Targeting [KPRK20]. Targets [SAL19]. Task [APS22, CCM^{+}16, CMAP22, DHD^{+}14, GTT^{+}16, KKAR16, LLC22, MPPS18, RHC15, SN17, SDH^{+}15, ZCQ^{+}19, ZWY17, CG14, LMJ^{+}13b, VTN13, ZYCZ10, APS22]. Task-Parallel [APS22, DHD^{+}14, MPPS18, SN17]. Task-RM [APS22]. Task-stealing [ZCQ^{+}19]. Tasks [DT17, MKKE15, PVS^{+}17, PWPD19, ZGC^{+}12, PWPD19]. TEA [HLZ^{+}24]. Technique [BSG23, HNKK17, PGB16, XT09]. Techniques [ATGN^{+}13, DJC16, HAC13, VZS^{+}18, YMM^{+}15, MMdS06, MG12, RCG^{+}10a]. technologies [WLZ^{+}10]. Technology [WZX^{+}24, NED^{+}13, RWY13]. Temperature [CS21, MBLK24, SSS^{+}04, MSF^{+}07]. Temperature-aware [SSS^{+}04]. temperature-constrained [MSF^{+}07]. Template [UJW15]. Temporal [HLZ^{+}24, TKJ13]. Temporal-based [TKJ13]. Tenant [FCD^{+}17, KLA^{+}19]. Tensor [ACGH24, BKS^{+}22, GGK18, SMS23, ZPH^{+}23, WZWW23]. Tensors [LY24]. tenure [RBM10]. TEP [LP17]. Test [KPM21, SV05]. Test-pattern [KPM21]. Tetris [XT09]. Tetris-XL [XT09]. Thefts [GAH22]. Their [SKS23, ZG05]. Theory [YDL^{+}17, YDS^{+}19]. Thermal [LMMM08, MRK^{+}22, SKP^{+}22, CK11, WA08, ZYCZ10]. Think [TIP^{+}23]. Thread [CDPN16, DSR15, HAM^{+}20, LMZ18, LWL18, LYH16, MGI15, PGB12, RCG^{+}10a, SF18, TAB^{+}21, YBSY19, BTS10, CCPG13, DEE13, GPL^{+}05, LHZ13, MSF^{+}07]. Thread-Aware [LYH16]. Thread-Data [LWL18]. Thread-Level [LMZ18, MGI15, YBSY19, GPL^{+}05]. Thread-management [RCG^{+}10a]. Threaded [GVT^{+}17]. Threading [KS16, TCS20]. Threading-Based [KS16]. Threads [BAZ^{+}19, GB06, LZ12, ZSCM08]. Three [VFW16]. Threshold [HK14]. Throughput [CSW^{+}23, EMR14, KCA^{+}13, SAG22, BKA13, BTS10, OGK^{+}12, TBC^{+}12]. throughput-oriented [BTS10]. throughput/watt [TBC^{+}12]. Tiered [CWMC16]. Tightly [Abd20]. Tile [MBY13, MMGS21]. Tiled [KPP^{+}15, MFSK24, SYE19, ZCF18, CC13]. Tiled-MapReduce [CC13]. Tiles [ZC20]. Tiling [CC13, SHS^{+}20, XLH^{+}24, ZGP15, BCVT13]. Time [BC13, BNS^{+}21, CSF^{+}20, CCL^{+}24, CEP^{+}16, DPBI^{+}19, HZN^{+}22, KE15, KTAE16, LWW^{+}24, Nas13, PKPM19, SEF^{+}19, WZY^{+}24b, CCD12, GK13, KHL^{+}13, LTG12, LMCV13, RGG^{+}12, ZGC^{+}12]. Time- [BC13, Nas13]. Time-Aware [WZY^{+}24b]. time-critical [RGG^{+}12]. time-series [LTG12]. timekeeping [WM11]. Times [ACGH24]. timestamp [RLS13]. timestamp-based [RLS13]. Timing [AGG22, AMS23, BB21, Bis21, JFK20, LAS^{+}13, LFK19]. Tiny [LY24]. TL [ZGC^{+}12]. TL-plane-based [ZGC^{+}12]. TLB [CTY^{+}23, DHX^{+}22, JED19, LMJ13a, LBM13]. TLB-pilot [DHX^{+}22]. TLBs [LBM13]. TLC [PM17]. TLP [LMZ18, SNL^{+}04]. TNT [RKL23]. Token [RBM10]. token-counting [RBM10]. Tokens [ZFL18]. TokenSmart [SSS^{+}23]. Tolerance [AAI^{+}16, RCV^{+}05]. Tolerant [DSH^{+}18, HAM17, LCC11]. Tolerating [KWCL09, YLTL04]. Tomasulo [WLZ^{+}13].

Tomography [MMT^{+}12]. Tool [FPMR21, GDL16, JLL^{+}23, MPW^{+}17, PD17]. Toolchain [SKP^{+}22]. Tools [BKM^{+}17, UDLD20]. Topological [CVB15, KKM^{+}13]. Topologies [DJC16, YRGES^{+}19]. Topology [DHD^{+}14]. Topology-Aware [DHD^{+}14]. TornadoNoC [LNLK13]. Trace [HWM14, XDXL19, CWS06, HCC^{+}14, SWH09]. trace-based [HWM14]. Traces [HEMK17, PDCS24, SLJ^{+}18, TG07, ZG05]. Tracing [HWL^{+}19, HCC^{+}14]. Tracker [CSSU21]. Tracking [LLRC17, MMT^{+}12, KHL^{+}13, VTN13]. Tracks [SKS23]. Trade [ABK21, AVG12]. Trade-off [ABK21, AVG12]. Tradeoffs [GPL^{+}05, MAY23]. Trading [NMPS22]. Traffic [SLH^{+}20, FQRG13, LYYB07]. Training [WZWW23, XDW^{+}23]. Tranquilizer [PGB12]. Transaction [FYI^{+}24, LZS^{+}24, ZCCD16, SSU^{+}13]. Transactional [DHK18, DD16, GMGZP14, CDM^{+}22, IPSD21, NZ15, PD17, RLS15, VSDL16, ZZB^{+}19, ATGN^{+}13, RLS13, SSU^{+}13, TGAG^{+}12, WKCS12, YJTF13]. Transactions [DD16, LDC15, SSU^{+}13]. Transcendental [SSRS15]. Transfer [HHC^{+}16, TMSR23]. transfers [STLM12]. Transformation [BDB^{+}20, CLA^{+}19, JSL13]. Transformations [RAF22, TPB24, BCVN10, RCG^{+}10b, SLM12]. Transformer [DJZ^{+}23]. transition [CW13]. Transitioning [ZNX^{+}24, HWM14]. transitions [SW13]. Translation [CTY^{+}23, HWL^{+}19, JED19, KPM21, LHW^{+}19, TKKM15, ZCDD23, HWH^{+}11, LWH11, LMJ13a]. Translator [SHY14, HLC10]. Translators [DGGL16, XYY^{+}24, GHS12]. Transparency [GKCE17]. Transparent [MSFC21, RVKP19, ZHS^{+}19]. Transport [ÄJE^{+}16]. transpose [GS12]. transpose-free [GS12]. Traversal [LZZ^{+}22, RMA14, WDK^{+}24]. Traversals [SSS^{+}24]. Traversing [RKL23]. Tree [HFL^{+}23, LHS^{+}24, SSS^{+}24, WDK^{+}24, WX24, ZX19, CDPD13, DZL^{+}24, PRMH13, XZW^{+}22]. Trees [JGSM15, BRSJG12]. Triangle [CNAA^{+}22]. Triangular [BSL17]. Triggered [ÄJE^{+}16, JHHM21, YXS^{+}22]. Triple [HBPH24, LP17]. TRIPS [SNL^{+}04]. Trusted [PKL^{+}24]. TSV [NRQ16a]. Tumbler [PGB16]. Tunable [MGSH16, ZNX^{+}24]. Tuning [CG15a, JGSM15, JA14, LL22, MGI15, WG17, XFS^{+}19, APG^{+}23, WKCS12]. Turbo [KH18]. Turn [PL23, AGVO05]. Turn-based [PL23]. turn-off [AGVO05]. Two [CWMC16, JYE^{+}16, LIS20]. Two-Level [JYE^{+}16, LIS20]. Two-Tiered [CWMC16]. Tyche [XHL^{+}24]. type [BR13, JML^{+}20]. Types [PD17].

ULEEN [SAM^{+}23]. Ultra [ESB^{+}20, SAM^{+}23, ZML^{+}24]. Ultra-Efficient [ZML^{+}24]. Ultra-low-energy [SAM^{+}23]. Ultra-Low-Power [ESB^{+}20]. UMH [ZSM^{+}16]. Understanding [EPAG16, LS10, MMT^{+}12, RS21, VE13]. Unified [ACG24, LSH^{+}23, TG07, WWL^{+}21, ZSM^{+}16, YXK^{+}12, KRHK16]. Unified-memory-efficient [WWL^{+}21]. Uniform [HK14]. Unit [FLW^{+}24, WZW^{+}24]. Units [DZL^{+}24, GG18, GáSÁ^{+}16, SEF^{+}19, GÁSÁ^{+}13, HVJ06, YCCY11]. unloading [ZK05]. Unnecessary [TMSR23]. Unreliable [PVA^{+}17, SQZK20]. Unstructured [TPB24]. Unsynchronized [DSR15]. UPC [SKAEG16]. update [LZYZ09]. update-conscious [LZYZ09]. Upon [YXS^{+}22]. usage [VS11]. Use [SW17a]. Useful [SAT20]. User [KKAR16, PSL^{+}23, ZHS^{+}19]. User-Assisted [KKAR16]. User-driven [PSL^{+}23]. User-Transparent [ZHS^{+}19]. uses [GB06]. Using [AZG17, AMP^{+}16, ABP^{+}17, BSL17, BAZ^{+}19, CCL^{+}13, DKB^{+}20, DAKK19, ESR^{+}15, EZYA23, FDF^{+}14, FSP^{+}23, GáSÁ^{+}16, GR15, CDM^{+}22, JGSM15, KR19, Lou19, MLB^{+}23, RGK^{+}23, RLBBN15, SSH19, SYX^{+}15, SGM^{+}22, SPS17, SPS12, SHS^{+}20, SSH^{+}13, SSRS15, TPB24, UJW15, WO13, YGB21, ZLYW18, ASK13, BZS13, CAMJ15, CGP23, DDU12, DWDS13, DXMJ11, DJB13, EE11, HVJ06, JSH09, JSM^{+}04, KKM^{+}13, LZM14, MG13, RCV^{+}12, SHLM14, SWH09, SSR13, TTS19, YRGES^{+}19, YCCY11, YCA18, ZHD^{+}04, CST^{+}06]. Utility [LDL22, PDCS24, PB15]. Utility-Aware [PDCS24]. Utility-Driven [PB15]. Utility-of-Allocation [LDL22]. Utilization [CAGS17, FLW^{+}24, LWF^{+}16, SKKB18, TZK18, VZS^{+}18, YWXW12, ZCCD16, XCC^{+}13]. Utilizing [TBC^{+}12, KCP13]. UVMs [KRHK16].

V [JHQ23, MPU^{+}23, TGRK21]. v2 [DZSL20]. Value [CSSU21, EPS17, GAM12, KS21, OAM19, YPT^{+}16, CST^{+}06]. Value-next [OAM19]. Valued [RSU^{+}20]. variability [LYYB07]. Variable [MY16, NB13]. variation [CK11, PGB12, XL07]. variations [KWCL09]. Vector [ASH20, HBPH24, MPU^{+}23, SPH^{+}17, YAG^{+}16, ZLWS24]. Vector-Scalar [SPH^{+}17]. Vectorization [AMG16, RWY13, SPS12]. Vectorized [TPB24]. Vectorizing [SAT20]. Vectors [ICEDR24, SL09]. Verification [MMGS21]. Versatility [SJV08]. versioning [NTG13]. versus [SCEG08]. VGRIS [QYZ^{+}14]. via [ADGA20, CHD^{+}23, DSH^{+}18, FBC^{+}22, IMS^{+}08, JHH^{+}23, LFX09, LGH24, MNSC16, RCG^{+}10b, SYE19, SCFD22, WGL^{+}24, XGD^{+}23, XHJY17, XDXL19, ZYCZ10]. viable [PI12]. victim [VSP^{+}12]. Video [CAMJ15, HHW^{+}22]. Virtual [ACG24, BSSS14, HWJ^{+}15, KBB^{+}14, KRHK16, SCEG08, JA14, LHC^{+}24, VED07, WHV^{+}13, YZ08]. Virtualization [HHC^{+}16, SWF16, WWH^{+}16, DCP^{+}12]. Virtualized [QYZ^{+}14, WWWL13]. Virtualizing [MTPK24, WFKL10]. Virtually [RFD13]. Visible [KBB^{+}14]. Visual [ZHB18]. Visualizing [MMT^{+}12]. Vitruvius [MPU^{+}23]. VLIW [CPP08, GKP14, LKL^{+}13, LDG^{+}13, LZM14, PI12, TC07, XL07, XT09]. VLIW-based [CPP08]. VM [YKM17, ZLWS24]. Volatile [RTK15, WSJ^{+}21, YXS^{+}22, AEE^{+}19, WDXJ14]. Voltage [APBR16, RCG^{+}10b, XMM04]. Voltages [HK14]. vs [LMZ18, SV05]. VSim [RPE12]. Vulnerability [TS15, WAST16, LKL^{+}13]. Vulnerable [WWY^{+}24].

WA [LHS^{+}24]. WA-Zone [LHS^{+}24]. WADE [WSC^{+}13]. WaFFLe [CS21]. wakeup [YCCY11]. Walk [HLZ^{+}24]. Walker [TCR^{+}22]. warmup [HS05]. Warp [HLSK22, FSYA09]. Warp-level [HLSK22]. watt [TBC^{+}12]. Wave [CAY^{+}18]. Wavefront [GGYK19]. Way [LMZ18, ZVYN05]. way-halting [ZVYN05]. Ways [CS21, KBB^{+}14]. WC [ZWHM05]. WCET [DBH16, KTAE16, ZWHM05, ZWY17]. WCET-Aware [ZWY17]. WCETs [Lou19]. Wear [JDZ^{+}13, LHS^{+}24]. Wear-Aware [LHS^{+}24]. Wear-Leveling [JDZ^{+}13]. Weaving [PBCB22]. Web [PCM16]. Weight [GG18]. Weight-Sharing [GG18]. weighting [VS11]. WENO [CAY^{+}18]. while [GBD21]. Whole [ZG05]. Wide [MMS15, PI12]. wide-issue [PI12]. Width [SMKH15, RPS06]. width-partitioned [RPS06]. window [VS11]. Winograd [WLJ^{+}24, XLH^{+}24]. Winols [XLH^{+}24]. wins [ATGN^{+}13]. WIPE [WDS^{+}24]. wire [RKL23]. wires [IWP^{+}04]. wise [ZBC^{+}22]. Withdrawal [SAT20]. within [BCVN10]. Without [LHC^{+}17, RLS15, Bis21, KRHK16]. Work [CG15a, CMAP22, HLSW17, MN24]. Work-Queue [HLSW17]. Workload [KKL^{+}24, MBLK24, WLWB19, AVG12, CG14]. Workload-Aware [MBLK24, CG14]. Workloads [Akr21, DVG^{+}23, GVT^{+}17, LZW23, LYH16, SLJ^{+}18, SJC^{+}21, vdVSAAS20, DWDS13, JEBJ08, LTG12, LHC^{+}24, WA08]. Works [LKV12]. world [DJZ^{+}23]. worst [AFD12]. worst-case [AFD12]. Write [LLW^{+}22, LWF^{+}16, RJSA18, RLS15, WDS^{+}24, DZC^{+}13, JLCR13]. Write-Friendly [LLW^{+}22]. Write-Optimized [WDS^{+}24]. Writeback [WSC^{+}13, ZDC^{+}12]. Writeback-aware [WSC^{+}13, ZDC^{+}12]. WSNs [LZYZ09].

X [QSZ^{+}21, TCS20]. X10 [TN20]. x86 [CCD12, LHW^{+}19, ZGX22]. x86-64 [ZGX22]. XEngine [SMS23]. XL [XT09, YGB21]. xMeta [CKL^{+}24].

YaConv [KDMA23].

Zero [RSU^{+}20]. Zeroploit [RSU^{+}20]. ZNS [LHS^{+}24]. Zone [LHS^{+}24, LHS^{+}24]. Zoned [HFL^{+}23].

References

Akturk:2016:ABN

[AAI^{+}16] Ismail Akturk, Riad Akram, Mohammad Majharul Islam, Abdullah Muzahid, and Ulya R. Karpuzcu. Accuracy bugs: a new class of concurrency bugs to exploit algorithmic noise tolerance. ACM Transactions on Architecture and Code Optimization, 13(4): 48:1–48:??, December 2016.

CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Andreetta:2016:FPF

[ABB^{+}16] Christian Andreetta, Vivien Bégot, Jost Berthold, Martin Elsman, Fritz Henglein, Troels Henriksen, Maj-Britt Nordfang, and Cosmin E. Oancea. FinPar: a parallel financial benchmark. ACM Transactions on Architecture and Code Optimization, 13(2): 18:1–18:??, June 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Abdelrahman:2020:CSH

[Abd20] Tarek S. Abdelrahman. Cooperative software-hardware acceleration of K-means on a tightly coupled CPU–FPGA system. ACM Transactions on Architecture and Code Optimization, 17(3):20:1–20:24, August 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3406114.

Abera:2021:PET

[ABK21] Solomon Abera, M. Balakrishnan, and Anshul Kumar. Performance-energy trade-off in modern CMPs. ACM Transactions on Architecture and Code Optimization, 18(1):3:1– 3:26, January 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3427092.

Ashouri:2017:MMC

[ABP^{+}17] Amir H. Ashouri, Andrea Bignoli, Gianluca Palermo, Cristina Silvano, Sameer Kulkarni, and John Cavazos. MiCOMP: Mitigating the compiler phase-ordering problem using optimization subsequences and machine learning. ACM Transactions on Architecture and Code Optimization, 14(3):29:1–29:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Acacio:2016:LDR

[Aca16] Manuel Acacio. List of distinguished reviewers ACM TACO 2014. ACM Transactions on Architecture and Code Optimization, 13(3):31:1–31:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Andujar:2019:PPA

[ACA^{+}19] Franciso J. Andújar, Salvador Coll, Marina Alonso, Pedro López, and Juan-Miguel Martínez. POWAR: Poweraware routing in HPC networks with on/off links. ACM Transactions on Architecture and Code Optimization, 15(4): 61:1–61:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Allen:2024:FGQ

[ACG24] Tyler Allen, Bennett Cooper, and Rong Ge. Fine-grain quantitative analysis of demand paging in unified virtual memory. ACM Transactions on Architecture and Code Optimization, 21(1):14:1–14:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632953.

Ahmad:2024:EDL

[ACGH24] Khalid Ahmad, Cris Cecka, Michael Garland, and Mary Hall. Exploring data layout for sparse tensor times dense matrix on GPUs. ACM Transactions on Architecture and Code Optimization, 21(1): 20:1–20:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3633462.

Aleta:2004:RCC

[ACGK04] Alex Aletà, Josep M. Codina, Antonio González, and David Kaeli. Removing communications in clustered microarchitectures through instruction replication. ACM Transactions on Architecture and Code Optimization, 1(2): 127–151, June 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Adamek:2020:GFC

[ADGA20] Karel Adámek, Sofia Dimoudi, Mike Giles, and Wesley Armour. GPU fast convolution via the overlap-and-save method in shared memory.

ACM Transactions on Architecture and Code Optimization, 17(3):18:1–18:20, August 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3394116.

Alshboul:2019:ECR

[AEE^{+}19] Mohammad Alshboul, Hussein Elnawawy, Reem Elkhouly, Keiji Kimura, James Tuck, and Yan Solihin. Efficient checkpointing with recompute scheme for non-volatile main memory. ACM Transactions on Architecture and Code Optimization, 16(2):18:1–18:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Adileh:2016:MHP

[AEJE16] Almutaz Adileh, Stijn Eyerman, Aamer Jaleel, and Lieven Eeckhout. Maximizing heterogeneous processor performance under power constraints. ACM Transactions on Architecture and Code Optimization, 13(3): 29:1–29:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Andrade:2007:PAA

[AFD07] Diego Andrade, Basilio B.

Fraguela, and Ramón Doallo. Precise automatable analytical modeling of the cache behavior of codes with indirections. ACM Transactions on Architecture and Code Optimization, 4(3):16:1–16:??, September 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Andrade:2012:SAW

[AFD12] Diego Andrade, Basilio B. Fraguela, and Ramón Doallo. Static analysis of the worstcase memory performance for irregular codes with indirections. ACM Transactions on Architecture and Code Optimization, 9(3):20:1–20:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Alam:2021:LPL

[AGG21] Syed Asad Alam, James Garland, and David Gregg. Lowprecision logarithmic number systems: Beyond base-2. ACM Transactions on Architecture and Code Optimization, 18(4): 47:1–47:25, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3461699.

Athanasios:2022:LPN

[AGG22] Tziouvaras Athanasios, Dimitriou Georgios, and Stamoulis Georgios. Low-power neardata instruction execution leveraging opcode-based timing analysis. ACM Transactions on Architecture and Code Optimization, 19(2): 19:1–19:26, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3504005.

Albericio:2012:ALC

[AGI^{+}12] Jorge Albericio, Rubén Gran, Pablo Ibáñez, Víctor Viñals, and Jose María Llabería. ABS: a low-cost adaptive controller for prefetching in a banked shared last-level cache. ACM Transactions on Architecture and Code Optimization, 8(4): 19:1–19:??, January 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Abella:2005:ISP

[AGVO05] Jaume Abella, Antonio González, Xavier Vera, and Michael F. P. O’Boyle. IATAC: a smart predictor to turn-off L2 cache lines. ACM Transactions on Architecture and Code Optimization, 2(1):55–77, March 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Azriel:2019:MSP

[AHA^{+}19] Leonid Azriel, Lukas Humbel, Reto Achermann, Alex Richardson, Moritz Hoffmann, Avi Mendelson, Timothy Roscoe, Robert N. M. Watson, Paolo Faraboschi, and Dejan Milojicic. Memory-side protection with a capability enforcement co-processor. ACM Transactions on Architecture and Code Optimization, 16(1): 5:1–5:??, March 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Albericio:2013:ERL

[AIVL13] Jorge Albericio, Pablo Ibáñez, Víctor Viñals, and Jose María Llabería. Exploiting reuse locality on inclusive shared lastlevel caches. ACM Transactions on Architecture and Code Optimization, 9(4):38:1–38:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Aijo:2016:ILP

[ÄJE^{+}16] Tomi Äijö, Pekka Jääskeläinen, Tapio Elomaa, Heikki Kultala, and Jarmo Takala. Integer linear programming-based scheduling for transport triggered architectures. ACM Transactions on Architecture and Code Optimization, 12(4): 59:1–59:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ahn:2012:ISE

[AJK^{+}12] Jung Ho Ahn, Norman P. Jouppi, Christos Kozyrakis, Jacob Leverich, and Robert S. Schreiber. Improving system energy efficiency with memory rank subsetting. ACM Transactions on Architecture and Code Optimization, 9(1): 4:1–4:??, March 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Alves:2021:EAP

[AKBS21] Ricardo Alves, Stefanos Kaxiras, and David Black-Schaffer. Early address prediction: Efficient pipeline prefetch and reuse. ACM Transactions on Architecture and Code Optimization, 18(3):39:1–39:22, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458883.

Akram:2021:PEI

[Akr21] Shoaib Akram. Performance evaluation of Intel Optane memory for managed workloads. ACM Transactions on Architecture and Code Optimization, 18(3):29:1–29:26, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3451342.

Anderson:2016:AVI

[AMG16] Andrew Anderson, Avinash Malik, and David Gregg. Automatic vectorization of interleaved data revisited. ACM Transactions on Architecture and Code Optimization, 12(4): 50:1–50:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ashouri:2016:CCA

[AMP^{+}16] Amir Hossein Ashouri, Giovanni Mariani, Gianluca Palermo, Eunjung Park, John Cavazos, and Cristina Silvano. COBAYN: Compiler autotuning framework using Bayesian networks. ACM Transactions on Architecture and Code Optimization, 13(2):21:1–21:??, June 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Azhar:2023:ARR

[AMS23] Muhammad Waqar Azhar, Madhavan Manivannan, and Per Stenström. Approx-RM: Reducing energy on heterogeneous multicore processors under accuracy and timing constraints. ACM Transactions on Architecture and Code Optimization, 20(3):44:1–44:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3605214.

Anonymous:2013:LDR

[Ano13a]  Anonymous.  List of dis-

tinguished reviewers ACM TACO. ACM Transactions on Architecture and Code Optimization, 10(4):65:1–65:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Anonymous:2013:TR

[Ano13b] Anonymous. TACO reviewers 2012. ACM Transactions on Architecture and Code Optimization, 10(3):9:1–9:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Anonymous:2015:LDR

[Ano15] Anonymous. List of distinguished reviewers ACM TACO 2014. ACM Transactions on Architecture and Code Optimization, 11(4):68:1–68:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Alsop:2022:CFG

[ANS^{+}22] Johnathan Alsop, Weon Taek Na, Matthew D. Sinclair, Samuel Grayson, and Sarita Adve. A case for fine-grain coherence specialization in heterogeneous systems. ACM Transactions on Architecture and Code Optimization, 19(3): 41:1–41:26, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3530819.

Alias:2017:OAC

[AP17] Christophe Alias and Alexandru Plesco. Optimizing affine control with semantic factorizations. ACM Transactions on Architecture and Code Optimization, 14(4):52:1–52:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ardestani:2016:MMV

[APBR16] Ehsan K. Ardestani, Rafael Trapani Possignolo, Jose Luis Briz, and Jose Renau. Managing mismatches in voltage stacking with CoreUnfolding. ACM Transactions on Architecture and Code Optimization, 12(4):43:1–43:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Abad:2013:LLE

[APG13] Pablo Abad, Valentin Puente, and Jose-Angel Gregorio. LIGERO: a light but efficient router conceived for cachecoherent chip multiprocessors. ACM Transactions on Architecture and Code Optimization, 9(4):37:1–37:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Anastasiadis:2023:PPA

[APG^{+}23] Petros Anastasiadis, Nikela Papadopoulou, Georgios Goumas, Nectarios Koziris, Dennis Hoppe, and Li Zhong. PAR-ALiA: a performance aware runtime for auto-tuning linear algebra on heterogeneous systems. ACM Transactions on Architecture and Code Optimization, 20(4):52:1–52:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624569.

Azhar:2022:TRR

[APS22] M. Waqar Azhar, Miquel Pericàs, and Per Stenström. Task-RM: a resource manager for energy reduction in task-parallel applications under quality of service constraints. ACM Transactions on Architecture and Code Optimization, 19(1):11:1–11:26, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494537.

Akkary:2004:ARE

[ARS04] Haitham Akkary, Ravi Rajwar, and Srikanth T. Srinivasan. An analysis of a resource efficient checkpoint architecture. ACM Transactions on Architecture and Code Optimization, 1(4):418–444, December 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Antao:2013:CFA

[AS13] Samuel Antão and Leonel Sousa. The CRNS framework and its application to programmable and reconfigurable cryptography. ACM Transactions on Architecture and Code Optimization, 9(4):33:1–33:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ahmad:2020:DDM

[ASH20] Khalid Ahmad, Hari Sundar, and Mary Hall. Data-driven mixed precision sparse matrix vector multiplication for GPUs. ACM Transactions on Architecture and Code Optimization, 16(4):51:1–51:24, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3371275.

Ahn:2013:SHR

[ASK13] Jung Ho Ahn, Young Hoon Son, and John Kim. Scalable high-radix router microarchitecture using a network switch organization. ACM Transactions on Architecture and Code Optimization, 10(3): 17:1–17:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Anbar:2016:EHL

[ASK^{+}16] Ahmad Anbar, Olivier Serres, Engin Kayraklioglu, Abdel-Hameed A. Badawy, and Tarek El-Ghazawi. Exploiting hierarchical locality in deep parallel architectures. ACM Transactions on Architecture and Code Optimization, 13(2): 16:1–16:??, June 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Azhar:2017:SQS

[ASP17] M. Waqar Azhar, Per Stenström, and Vassilis Papaefstathiou. SLOOP: QoSsupervised loop execution to reduce energy on heterogeneous architectures. ACM Transactions on Architecture and Code Optimization, 14(4): 41:1–41:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Angerd:2017:FAC

[ASS17] Alexandra Angerd, Erik Sintorn, and Per Stenström. A framework for automated and controlled floating-point accuracy reduction in graphics applications on GPUs. ACM Transactions on Architecture and Code Optimization, 14(4): 46:1–46:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Akram:2016:BPG

[ASV^{+}16] Shoaib Akram, Jennifer B.

Sartor, Kenzo Van Craeynest, Wim Heirman, and Lieven Eeckhout. Boosting the priority of garbage: Scheduling collection on heterogeneous multicore processors. ACM Transactions on Architecture and Code Optimization, 13(1):4:1– 4:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Armejach:2013:TIP

[ATGN^{+}13] Adrià Armejach, Ruben Titos-Gil, Anurag Negi, Osman S. Unsal, and Adrián Cristal. Techniques to improve performance in requester-wins hardware transactional memory. ACM Transactions on Architecture and Code Optimization, 10(4):42:1–42:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Abbasi:2012:TSW

[AVG12] Zahra Abbasi, Georgios Varsamopoulos, and Sandeep K. S. Gupta. TACOMA: Server and workload management in Internet data centers considering cooling-computing power trade-off and energy proportionality. ACM Transactions on Architecture and Code Optimization, 9(2):11:1–11:??, June 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Amme:2007:SBM

[AvRF07] Wolfram Amme, Jeffery von Ronne, and Michael Franz. SSA-based mobile code: Implementation and empirical evaluation. ACM Transactions on Architecture and Code Optimization, 4(2):13:1–13:??, June 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ahn:2016:AEE

[AYC16] Junwhan Ahn, Sungjoo Yoo, and Kiyoung Choi. AIM: Energy-efficient aggregation inside the memory hierarchy. ACM Transactions on Architecture and Code Optimization, 13(4):34:1–34:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ao:2018:POH

[AYL^{+}18] Yulong Ao, Chao Yang, Fangfang Liu, Wanwang Yin, Lijuan Jiang, and Qiao Sun. Performance optimization of the HPCG benchmark on the Sunway TaihuLight Supercomputer. ACM Transactions on Architecture and Code Optimization, 15(1):11:1–11:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Arteaga:2017:GFG

[AZG17] Jaime Arteaga, Stéphane Zuckerman, and Guang R. Gao. Generating fine-grain multithreaded applications using a multigrain approach. ACM Transactions on Architecture and Code Optimization, 14(4):47:1–47:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Budhkar:2019:AMD

[BAZ^{+}19] Prerna Budhkar, Ildar Absalyamov, Vasileios Zois, Skyler Windh, Walid A. Najjar, and Vassilis J. Tsotras. Accelerating in-memory database selections using latency masking hardware threads. ACM Transactions on Architecture and Code Optimization, 16(2):13:1–13:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bai:2004:LPO

[BB04] Yu Bai and R. Iris Bahar.

A low-power in-order/out-of-order issue queue. ACM Transactions on Architecture and Code Optimization, 1(2):152– 179, June 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Benz:2021:SAP

[BB21] Joscha Benz and Oliver Bringmann. Scenario-aware program specialization for timing predictability. ACM Transactions on Architecture and Code Optimization, 18(4): 54:1–54:26, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3473333.

Belviranli:2013:DSS

[BBG13] Mehmet E. Belviranli, Laxmi N. Bhuyan, and Rajiv Gupta. A dynamic self-scheduling scheme for heterogeneous multiprocessor architectures. ACM Transactions on Architecture and Code Optimization, 9(4):57:1–57:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Becchi:2013:DTS

[BC13] Michela Becchi and Patrick Crowley. A-DFA: a time- and space-efficient DFA compression algorithm for fast regular expression evaluation. ACM Transactions on Architecture and Code Optimization, 10(1): 4:1–4:26, April 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Belleville:2019:ASP

[BCHC19] Nicolas Belleville, Damien Couroussé, Karine Heydemann, and Henri-Pierre Charles.

Automated software protection for the masses against side-channel attacks. ACM Transactions on Architecture and Code Optimization, 15(4): 47:1–47:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bhattacharjee:2011:PLC

[BCM11] Abhishek Bhattacharjee, Gilberto Contreras, and Margaret Martonosi. Parallelization libraries: Characterizing and reducing overheads. ACM Transactions on Architecture and Code Optimization, 8(1): 5:1–5:??, April 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Buyukkurt:2010:IHL

[BCVN10] Betul Buyukkurt, John Cortes, Jason Villarreal, and Walid A. Najjar. Impact of high-level transformations within the ROCCC framework. ACM Transactions on Architecture and Code Optimization, 7(4): 17:1–17:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Baghdadi:2013:ILT

[BCVT13] Riyadh Baghdadi, Albert Cohen, Sven Verdoolaege, and Konrad Trifunović. Improved loop tiling based on the removal of spurious false dependences. ACM Transactions on Architecture and Code Optimization, 9(4):52:1–52:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Briggs:2020:FRT

[BDB^{+}20] Ian Briggs, Arnab Das, Mark Baranowski, Vishal Sharma, Sriram Krishnamoorthy, Zvonimir Rakamarić, and Ganesh Gopalakrishnan. FailAmp: Relativization transformation for soft error detection in structured address generation. ACM Transactions on Architecture and Code Optimization, 16(4):50:1–50:21, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3369381.

Breughe:2013:SRB

[BE13] Maximilien B. Breughe and Lieven Eeckhout. Selecting representative benchmark inputs for exploring microprocessor design spaces. ACM Transactions on Architecture and Code Optimization, 10(4): 37:1–37:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Breugh:2015:MAM

[BEE15] Maximilien B. Breugh, Stijn Eyerman, and Lieven Eeckhout. Mechanistic analytical modeling of superscalar inorder processor performance. ACM Transactions on Architecture and Code Optimization, 11(4):50:1–50:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Balasubramanian:2015:EGL

[BGG^{+}15] Raghuraman Balasubramanian, Vinay Gangadhar, Ziliang Guo, Chen-Han Ho, Cherin Joseph, Jaikrishnan Menon, Mario Paulo Drumond, Robin Paul, Sharath Prasad, Pradip Valathol, and Karthikeyan Sankaralingam. Enabling GPGPU low-level hardware explorations with MIAOW: an open-source RTL implementation of a GPGPU. ACM Transactions on Architecture and Code Optimization, 12(2):21:1–21:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bao:2016:SDF

[BHC^{+}16] Wenlei Bao, Changwan Hong, Sudheer Chunduri, Sriram Krishnamoorthy, Louis-Noël Pouchet, Fabrice Rastello, and P. Sadayappan. Static and dynamic frequency scaling on multicore CPUs. ACM Transactions on Architecture and Code Optimization, 13(4): 51:1–51:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Borbon:2021:APB

[BHWN21] Jose M. Rodriguez Borbon, Junjie Huang, Bryan M. Wong, and Walid Najjar. Acceleration of parallel-blocked QR decomposition of tall-and-skinny matrices on FPGAs. ACM Transactions on Architecture and Code Optimization, 18(3):27:1–27:25, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447775.

Bilas:2019:LDR

[Bil19] Angelos Bilas. List of 2018 distinguished reviewers ACM TACO. ACM Transactions on Architecture and Code Optimization, 15(4):69:1–69:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Biswas:2021:CSI

[Bis21] Arnab Kumar Biswas. Cryptographic software IP protection without compromising performance or timing side-channel leakage. ACM Transactions on Architecture and Code Optimization, 18(2):20:1–20:20, March 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3443707.

Benatia:2018:BSM

[BJWS18] Akrem Benatia, Weixing Ji, Yizhuo Wang, and Feng Shi. BestSF: a sparse meta-format for optimizing SpMV on GPU. ACM Transactions on Architecture and Code Optimization, 15(3):29:1–29:??, October 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bakhoda:2013:DCN

[BKA13] Ali Bakhoda, John Kim, and Tor M. Aamodt. Designing on-chip networks for throughput accelerators. ACM Transactions on Architecture and Code Optimization, 10(3): 21:1–21:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Balasubramonian:2017:CNT

[BKM^{+}17] Rajeev Balasubramonian, Andrew B. Kahng, Naveen Muralimanohar, Ali Shafiee, and Vaishnav Srinivas. CACTI 7: New tools for interconnect exploration in innovative off-chip memories. ACM Transactions on Architecture and Code Optimization, 14(2): 14:1–14:??, July 2017. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bik:2022:CSS

[BKS^{+}22] Aart Bik, Penporn Koanantakool, Tatiana Shpeisman, Nicolas Vasilache, Bixia Zheng, and Fredrik Kjolstad. Compiler support for sparse tensor computations in MLIR. ACM Transactions on Architecture and Code Optimization, 19(4): 50:1–50:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3544559.

Braun:2021:SMP

[BNS^{+}21] Lorenz Braun, Sotirios Nikas, Chen Song, Vincent Heuveline, and Holger Fröning. A simple model for portable and fast prediction of execution time and power consumption of GPU kernels. ACM Transactions on Architecture and Code Optimization, 18(1):7:1–7:25, January 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3431731.

Benmeziane:2023:MOH

[BOEN23] Hadjer Benmeziane, Hamza Ouarnoughi, Kaoutar El Maghraoui, and Smail Niar. Multi-objective hardware-aware neural architecture search with Pareto rank-preserving surrogate models. ACM Transactions on Architecture and Code Optimization, 20(2): 29:1–29:??, June 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3579853.

BenAsher:2013:HTL

[BR13] Yosi Ben Asher and Nadav Rotem. Hybrid type legalization for a sparse SIMD instruction set. ACM Transactions on Architecture and Code Optimization, 10(3):11:1–11:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bahmann:2015:PRC

[BRJM15] Helge Bahmann, Nico Reissmann, Magnus Jahre, and Jan Christian Meyer. Perfect reconstructability of control flow from demand dependence graphs. ACM Transactions on Architecture and Code Optimization, 11(4):66:1–66:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bogdanski:2012:SFC

[BRSJG12] Bartosz Bogdanski, Sven-Arne Reinemo, Frank Olaf Sem-Jacobsen, and Ernst Gunnar Gran. sFtree: a fully connected and deadlock-free switch-to-switch routing algorithm for fat-trees. ACM Transactions on Architecture and Code Optimization, 8(4): 55:1–55:??, January 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Badri:2023:MPE

[BSG23] Satya Jaswanth Badri, Mukesh Saini, and Neeraj Goel. Mapi-Pro: an energy efficient memory mapping technique for intermittent computing. ACM Transactions on Architecture and Code Optimization, 20(4): 58:1–58:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3629524.

Baroudi:2017:OTB

[BSL17] Toufik Baroudi, Rachid Seghir, and Vincent Loechner. Optimization of triangular and banded matrix operations using 2 d-packed layouts. ACM Transactions on Architecture and Code Optimization, 14(4): 55:1–55:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bower:2007:ODH

[BSO07] Fred A. Bower, Daniel J. Sorin, and Sule Ozev. Online diagnosis of hard faults in microprocessors. ACM Transactions on Architecture and Code Optimization, 4(2):8:1–8:??, June 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bartolini:2014:AFG

[BSSS14] Davide B. Bartolini, Filippo Sironi, Donatella Sciuto, and Marco D. Santambrogio. Automated fine-grained CPU provisioning for virtual machines. ACM Transactions on Architecture and Code Optimization, 11(3):27:1–27:??, October 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bardizbanyan:2013:DPD

[BSWLE13] Alen Bardizbanyan, Magnus Själander, David Whalley, and Per Larsson-Edefors. Designing a practical data filter cache to improve both energy efficiency and performance. ACM Transactions on Architecture and Code Optimization, 10(4): 54:1–54:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Boyer:2010:FBP

[BTS10] Michael Boyer, David Tarjan, and Kevin Skadron. Federation: Boosting per-thread performance of throughputoriented manycore architectures. ACM Transactions on Architecture and Code Optimization, 7(4):19:1–19:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bayrak:2012:AII

[BVIB12] Ali Galip Bayrak, Nikola Velickovic, Paolo Ienne, and Wayne Burleson. An architectureindependent instruction shuffler to protect against sidechannel attacks. ACM Transactions on Architecture and Code Optimization, 8(4):20:1– 20:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bruintjes:2012:SLA

[BWG^{+}12] Tom M. Bruintjes, Karel H. G.

Walters, Sabih H. Gerez, Bert Molenkamp, and Gerard J. M. Smit. Sabrewing: a lightweight architecture for combined floatingpoint and integer arithmetic. ACM Transactions on Architecture and Code Optimization, 8(4):41:1–41:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bentley:2006:IAB

[BWLR06] Chris Bentley, Scott A. Watterson, David K. Lowenthal, and Barry Rountree. Implicit array bounds checking on 64bit architectures. ACM Transactions on Architecture and Code Optimization, 3(4):502– 527, December 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Barik:2013:DNS

[BZS13] Rajkishore Barik, Jisheng Zhao, and Vivek Sarkar. A decoupled non-SSA global register allocation using bipartite liveness graphs. ACM Transactions on Architecture and Code Optimization, 10(4): 63:1–63:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2011:HAM

[CA11] Xi E. Chen and Tor M. Aamodt. Hybrid analytical modeling of pending cache hits, data prefetching, and MSHRs. ACM Transactions on Architecture and Code Optimization, 8(3):10:1–10:??, October 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chaudhuri:2017:MSC

[CAGS17] Mainak Chaudhuri, Mukesh Agrawal, Jayesh Gaur, and Sreenivas Subramoney. Microsector cache: Improving space utilization in sectored DRAM caches. ACM Transactions on Architecture and Code Optimization, 14(1):7:1–7:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chi:2015:LPH

[CAMJ15] Chi Ching Chi, Mauricio Alvarez-Mesa, and Ben Juurlink. Low-power highefficiency video decoding using general-purpose processors. ACM Transactions on Architecture and Code Optimization, 11(4):56:1–56:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cai:2018:ESH

[CAY^{+}18] Ying Cai, Yulong Ao, Chao Yang, Wenjing Ma, and Haitao Zhao. Extreme-scale highorder WENO simulations of 3-D detonation wave with 10 million cores. ACM Transactions on Architecture and Code Optimization, 15(2):26:1–26:??, June 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Colombet:2015:SOS

[CBD15] Quentin Colombet, Florian Brandner, and Alain Darte. Studying optimal spilling in the light of SSA. ACM Transactions on Architecture and Code Optimization, 11(4):

47:1–47:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2013:TME

[CC13] Rong Chen and Haibo Chen. Tiled-MapReduce: Efficient and flexible MapReduce processing on multicore with tiling. ACM Transactions on Architecture and Code Optimization, 10(1):3:1–3:??, April 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2018:ESE

[CC18] Kuan-Chung Chen and Chung-Ho Chen. Enabling SIMT execution model on homogeneous multi-core system. ACM Transactions on Architecture and Code Optimization, 15(1): 6:1–6:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cherubin:2020:DPA

[CCCA20] Stefano Cherubin, Daniele Cattaneo, Michele Chiari, and Giovanni Agosta. Dynamic precision autotuning with TAFFO. ACM Transactions on Architecture and Code Optimization, 17(2):10:1–10:26, June 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3388785.

Cleemput:2012:CMT

[CCD12] Jeroen V. Cleemput, Bart Coppens, and Bjorn De Sutter. Compiler mitigations for time attacks on modern x86 processors. ACM Transactions on Architecture and Code Optimization, 8(4):23:1–23:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2013:DRU

[CCL^{+}13] Yunji Chen, Tianshi Chen, Ling Li, Ruiyang Wu, Daofu Liu, and Weiwu Hu. Deterministic replay using global clock. ACM Transactions on Architecture and Code Optimization, 10(1):1:1–1:??, April 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2024:DOO

[CCL^{+}24] Peng Chen, Hui Chen, Weichen Liu, Linbo Long, Wanli Chang, and Nan Guan. DAG-Order: an order-based dynamic DAG scheduling for real-time networks-on-chip. ACM Transactions on Architecture and Code Optimization, 21(1):2:1–2:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3631527.

Chasapis:2016:PEI

[CCM^{+}16] Dimitrios Chasapis, Marc Casas, Miquel Moretó, Raul Vidal, Eduard Ayguadé, Jesús Labarta, and Mateo Valero. PARSECSs: Evaluating the impact of task parallelism in the PARSEC benchmark suite. ACM Transactions on Architecture and Code Optimization, 12(4):41:1–41:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cleary:2013:FAT

[CCPG13] Jimmy Cleary, Owen Callanan, Mark Purcell, and David Gregg. Fast asymmetric thread synchronization. ACM Transactions on Architecture and Code Optimization, 9(4): 27:1–27:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2013:CME

[CCZ13] Long Chen, Yanan Cao, and Zhao Zhang. E^{3}CC: a memory error protection scheme with novel address mapping for subranked and low-power memories. ACM Transactions on Architecture and Code Optimization, 10(4):32:1–32:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Coppens:2013:FDB

[CDM13] Bart Coppens, Bjorn De Sutter, and Jonas Maebe. Feedback-driven binary code diversification to the special issue on high-performance embedded architectures and compilers. ACM Transactions on Architecture and Code Optimization, 9(4):24:1–24:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Honorio:2022:UBE

[CDM^{+}22] Bruno Chinelato Honorio, João P. L. De Carvalho, Catalina Munoz Morales, Alexandro Baldassin, and Guido Araujo. Using barrier elision to improve transactional code generation. ACM Transactions on Architecture and Code Optimization, 19(3): 46:1–46:23, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3533318.

Chrysos:2013:HCP

[CDPD13] Grigorios Chrysos, Panagiotis Dagritzikos, Ioannis Papaefstathiou, and Apostolos Dollas. HC-CART: a parallel system implementation of data mining classification and regression tree (CART) algorithm on a multi-FPGA system. ACM Transactions on Architecture and Code Optimization, 9(4): 47:1–47:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cruz:2016:HAT

[CDPN16] Eduardo H. M. Cruz, Matthias Diener, Laércio L. Pilla, and Philippe O. A. Navaux. Hardware-assisted thread and data mapping in hierarchical multicore architectures. ACM Transactions on Architecture and Code Optimization, 13(3): 28:1–28:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chrysanthou:2016:ORT

[CEP^{+}16] Kypros Chrysanthou, Panayiotis Englezakis, Andreas Prodromou, Andreas Panteli, Chrysostomos Nicopoulos, Yiannakis Sazeides, and Giorgos Dimitrakopoulos. An online and real-time fault detection and localization mechanism for network-on-chip architectures. ACM Transactions on Architecture and Code Optimization, 13(2): 22:1–22:??, June 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2012:DIO

[CFH^{+}12] Yang Chen, Shuangde Fang, Yuanjie Huang, Lieven Eeckhout, Grigori Fursin, Olivier Temam, and Chengyong Wu. Deconstructing iterative optimization. ACM Transactions on Architecture and Code Optimization, 9(3):21:1–21:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2014:AWA

[CG14] Quan Chen and Minyi Guo. Adaptive workload-aware task scheduling for single-ISA asymmetric multicore architectures. ACM Transactions on Architecture and Code Optimization, 11(1):8:1–8:25, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2015:LAW

[CG15a] Quan Chen and Minyi Guo.

Locality-aware work stealing based on online profiling and auto-tuning for multisocket multicore architectures. ACM Transactions on Architecture and Code Optimization, 12(2): 22:1–22:??, July 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cilardo:2015:IMM

[CG15b] Alessandro Cilardo and Luca Gallo. Improving multibank memory access parallelism with lattice-based partitioning. ACM Transactions on Architecture and Code Optimization, 11(4):45:1–45:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Choudhury:2023:FAI

[CGP23] Ziaul Choudhury, Anish Gulati, and Suresh Purini. Flow-Pix: Accelerating image processing pipelines on an FPGA overlay using a domain specific compiler. ACM Transactions on Architecture and Code Optimization, 20(4):60:1–60:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3629523.

Choi:2006:ORR

[CH06] Yoonseo Choi and Hwansoo Han. Optimal register reassignment for register stack overflow minimization. ACM Transactions on Architecture and Code Optimization, 3(1): 90–114, March 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2023:LFH

[CHD^{+}23] Zhangyu Chen, Yu Hua, Luochangqi Ding, Bo Ding, Pengfei Zuo, and Xue Liu. Lock-free high-performance hashing for persistent memory via PM-aware holistic optimization. ACM Transactions on Architecture and Code Optimization, 20(1):5:1–5:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3561651.

Carlson:2014:EHL

[CHE^{+}14] Trevor E. Carlson, Wim Heirman, Stijn Eyerman, Ibrahim Hur, and Lieven Eeckhout. An evaluation of high-level mechanistic core models. ACM Transactions on Architecture and Code Optimization, 11(3): 28:1–28:??, October 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Coelho:2013:ACI

[CI13] Fabien Coelho and François Irigoin. API compilation for image hardware accelerators. ACM Transactions on Architecture and Code Optimization, 9(4):49:1–49:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cher:2011:EEC

[CK11] Chen-Yong Cher and Eren Kursun. Exploring the effects of on-chip thermal variation on high-performance multicore architectures. ACM Transactions on Architecture and Code Optimization, 8(1):2:1– 2:??, April 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2024:XSH

[CKL^{+}24] Yan Chen, Qiwen Ke, Huiba Li, Yongwei Wu, and Yiming Zhang. xMeta: SSD-HDD-hybrid optimization for metadata maintenance of cloudscale object storage. ACM Transactions on Architecture and Code Optimization, 21(2): 40:1–40:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3652606.

Chatarasi:2022:MDC

[CKP^{+}22] Prasanth Chatarasi, Hyoukjun Kwon, Angshuman Parashar, Michael Pellauer, Tushar Krishna, and Vivek Sarkar. Marvel: a data-centric approach for mapping deep learning operators on spatial accelerators. ACM Transactions on Architecture and Code Optimization, 19(1):6:1–6:26, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485137.

Cha:2019:MDC

[CKPH19] Sanghoon Cha, Bokyeong Kim, Chang Hyun Park, and Jaehyuk Huh. Morphable DRAM cache design for hybrid memory systems. ACM Transactions on Architecture and Code Optimization, 16(3): 31:1–31:??, July 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chikin:2019:MAA

[CLA^{+}19] Artem Chikin, Taylor Lloyd, José Nelson Amaral, Ettore Tiotto, and Muhammad Usman. Memory-access-aware safety and profitability analysis for transformation of accelerator-bound OpenMP loops. ACM Transactions on Architecture and Code Optimization, 16(3):30:1–30:??, July 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2022:EEE

[CMAP22] Jing Chen, Madhavan Manivannan, Mustafa Abduljabbar, and Miquel Pericàs. ERASE: Energy efficient task mapping and resource management for work stealing runtimes. ACM Transactions on Architecture and Code Optimization, 19(2): 27:1–27:29, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3510422.

Corbalan-Navarro:2022:TDO

[CNAA^{+}22] David Corbalán-Navarro, Juan L.

Aragón, Martí Anglada, Joan-Manuel Parcerisa, and Antonio González. Triangle dropping: an occluded-geometry predictor for energy-efficient mobile GPUs. ACM Transactions on Architecture and Code Optimization, 19(3):39:1–39:20, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3527861.

C:2016:FGM

[CNS16a] Unnikrishnan C, Rupesh Nasre, and Y. N. Srikant. Falcon: a graph manipulation language for heterogeneous systems. ACM Transactions on Architecture and Code Optimization, 12(4):54:1–54:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cattaneo:2016:HAI

[CNS^{+}16b] Riccardo Cattaneo, Giuseppe Natale, Carlo Sicignano, Donatella Sciuto, and Marco Domenico Santambrogio. On how to accelerate iterative stencil loops: a scalable streaming-based approach. ACM Transactions on Architecture and Code Optimization, 12(4):53:1–53:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Constantinides:2007:ARC

[CPB^{+}07] Kypros Constantinides, Stephen Plaza, Jason Blome, Valeria Bertacco, Scott Mahlke, Todd Austin, Bin Zhang, and Michael Orshansky. Architecting a reliable CMP switch architecture. ACM Transactions on Architecture and Code Optimization, 4(1):2:1– 2:37, March 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Carle:2014:PAM

[CPB14] Thomas Carle and Dumitru Potop-Butucaru. Predicateaware, makespan-preserving software pipelining of scheduling tables. ACM Transactions on Architecture and Code Optimization, 11(1):12:1–12:26, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/2579676.

Charitopoulos:2021:MDC

[CPG21] George Charitopoulos, Dionisios N. Pnevmatikatos, and Georgi Gaydadjiev. MC-DeF: Creating customized CGRAs for dataflow applications. ACM Transactions on Architecture and Code Optimization, 18(3):26:1–26:25, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447970.

Catania:2008:RCM

[CPP08] Vincenzo Catania, Maurizio Palesi, and Davide Patti. Reducing complexity of multiobjective design space exploration in VLIW-based embedded systems. ACM Transactions on Architecture and Code Optimization, 5(2):11:1–11:??, August 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cheng:2015:ECS

[CPS^{+}15] Hsiang-Yun Cheng, Matt Poremba, Narges Shahidi, Ivan Stalev, Mary Jane Irwin, Mahmut Kandemir, Jack Sampson, and Yuan Xie. EECache: a comprehensive study on the architectural design for energy-efficient lastlevel caches in chip multiprocessors. ACM Transactions on Architecture and Code Optimization, 12(2):17:1–17:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Canesche:2024:DSA

[CRBP24] Michael Canesche, Vanderson Rosário, Edson Borin, and Fernando Quintão Pereira. The droplet search algorithm for kernel scheduling. ACM Transactions on Architecture and Code Optimization, 21(2): 35:1–35:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3650109.

Cilasun:2021:SNN

[CRC^{+}21] Hüsrev Cilasun, Salonik Resch, Zamshed I. Chowdhury, Erin Olson, Masoud Zabihi, Zhengyang Zhao, Thomas Peterson, Keshab K. Parhi, Jian-Ping Wang, Sachin S. Sapatnekar, and Ulya R. Karpuzcu. Spiking neural networks in spintronic computational RAM. ACM Transactions on Architecture and Code Optimization, 18(4):59:1–59:21, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3475963.

Chhabra:2009:MSP

[CRSP09] Siddhartha Chhabra, Brian Rogers, Yan Solihin, and Milos Prvulovic. Making secure processors OS- and performancefriendly. ACM Transactions on Architecture and Code Optimization, 5(4):16:1–16:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2010:HSF

[CS10] Zhong-Ho Chen and Alvin W. Y. Su. A hardware/ software framework for instruction and data scratchpad memory allocation. ACM Transactions on Architecture and Code Optimization, 7(1): 2:1–2:??, April 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2013:PGF

[CS13] Doris Chen and Deshanand Singh. Profile-guided floatingto fixed-point conversion for hybrid FPGA-processor applications. ACM Transactions on Architecture and Code Optimization, 9(4):43:1–43:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chakraborty:2021:WGC

[CS21] Shounak Chakraborty and Magnus Själander. WaF-FLe: Gated cache-ways with per-core fine-grained DVFS for reduced on-chip temperature and leakage consumption. ACM Transactions on Architecture and Code Optimization, 18(4):55:1–55:25, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3471908.

Cerina:2020:EDO

[CSF^{+}20] L. Cerina, M. D. Santambrogio, G. Franco, C. Gallicchio, and A. Micheli. EchoBay: Design and optimization of echo state networks under memory and time constraints. ACM Transactions on Architecture and Code Optimization, 17(3): 22:1–22:24, August 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3404993.

Crago:2019:EMA

[CSK19] Neal C. Crago, Mark Stephenson, and Stephen W. Keckler. Exposing memory access patterns to improve instruction and memory efficiency in GPUs. ACM Transactions on Architecture and Code Optimization, 15(4):45:1–45:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Choudhury:2022:FOC

[CSRP22] Ziaul Choudhury, Shashwat Shrivastava, Lavanya Ramapantulu, and Suresh Purini. An FPGA overlay for CNN inference with fine-grained flexible parallelism. ACM Transactions on Architecture and Code Optimization, 19(3): 34:1–34:26, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3519598.

Cavus:2021:FKV

[CSSU21] Mustafa Cavus, Mohammed Shatnawi, Resit Sendag, and Augustus K. Uht. Fast key-value lookups with node tracker. ACM Transactions on Architecture and Code Optimization, 18(3):34:1–34:26, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3452099.

Ceze:2006:CUC

[CST^{+}06] Luis Ceze, Karin Strauss, James Tuck, Josep Torrellas, and Jose Renau. CAVA: Using checkpoint-assisted value prediction to hide L2 misses. ACM Transactions on Architecture and Code Optimization, 3(2):182–208, June 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cristal:2004:TKI

[CSVM04] Adrián Cristal, Oliverio J.

Santana, Mateo Valero, and José F. Martínez. Toward kiloinstruction processors. ACM Transactions on Architecture and Code Optimization, 1(4): 389–417, December 2004. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2023:JOJ

[CSW^{+}23] Ruobing Chen, Haosen Shi, Jinping Wu, Yusen Li, Xiaoguang Liu, and Gang Wang. Jointly optimizing job assignment and resource partitioning for improving system throughput in cloud datacenters. ACM Transactions on Architecture and Code Optimization, 20(3): 34:1–34:??, September 2023.

CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3593055.

Cavus:2020:IPI

[CSY20] Mustafa Cavus, Resit Sendag, and Joshua J. Yi. Informed prefetching for indirect memory accesses. ACM Transactions on Architecture and Code Optimization, 17(1):4:1–4:29, March 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3374216.

Calder:2004:I

[CT04] Brad Calder and Dean Tullsen. Introduction. ACM Transactions on Architecture and Code Optimization, 1(1):1–2, March 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Calder:2005:I

[CT05] Brad Calder and Dean Tullsen. Introduction. ACM Transactions on Architecture and Code Optimization, 2(1):1–2, March 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Calder:2006:I

[CT06] Brad Calder and Dean Tullsen. Introduction. ACM Transactions on Architecture and Code Optimization, 3(1):1–2, March 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Calder:2007:I

[CT07] Brad Calder and Dean Tullsen.

Introduction. ACM Transactions on Architecture and Code Optimization, 4(1):1:1, March 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Calder:2008:E

[CT08] Brad Calder and Dean Tullsen.

Editorial. ACM Transactions on Architecture and Code Optimization, 5(1):1:1–1:??, May 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2023:FFA

[CTY^{+}23] Dongwei Chen, Dong Tong, Chun Yang, Jiangfang Yi, and Xu Cheng. FlexPointer: Fast address translation based on range TLB and tagged pointers. ACM Transactions on Architecture and Code Optimization, 20(2):30:1–30:??, June 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579854.

Camarero:2015:TCH

[CVB15] Cristóbal Camarero, Enrique Vallejo, and Ramón Beivide. Topological characterization of Hamming and dragonfly networks and its implications on routing. ACM Transactions on Architecture and Code Optimization, 11(4):39:1–39:??, January 2015. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2013:EMT

[CW13] Chien-Chi Chen and Sheng-De Wang. An efficient multicharacter transition stringmatching engine based on the Aho–Corasick algorithm. ACM Transactions on Architecture and Code Optimization, 10(4):25:1–25:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Crandall:2006:MAS

[CWC06] Jedidiah R. Crandall, S. Felix Wu, and Frederic T. Chong. Minos: Architectural support for protecting control data. ACM Transactions on Architecture and Code Optimization, 3(4):359–389, December 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cui:2013:LCA

[CWCS13] Yan Cui, Yingxin Wang, Yu Chen, and Yuanchun Shi. Lock-contention-aware scheduler: a scalable and energyefficient method for addressing scalability collapse on multicore systems. ACM Transactions on Architecture and Code Optimization, 9(4):44:1–44:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chacon:2024:CAC

[CWK^{+}24] Gino A. Chacon, Charles Williams, Johann Knechtel, Ozgur Sinanoglu, Paul V. Gratz, and Vassos Soteriou. Coherence attacks and countermeasures in interposerbased chiplet systems. ACM Transactions on Architecture and Code Optimization, 21(2): 23:1–23:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3633461.

Chen:2016:RER

[CWMC16] Hsing-Min Chen, Carole-Jean Wu, Trevor Mudge, and Chaitali Chakrabarti. RATT-ECC: Rate adaptive two-tiered error correction codes for reliable 3D die-stacked memory. ACM Transactions on Architecture and Code Optimization, 13(3):24:1–24:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Co:2006:ETC

[CWS06] Michele Co, Dee A. B. Weikle, and Kevin Skadron. Evaluating trace cache energy efficiency. ACM Transactions on Architecture and Code Optimization, 3(4):450–476, December 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2016:IDO

[CWW^{+}16] Wenjie Chen, Zhibin Wang, Qin Wu, Jiuzhen Liang, and Zhilei Chai. Implementing dense optical flow computation on a heterogeneous FPGA SoC in C. ACM Transactions on Architecture and Code Optimization, 13(3):25:1–25:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cui:2012:EPO

[CXW^{+}12] Huimin Cui, Jingling Xue, Lei Wang, Yang Yang, Xiaobing Feng, and Dongrui Fan. Extendable pattern-oriented optimization directives. ACM Transactions on Architecture and Code Optimization, 9(3): 14:1–14:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cui:2013:LOC

[CYXF13] Huimin Cui, Qing Yi, Jingling Xue, and Xiaobing Feng. Layout-oblivious compiler optimization for matrix computations. ACM Transactions on Architecture and Code Optimization, 9(4):35:1–35:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2007:CRL

[CZ07] Yu Chen and Fuxin Zhang. Code reordering on limited branch offset. ACM Transactions on Architecture and Code Optimization, 4(2):10:1–10:??, June 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chelini:2020:DLT

[CZGC20] Lorenzo Chelini, Oleksandr Zinenko, Tobias Grosser, and Henk Corporaal. Declarative loop tactics for domainspecific optimization. ACM Transactions on Architecture and Code Optimization, 16(4): 55:1–55:25, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3372266.

Das:2016:RDB

[DAD16] Subhasis Das, Tor M. Aamodt, and William J. Dally. Reuse distance-based probabilistic cache replacement. ACM Transactions on Architecture and Code Optimization, 12(4): 33:1–33:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Dogan:2019:ASU

[DAKK19] Halit Dogan, Masab Ahmad, Brian Kahne, and Omer Khan. Accelerating synchronization using moving compute to data model at 1,000-core multicore scale. ACM Transactions on Architecture and Code Optimization, 16(1):4:1–4:??, March 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

DeOliveiraCastro:2015:CLB

[DAP^{+}15] Pablo De Oliveira Castro, Chadi Akel, Eric Petit, Mihail Popov, and William Jalby. CERE: LLVM-based Codelet Extractor and REplayer for piecewise benchmarking and optimization. ACM Transactions on Architecture and Code Optimization, 12(1):6:1– 6:??, April 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Damschen:2016:EWP

[DBH16] Marvin Damschen, Lars Bauer, and Jörg Henkel. Extending the WCET problem to optimize for runtimereconfigurable processors. ACM Transactions on Architecture and Code Optimization, 13(4): 45:1–45:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Dou:2007:CCM

[DC07] Jialin Dou and Marcelo Cintra. A compiler cost model for speculative parallelization. ACM Transactions on Architecture and Code Optimization, 4(2):12:1–12:??, June 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ding:2022:CCA

[DCL^{+}22] Chen Ding, Dong Chen, Fangzhou Liu, Benjamin Reber, and Wesley Smith. CARL: Compiler assigned reference leasing. ACM Transactions on Architecture and Code Optimization, 19(1):15:1–15:28, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498730.

Dong:2012:RAE

[DCP^{+}12] Yaozu Dong, Yu Chen, Zhenhao Pan, Jinquan Dai, and Yunhong Jiang. ReNIC: Architectural extension to SR-IOV I/O virtualization for efficient replication. ACM Transactions on Architecture and Code Optimization, 8(4):40:1–40:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Do:2016:PEH

[DD16] Sang Wook Stephen Do and Michel Dubois. Power efficient hardware transactional memory: Dynamic issue of transactions. ACM Transactions on Architecture and Code Optimization, 13(1):9:1–9:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

DeSensi:2017:BPP

[DDT^{+}17] Daniele De Sensi, Tiziano De Matteis, Massimo Torquati, Gabriele Mencagli, and Marco Danelutto. Bringing parallel patterns out of the corner: The P^{3}ARSEC benchmark suite. ACM Transactions on Architecture and Code Optimization, 14(4):33:1–33:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Das:2012:ELC

[DDU12] Dibyendu Das, B. Dupont De Dinechin, and Ramakrishna Upadrasta. Efficient liveness computation using merge sets and DJ-graphs. ACM Transactions on Architecture and Code Optimization, 8(4):27:1–27:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

DuBois:2013:PTC

[DEE13] Kristof Du Bois, Stijn Eyerman, and Lieven Eeckhout. Per-thread cycle accounting in multicore processors. ACM Transactions on Architecture and Code Optimization, 9(4): 29:1–29:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Dantras:2016:OIB

[DGGL16] Amanieu D’antras, Cosmin Gorgovan, Jim Garside, and Mikel Luján. Optimizing indirect branches in dynamic binary translators. ACM Transactions on Architecture and Code Optimization, 13(1):7:1– 7:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Carlo:2014:FAA

[DGI^{+}14] Stefano Di Carlo, Salvatore Galfano, Marco Indaco, Paolo Prinetto, Davide Bertozzi, Piero Olivo, and Cristian Zambelli. FLARES: an aging aware algorithm to autonomously adapt the error correction capability in NAND flash memories. ACM Transactions on Architecture and Code Optimization, 11(3):26:1–26:??, October 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Demir:2016:EPP

[DH16] Yigit Demir and Nikos Hardavellas. Energy-proportional photonic interconnects. ACM Transactions on Architecture and Code Optimization, 13(4): 54:1–54:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Diouf:2013:DLM

[DHC^{+}13] Boubacar Diouf, Can Hantas, Albert Cohen, Özcan Özturk, and Jens Palsberg. A decoupled local memory allocator. ACM Transactions on Architecture and Code Optimization, 9(4):34:1–34:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Drebes:2014:TAD

[DHD^{+}14] Andi Drebes, Karine Heydemann, Nathalie Drach, Antoniu Pop, and Albert Cohen. Topology-aware and dependence-aware scheduling and memory allocation for task-parallel languages. ACM Transactions on Architecture and Code Optimization, 11(3): 30:1–30:??, October 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Dice:2018:IPH

[DHK18] Dave Dice, Maurice Herlihy, and Alex Kogan. Improving parallelism in hardware transactional memory. ACM Transactions on Architecture and Code Optimization, 15(1):9:1– 9:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Di:2022:TPM

[DHX^{+}22] Bang Di, Daokun Hu, Zhen Xie, Jianhua Sun, Hao Chen, Jinkui Ren, and Dong Li. TLB-pilot: Mitigating TLB contention attack on GPUs with microarchitecture-aware scheduling. ACM Transactions on Architecture and Code Optimization, 19(1):9:1–9:23, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491218.

Dubach:2013:DMA

[DJB13] Christophe Dubach, Timothy M. Jones, and Edwin V. Bonilla. Dynamic microarchitectural adaptation using machine learning. ACM Transactions on Architecture and Code Optimization, 10(4): 31:1–31:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Dsouza:2016:IMS

[DJC16] Sandeep D’souza, Soumya J., and Santanu Chattopadhyay. Integrated mapping and synthesis techniques for networkon-chip topologies with express channels. ACM Transactions on Architecture and Code Optimization, 12(4): 40:1–40:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Domnitser:2012:NMC

[DJL^{+}12] Leonid Domnitser, Aamer Jaleel, Jason Loew, Nael Abu-Ghazaleh, and Dmitry Ponomarev. Non-monopolizable caches: Low-complexity mitigation of cache side channel attacks. ACM Transactions on Architecture and Code Optimization, 8(4):35:1–35:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Dong:2013:CAC

[DJX13] Xiangyu Dong, Norman P.

Jouppi, and Yuan Xie. A circuit-architecture co-optimization framework for exploring nonvolatile memory hierarchies. ACM Transactions on Architecture and Code Optimization, 10(4):23:1–23:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Du:2023:ICM

[DJZ^{+}23] Jiangsu Du, Jiazhi Jiang, Jiang Zheng, Hongbin Zhang, Dan Huang, and Yutong Lu. Improving computation and memory efficiency for realworld Transformer inference on GPUs. ACM Transactions on Architecture and Code Optimization, 20(4):46:1–46:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617689.

Das:2020:FER

[DKB^{+}20] Arnab Das, Sriram Krishnamoorthy, Ian Briggs, Ganesh Gopalakrishnan, and Ramakrishna Tipireddy. FPDetect: Efficient reasoning about stencil programs using selective direct evaluation. ACM Transactions on Architecture and Code Optimization, 17(3): 19:1–19:27, August 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3402451.

Carvalho:2021:KRN

[DKK^{+}21] João P. L. De Carvalho, Braedy Kuzma, Ivan Korostelev, José Nelson Amaral, Christopher Barton, José Moreira, and Guido Araujo. KernelFaRer: Replacing nativecode idioms with high-performance library calls. ACM Transactions on Architecture and Code Optimization, 18(3): 38:1–38:22, June 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3459010.

Dang:2022:LAP

[DLS22] Dharanidhar Dang, Bill Lin, and Debashis Sahoo. Lite-CON: an all-photonic neuromorphic accelerator for energy-efficient deep learning. ACM Transactions on Architecture and Code Optimization, 19(3):43:1–43:22, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531226.

Dolan:2013:CSL

[DMG13] Stephen Dolan, Servesh Muralidharan, and David Gregg. Compiler support for lightweight context switching. ACM Transactions on Architecture and Code Optimization, 9(4): 36:1–36:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Dardaillon:2016:NCF

[DMR^{+}16] Mickaël Dardaillon, Kevin Marquet, Tanguy Risset, Jérôme Martin, and Henri-Pierre Charles. A new compilation flow for software-defined radio applications on heterogeneous MPSoCs. ACM Transactions on Architecture and Code Optimization, 13(2): 19:1–19:??, June 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Deng:2024:FPE

[DNS^{+}24] Bobin Deng, Bhargava Nadendla, Kun Suo, Yixin Xie, and Dan Chia-Tien Lo. Fixed-point encoding and architecture exploration for residue number systems. ACM Transactions on Architecture and Code Optimization, 21(3):53:1–53:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3664923.

Dublish:2016:CCG

[DNT16] Saumay Dublish, Vijay Nagarajan, and Nigel Topham. Cooperative caching for GPUs. ACM Transactions on Architecture and Code Optimization, 13(4):39:1–39:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Didier:2019:CCP

[DPBI^{+}19] Keryan Didier, Dumitru Potop-Butucaru, Guillaume Iooss, Albert Cohen, Jean Souyris, Philippe Baufreton, and Amaury Graillat. Correct-by-construction parallelization of hard realtime avionics applications on off-the-shelf predictable hardware. ACM Transactions on Architecture and Code Optimization, 16(3):24:1–24:??, July 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Du:2024:FDR

[DQCK24] Haitao Du, Yuhan Qin, Song Chen, and Yi Kang. FASA-DRAM: Reducing DRAM latency with destructive activation and delayed restoration. ACM Transactions on Architecture and Code Optimization, 21(2):34:1–34:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3649455.

Davari:2015:EGA

[DRHK15] Mahdad Davari, Alberto Ros, Erik Hagersten, and Stefanos Kaxiras. The effects of granularity and adaptivity on private/shared classification for coherence. ACM Transactions on Architecture and Code Optimization, 12(3):26:1–26:??, October 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Demme:2012:AGC

[DS12] John Demme and Simha Sethumadhavan. Approximate graph clustering for program characterization. ACM Transactions on Architecture and Code Optimization, 8(4):21:1– 21:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Deniz:2016:MGM

[DS16] Etem Deniz and Alper Sen.

MINIME-GPU: Multicore benchmark synthesizer for GPUs.

ACM Transactions on Architecture and Code Optimization, 12(4):34:1–34:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Deng:2018:EML

[DSH^{+}18] Bobin Deng, Sriseshan Srikanth, Eric R. Hein, Thomas M. Conte, Erik Debenedictis, Jeanine Cook, and Michael P. Frank. Extending Moore’s Law via computationally errortolerant computing. ACM Transactions on Architecture and Code Optimization, 15(1): 8:1–8:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Debrunner:2019:AAK

[DSK19] Thomas Debrunner, Sajad Saeedi, and Paul H. J. Kelly. AUKE: Automatic kernel code generation for an analogue SIMD focal-plane sensor-processor array. ACM Transactions on Architecture and Code Optimization, 15(4): 59:1–59:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Das:2015:SBP

[DSR15] Madan Das, Gabriel Southern, and Jose Renau. Sectionbased program analysis to reduce overhead of detecting unsynchronized thread communication. ACM Transactions on Architecture and Code Optimization, 12(2):23:1–23:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Diavastos:2017:SLR

[DT17] Andreas Diavastos and Pedro Trancoso. SWITCHES: a lightweight runtime for dataflow execution of tasks on many-cores. ACM Transactions on Architecture and Code Optimization, 14(3): 31:1–31:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

DeSensi:2016:RAP

[DTD16] Daniele De Sensi, Massimo Torquati, and Marco Danelutto. A reconfiguration algorithm for power-aware parallel applications. ACM Transactions on Architecture and Code Optimization, 13(4): 43:1–43:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Domke:2023:LPQ

[DVG^{+}23] Jens Domke, Emil Vatai, Balazs Gerofi, Yuetsu Kodama, Mohamed Wahib, Artur Podobas, Sparsh Mittal, Miquel Pericàs, Lingqi Zhang, Peng Chen, Aleksandr Drozd, and Satoshi Matsuoka. At the locus of performance: Quantifying the effects of copious 3Dstacked cache on HPC workloads. ACM Transactions on Architecture and Code Optimization, 20(4):57:1–57:??, December 2023. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3629520.

Dey:2013:RMD

[DWDS13] Tanima Dey, Wei Wang, Jack W. Davidson, and Mary Lou Soffa. ReSense: Mapping dynamic workloads of colocated multithreaded applications using resource sensitivity. ACM Transactions on Architecture and Code Optimization, 10(4):41:1–41:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Du:2023:FOS

[DWF^{+}23] Jingwen Du, Fang Wang, Dan Feng, Changchen Gan, Yuchao Cao, Xiaomin Zou, and Fan Li. Fast one-sided RDMA-based state machine replication for disaggregated memory. ACM Transactions on Architecture and Code Optimization, 20(2): 31:1–31:??, June 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3587096.

Dong:2011:HCU

[DXMJ11] Xiangyu Dong, Yuan Xie, Naveen Muralimanohar, and Norman P. Jouppi. Hybrid checkpointing using emerging nonvolatile memories for future exascale systems. ACM Transactions on Architecture and Code Optimization, 8(2): 6:1–6:??, July 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Du:2013:DCC

[DZC^{+}13] Yu Du, Miao Zhou, Bruce Childers, Rami Melhem, and Daniel Mossé. Delta-compressed caching for overcoming the write bandwidth limitation of hybrid main memory. ACM Transactions on Architecture and Code Optimization, 9(4): 55:1–55:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ding:2024:DCE

[DZL^{+}24] Chen Ding, Jian Zhou, Kai Lu, Sicen Li, Yiqin Xiong, Jiguang Wan, and Ling Zhan. D^{2}Comp: Efficient offload of LSM-tree compaction with data processing units on disaggregated storage. ACM Transactions on Architecture and Code Optimization, 21(3): 46:1–46:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3656584.

Daruwalla:2020:BVC

[DZSL20] Kyle Daruwalla, Heng Zhuo, Rohit Shukla, and Mikko Lipasti. BitSAD v2: Compiler optimization and analysis for bitstream computing. ACM Transactions on Architecture and Code Optimization, 16(4): 43:1–43:25, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3364999.

Elkhouly:2020:CSC

[EAH^{+}20] Reem Elkhouly, Mohammad Alshboul, Akihiro Hayashi, Yan Solihin, and Keiji Kimura. Compiler-support for critical data persistence in NVM. ACM Transactions on Architecture and Code Optimization, 16(4):54:1–54:25, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3371236.

Eyerman:2009:MLP

[EE09] Stijn Eyerman and Lieven Eeckhout. Memory-level parallelism aware fetch policies for simultaneous multithreading processors. ACM Transactions on Architecture and Code Optimization, 6(1):3:1– 3:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Eyerman:2011:FGD

[EE11] Stijn Eyerman and Lieven Eeckhout. Fine-grained DVFS using on-chip regulators. ACM Transactions on Architecture and Code Optimization, 8(1): 1:1–1:??, April 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Eyerman:2012:PMJ

[EE12] Stijn Eyerman and Lieven Eeckhout. Probabilistic modeling for job symbiosis scheduling on SMT processors. ACM Transactions on Architecture and Code Optimization, 9(2): 7:1–7:??, June 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Eris:2023:PRF

[ELE^{+}23] Furkan Eris, Marcia Louis, Kubra Eris, José Abellán, and Ajay Joshi. Puppeteer: a random forest based manager for hardware prefetchers across the memory hierarchy. ACM Transactions on Architecture and Code Optimization, 20(1): 19:1–19:??, March 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3570304.

Eyerman:2014:MTM

[EMR14] Stijn Eyerman, Pierre Michaud, and Wouter Rogiest. Multiprogram throughput metrics: a systematic approach. ACM Transactions on Architecture and Code Optimization, 11(3): 34:1–34:??, October 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Evtyushkin:2016:UMC

[EPAG16] Dmitry Evtyushkin, Dmitry Ponomarev, and Nael Abu-Ghazaleh. Understanding and mitigating covert channels through branch predictors. ACM Transactions on Architecture and Code Optimization, 13(1):10:1–10:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Endo:2017:IBV

[EPS17] Fernando A. Endo, Arthur Perais, and André Seznec. On the interactions between value prediction and compiler optimizations in the context of EOLE. ACM Transactions on Architecture and Code Optimization, 14(2):18:1–18:??, July 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ejaz:2018:DDD

[EPS18] Ahsen Ejaz, Vassilios Papaefstathiou, and Ioannis Sourdis. DDRNoC: Dual datarate network-on-chip. ACM Transactions on Architecture and Code Optimization, 15(2): 25:1–25:??, June 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Elwell:2016:RMP

[ERAG^{+}16] Jesse Elwell, Ryan Riley, Nael Abu-Ghazaleh, Dmitry Ponomarev, and Iliano Cervesato. Rethinking memory permissions for protection against cross-layer attacks. ACM Transactions on Architecture and Code Optimization, 12(4): 56:1–56:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Erdem:2020:RDS

[ESB^{+}20] Ahmet Erdem, Cristina Silvano, Thomas Boesch, Andrea Carlo Ornstein, Surinder-Pal Singh, and Giuseppe Desoli. Runtime design space exploration and mapping of DCNNs for the ultra-low-power Orlando SoC. ACM Transactions on Architecture and Code Optimization, 17(2): 11:1–11:25, June 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3379933.

Elango:2015:URM

[ESR^{+}15] Venmugil Elango, Naser Sedaghati, Fabrice Rastello, Louis-Noël Pouchet, J. Ramanujam, Radu Teodorescu, and P. Sadayappan. On using the roofline model with lower bounds on data movement. ACM Transactions on Architecture and Code Optimization, 11(4): 67:1–67:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Espindola:2023:SMR

[EZYA23] Vinicius Espindola, Luciano Zago, Hervé Yviquel, and Guido Araujo. Source matching and rewriting for MLIR using string-based automata. ACM Transactions on Architecture and Code Optimization, 20(2):22:1–22:??, June 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571283.

Fu:2022:GDS

[FBC^{+}22] Yaosheng Fu, Evgeny Bolotin, Niladrish Chatterjee, David Nellans, and Stephen W. Keckler. GPU domain specialization via composable onpackage architecture. ACM Transactions on Architecture and Code Optimization, 19(1): 4:1–4:23, March 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3484505.

Fields:2004:ICS

[FBHN04] Brian A. Fields, Rastislav Bodik, Mark D. Hill, and Chris J. Newburn. Interaction cost and shotgun profiling. ACM Transactions on Architecture and Code Optimization, 1(3):272–304, September 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Fowers:2013:PEC

[FBWS13] Jeremy Fowers, Greg Brown, John Wernsing, and Greg Stitt. A performance and energy comparison of convolution on GPUs, FPGAs, and multicore processors. ACM Transactions on Architecture and Code Optimization, 9(4): 25:1–25:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ferroni:2017:PCM

[FCD^{+}17] Matteo Ferroni, Andrea Corna, Andrea Damiani, Rolando Brondolin, Juan A. Colmenares, Steven Hofmeyr, John D. Kubiatowicz, and Marco D. Santambrogio. Power consumption models for multitenant server infrastructures. ACM Transactions on Architecture and Code Optimization, 14(4):38:1–38:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Santos:2024:AIC

[FCVR24] Fernando Fernandes Dos Santos, Luigi Carro, Flavio Vella, and Paolo Rech. Assessing the impact of compiler optimizations on GPUs reliability. ACM Transactions on Architecture and Code Optimization, 21(2):26:1–26:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638249.

Fang:2014:PPA

[FDF^{+}14] Shuangde Fang, Zidong Du, Yuntan Fang, Yuanjie Huang, Yang Chen, Lieven Eeckhout, Olivier Temam, Huawei Li, Yunji Chen, and Chengyong Wu. Performance portability across heterogeneous SoCs using a generalized library-based approach. ACM Transactions on Architecture and Code Optimization, 11(2):21:1–21:??, June 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Fauzia:2013:BRD

[FER^{+}13] Naznin Fauzia, Venmugil Elango, Mahesh Ravishankar, J. Ramanujam, Fabrice Rastello, Atanas Rountev, Louis-Noël Pouchet, and P. Sadayappan. Beyond reuse distance analysis: Dynamic analysis for characterization of data locality potential. ACM Transactions on Architecture and Code Optimization, 10(4):53:1–53:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Feng:2012:PPL

[FLG12] Min Feng, Changhui Lin, and Rajiv Gupta. PLDS: Partitioning linked data structures for parallelism. ACM Transactions on Architecture and Code Optimization, 8(4):38:1–38:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Fan:2024:IUD

[FLW^{+}24] Zhihua Fan, Wenming Li, Zhen Wang, Yu Yang, Xiaochun Ye, Dongrui Fan, Ninghui Sun, and Xuejun An. Improving utilization of dataflow unit for multi-batch processing. ACM Transactions on Architecture and Code Optimization, 21(1):17:1–17:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3637906.

Fang:2015:MMD

[FMY^{+}15] Zhenman Fang, Sanyam Mehta, Pen-Chung Yew, Antonia Zhai, James Greensky, Gautham Beeraka, and Binyu Zang. Measuring microarchitectural details of multi- and many-core memory systems through microbenchmarking. ACM Transactions on Architecture and Code Optimization, 11(4):55:1–55:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Firtina:2024:AAP

[FPK^{+}24] Can Firtina, Kamlesh Pillai, Gurpreet S. Kalsi, Bharathwaj Suresh, Damla Senol Cali, Jeremie S. Kim, Taha Shahroodi, Meryem Banu Cavlak, Joël Lindegger, Mohammed Alser, Juan Gómez Luna, Sreenivas Subramoney, and Onur Mutlu. ApHMM: Accelerating profile hidden Markov models for fast and energy-efficient genome analysis. ACM Transactions on Architecture and Code Optimization, 21(1):19:1–19:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632950.

France-Pillois:2021:NIT

[FPMR21] Maxime France-Pillois, Jérôme Martin, and Frédéric Rousseau. A non-intrusive tool chain to optimize MPSoC end-to-end systems. ACM Transactions on Architecture and Code Optimization, 18(2):21:1–21:22, March 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3445030.

Fedorov:2013:AAL

[FQRG13] Viacheslav V. Fedorov, Sheng Qiu, A. L. Narasimha Reddy, and Paul V. Gratz. ARI: Adaptive LLC-memory traffic management. ACM Transactions on Architecture and Code Optimization, 10(4): 46:1–46:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ferrari:2023:ADC

[FSP^{+}23] Victor Ferrari, Rafael Sousa, Marcio Pereira, João P. L. De Carvalho, José Nelson Amaral, José Moreira, and Guido Araujo. Advancing direct convolution using convolution slicing optimization and ISA extensions. ACM Transactions on Architecture and Code Optimization, 20(4):54:1–54:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3625004.

Falahati:2024:CCD

[FSX^{+}24] Hajar Falahati, Mohammad Sadrosadati, Qiumin Xu, Juan Gómez-Luna, Banafsheh Saber Latibari, Hyeran Jeon, Shaahin Hesaabi, Hamid Sarbazi-Azad, Onur Mutlu, Murali Annavaram, and Masoud Pedram. Cross-core data sharing for energy-efficient GPUs. ACM Transactions on Architecture and Code Optimization, 21(3):42:1–42:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3653019.

Fung:2009:DWF

[FSYA09] Wilson W. L. Fung, Ivan Sham, George Yuan, and Tor M. Aamodt. Dynamic warp formation: Efficient MIMD control flow on SIMD graphics hardware. ACM Transactions on Architecture and Code Optimization, 6(2): 7:1–7:??, June 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Fursin:2010:COP

[FT10] Grigori Fursin and Olivier Temam. Collective optimization: a practical collaborative approach. ACM Transactions on Architecture and Code Optimization, 7(4):20:1–20:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Feng:2011:DAD

[FTLG11] Min Feng, Chen Tian, Changhui Lin, and Rajiv Gupta. Dynamic access distance driven cache replacement. ACM Transactions on Architecture and Code Optimization, 8(3): 14:1–14:??, October 2011. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Fernandes:2016:EHO

[FWJ^{+}16] Fernando Fernandes, Lucas Weigel, Claudio Jung, Philippe Navaux, Luigi Carro, and Paolo Rech. Evaluation of histogram of oriented gradients soft errors criticality for automotive applications. ACM Transactions on Architecture and Code Optimization, 13(4): 38:1–38:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Fang:2015:PIO

[FXC^{+}15] Shuangde Fang, Wenwen Xu, Yang Chen, Lieven Eeckhout, Olivier Temam, Yunji Chen, Chengyong Wu, and Xiaobing Feng. Practical iterative optimization for the data center. ACM Transactions on Architecture and Code Optimization, 12(2):15:1–15:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Fan:2024:QPQ

[FYI^{+}24] Hao Fan, Yiliang Ye, Shadi Ibrahim, Zhuo Huang, Xingru Li, Weibin Xue, Song Wu, Chen Yu, Xuanhua Shi, and Hai Jin. QoS-pro: a QoSenhanced transaction processing framework for shared SSDs. ACM Transactions on Architecture and Code Optimization, 21(1):6:1–6:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632955.

Gomes:2022:CCA

[GAH22] Cesar Gomes, Maziar Amiraski, and Mark Hempstead. CASHT: Contention analysis in shared hierarchies with thefts. ACM Transactions on Architecture and Code Optimization, 19(1):12:1–12:27, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494538.

Ghandour:2012:LSB

[GAM12] Walid J. Ghandour, Haitham Akkary, and Wes Masri. Leveraging strength-based dynamic information flow analysis to enhance data value prediction. ACM Transactions on Architecture and Code Optimization, 9(1):1:1–1:??, March 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gonzalez-Alvarez:2013:AAD

[GÁSÁ^{+}13] Cecilia González-Álvarez, Jennifer B. Sartor, Carlos Álvarez, Daniel Jiménez-González, and Lieven Eeckhout. Accelerating an application domain with specialized functional units.

ACM Transactions on Architecture and Code Optimization, 10(4):47:1–47:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gonzalez-alvarez:2016:MEF

[GáSÁ^{+}16] Cecilia González-álvarez, Jennifer B. Sartor, Carlos Álvarez, Daniel Jiménez-González, and Lieven Eeckhout. MInGLE: an efficient framework for domain acceleration using low-power specialized functional units. ACM Transactions on Architecture and Code Optimization, 13(2):17:1–17:??, June 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ganusov:2006:FEP

[GB06] Ilya Ganusov and Martin Burtscher. Future execution: a prefetching mechanism that uses multiple cores to speed up single threads. ACM Transactions on Architecture and Code Optimization, 3(4):424– 449, December 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gottscho:2015:DDP

[GBD^{+}15] Mark Gottscho, Abbas BanaiyanMofrad, Nikil Dutt, Alex Nicolau, and Puneet Gupta. DPCS: Dynamic power/capacity scaling for SRAM caches in the nanoscale era. ACM Transactions on Architecture and Code Optimization, 12(3):27:1–27:??, October 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Goswami:2021:TES

[GBD21] Kaustav Goswami, Dip Sankar Banerjee, and Shirshendu Das. Towards enhanced system efficiency while mitigating row hammer. ACM Transactions on Architecture and Code Optimization, 18(4):40:1–40:26, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458749.

Gorgovan:2016:MLO

[GDL16] Cosmin Gorgovan, Amanieu D’antras, and Mikel Luján. MAMBO: a low-overhead dynamic binary modification tool for ARM. ACM Transactions on Architecture and Code Optimization, 13(1):14:1–14:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gracia:2014:RLN

[GFD^{+}14] Darío Suárez Gracia, Alexandra Ferrerón, Luis Montesano Del Campo, Teresa Monreal Arnal, and Víctor Viñals Yúfera. Revisiting LP–NUCA energy consumption: Cache access policies and adaptive block dropping. ACM Transactions on Architecture and Code Optimization, 11(2): 19:1–19:??, June 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Garland:2018:LCM

[GG18] James Garland and David Gregg. Low complexity multiply-accumulate units for convolutional neural networks with weight-sharing. ACM Transactions on Architecture and Code Optimization, 15(3): 31:1–31:??, October 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Garcia-Guirado:2012:DDA

[GGFPRG12] Antonio García-Guirado, Ricardo Fernández-Pascual, Alberto Ros, and José M. García. DAPSCO: Distanceaware partially shared cache organization. ACM Transactions on Architecture and Code Optimization, 8(4):25:1–25:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gareev:2018:HPG

[GGK18] Roman Gareev, Tobias Grosser, and Michael Kruse. Highperformance generalized tensor operations: a compileroriented approach. ACM Transactions on Architecture and Code Optimization, 15(3): 34:1–34:??, October 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ganser:2017:ISO

[GGS^{+}17] Stefan Ganser, Armin Grösslinger, Norbert Siegmund, Sven Apel, and Christian Lengauer. Iterative schedule optimization for parallelization in the polyhedron model. ACM Transactions on Architecture and Code Optimization, 14(3): 23:1–23:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ganser:2019:SIP

[GGS^{+}19] Stefan Ganser, Armin Größlinger, Norbert Siegmund, Sven Apel, and Christian Lengauer. Speeding up iterative polyhedral schedule optimization with surrogate performance models. ACM Transactions on Architecture and Code Optimization, 15(4):56:1–56:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gong:2019:HAG

[GGYK19] Xun Gong, Xiang Gong, Leiming Yu, and David Kaeli. HAWS: Accelerating GPU wavefront execution through selective out-of-order execution. ACM Transactions on Architecture and Code Optimization, 16(2):15:1–15:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gaster:2015:HRA

[GHH15] Benedict R. Gaster, Derek Hower, and Lee Howes. HRFrelaxed: Adapting HRF to the complexities of industrial heterogeneous memory models. ACM Transactions on Architecture and Code Optimization, 12(1):7:1–7:??, April 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Guha:2012:MOD

[GHS12] Apala Guha, Kim Hazelwood, and Mary Lou Soffa. Memory optimization of dynamic binary translators for embedded systems. ACM Transactions on Architecture and Code Optimization, 9(3):22:1–22:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gerards:2013:ODD

[GK13] Marco E. T. Gerards and Jan Kuper. Optimal DPM and DVFS for frame-based realtime systems. ACM Transactions on Architecture and Code Optimization, 9(4):41:1–41:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Georgiou:2017:ETD

[GKCE17] Kyriakos Georgiou, Steve Kerrison, Zbigniew Chamski, and Kerstin Eder. Energy transparency for deeply embedded programs. ACM Transactions on Architecture and Code Optimization, 14(1):8:1– 8:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Goel:2014:SPR

[GKP14] Neeraj Goel, Anshul Kumar, and Preeti Ranjan Panda. Shared-port register file architecture for low-energy VLIW processors. ACM Transactions on Architecture and Code Optimization, 11(1):1:1–1:32, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gondimalla:2023:OOD

[GLTV23] Ashish Gondimalla, Jianqiao Liu, Mithuna Thottethodi, and T. N. Vijaykumar. Occam: Optimal data reuse for convolutional neural networks. ACM Transactions on Architecture and Code Optimization, 20(1): 12:1–12:??, March 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3566052.

Gonzalez-Mesa:2014:ETM

[GMGZP14] M. A. Gonzalez-Mesa, Eladio Gutierrez, Emilio L. Zapata, and Oscar Plata. Effective transactional memory execution management for improved concurrency. ACM Transactions on Architecture and Code Optimization, 11(3): 24:1–24:??, October 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gabor:2009:SLA

[GMW09] Ron Gabor, Avi Mendelson, and Shlomo Weiss. Service level agreement for multithreaded processors. ACM Transactions on Architecture and Code Optimization, 6(2): 6:1–6:??, June 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gysi:2021:DSM

[GMZ^{+}21] Tobias Gysi, Christoph Müller, Oleksandr Zinenko, Stephan Herhut, Eddie Davis, Tobias Wicky, Oliver Fuhrer, Torsten Hoefler, and Tobias Grosser. Domain-specific multi-level IR rewriting for GPU: The Open Earth compiler for GPUaccelerated climate simulation. ACM Transactions on Architecture and Code Optimization, 18(4):51:1–51:23, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3469030.

Guo:2008:EHC

[GNB08] Zhi Guo, Walid Najjar, and Betul Buyukkurt. Efficient hardware code generation for FPGAs. ACM Transactions on Architecture and Code Optimization, 5(1):6:1–6:??, May 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Garzaran:2005:TBS

[GPL^{+}05] María Jesús Garzarán, Milos Prvulovic, José María Llabería, Víctor Viñals, Lawrence Rauchwerger, and Josep Torrellas. Tradeoffs in buffering speculative memory state for thread-level speculation in multiprocessors. ACM Transactions on Architecture and Code Optimization, 2(3):247– 279, September 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Grigorian:2015:ADA

[GR15] Beayna Grigorian and Glenn Reinman. Accelerating divergent applications on SIMD architectures using neural networks. ACM Transactions on Architecture and Code Optimization, 12(1):2:1–2:??, April 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Geraci:2012:TFP

[GS12] James R. Geraci and Sharon M.

Sacco. A transpose-free inplace SIMD optimized FFT. ACM Transactions on Architecture and Code Optimization, 9(3):23:1–23:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Goens:2017:SSS

[GSC17] Andrés Goens, Sergio Siccha, and Jeronimo Castrillon. Symmetry in software synthesis. ACM Transactions on Architecture and Code Optimization, 14(2):20:1–20:??, July 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Guo:2010:QSS

[GSZI10] Fei Guo, Yan Solihin, Li Zhao, and Ravishankar Iyer. Quality of service shared cache management in chip multiprocessor architecture. ACM Transactions on Architecture and Code Optimization, 7(3):14:1–14:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gerzhoy:2020:NMS

[GSZY20] Daniel Gerzhoy, Xiaowu Sun, Michael Zuzak, and Donald Yeung. Nested MIMD–SIMD parallelization for heterogeneous microprocessors. ACM Transactions on Architecture and Code Optimization, 16(4): 48:1–48:27, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3368304.

Gaspar:2016:FAG

[GTT^{+}16] Francisco Gaspar, Luis Taniça, Pedro Tomás, Aleksandar Ilic, and Leonel Sousa. A framework for application-guided task management on heterogeneous embedded systems. ACM Transactions on Architecture and Code Optimization, 12(4):42:1–42:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Georgakoudis:2017:SSA

[GVT^{+}17] Giorgis Georgakoudis, Hans Vandierendonck, Peter Thoman, Bronis R. De Supinski, Thomas Fahringer, and Dimitrios S. Nikolopoulos. SCALO: Scalabilityaware parallelism orchestration for multi-threaded workloads. ACM Transactions on Architecture and Code Optimization, 14(4):54:1–54:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Golander:2008:HMP

[GW08] Amit Golander and Shlomo Weiss. Hiding the misprediction penalty of a resourceefficient high-performance processor. ACM Transactions on Architecture and Code Optimization, 4(4):6:1–6:??, January 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Golander:2009:CAR

[GW09] Amit Golander and Shlomo Weiss. Checkpoint allocation and release. ACM Transactions on Architecture and Code Optimization, 6(3):10:1–10:??, September 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gabor:2007:FES

[GWM07] Ron Gabor, Shlomo Weiss, and Avi Mendelson. Fairness enforcement in switch on event multithreading. ACM Transactions on Architecture and Code Optimization, 4(3):15:1– 15:??, September 2007. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gavin:2013:RIF

[GWS13] Peter Gavin, David Whalley, and Magnus Själander. Reducing instruction fetch energy in multi-issue processors. ACM Transactions on Architecture and Code Optimization, 10(4):64:1–64:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gao:2022:ACM

[GWZ22] Lan Gao, Jing Wang, and Weigong Zhang. Adaptive contention management for fine-grained synchronization on commodity GPUs. ACM Transactions on Architecture and Code Optimization, 19(4): 58:1–58:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3547301.

Han:2013:PEP

[HAC13] Kyuseung Han, Junwhan Ahn, and Kiyoung Choi. Powerefficient predication techniques for acceleration of control flow execution on CGRA. ACM Transactions on Architecture and Code Optimization, 10(2): 8:1–8:??, May 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hasenplaugh:2012:GBC

[HAJ^{+}12] William Hasenplaugh, Pritpal S. Ahuja, Aamer Jaleel, Simon Steely, Jr., and Joel Emer. The gradient-based cache partitioning algorithm. ACM Transactions on Architecture and Code Optimization, 8(4):44:1–44:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ham:2017:DDS

[HAM17] Tae Jun Ham, Juan L. Aragón, and Margaret Martonosi. Decoupling data supply from computation for latency-tolerant communication in heterogeneous architectures. ACM Transactions on Architecture and Code Optimization, 14(2): 16:1–16:??, July 2017. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ham:2019:EDS

[HAM19] Tae Jun Ham, Juan L. Aragón, and Margaret Martonosi. Efficient data supply for parallel heterogeneous architectures. ACM Transactions on Architecture and Code Optimization, 16(2):9:1–9:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Huzaifa:2020:IKR

[HAM^{+}20] Muhammad Huzaifa, Johnathan Alsop, Abdulrahman Mahmoud, Giordano Salvador, Matthew D. Sinclair, and Sarita V. Adve. Interkernel reuse-aware thread block scheduling. ACM Transactions on Architecture and Code Optimization, 17(3): 24:1–24:27, August 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3406538.

Hoseinzadeh:2016:SSP

[HASA16] Morteza Hoseinzadeh, Mohammad Arjomand, and Hamid Sarbazi-Azad. SPCM: The striped phase change memory. ACM Transactions on Architecture and Code Optimization, 12(4):38:1–38:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hwang:2024:CTM

[HBPH24] Soojin Hwang, Daehyeon Baek, Jongse Park, and Jaehyuk Huh. Cerberus: Triple mode acceleration of sparse matrix and vector multiplication. ACM Transactions on Architecture and Code Optimization, 21(2):38:1–38:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3653020.

Huang:2014:HHH

[HCC^{+}14] Yongbing Huang, Licheng Chen, Zehan Cui, Yuan Ruan, Yungang Bao, Mingyu Chen, and Ninghui Sun. HMTT: a hybrid hardware/software tracing system for bridging the DRAM access trace’s semantic gap. ACM Transactions on Architecture and Code Optimization, 11(1):7:1–7:25, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ho:2021:GFD

[HDW21] Nhut-Minh Ho, Himeshi De Silva, and Weng-Fai Wong. GRAM: a framework for dynamically mixing precisions in GPU applications. ACM Transactions on Architecture and Code Optimization, 18(2): 19:1–19:24, March 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3441830.

Heirman:2021:ASE

[HEDH21] Wim Heirman, Stijn Eyerman, Kristof Du Bois, and Ibrahim Hur. Automatic sublining for efficient sparse memory accesses. ACM Transactions on Architecture and Code Optimization, 18(3):33:1–33:23, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3452141.

Hohenauer:2009:SOF

[HEL^{+}09] Manuel Hohenauer, Felix Engel, Rainer Leupers, Gerd Ascheid, and Heinrich Meyr. A SIMD optimization framework for retargetable compilers. ACM Transactions on Architecture and Code Optimization, 6(1):2:1–2:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hroub:2017:EGC

[HEMK17] Ayman Hroub, M. E. S. Elrabaa, M. F. Mudawar, and A. Khayyat. Efficient generation of compact execution traces for multicore architectural simulations. ACM Transactions on Architecture and Code Optimization, 14(3): 27:1–27:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Huang:2023:STE

[HFL^{+}23] Dong Huang, Dan Feng, Qiankun Liu, Bo Ding, Wei Zhao, Xueliang Wei, and Wei Tong. SplitZNS: Towards an efficient LSM-tree on zoned namespace SSDs. ACM Transactions on Architecture and Code Optimization, 20(3): 45:1–45:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3608476.

Hong:2016:OCT

[HHC^{+}16] Ding-Yong Hong, Chun-Chen Hsu, Cheng-Yi Chou, Wei-Chung Hsu, Pangfeng Liu, and Jan-Jan Wu. Optimizing control transfer and memory virtualization in full system emulators. ACM Transactions on Architecture and Code Optimization, 12(4):47:1–47:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Huang:2022:AVC

[HHW^{+}22] Horng-Ruey Huang, Ding-Yong Hong, Jan-Jan Wu, Kung-Fu Chen, Pangfeng Liu, and Wei-Chung Hsu. Accelerating video captioning on heterogeneous system architectures. ACM Transactions on Architecture and Code Optimization, 19(3):38:1–38:25, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3527609.

Hijaz:2014:NLN

[HK14] Farrukh Hijaz and Omer Khan. NUCA-L1: a nonuniform access latency level-1 cache architecture for multicores operating at nearthreshold voltages. ACM Transactions on Architecture and Code Optimization, 11(3): 29:1–29:??, October 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hadjilambrou:2019:CCO

[HKA^{+}19] Zacharias Hadjilambrou, Marios Kleanthous, Georgia Antoniou, Antoni Portero, and Yiannakis Sazeides. Comprehensive characterization of an open source document search engine. ACM Transactions on Architecture and Code Optimization, 16(2):19:1–19:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hwang:2007:SSA

[HL07] Yuan-Shin Hwang and Jia-Jhe Li. Snug set-associative caches: Reducing leakage power of instruction and data caches with no performance penalties. ACM Transactions on Architecture and Code Optimization, 4(1):6:1–6:28, March 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hwang:2010:DCR

[HLC10] Yuan-Shin Hwang, Tzong-Yen Lin, and Rong-Guey Chang. DisIRer: Converting a retargetable compiler into a multiplatform binary translator. ACM Transactions on Architecture and Code Optimization, 7(4):18:1–18:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hagiescu:2013:GCG

[HLR^{+}13] Andrei Hagiescu, Bing Liu, R. Ramanathan, Sucheendra K. Palaniappan, Zheng Cui, Bipasa Chattopadhyay, P. S. Thiagarajan, and Weng-Fai Wong. GPU code generation for ODE-based applications with phased shareddata access patterns. ACM Transactions on Architecture and Code Optimization, 10(4): 55:1–55:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Han:2022:CEC

[HLSK22] Ruobing Han, Jaewon Lee, Jaewoong Sim, and Hyesoon Kim. COX : Exposing CUDA warp-level functions to CPUs. ACM Transactions on Architecture and Code Optimization, 19(4):59:1–59:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3554736.

Huang:2017:IEG

[HLSW17] Libo Huang, Yashuai Lü, Li Shen, and Zhiying Wang. Improving the efficiency of GPGPU work-queue through data awareness. ACM Transactions on Architecture and Code Optimization, 14(4): 45:1–45:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Huan:2024:TNT

[HLZ^{+}24] Chengying Huan, Yongchao Liu, Heng Zhang, Shuaiwen Song, Santosh Pandey, Shiyang Chen, Xiangfei Fang, Yue Jin, Baptiste Lepers, Yanjun Wu, and Hang Liu. TEA+: a novel temporal graph random walk engine with hybrid storage architecture. ACM Transactions on Architecture and Code Optimization, 21(2):

37:1–37:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3652604.

Holey:2015:PEC

[HMYZ15] Anup Holey, Vineeth Mekkat, Pen-Chung Yew, and Antonia Zhai. Performance-energy considerations for shared cache management in a heterogeneous multicore processor. ACM Transactions on Architecture and Code Optimization, 12(1):3:1–3:??, April 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hadidi:2017:CCA

[HNKK17] Ramyad Hadidi, Lifeng Nai, Hyojong Kim, and Hyesoon Kim. CAIRO: a compilerassisted technique for enabling instruction-level offloading of processing-in-memory. ACM Transactions on Architecture and Code Optimization, 14(4): 48:1–48:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hartstein:2004:OPD

[HP04] A. Hartstein and Thomas R. Puzak. The optimum pipeline depth considering both power and performance. ACM Transactions on Architecture and Code Optimization, 1(4):369– 388, December 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hassan:2021:RCM

[HPBS21] Muhammad Hassan, Chang Hyun Park, and David Black-Schaffer. A reusable characterization of the memory system behavior of SPEC2017 and SPEC2006. ACM Transactions on Architecture and Code Optimization, 18(2): 24:1–24:20, March 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3446200.

Haskins:2005:AWS

[HS05] John W. Haskins, Jr. and Kevin Skadron. Accelerated warmup for sampled microarchitecture simulation. ACM Transactions on Architecture and Code Optimization, 2(1): 78–108, March 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hazelwood:2006:MBC

[HS06] Kim Hazelwood and Michael D.

Smith. Managing bounded code caches in dynamic binary optimization systems. ACM Transactions on Architecture and Code Optimization, 3(3): 263–294, September 2006. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hu:2006:EMM

[HVJ06] Shiwen Hu, Madhavi Valluri, and Lizy Kurian John. Effective management of multiple configurable units using dynamic optimization. ACM Transactions on Architecture and Code Optimization, 3(4): 477–501, December 2006. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hiser:2011:EIB

[HWH^{+}11] Jason D. Hiser, Daniel W. Williams, Wei Hu, Jack W. Davidson, Jason Mars, and Bruce R. Childers. Evaluating indirect branch handling mechanisms in software dynamic translation systems. ACM Transactions on Architecture and Code Optimization, 8(2):9:1–9:??, July 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

He:2015:IHF

[HWJ^{+}15] Dan He, Fang Wang, Hong Jiang, Dan Feng, Jing Ning Liu, Wei Tong, and Zheng Zhang. Improving hybrid FTL by fully exploiting internal SSD parallelism with virtual blocks. ACM Transactions on Architecture and Code Optimization, 11(4):43:1–43:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hong:2019:PTG

[HWL^{+}19] Ding-Yong Hong, Jan-Jan Wu, Yu-Ping Liu, Sheng-Yu Fu, and Wei-Chung Hsu. Processor-tracing guided region formation in dynamic binary translation. ACM Transactions on Architecture and Code Optimization, 15(4): 52:1–52:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Haubl:2014:TTE

[HWM14] Christian Häubl, Christian Wimmer, and Hanspeter Mössenböck. Trace transitioning and exception handling in a trace-based JIT compiler for Java. ACM Transactions on Architecture and Code Optimization, 11(1):6:1–6:26, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Huang:2013:ACM

[HWX^{+}13] Libo Huang, Zhiying Wang, Nong Xiao, Yongwen Wang, and Qiang Dou. Adaptive communication mechanism for accelerating MPI functions in NoC-based multicore processors. ACM Transactions on Architecture and Code Optimization, 10(3):18:1–18:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Haj-Yihia:2015:CDP

[HYBR^{+}15] Jawad Haj-Yihia, Yosi Ben Asher, Efraim Rotem, Ahmad Yasin, and Ran Ginosar. Compiler-directed power management for superscalars. ACM Transactions on Architecture and Code Optimization, 11(4):48:1–48:??, January 2015. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic).

Haj-Yihia:2016:FGP

[HYYBM16] Jawad Haj-Yihia, Ahmad Yasin, Yosi Ben Asher, and Avi Mendelson. Fine-grain power breakdown of modern out-of-order cores and its implications on Skylake-based systems. ACM Transactions on Architecture and Code Optimization, 13(4):56:1–56:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

He:2023:DLS

[HZC^{+}23] Bowen He, Xiao Zheng, Yuan Chen, Weinan Li, Yajin Zhou, Xin Long, Pengcheng Zhang, Xiaowei Lu, Linquan Jiang, Qiang Liu, Dennis Cai, and Xiantao Zhang. DxPU: Largescale disaggregated GPU pools in the datacenter. ACM Transactions on Architecture and Code Optimization, 20(4): 55:1–55:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3617995.

Hartley:2022:JTC

[HZN^{+}22] Tim Hartley, Foivos S. Zakkak, Andy Nisbet, Christos Kotselidis, and Mikel Luján. Just-in-time compilation on ARM — a closer look at callsite code consistency. ACM Transactions on Architecture and Code Optimization, 19(4):

54:1–54:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3546568.

Isaac-Chassande:2024:DHA

[ICEDR24] Valentin Isaac-Chassande, Adrian Evans, Yves Durand, and Frédéric Rousseau. Dedicated hardware accelerators for processing of sparse matrices and vectors: a survey. ACM Transactions on Architecture and Code Optimization, 21(2): 27:1–27:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3640542.

Ipek:2008:EAD

[IMS^{+}08] Engin Ipek, Sally A. McKee, Karan Singh, Rich Caruana, Bronis R. de Supinski, and Martin Schulz. Efficient architectural design space exploration via predictive modeling. ACM Transactions on Architecture and Code Optimization, 4(4):1:1–1:??, January 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Izadpanah:2021:PPT

[IPSD21] Ramin Izadpanah, Christina Peterson, Yan Solihin, and Damian Dechev. PETRA: Persistent transactional nonblocking linked data structures. ACM Transactions on Architecture and Code Optimization, 18(2):23:1–23:26, March 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3446391.

Isailovic:2004:DCQ

[IWP^{+}04] Nemanja Isailovic, Mark Whitney, Yatish Patel, John Kubiatowicz, Dean Copsey, Frederic T. Chong, Isaac L. Chuang, and Mark Oskin. Datapath and control for quantum wires. ACM Transactions on Architecture and Code Optimization, 1(1):34– 61, March 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jothi:2014:TCF

[JA14] Komal Jothi and Haitham Akkary. Tuning the continual flow pipeline architecture with virtual register renaming. ACM Transactions on Architecture and Code Optimization, 11(1):11:1–11:27, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jatala:2017:SSG

[JAK17] Vishwesh Jatala, Jayvant Anantpur, and Amey Karkare. Scratchpad sharing in GPUs. ACM Transactions on Architecture and Code Optimization, 14(2):15:1–15:??, July 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jiang:2024:JRG

[JCG^{+}24] Zhang Jiang, Ying Chen, Xiaoli Gong, Jin Zhang, Wenwen Wang, and Pen-Chung Yew. JiuJITsu: Removing gadgets with safe register allocation for JIT code generation. ACM Transactions on Architecture and Code Optimization, 21(1): 3:1–3:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3631526.

Jiang:2013:HAC

[JDZ^{+}13] Lei Jiang, Yu Du, Bo Zhao, Youtao Zhang, Bruce R. Childers, and Jun Yang. Hardware-assisted cooperative integration of wear-leveling and salvaging for phase change memory. ACM Transactions on Architecture and Code Optimization, 10(2):7:1–7:??, May 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Joshi:2008:DEP

[JEBJ08] Ajay Joshi, Lieven Eeckhout, Robert H. Bell, Jr., and Lizy K. John. Distilling the essence of proprietary workloads into miniature benchmarks. ACM Transactions on Architecture and Code Optimization, 5(2):10:1–10:??, August 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jaleel:2019:DHP

[JED19] Aamer Jaleel, Eiman Ebrahimi, and Sam Duncan. DUCATI: High-performance address translation by extending TLB reach of GPU-accelerated systems. ACM Transactions on Architecture and Code Optimization, 16(1):6:1–6:??, March 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jiang:2020:EBC

[JFK20] Zhen Hang Jiang, Yunsi Fei, and David Kaeli. Exploiting bank conflict-based sidechannel timing leakage of GPUs. ACM Transactions on Architecture and Code Optimization, 16(4):42:1–42:24, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3361870.

Jia:2015:GPP

[JGSM15] Wenhao Jia, Elba Garza, Kelly A. Shaw, and Margaret Martonosi. GPU performance and power tuning using regression trees. ACM Transactions on Architecture and Code Optimization, 12(2):13:1–13:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jiang:2023:HMP

[JHH^{+}23] Jiazhi Jiang, Zijian Huang, Dan Huang, Jiangsu Du, Lin Chen, Ziguan Chen, and Yutong Lu. Hierarchical model parallelism for optimizing inference on many-core processor via decoupled 3D-CNN structure. ACM Transactions on Architecture and Code Optimization, 20(3):42:1–42:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3605149.

Jafri:2021:RTC

[JHHM21] Syed M. A. H. Jafri, Hasan Hassan, Ahmed Hemani, and Onur Mutlu. Refresh triggered computation: Improving the energy efficiency of convolutional neural network accelerators. ACM Transactions on Architecture and Code Optimization, 18(1):2:1–2:29, January 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3417708.

Jin:2023:SBS

[JHQ23] Hai Jin, Zhuo He, and Weizhong Qiang. SpecTerminator: Blocking speculative side channels based on instruction classes on RISC-V. ACM Transactions on Architecture and Code Optimization, 20(1): 15:1–15:??, March 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3566053.

Jimenez:2009:GNB

[Jim09] Daniel A. Jiménez. Generalizing neural branch prediction. ACM Transactions on Architecture and Code Optimization, 5(4):17:1–17:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jantz:2013:ESM

[JK13] Michael R. Jantz and Prasad A. Kulkarni. Exploring single and multilevel JIT compilation policy for modern machines 1. ACM Transactions on Architecture and Code Optimization, 10(4):22:1–22:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jensen:2017:ILD

[JK17] Nicklas Bo Jensen and Sven Karlsson. Improving loop dependence analysis. ACM Transactions on Architecture and Code Optimization, 14(3): 22:1–22:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jeon:2013:RDR

[JLCR13] Myeongjae Jeon, Conglong Li, Alan L. Cox, and Scott Rixner. Reducing DRAM row activations with eager read/write clustering. ACM Transactions on Architecture and Code Optimization, 10(4):43:1–43:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jang:2012:ACO

[JLER12] Choonki Jang, Jaejin Lee, Bernhard Egger, and Soojung Ryu. Automatic code overlay generation and partially redundant code fetch elimination. ACM Transactions on Architecture and Code Optimization, 9(2):10:1–10:??, June 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jin:2018:LCM

[JLJ^{+}18a] Hai Jin, Bo Liu, Wenbin Jiang, Yang Ma, Xuanhua Shi, Bingsheng He, and Shaofeng Zhao. Layer-centric memory reuse and data migration for extreme-scale deep learning on many-core architectures. ACM Transactions on Architecture and Code Optimization, 15(3): 37:1–37:??, October 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jo:2018:DSD

[JLJ^{+}18b] Jae-Eon Jo, Gyu-Hyeon Lee, Hanhwi Jang, Jaewon Lee, Mohammadamin Ajdari, and Jangwoo Kim. DiagSim: Systematically diagnosing simulators for healthy simulations. ACM Transactions on Architecture and Code Optimization, 15(1):4:1–4:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jin:2023:CTC

[JLL^{+}23] Hai Jin, Bo Lei, Haikun Liu, Xiaofei Liao, Zhuohui Duan, Chencheng Ye, and Yu Zhang. A compilation tool for computation offloading in ReRAM-based CIM architectures. ACM Transactions on Architecture and Code Optimization, 20(4):47:1–47:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617686.

Jiang:2020:LLA

[JML^{+}20] Wenbin Jiang, Yang Ma, Bo Liu, Haikun Liu, Bing Bing Zhou, Jian Zhu, Song Wu, and Hai Jin. Layup: Layeradaptive and multi-type intermediateoriented memory optimization for GPU-based CNNs. ACM Transactions on Architecture and Code Optimization, 16(4): 39:1–39:23, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jones:2009:EER

[JOA^{+}09a] Timothy M. Jones, Michael F. P. O’Boyle, Jaume Abella, Antonio González, and Oğuz Ergin. Energy-efficient register caching with compiler assistance. ACM Transactions on Architecture and Code Optimization, 6(4):13:1–13:??, October 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jones:2009:ELE

[JOA^{+}09b] Timothy M. Jones, Michael F. P. O’Boyle, Jaume Abella, Antonio González, and Oğuz Ergin. Exploring the limits of early register release: Exploiting compiler analysis. ACM Transactions on Architecture and Code Optimization, 6(3): 12:1–12:??, September 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jellum:2022:SSA

[JOB^{+}22] Erling Jellum, Milica Orlandić, Edmund Brekke, Tor Johansen, and Torleiv Bryne. Solving sparse assignment problems on FPGAs. ACM Transactions on Architecture and Code Optimization, 19(4): 55:1–55:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3546072.

Jain:2017:CMA

[JPS17] Rahul Jain, Preeti Ranjan Panda, and Sreenivas Subramoney. Cooperative multiagent reinforcement learningbased co-optimization of cores, caches, and on-chip network. ACM Transactions on Architecture and Code Optimization, 14(4):32:1–32:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jung:2021:PPB

[JRH21] Tina Jung, Fabian Ritter, and Sebastian Hack. PICO: a Presburger in-bounds check optimization for compiler-based memory safety instrumentations. ACM Transactions on Architecture and Code Optimization, 18(4):45:1–45:27, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460434.

Jantz:2016:IIP

[JRK16] Michael R. Jantz, Forrest J. Robinson, and Prasad A. Kulkarni. Impact of intrinsic profiling limitations on effectiveness of adaptive optimizations. ACM Transactions on Architecture and Code Optimization, 13(4):44:1–44:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jeon:2009:AAP

[JSH09] Jinseong Jeon, Keoncheol Shin, and Hwansoo Han. Abstracting access patterns of dynamic memory using regular expressions. ACM Transactions on Architecture and Code Optimization, 5(4):18:1–18:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jeong:2013:EET

[JSL13] Yeonghun Jeong, Seongseok Seo, and Jongeun Lee. Evaluatorexecutor transformation for efficient pipelining of loops with conditionals. ACM Transactions on Architecture and Code Optimization, 10(4): 62:1–62:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Juang:2004:IBP

[JSM^{+}04] Philo Juang, Kevin Skadron, Margaret Martonosi, Zhigang Hu, Douglas W. Clark, Philip W. Diodato, and Stefanos Kaxiras. Implementing branch-predictor decay using quasi-static memory cells. ACM Transactions on Architecture and Code Optimization, 1(2):180–219, June 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jiang:2016:TLH

[JYE^{+}16] Chuntao Jiang, Zhibin Yu, Lieven Eeckhout, Hai Jin, Xiaofei Liao, and Chengzhong Xu. Two-level hybrid sampled simulation of multithreaded applications. ACM Transactions on Architecture and Code Optimization, 12(4): 39:1–39:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jiang:2013:PAP

[JYJ^{+}13] Chuntao Jiang, Zhibin Yu, Hai Jin, Chengzhong Xu, Lieven Eeckhout, Wim Heirman, Trevor E. Carlson, and Xiaofei Liao. PCantorSim: Accelerating parallel architecture simulation through fractal-based sampling. ACM Transactions on Architecture and Code Optimization, 10(4): 49:1–49:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jiang:2020:EHE

[JYM20] Lijuan Jiang, Chao Yang, and Wenjing Ma. Enabling highly efficient batched matrix multiplications on SW26010 manycore processor. ACM Transactions on Architecture and Code Optimization, 17(1):3:1–3:23, March 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3378176.

Jahanshahi:2022:PQA

[JYW22] Ali Jahanshahi, Nanpeng Yu, and Daniel Wong. PowerMorph: QoS-aware server power reshaping for data center regulation service. ACM Transactions on Architecture and Code Optimization, 19(3): 36:1–36:27, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524129.

Jia:2022:GND

[JZY^{+}22] Xingguo Jia, Jin Zhang, Boshi Yu, Xingyue Qian, Zhengwei Qi, and Haibing Guan. GiantVM: a novel distributed hypervisor for resource aggregation with DSM-aware optimizations. ACM Transactions on Architecture and Code Optimization, 19(2):20:1–20:27, June 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505251.

Kumar:2022:DAS

[KABS22] Rakesh Kumar, Mehdi Alipour, and David Black-Schaffer. Dependence-aware slice execution to boost MLP in sliceout-of-order cores. ACM Transactions on Architecture and Code Optimization, 19(2): 25:1–25:28, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3506704.

Komuravelli:2015:RCH

[KAC15] Rakesh Komuravelli, Sarita V.

Adve, and Ching-Tsun Chou. Revisiting the complexity of hardware cache coherence and some implications. ACM Transactions on Architecture and Code Optimization, 11(4): 37:1–37:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kim:2018:BEE

[KAC^{+}18] Namhyung Kim, Junwhan Ahn, Kiyoung Choi, Daniel Sanchez, Donghoon Yoo, and Soojung Ryu. Benzene: an energy-efficient distributed hybrid cache architecture for manycore systems. ACM Transactions on Architecture and Code Optimization, 15(1):

10:1–10:??, April 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kaeli:2020:EME

[Kae20] Dave Kaeli. Editorial: a message from the Editor-in-Chief. ACM Transactions on Architecture and Code Optimization, 17(3):16:1–16:2, August 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3409369.

Kluter:2014:VWL

[KBB^{+}14] Theo Kluter, Samuel Burri, Philip Brisk, Edoardo Charbon, and Paolo Ienne. Virtual ways: Low-cost coherence for instruction set extensions with architecturally visible storage. ACM Transactions on Architecture and Code Optimization, 11(2):15:1–15:26, July 2014. ISSN 1544-3566 (print), 1544-3973 (electronic).

Khan:2013:SBA

[KBR^{+}13] Malik Khan, Protonu Basu, Gabe Rudy, Mary Hall, Chun Chen, and Jacqueline Chame. A script-based autotuning compiler system to generate high-performance CUDA code. ACM Transactions on Architecture and Code Optimization, 9(4):31:1–31:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kritikakou:2013:NOM

[KCA^{+}13] Angeliki Kritikakou, Francky Catthoor, George S. Athanasiou, Vasilios Kelefouras, and Costas Goutis. Near-optimal microprocessor and accelerators codesign with latency and throughput constraints. ACM Transactions on Architecture and Code Optimization, 10(2): 6:1–6:??, May 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kritikakou:2014:SNO

[KCKG14] Angeliki Kritikakou, Francky Catthoor, Vasilios Kelefouras, and Costas Goutis. A scalable and near-optimal representation of access schemes for memory management. ACM Transactions on Architecture and Code Optimization, 11 (1):13:1–13:25, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kim:2013:FMS

[KCP13] Wonsub Kim, Yoonseo Choi, and Haewoo Park. Fast modulo scheduler utilizing patternized routes for coarsegrained reconfigurable architectures. ACM Transactions on Architecture and Code Optimization, 10(4):58:1–58:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Korostelev:2023:YCL

[KDMA23] Ivan Korostelev, João P. L. De Carvalho, José Moreira, and José Nelson Amaral. YaConv: Convolution with low cache footprint. ACM Transactions on Architecture and Code Optimization, 20(1):18:1–18:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570305.

Kafshdooz:2015:DSS

[KE15] Morteza Mohajjel Kafshdooz and Alireza Ejlali. Dynamic shared SPM reuse for real-time multicore embedded systems. ACM Transactions on Architecture and Code Optimization, 12(2):12:1–12:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kayraklioglu:2018:LLA

[KFEG18] Engin Kayraklioglu, Michael P. Ferguson, and Tarek El-Ghazawi. LAPPS: Localityaware productive prefetching support for PGAS. ACM Transactions on Architecture and Code Optimization, 15(3): 28:1–28:??, October 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Koraei:2020:DSS

[KFJ20] Mostafa Koraei, Omid Fatemi, and Magnus Jahre. DCMI: a scalable strategy for accelerating iterative stencil loops on FPGAs. ACM Transactions on Architecture and Code Optimization, 16(4):36:1–36:24, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kourtis:2010:ECO

[KGK10] Kornilios Kourtis, Georgios Goumas, and Nectarios Koziris. Exploiting compression opportunities to improve SpMxV performance on shared memory systems. ACM Transactions on Architecture and Code Optimization, 7(3):16:1–16:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kondguli:2018:CME

[KH18] Sushant Kondguli and Michael Huang. A case for a more effective, power-efficient turbo boosting. ACM Transactions on Architecture and Code Optimization, 15(1):5:1–5:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Khan:2020:SMS

[KHB^{+}20] Asif Ali Khan, Fazal Hameed, Robin Bläsing, Stuart S. P. Parkin, and Jeronimo Castrillon. ShiftsReduce: Minimizing shifts in Racetrack Memory 4.0. ACM Transactions on Architecture and Code Optimization, 16(4):56:1–56:23, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3372489.

Kerschbaumer:2013:IFT

[KHL^{+}13] Christoph Kerschbaumer, Eric Hennigan, Per Larsen, Stefan Brunthaler, and Michael Franz. Information flow tracking meets just-in-time compilation. ACM Transactions on Architecture and Code Optimization, 10(4):38:1–38:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kim:2018:CEC

[KHN^{+}18] Hyojong Kim, Ramyad Hadidi, Lifeng Nai, Hyesoon Kim, Nuwan Jayasena, Yasuko Eckert, Onur Kayiran, and Gabriel Loh. CODA: Enabling co-location of computation and data for multiple GPU systems. ACM Transactions on Architecture and Code Optimization, 15(3): 32:1–32:??, October 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kaitoua:2014:HED

[KHS^{+}14] Abdulrahman Kaitoua, Hazem Hajj, Mazen A. R. Saghir, Hassan Artail, Haitham Akkary, Mariette Awad, Mageda Sharafeddine, and Khaleel Mershad. Hadoop extensions for distributed computing on reconfigurable active SSD clusters. ACM Transactions on Architecture and Code Optimization, 11(2):22:1–22:??, June 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kulkarni:2005:FES

[KHW^{+}05] Prasad A. Kulkarni, Stephen R.

Hines, David B. Whalley, Jason D. Hiser, Jack W. Davidson, and Douglas L. Jones. Fast and efficient searches for effective optimization-phase sequences. ACM Transactions on Architecture and Code Optimization, 2(2):165–198, June 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kanuparthi:2015:RIC

[KK15] Arun Kanuparthi and Ramesh Karri. Reliable integrity checking in multicore processors. ACM Transactions on Architecture and Code Optimization, 12(2):10:1–10:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kurt:2016:UAS

[KKAR16] Mehmet Can Kurt, Sriram Krishnamoorthy, Gagan Agrawal, and Bin Ren. Userassisted store recycling for dynamic task graph schedulers. ACM Transactions on Architecture and Code Optimization, 13(4):55:1–55:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kang:2024:IAG

[KKL^{+}24] Seokwon Kang, Jongbin Kim, Gyeongyong Lee, Jeongmyung Lee, Jiwon Seo, Hyungsoo Jung, Yong Ho Song, and Yongjun Park. ISP Agent: a generalized instorage-processing workload offloading framework by providing multiple optimization opportunities. ACM Transactions on Architecture and Code Optimization, 21(1): 11:1–11:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3632951.

Kawahito:2013:IRF

[KKM^{+}13] Motohiro Kawahito, Hideaki Komatsu, Takao Moriyama, Hiroshi Inoue, and Toshio Nakatani. Idiom recognition framework using topological embedding. ACM Transactions on Architecture and Code Optimization, 10(3): 13:1–13:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Karimi:2015:MMA

[KKW^{+}15] Naghmeh Karimi, Arun Karthik Kanuparthi, Xueyang Wang, Ozgur Sinanoglu, and Ramesh Karri. MAGIC: Malicious aging in circuits/cores. ACM Transactions on Architecture and Code Optimization, 12(1): 5:1–5:??, April 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kronawitter:2019:PSS

[KL19] Stefan Kronawitter and Christian Lengauer. Polyhedral search space exploration in the ExaStencils code generator. ACM Transactions on Architecture and Code Optimization, 15(4):40:1–40:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kannan:2019:CIE

[KLA^{+}19] Ram Srivatsa Kannan, Michael Laurenzano, Jeongseob Ahn, Jason Mars, and Lingjia Tang. Caliper: Interference estimator for multi-tenant environments sharing architectural resources. ACM Transactions on Architecture and Code Optimization, 16(3):22:1–22:??, July 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kim:2012:IPN

[KLMP12] Yongjoo Kim, Jongeun Lee, Toan X. Mai, and Yunheung Paek. Improving performance of nested loops on reconfigurable array processors. ACM Transactions on Architecture and Code Optimization, 8(4): 32:1–32:??, January 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kundan:2022:PAP

[KMAK22] Shivam Kundan, Theodoros Marinakis, Iraklis Anagnostopoulos, and Dimitri Kagaris. A pressure-aware policy for contention minimization on multicore systems. ACM Transactions on Architecture and Code Optimization, 19(3):

40:1–40:26, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3524616.

Kumar:2014:EPG

[KMG14] Rakesh Kumar, Alejandro Martínez, and Antonio González. Efficient power gating of SIMD accelerators through dynamic selective devectorization in an HW/SW codesigned environment. ACM Transactions on Architecture and Code Optimization, 11(3):25:1–25:??, October 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kicherer:2012:SPA

[KNBK12] Mario Kicherer, Fabian Nowak, Rainer Buchty, and Wolfgang Karl. Seamlessly portable applications: Managing the diversity of modern heterogeneous systems. ACM Transactions on Architecture and Code Optimization, 8(4):42:1–42:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kanakagiri:2017:MMD

[KPM17] Raghavendra Kanakagiri, Biswabandan Panda, and Madhu Mutyam. MBZip: Multiblock data compression. ACM Transactions on Architecture and Code Optimization, 14(4): 42:1–42:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kim:2021:IRA

[KPM21] Minsu Kim, Jeong-Keun Park, and Soo-Mook Moon. Irregular register allocation for translation of test-pattern programs. ACM Transactions on Architecture and Code Optimization, 18(1):5:1–5:23, January 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3427378.

Kong:2015:CRF

[KPP^{+}15] Martin Kong, Antoniu Pop, Louis-Noël Pouchet, R. Govindarajan, Albert Cohen, and P. Sadayappan. Compiler/ runtime framework for dynamic dataflow parallelization of tiled programs. ACM Transactions on Architecture and Code Optimization, 11(4): 61:1–61:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kaushik:2021:GHP

[KPP21] Anirudh Mohan Kaushik, Gennady Pekhimenko, and Hiren Patel. Gretch: a hardware prefetcher for graph analytics. ACM Transactions on Architecture and Code Optimization, 18(2):18:1–18:25, March 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3439803.

Kalra:2020:ACB

[KPRK20] Charu Kalra, Fritz Previlon, Norm Rubin, and David Kaeli. ArmorAll: Compiler-based resilience targeting GPU applications. ACM Transactions on Architecture and Code Optimization, 17(2):9:1–9:24, June 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3382132.

Kiani:2019:ECP

[KR19] Mohsen Kiani and Amir Rajabzadeh. Efficient cache performance modeling in GPUs using reuse distance analysis. ACM Transactions on Architecture and Code Optimization, 15(4):58:1–58:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Koukos:2016:BHU

[KRHK16] Konstantinos Koukos, Alberto Ros, Erik Hagersten, and Stefanos Kaxiras. Building heterogeneous Unified Virtual Memories (UVMs) without the overhead. ACM Transactions on Architecture and Code Optimization, 13(1):1:1–1:22, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kleanthous:2011:CMD

[KS11] Marios Kleanthous and Yiannakis Sazeides. CATCH: a mechanism for dynamically detecting cache-content-duplication in instruction caches. ACM Transactions on Architecture and Code Optimization, 8(3):11:1–11:??, October 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kalayappan:2016:FRT

[KS16] Rajshekar Kalayappan and Smruti R. Sarangi. Fluid-Check: a redundant threadingbased approach for reliable execution in manycore processors. ACM Transactions on Architecture and Code Optimization, 12(4):55:1–55:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kalaitzidis:2021:LVE

[KS21] Kleovoulos Kalaitzidis and André Seznec. Leveraging value equality prediction for value speculation. ACM Transactions on Architecture and Code Optimization, 18(1): 13:1–13:20, January 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3436821.

Kafshdooz:2016:CTO

[KTAE16] Morteza Mohajjel Kafshdooz, Mohammadkazem Taram, Sepehr Assadi, and Alireza Ejlali. A compile-time optimization method for WCET reduction in real-time embedded systems through block formation. ACM Transactions on Architecture and Code Optimization, 12(4):66:1–66:25, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Krolik:2023:RFQ

[KVH23] Alexander Krolik, Clark Verbrugge, and Laurie Hendren. rNdN: Fast query compilation for NVIDIA GPUs. ACM Transactions on Architecture and Code Optimization, 20(3): 41:1–41:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3603503.

Koh:2009:TPV

[KWCL09] Cheng-Kok Koh, Weng-Fai Wong, Yiran Chen, and Hai Li. Tolerating process variations in large, set-associative caches: The buddy cache. ACM Transactions on Architecture and Code Optimization, 6(2):8:1– 8:??, June 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kotzmann:2008:DJH

[KWM^{+}08] Thomas Kotzmann, Christian Wimmer, Hanspeter Mössenböck, Thomas Rodriguez, Kenneth Russell, and David Cox. Design of the Java HotSpot^{T M} client compiler for Java 6. ACM Transactions on Architecture and Code Optimization, 5(1):7:1–7:??, May 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kulkarni:2009:PEO

[KWTD09] Prasad A. Kulkarni, David B.

Whalley, Gary S. Tyson, and Jack W. Davidson. Practical exhaustive optimization phase order exploration and evaluation. ACM Transactions on Architecture and Code Optimization, 6(1):1:1–1:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lucas:2015:SSS

[LAAMJ15] Jan Lucas, Michael Andersch, Mauricio Alvarez-Mesa, and Ben Juurlink. Spatiotemporal SIMT and scalarization for improving GPU efficiency. ACM Transactions on Architecture and Code Optimization, 12(3): 32:1–32:??, October 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Leverich:2008:CEM

[LAS^{+}08] Jacob Leverich, Hideho Arakida, Alex Solomatnikov, Amin Firoozshahian, Mark Horowitz, and Christos Kozyrakis. Comparative evaluation of memory models for chip multiprocessors. ACM Transactions on Architecture and Code Optimization, 5(3):12:1–12:??, November 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2013:MFM

[LAS^{+}13] Sheng Li, Jung Ho Ahn, Richard D. Strong, Jay B. Brockman, Dean M. Tullsen, and Norman P. Jouppi. The McPAT framework for multicore and manycore architectures: Simultaneously modeling power, area, and timing. ACM Transactions on Architecture and Code Optimization, 10(1):5:1–5:??, April 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lee:2010:AIC

[LB10] Benjamin C. Lee and David Brooks. Applied inference: Case studies in microarchitectural design. ACM Transactions on Architecture and Code Optimization, 7(2):8:1– 8:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2005:ABT

[LBJ05] Tao Li, Ravi Bhargava, and Lizy Kurian John. Adapting branch-target buffer to improve the target predictability of Java code. ACM Transactions on Architecture and Code Optimization, 2(2):109– 130, June 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lustig:2013:TIC

[LBM13] Daniel Lustig, Abhishek Bhattacharjee, and Margaret Martonosi.

TLB improvements for chip multiprocessors: Inter-core cooperative prefetchers and shared last-level TLBs. ACM Transactions on Architecture and Code Optimization, 10(1): 2:1–2:??, April 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Leather:2014:AFG

[LBO14] Hugh Leather, Edwin Bonilla, and Michael O’Boyle. Automatic feature generation for machine learning–based optimising compilation. ACM Transactions on Architecture and Code Optimization, 11 (1):14:1–14:32, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lee:2011:DDE

[LCC11] Hyunjin Lee, Sangyeun Cho, and Bruce R. Childers. DEF-CAM: a design and evaluation framework for defect-tolerant cache memories. ACM Transactions on Architecture and Code Optimization, 8(3):17:1– 17:??, October 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lin:2004:CFS

[LCH^{+}04] Jin Lin, Tong Chen, Wei-Chung Hsu, Pen-Chung Yew, Roy Dz-Ching Ju, Tin-Fook Ngai, and Sun Chan. A compiler framework for speculative optimizations. ACM Transactions on Architecture and Code Optimization, 1(3):247– 271, September 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2014:BBS

[LCL^{+}14] Lei Liu, Zehan Cui, Yong Li, Yungang Bao, Mingyu Chen, and Chengyong Wu. BPM/BPM+: Software-based dynamic memory partitioning mechanisms for mitigating DRAM bank-/channel-level interferences in multicore systems. ACM Transactions on Architecture and Code Optimization, 11(1):5:1–5:28, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Labini:2021:APM

[LCP^{+}21] Paolo Sylos Labini, Marco Cianfriglia, Damiano Perri, Osvaldo Gervasi, Grigori Fursin, Anton Lokhmotov, Cedric Nugteren, Bruno Carpentieri, Fabiana Zollo, and Flavio Vella. On the anatomy of predictive models for accelerating GPU convolution kernels and beyond. ACM Transactions on Architecture and Code Optimization, 18(1): 16:1–16:24, January 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3434402.

Lee:2019:SLS

[LCS^{+}19] Matthew Kay Fei Lee, Yingnan Cui, Thannirmalai Somu, Tao Luo, Jun Zhou, Wai Teng Tang, Weng-Fai Wong, and Rick Siow Mong Goh. A system-level simulator for RRAM-based neuromorphic computing chips. ACM Transactions on Architecture and Code Optimization, 15(4): 64:1–64:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2023:GGM

[LCW^{+}23] Shiyi Li, Qiang Cao, Shenggang Wan, Wen Xia, and Changsheng Xie. gPPM: a generalized matrix operation and parallel algorithm to accelerate the encoding/decoding process of erasure codes. ACM Transactions on Architecture and Code Optimization, 20(4): 51:1–51:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3625005.

Li:2024:COL

[LCW^{+}24] Qiao Li, Yu Chen, Guanyu Wu, Yajuan Du, Min Ye, Xinbiao Gan, Jie Zhang, Zhirong Shen, Jiwu Shu, and Chun Xue. Characterizing and optimizing LDPC performance on 3D NAND flash memories. ACM Transactions on Architecture and Code Optimization, 21(3):62:1–62:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663478.

Liu:2024:QHQ

[LD24] Lei Liu and Xinglei Dou. QuCloud+: a holistic qubit mapping scheme for single/ multi-programming on 2D/ 3D NISQ quantum computers. ACM Transactions on Architecture and Code Optimization, 21(1):9:1–9:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3631525.

Litz:2015:ECA

[LDC15] Heiner Litz, Ricardo J. Dias, and David R. Cheriton. Efficient correction of anomalies in snapshot isolation transactions. ACM Transactions on Architecture and Code Optimization, 11(4):65:1–65:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lei:2013:VCI

[LDG^{+}13] Yuanwu Lei, Yong Dou, Lei Guo, Jinbo Xu, Jie Zhou, Yazhuo Dong, and Hongjian Li. VLIW coprocessor for IEEE-754 quadruple-precision elementary functions. ACM Transactions on Architecture and Code Optimization, 10(3): 12:1–12:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2022:MRM

[LDL22] Cunlu Li, Dezun Dong, and Xiangke Liao. MUA-Router:

Maximizing the utility-of-allocation for on-chip pipelining routers. ACM Transactions on Architecture and Code Optimization, 19(3): 33:1–33:23, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3519027.

Lin:2019:CCC

[LDMZ19] Zhen Lin, Hongwen Dai, Michael Mantor, and Huiyang Zhou. Coordinated CTA combination and bandwidth partitioning for GPU concurrent kernel execution. ACM Transactions on Architecture and Code Optimization, 16(3): 23:1–23:??, July 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2021:CHC

[LDY^{+}21] Cunlu Li, Dezun Dong, Shazhou Yang, Xiangke Liao, Guangyu Sun, and Yongheng Liu. CIB-HIER: Centralized input buffer design in hierarchical high-radix routers. ACM Transactions on Architecture and Code Optimization, 18(4):50:1–50:21, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468062.

Lee:2016:ACS

[Lee16] Byeongcheol Lee. Adaptive correction of sampling bias in dynamic call graphs. ACM Transactions on Architecture and Code Optimization, 12(4): 45:1–45:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lutz:2013:PAF

[LFC13] Thibaut Lutz, Christian Fensch, and Murray Cole. PAR-TANS: an autotuning framework for stencil computation on multi-GPU systems. ACM Transactions on Architecture and Code Optimization, 9(4): 59:1–59:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Luo:2019:SCT

[LFK19] Chao Luo, Yunsi Fei, and David Kaeli. Side-channel timing attack of RSA on a GPU. ACM Transactions on Architecture and Code Optimization, 16(3):32:1–32:??, August 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3341729.

Li:2009:CDS

[LFX09] Lian Li, Hui Feng, and Jingling Xue. Compiler-directed scratchpad memory management via graph coloring. ACM Transactions on Architecture and Code Optimization, 6(3): 9:1–9:??, September 2009. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2007:CCE

[LGAZ07] Xiaodong Li, Ritu Gupta, Sarita V. Adve, and Yuanyuan Zhou. Cross-component energy management: Joint adaptation of processor and memory. ACM Transactions on Architecture and Code Optimization, 4(3):14:1–14:??, September 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lu:2021:GAG

[LGH^{+}21] Yashuai Lü, Hui Guo, Libo Huang, Qi Yu, Li Shen, Nong Xiao, and Zhiying Wang. GraphPEG: Accelerating graph processing on GPUs. ACM Transactions on Architecture and Code Optimization, 18(3):30:1–30:24, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3450440.

Liu:2024:ECP

[LGH24] Tong-Yu Liu, Jianmei Guo, and Bo Huang. Efficient crossplatform multiplexing of hardware performance counters via adaptive grouping. ACM Transactions on Architecture and Code Optimization, 21(1): 8:1–8:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3629525.

Lee:2016:SML

[LGP^{+}16] Donghyuk Lee, Saugata Ghose, Gennady Pekhimenko, Samira Khan, and Onur Mutlu. Simultaneous multi-layer access: Improving 3D-stacked memory bandwidth at low cost. ACM Transactions on Architecture and Code Optimization, 12(4): 63:1–63:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2017:LLO

[LHC^{+}17] Pengcheng Li, Xiaoyu Hu, Dong Chen, Jacob Brock, Hao Luo, Eddy Z. Zhang, and Chen Ding. LD: Low-overhead GPU race detection without access monitoring. ACM Transactions on Architecture and Code Optimization, 14(1):9:1–9:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2024:IAS

[LHC^{+}24] Qunyou Liu, Darong Huang, Luis Costero, Marina Zapater, and David Atienza. Intermediate address space: virtual memory optimization of heterogeneous architectures for cache-resident workloads. ACM Transactions on Architecture and Code Optimization, 21(3):50:1–50:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3659207.

Long:2024:WZW

[LHS^{+}24] Linbo Long, Shuiyong He, Jingcheng Shen, Renping Liu, Zhenhua Tan, Congming Gao, Duo Liu, Kan Zhong, and Yi Jiang. WA-Zone: Wearaware zone management optimization for LSM-tree on ZNS SSDs. ACM Transactions on Architecture and Code Optimization, 21(1):16:1–16:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3637488.

Liu:2019:ESA

[LHW^{+}19] Yu-Ping Liu, Ding-Yong Hong, Jan-Jan Wu, Sheng-Yu Fu, and Wei-Chung Hsu. Exploiting SIMD asymmetry in ARM-to-x86 dynamic binary translation. ACM Transactions on Architecture and Code Optimization, 16(1):2:1–2:??, March 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lyons:2012:ASS

[LHWB12] Michael J. Lyons, Mark Hempstead, Gu-Yeon Wei, and David Brooks. The accelerator store: a shared memory framework for accelerator-based systems. ACM Transactions on Architecture and Code Optimization, 8(4):48:1–48:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lin:2006:RCG

[LHY^{+}06] Jin Lin, Wei-Chung Hsu, Pen-Chung Yew, Roy Dz-Ching Ju, and Tin-Fook Ngai. Recovery code generation for general speculative optimizations. ACM Transactions on Architecture and Code Optimization, 3(1):67–89, March 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Luo:2013:DIH

[LHZ13] Yangchun Luo, Wei-Chung Hsu, and Antonia Zhai. The design and implementation of heterogeneous multicore systems for energy-efficient speculative thread execution. ACM Transactions on Architecture and Code Optimization, 10(4): 26:1–26:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lee:2020:SBP

[LIS20] Jaekyu Lee, Yasuo Ishii, and Dam Sunwoo. Securing branch predictors with two-level encryption. ACM Transactions on Architecture and Code Optimization, 17(3):21:1–21:25, August 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3404189.

Lira:2012:MPA

[LJMG12] Javier Lira, Timothy M. Jones, Carlos Molina, and Antonio González. The migration prefetcher: Anticipating data promotion in dynamic NUCA caches. ACM Transactions on Architecture and Code Optimization, 8(4):45:1–45:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2024:GDA

[LK24] Junkaixuan Li and Yi Kang.

GraphSER: Distance-aware stream-based edge repartition for many-core systems. ACM Transactions on Architecture and Code Optimization, 21(3): 48:1–48:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3661998.

Lee:2013:DCD

[LKL^{+}13] Jongwon Lee, Yohan Ko, Kyoungwoo Lee, Jonghee M. Youn, and Yunheung Paek. Dynamic code duplication with vulnerability awareness for soft error detection on VLIW architectures. ACM Transactions on Architecture and Code Optimization, 9(4): 48:1–48:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lee:2012:WPW

[LKV12] Jaekyu Lee, Hyesoon Kim, and Richard Vuduc. When prefetching works, when it doesn’t, and why. ACM Transactions on Architecture and Code Optimization, 9(1):2:1– 2:??, March 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2022:PEP

[LL22] Yuhao Li and Benjamin C. Lee. Phronesis: Efficient performance modeling for highdimensional configuration tuning. ACM Transactions on Architecture and Code Optimization, 19(4):56:1–56:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3546868.

Lee:2022:SBC

[LLC22] Daeyeal Lee, Bill Lin, and Chung-Kuan Cheng. SMTbased contention-free task mapping and scheduling on 2D/3D SMART NoC with mixed dimension-order routing. ACM Transactions on Architecture and Code Optimization, 19(1):5:1–5:21, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487018.

Liu:2022:ICO

[LLLW22] Hongzhi Liu, Jie Luo, Ying Li, and Zhonghai Wu. Iterative compilation optimization based on metric learning and collaborative filtering. ACM Transactions on Architecture and Code Optimization, 19(1): 2:1–2:25, March 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3480250.

Lee:2017:DBT

[LLRC17] Dongwoo Lee, Sangheon Lee, Soojung Ryu, and Kiyoung Choi. Dirty-block tracking in a direct-mapped DRAM cache with self-balancing dispatch. ACM Transactions on Architecture and Code Optimization, 14(2):11:1–11:??, July 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Luinaud:2023:SAD

[LLS23] Thomas Luinaud, J. M. Pierre Langlois, and Yvon Savaria. Symbolic analysis for data plane programs specialization. ACM Transactions on Architecture and Code Optimization, 20(1):1:1–1:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3557727.

Lei:2022:SEW

[LLW^{+}22] Mengya Lei, Fan Li, Fang Wang, Dan Feng, Xiaomin Zou, and Renzhi Xiao. Sec-NVM: an efficient and writefriendly metadata crash consistency scheme for secure NVM. ACM Transactions on Architecture and Code Optimization, 19(1):8:1–8:26, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3488724.

Li:2005:PPC

[LM05] Jian Li and José F. Martínez. Power-performance considerations of parallel computing on chip multiprocessors. ACM Transactions on Architecture and Code Optimization, 2(4): 397–422, December 2005. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2016:SEA

[LMA^{+}16] Qixiao Liu, Miquel Moreto, Jaume Abella, Francisco J. Cazorla, Daniel A. Jimenez, and Mateo Valero. Sensible energy accounting with abstract metering for multicore systems. ACM Transactions on Architecture and Code Optimization, 12(4):60:1–60:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Luque:2013:FCT

[LMCV13] Carlos Luque, Miquel Moreto, Francisco J. Cazorla, and Mateo Valero. Fair CPU time accounting in CMP+SMT processors. ACM Transactions on Architecture and Code Optimization, 9(4):50:1–50:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2013:PTL

[LMJ13a] Yong Li, Rami Melhem, and Alex K. Jones. PS-TLB: Leveraging page classification information for fast, scalable and efficient translation for future CMPs. ACM Transactions on Architecture and Code Optimization, 9(4):28:1–28:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2013:HSA

[LMJ^{+}13b] Qixiao Liu, Miquel Moreto, Victor Jimenez, Jaume Abella, Francisco J. Cazorla, and Mateo Valero. Hardware support for accurate per-task energy metering in multicore systems. ACM Transactions on Architecture and Code Optimization, 10(4):34:1–34:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Long:2008:TMM

[LMMM08] Jieyi Long, Seda Ogrenci Memik, Gokhan Memik, and Rajarshi Mukherjee. Thermal monitoring mechanisms for chip multiprocessors. ACM Transactions on Architecture and Code Optimization, 5(2): 9:1–9:??, August 2008. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lee:2018:IEE

[LMSE18] Hochan Lee, Mansureh S.

Moghaddam, Dongkwan Suh, and Bernhard Egger. Improving energy efficiency of coarsegrain reconfigurable arrays through modulo schedule compression/decompression. ACM Transactions on Architecture and Code Optimization, 15(1): 1:1–1:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lin:2018:GPV

[LMZ18] Zhen Lin, Michael Mantor, and Huiyang Zhou. GPU performance vs. thread-level parallelism: Scalability analysis and a novel way to improve TLP. ACM Transactions on Architecture and Code Optimization, 15(1):15:1–15:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lakshminarasimhan:2022:FSC

[LNFE22] Kartik Lakshminarasimhan, Ajeya Naithani, Josué Feliu, and Lieven Eeckhout. The forward slice core: a high-performance, yet lowcomplexity microarchitecture. ACM Transactions on Architecture and Code Optimization, 19(2):17:1–17:25, June 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3499424.

Lee:2013:TLS

[LNLK13] Junghee Lee, Chrysostomos Nicopoulos, Hyung Gyu Lee, and Jongman Kim. TornadoNoC: a lightweight and scalable on-chip network architecture for the many-core era. ACM Transactions on Architecture and Code Optimization, 10(4):56:1–56:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Louise:2019:FST

[Lou19] Stephane Louise. A first step toward using quantum computing for low-level WCETs estimations. ACM Transactions on Architecture and Code Optimization, 16(3): 29:1–29:??, July 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lim:2017:TEP

[LP17] Hongyeol Lim and Giho Park. Triple Engine Processor (TEP): a heterogeneous near-memory processor for diverse kernel operations. ACM Transactions on Architecture and Code Optimization, 14(4): 49:1–49:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2012:DQM

[LPZI12] Bin Li, Li-Shiuan Peh, Li Zhao, and Ravi Iyer. Dynamic QoS management for chip multiprocessors. ACM Transactions on Architecture and Code Optimization, 9(3):17:1–17:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lotfi:2015:AAC

[LRBG15] Atieh Lotfi, Abbas Rahimi, Luca Benini, and Rajesh K. Gupta. Aging-aware compilation for GP-GPUs. ACM Transactions on Architecture and Code Optimization, 12(2): 24:1–24:??, July 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2010:UBI

[LS10] Fang Liu and Yan Solihin. Understanding the behavior and implications of context switch misses. ACM Transactions on Architecture and Code Optimization, 7(4):21:1–21:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lin:2015:SSE

[LSC^{+}15] Chung-Hsiang Lin, De-Yu Shen, Yi-Jung Chen, Chia-Lin Yang, and Cheng-Yuan Michael Wang. SECRET: a selective error correction framework for refresh energy reduction in DRAMs. ACM Transactions on Architecture and Code Optimization, 12(2):19:1–19:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2023:UBC

[LSH^{+}23] Qiaoyi Liu, Jeff Setter, Dillon Huff, Maxwell Strange, Kathleen Feng, Mark Horowitz, Priyanka Raina, and Fredrik Kjolstad. Unified buffer: Compiling image processing and machine learning applications to push-memory accelerators. ACM Transactions on Architecture and Code Optimization, 20(2):26:1–26:??, June 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572908.

Li:2020:DCP

[LSL20] Yuhao Li, Dan Sun, and Benjamin C. Lee. Dynamic colocation policies with reinforcement learning. ACM Transactions on Architecture and Code Optimization, 17(1):1:1–1:25, March 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3375714.

Lee:2013:APF

[LT13] Sanghoon Lee and James Tuck. Automatic parallelization of fine-grained metafunctions on a chip multiprocessor. ACM Transactions on Architecture and Code Optimization, 10(4):30:1–30:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Leben:2019:PCM

[LT19] Jakob Leben and George Tzanetakis. Polyhedral compilation for multi-dimensional stream processing. ACM Transactions on Architecture and Code Optimization, 16(3):

27:1–27:??, July 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lewis:2012:REC

[LTG12] Adam Wade Lewis, Nian-Feng Tzeng, and Soumik Ghosh. Runtime energy consumption estimation for server workloads based on chaotic timeseries approximation. ACM Transactions on Architecture and Code Optimization, 9(3): 15:1–15:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liao:2016:DPM

[LTX16] Jianwei Liao, François Trahay, and Guoqiang Xiao. Dynamic process migration based on block access patterns occurring in storage servers. ACM Transactions on Architecture and Code Optimization, 13(2): 20:1–20:??, June 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Luporini:2015:CLO

[LVR^{+}15] Fabio Luporini, Ana Lucia Varbanescu, Florian Rathgeber, Gheorghe-Teodor Bercea, J. Ramanujam, David A. Ham, and Paul H. J. Kelly. Crossloop optimization of arithmetic intensity for finite element local assembly. ACM Transactions on Architecture and Code Optimization, 11(4): 57:1–57:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2022:AOM

[LWC^{+}22] Jiansong Li, Xueying Wang, Xiaobing Chen, Guangli Li, Xiao Dong, Peng Zhao, Xianzhi Yu, Yongxin Yang, Wei Cao, Lei Liu, and Xiaobing Feng. An applicationoblivious memory scheduling system for DNN accelerators. ACM Transactions on Architecture and Code Optimization, 19(4):47:1–47:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3535355.

Li:2016:MAP

[LWF^{+}16] Zheng Li, Fang Wang, Dan Feng, Yu Hua, Jingning Liu, and Wei Tong. MaxPB: Accelerating PCM write by maximizing the power budget utilization. ACM Transactions on Architecture and Code Optimization, 13(4):46:1–46:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2011:EEM

[LWH11] Jianjun Li, Chenggang Wu, and Wei-Chung Hsu. Efficient and effective misaligned data access handling in a dynamic binary translation system. ACM Transactions on Architecture and Code Optimization, 8(2):7:1–7:??, July 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lin:2018:GTD

[LWL18] Huanxin Lin, Cho-Li Wang, and Hongyuan Liu. On-GPU thread-data remapping for branch divergence reduction. ACM Transactions on Architecture and Code Optimization, 15(3):39:1–39:??, October 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2019:EGC

[LWS^{+}19] Bingchao Li, Jizeng Wei, Jizhou Sun, Murali Annavaram, and Nam Sung Kim. An efficient GPU cache architecture for applications with irregular memory access patterns. ACM Transactions on Architecture and Code Optimization, 16(3):20:1–20:??, July 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2024:SSR

[LWW^{+}24] Ke Liu, Kan Wu, Hua Wang, Ke Zhou, Peng Wang, Ji Zhang, and Cong Li. SLAP: Segmented reuse-time-label based admission policy for content delivery network caching. ACM Transactions on Architecture and Code Optimization, 21(2):32:1–32:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3646550.

Lankes:2012:BSP

[LWWH12] Andreas Lankes, Thomas Wild, Stefan Wallentowitz, and Andreas Herkersdorf. Benefits of selective packet discard in networks-on-chip. ACM Transactions on Architecture and Code Optimization, 9(2): 12:1–12:??, June 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lu:2016:AFB

[LY16] Zhonghai Lu and Yuan Yao.

Aggregate flow-based performance fairness in CMPs. ACM Transactions on Architecture and Code Optimization, 13(4): 53:1–53:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lee:2024:RRS

[LY24] Ching-Jui Lee and Tsung Tai Yeh. ReSA: Reconfigurable systolic array for multiple tiny DNN tensors. ACM Transactions on Architecture and Code Optimization, 21(3): 43:1–43:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3653363.

Liu:2016:TAA

[LYH16] Peng Liu, Jiyang Yu, and Michael C. Huang. Threadaware adaptive prefetcher on multicore systems: Improving the performance for multithreaded workloads. ACM Transactions on Architecture and Code Optimization, 13(1): 13:1–13:??, April 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lee:2015:NMD

[LYK^{+}15] Do-Heon Lee, Su-Kyung Yoon, Jung-Geun Kim, Charles C. Weems, and Shin-Dug Kim. A new memory-disk integrated system with HW optimizer. ACM Transactions on Architecture and Code Optimization, 12(2):11:1–11:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Luo:2024:CDB

[LYLS24] Longfei Luo, Dingcui Yu, Yina Lv, and Liang Shi. Critical data backup with hybrid flashbased consumer devices. ACM Transactions on Architecture and Code Optimization, 21(1): 1:1–1:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3631529.

Luo:2007:CNP

[LYYB07] Yan Luo, Jia Yu, Jun Yang, and Laxmi N. Bhuyan. Conserving network processor power consumption by exploiting traffic variability. ACM Transactions on Architecture and Code Optimization, 4(1): 4:1–4:26, March 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Luo:2012:DDS

[LZ12] Yangchun Luo and Antonia Zhai. Dynamically dispatching speculative threads to improve sequential execution. ACM Transactions on Architecture and Code Optimization, 9(3): 13:1–13:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2013:CCC

[LZL^{+}13] Yong Li, Yaojun Zhang, Hai LI, Yiran Chen, and Alex K. Jones. C1C: a configurable, compiler-guided STT-RAM L1 cache. ACM Transactions on Architecture and Code Optimization, 10(4):52:1–52:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liang:2014:DCC

[LZM14] Zhibin Liang, Wei Zhang, and Yung-Cheng Ma. Deadlineconstrained clustered scheduling for VLIW architectures using power-gated register files. ACM Transactions on Architecture and Code Optimization, 11(2):20:1–20:26, July 2014. ISSN 1544-3566 (print), 1544-3973 (electronic).

Lu:2024:SLL

[LZS^{+}24] Kai Lu, Siqi Zhao, Haikang Shan, Qiang Wei, Guokuan Li, Jiguang Wan, Ting Yao, Huatao Wu, and Daohui Wang. Scythe: a low-latency RDMAenabled distributed transaction system for disaggregated memory. ACM Transactions on Architecture and Code Optimization, 21(3):57:1–57:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3666004.

Liang:2023:QRC

[LZW23] Yi Liang, Shaokang Zeng, and Lei Wang. Quantifying resource contention of colocated workloads with the system-level entropy. ACM Transactions on Architecture and Code Optimization, 20(1): 10:1–10:??, March 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3563696.

Li:2009:TUC

[LZYZ09] Weijia Li, Youtao Zhang, Jun Yang, and Jiang Zheng. Towards update-conscious compilation for energy-efficient code dissemination in WSNs. ACM Transactions on Architecture and Code Optimization, 6(4):14:1–14:??, October 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2022:DAS

[LZZ^{+}22] Yiding Liu, Xingyao Zhang, Donglin Zhuang, Xin Fu, and Shuaiwen Song. DynamAP: Architectural support for dynamic graph traversal on the automata processor. ACM Transactions on Architecture and Code Optimization, 19(4): 60:1–60:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3556976.

Mehrara:2008:ESP

[MA08] Mojtaba Mehrara and Todd Austin. Exploiting selective placement for low-cost memory protection. ACM Transactions on Architecture and Code Optimization, 5(3):14:1–14:??, November 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mohammadi:2017:COE

[MAD17] Milad Mohammadi, Tor M.

Aamodt, and William J. Dally. CG-OoO: Energyefficient coarse-grain out-of-order execution near in-order energy with near out-of-order performance. ACM Transactions on Architecture and Code Optimization, 14(4): 39:1–39:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mysore:2008:FIP

[MAN^{+}08] Shashidhar Mysore, Banit Agrawal, Rodolfo Neuber, Timothy Sherwood, Nisheeth Shrivastava, and Subhash Suri. Formulating and implementing profiling over adaptive ranges. ACM Transactions on Architecture and Code Optimization, 5(1):2:1–2:??, May 2008.

CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mastoras:2023:DIN

[MAY23] Aristeidis Mastoras, Sotiris Anagnostidis, and Albert-Jan N. Yzelman. Design and implementation for nonblocking execution in GraphBLAS: Tradeoffs and performance. ACM Transactions on Architecture and Code Optimization, 20(1):6:1–6:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3561652.

Malits:2012:ELG

[MBKM12] Roman Malits, Evgeny Bolotin, Avinoam Kolodny, and Avi Mendelson. Exploring the limits of GPGPU scheduling in control flow bound applications. ACM Transactions on Architecture and Code Optimization, 8(4):29:1–29:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Min:2024:CCE

[MBLK24] Dongmoon Min, Ilkwon Byun, Gyu-Hyeon Lee, and Jangwoo Kim. CoolDC: a cost-effective immersion-cooled datacenter with workload-aware temperature scaling. ACM Transactions on Architecture and Code Optimization, 21(3): 51:1–51:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3664925.

Mehta:2013:TSS

[MBY13] Sanyam Mehta, Gautham Beeraka, and Pen-Chung Yew. Tile size selection revisited. ACM Transactions on Architecture and Code Optimization, 10(4):35:1–35:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Majumdar:2012:MPE

[MCB^{+}12] Abhinandan Majumdar, Srihari Cadambi, Michela Becchi, Srimat T. Chakradhar, and Hans Peter Graf. A massively parallel, energy efficient programmable accelerator for learning and classification. ACM Transactions on Architecture and Code Optimization, 9(1):6:1–6:??, March 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Matheou:2015:ASD

[ME15] George Matheou and Paraskevas Evripidou. Architectural support for data-driven execution. ACM Transactions on Architecture and Code Optimization, 11(4):52:1–52:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Matheou:2017:DDC

[ME17] George Matheou and Paraskevas Evripidou. Data-driven concurrency for high performance computing. ACM Transactions on Architecture and Code Optimization, 14(4): 53:1–53:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mummidi:2024:HES

[MFSK24] Chandra Sekhar Mummidi, Victor C. Ferreira, Sudarshan Srinivasan, and Sandip Kundu. Highly efficient selfchecking matrix multiplication on tiled AMX accelerators. ACM Transactions on Architecture and Code Optimization, 21(2):21:1–21:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3633332.

Mccandless:2012:CTI

[MG12] Jason Mccandless and David Gregg. Compiler techniques to improve dynamic branch prediction for indirect jump and call instructions. ACM Transactions on Architecture and Code Optimization, 8(4):24:1– 24:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Malik:2013:OSG

[MG13] Avinash Malik and David Gregg. Orchestrating stream graphs using model checking. ACM Transactions on Architecture and Code Optimization, 10(3):19:1–19:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mastoras:2019:ESE

[MG19] Aristeidis Mastoras and Thomas R.

Gross. Efficient and scalable execution of fine-grained dynamic linear pipelines. ACM Transactions on Architecture and Code Optimization, 16(2): 8:1–8:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mastoras:2020:CDL

[MG20] Aristeidis Mastoras and Thomas R.

Gross. Chunking for dynamic linear pipelines. ACM Transactions on Architecture and Code Optimization, 16(4): 44:1–44:25, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3363815.

Mendonca:2017:DAA

[MGA^{+}17] Gleison Mendonça, Breno Guimarães, Péricles Alves, Márcio Pereira, Guido Araújo, and Fernando Magno Quintão Pereira. DawnCC: Automatic annotation for data parallelism and offloading. ACM Transactions on Architecture and Code Optimization, 14(2): 13:1–13:??, July 2017. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Martinsen:2015:EPT

[MGI15] Jan Kasper Martinsen, Håkan Grahn, and Anders Isberg. The effects of parameter tuning in software thread-level speculation in JavaScript engines. ACM Transactions on Architecture and Code Optimization, 11(4):46:1–46:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Muralidharan:2016:DTN

[MGSH16] Saurav Muralidharan, Michael Garland, Albert Sidelnik, and Mary Hall. Designing a tunable nested data-parallel programming system. ACM Transactions on Architecture and Code Optimization, 13(4): 47:1–47:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Michaud:2016:SMF

[Mic16] Pierre Michaud. Some mathematical facts about optimal cache replacement. ACM Transactions on Architecture and Code Optimization, 13(4): 50:1–50:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Michaud:2018:ATL

[Mic18]  Pierre Michaud.  An al-

ternative TAGE-like conditional branch predictor. ACM Transactions on Architecture and Code Optimization, 15(3): 30:1–30:??, October 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mummidi:2023:AAC

[MK23] Chandra Sekhar Mummidi and Sandip Kundu. ACTION: Adaptive cache block migration in distributed cache architectures. ACM Transactions on Architecture and Code Optimization, 20(2):25:1–25:??, June 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572911.

Michelogiannakis:2022:CIR

[MKC^{+}22] George Michelogiannakis, Benjamin Klenk, Brandon Cook, Min Yee Teh, Madeleine Glick, Larry Dennison, Keren Bergman, and John Shalf. A case for intra-rack resource disaggregation in HPC. ACM Transactions on Architecture and Code Optimization, 19(2): 29:1–29:26, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3514245.

Melot:2015:FCS

[MKKE15] Nicolas Melot, Christoph Kessler, Jörg Keller, and Patrick Eitschberger. Fast crown scheduling heuristics for energy-efficient mapping and scaling of moldable streaming tasks on manycore systems.

ACM Transactions on Architecture and Code Optimization, 11(4):62:1–62:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Moolchandani:2022:PPP

[MKS22] Diksha Moolchandani, Anshul Kumar, and Smruti R. Sarangi. Performance and power prediction for concurrent execution on GPUs. ACM Transactions on Architecture and Code Optimization, 19(3): 35:1–35:27, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3522712.

Menard:2023:HPD

[MLB^{+}23] Christian Menard, Marten Lohstroh, Soroush Bateni, Matthew Chorlian, Arthur Deng, Peter Donovan, Clément Fournier, Shaokai Lin, Felix Suchert, Tassilo Tanneberger, Hokeun Kim, Jeronimo Castrillon, and Edward A. Lee. High-performance deterministic concurrency using Lingua Franca. ACM Transactions on Architecture and Code Optimization, 20(4):48:1–48:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617687.

Ma:2023:OFM

[MLC^{+}23] Wenjing Ma, Fangfang Liu, Daokun Chen, Qinglin Lu, Yi Hu, Hongsen Wang, and Xinhui Yuan. An optimized framework for matrix factorization on the new Sunway many-core platform. ACM Transactions on Architecture and Code Optimization, 20(2): 23:1–23:??, June 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3571856.

Marathe:2006:ACC

[MMdS06] Jaydeep Marathe, Frank Mueller, and Bronis R. de Supinski. Analysis of cachecoherence bottlenecks with hybrid hardware/software techniques. ACM Transactions on Architecture and Code Optimization, 3(4):390–423, December 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mettler:2021:DHM

[MMGS21] Marcel Mettler, Daniel Mueller-Gritschneder, and Ulf Schlichtmann. A distributed hardware monitoring system for runtime verification on multitile MPSoCs. ACM Transactions on Architecture and Code Optimization, 18(1):8:1–8:25, January 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3430699.

Mehrabi:2021:BOE

[MMLS21] Atefeh Mehrabi, Aninda Manocha, Benjamin C. Lee, and Daniel J.

Sorin. Bayesian optimization for efficient accelerator synthesis. ACM Transactions on Architecture and Code Optimization, 18(1):4:1–4:25, January 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3427377.

Michaud:2015:RCM

[MMS15] Pierre Michaud, Andrea Mondelli, and André Seznec. Revisiting clustered microarchitecture for future superscalar cores: a case for wide issue clusters. ACM Transactions on Architecture and Code Optimization, 12(3):28:1–28:??, October 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mazloom:2012:DTI

[MMT^{+}12] Bita Mazloom, Shashidhar Mysore, Mohit Tiwari, Banit Agrawal, and Tim Sherwood. Dataflow tomography: Information flow tracking for understanding and visualizing full systems. ACM Transactions on Architecture and Code Optimization, 9(1):3:1– 3:??, March 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mishra:2024:CHP

[MN24] Prasoon Mishra and V. Krishna Nandivada. COWS for high performance: Cost aware work stealing for irregular parallel loop. ACM Transactions on Architecture and Code Optimization, 21(1):12:1–12:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3633331.

Martins:2016:CBS

[MNC^{+}16] Luiz G. A. Martins, Ricardo Nobre, João M. P. Cardoso, Alexandre C. B. Delbem, and Eduardo Marques. Clusteringbased selection for the exploration of compiler optimization sequences. ACM Transactions on Architecture and Code Optimization, 13(1):8:1– 8:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mcpherson:2016:FPL

[MNSC16] Andrew J. Mcpherson, Vijay Nagarajan, Susmit Sarkar, and Marcelo Cintra. Fence placement for legacy data-race-free programs via synchronization read detection. ACM Transactions on Architecture and Code Optimization, 12(4): 46:1–46:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mattheakis:2013:SRM

[MP13] Pavlos M. Mattheakis and Ioannis Papaefstathiou. Significantly reducing MPI intercommunication latency and power overhead in both embedded and HPC systems. ACM Transactions on Architecture and Code Optimization, 9(4):51:1–51:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Michaud:2022:HHA

[MP22] Pierre Michaud and Anis Peysieux. HAIR: Halving the area of the integer register file with odd/even banking. ACM Transactions on Architecture and Code Optimization, 19(4): 51:1–51:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3544838.

Mpeis:2022:OIC

[MPHL22] Paschalis Mpeis, Pavlos Petoumenos, Kim Hazelwood, and Hugh Leather. Object intersection captures on interactive apps to drive a crowd-sourced replay-based compiler optimization. ACM Transactions on Architecture and Code Optimization, 19(3):32:1–32:25, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3517338.

Manivannan:2018:GDB

[MPPS18] Madhavan Manivannan, Miquel Pericás, Vassilis Papaefstathiou, and Per Stenström. Global dead-block management for task-parallel programs. ACM Transactions on Architecture and Code Optimization, 15(3):33:1–33:??, October 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Minervini:2023:VAE

[MPU^{+}23] Francesco Minervini, Oscar Palomar, Osman Unsal, Enrico Reggiani, Josue Quiroga, Joan Marimon, Carlos Rojas, Roger Figueras, Abraham Ruiz, Alberto Gonzalez, Jonnatan Mendoza, Ivan Vargas, César Hernandez, Joan Cabre, Lina Khoirunisya, Mustapha Bouhali, Julian Pavon, Francesc Moll, Mauro Olivieri, Mario Kovac, Mate Kovac, Leon Dragic, Mateo Valero, and Adrian Cristal. Vitruvius+: an area-efficient RISC-V decoupled vector coprocessor for high performance computing applications. ACM Transactions on Architecture and Code Optimization, 20(2): 28:1–28:??, June 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3575861.

Mukhanov:2017:AFG

[MPW^{+}17] Lev Mukhanov, Pavlos Petoumenos, Zheng Wang, Nikos Parasyris, Dimitrios S. Nikolopoulos, Bronis R. De Supinski, and Hugh Leather. ALEA: a fine-grained energy profiling tool. ACM Transactions on Architecture and Code Optimization, 14(1):1:1–1:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mettler:2022:FBA

[MRK^{+}22] Marcel Mettler, Martin Rapp, Heba Khdr, Daniel Mueller-Gritschneder, Jörg Henkel, and Ulf Schlichtmann. An FPGA-based approach to evaluate thermal and resource management strategies of many-core processors. ACM Transactions on Architecture and Code Optimization, 19(3): 31:1–31:24, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3516825.

Michaud:2007:STM

[MSF^{+}07] Pierre Michaud, André Seznec, Damien Fetis, Yiannakis Sazeides, and Theofanis Constantinou. A study of thread migration in temperatureconstrained multicores. ACM Transactions on Architecture and Code Optimization, 4(2): 9:1–9:??, June 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Metzger:2021:DHT

[MSFC21] Paul Metzger, Volker Seeker, Christian Fensch, and Murray Cole. Device hopping: Transparent mid-kernel runtime switching for heterogeneous systems. ACM Transactions on Architecture and Code Optimization, 18(4):57:1–57:25, December 2021. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3471909.

Meng:2005:ELL

[MSK05] Yan Meng, Timothy Sherwood, and Ryan Kastner. Exploring the limits of leakage power reduction in caches. ACM Transactions on Architecture and Code Optimization, 2(3):221–246, September 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Manocha:2021:GOD

[MST^{+}21] Aninda Manocha, Tyler Sorensen, Esin Tureci, Opeoluwa Matthews, Juan L. Aragón, and Margaret Martonosi. GraphAttack: Optimizing data supply for graph applications on in-order multicore architectures. ACM Transactions on Architecture and Code Optimization, 18(4): 53:1–53:26, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3469846.

Mbakoyiannis:2018:EPC

[MTK18] Dimitrios Mbakoyiannis, Othon Tomoutzoglou, and George Kornaros. Energy-performance considerations for data offloading to FPGA-based accelerators over PCIe. ACM Transactions on Architecture and Code Optimization, 15(1): 14:1–14:??, April 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Miliadis:2024:ASS

[MTPK24] Panagiotis Miliadis, Dimitris Theodoropoulos, Dionisios Pnevmatikatos, and Nectarios Koziris. Architectural support for sharing, isolating and virtualizing FPGA resources. ACM Transactions on Architecture and Code Optimization, 21(2):33:1–33:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3648475.

Mammadli:2019:AGD

[MWJ19] Rahim Mammadli, Felix Wolf, and Ali Jannesari. The art of getting deep neural networks in shape. ACM Transactions on Architecture and Code Optimization, 15(4): 62:1–62:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mehta:2016:VL

[MY16] Sanyam Mehta and Pen-Chung Yew. Variable liberalization. ACM Transactions on Architecture and Code Optimization, 13(3):23:1–23:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Morad:2015:GSP

[MYG15] Amir Morad, Leonid Yavits, and Ran Ginosar. GP–SIMD processing-in-memory. ACM Transactions on Architecture and Code Optimization, 11(4): 53:1–53:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Morad:2016:RGS

[MYKG16] Amir Morad, Leonid Yavits, Shahar Kvatinsky, and Ran Ginosar. Resistive GP-SIMD processing-in-memory. ACM Transactions on Architecture and Code Optimization, 12(4): 57:1–57:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nasre:2013:TSE

[Nas13] Rupesh Nasre. Time- and space-efficient flow-sensitive points-to analysis. ACM Transactions on Architecture and Code Optimization, 10(4): 39:1–39:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nandivada:2013:IBA

[NB13] V. Krishna Nandivada and Rajkishore Barik. Improved bitwidth-aware variable packing. ACM Transactions on Architecture and Code Optimization, 10(3):16:1–16:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nugteren:2015:BAS

[NC15] Cedric Nugteren and Henk Corporaal. Bones: an automatic skeleton-based C-to-CUDA compiler for GPUs. ACM Transactions on Architecture and Code Optimization, 11(4):35:1–35:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nugteren:2013:ASC

[NCC13] Cedric Nugteren, Pieter Custers, and Henk Corporaal. Algorithmic species: a classification of affine loop nests for parallel programming. ACM Transactions on Architecture and Code Optimization, 9(4):40:1–40:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nair:2014:RPD

[NCQ14] Prashant J. Nair, Chia-Chen Chou, and Moinuddin K. Qureshi. Refresh pausing in DRAM memory systems. ACM Transactions on Architecture and Code Optimization, 11(1):10:1–10:26, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/2579669.

Neill:2017:FAM

[NDP17] Richard Neill, Andi Drebes, and Antoniu Pop. Fuse: Accurate multiplexing of hardware performance counters across executions. ACM Transactions on Architecture and Code Optimization, 14(4):43:1–43:??, December 2017. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic).

Nuzman:2013:JTC

[NED^{+}13] Dorit Nuzman, Revital Eres, Sergei Dyshel, Marcel Zalmanovici, and Jose Castanos. JIT technology with C/C++: Feedback-directed dynamic recompilation for statically compiled languages. ACM Transactions on Architecture and Code Optimization, 10(4): 59:1–59:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Na:2016:JPC

[NKH16] Yeoul Na, Seon Wook Kim, and Youngsun Han. JavaScript parallelizing compiler for exploiting parallelism from dataparallel HTML5 applications. ACM Transactions on Architecture and Code Optimization, 12(4):64:1–64:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nagpurkar:2006:ERP

[NMKS06] Priya Nagpurkar, Hussam Mousa, Chandra Krintz, and Timothy Sherwood. Efficient remote profiling for resourceconstrained devices. ACM Transactions on Architecture and Code Optimization, 3(1): 35–66, March 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nejat:2022:CSM

[NMPS22] Mehrzad Nejat, Madhavan Manivannan, Miquel Pericàs, and Per Stenström. Cooperative slack management: Saving energy of multicore processors by trading performance slack between QoS-constrained applications. ACM Transactions on Architecture and Code Optimization, 19(2):21:1–21:27, June 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505559.

Nair:2016:CEP

[NRQ16a] Prashant J. Nair, David A. Roberts, and Moinuddin K. Qureshi. Citadel: Efficiently protecting stacked memory from TSV and large granularity failures. ACM Transactions on Architecture and Code Optimization, 12(4):49:1–49:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nair:2016:FFC

[NRQ16b] Prashant J. Nair, David A. Roberts, and Moinuddin K. Qureshi. FaultSim: a fast, configurable memory-reliability simulator for conventional and 3D-stacked systems. ACM Transactions on Architecture and Code Optimization, 12(4): 44:1–44:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Nematollahi:2021:ENN

[NSF^{+}21] Negin Nematollahi, Mohammad Sadrosadati, Hajar Falahati, Marzieh Barkhordar, Mario Paulo Drumond, Hamid Sarbazi-Azad, and Babak Falsafi. Efficient nearest-neighbor data sharing in GPUs. ACM Transactions on Architecture and Code Optimization, 18(1): 6:1–6:26, January 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3429981.

Negi:2013:SCF

[NTG13] Anurag Negi and Ruben Titos-Gil. SCIN-cache: Fast speculative versioning in multithreaded cores. ACM Transactions on Architecture and Code Optimization, 9(4):58:1–58:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Narayan:2022:AOC

[NTV^{+}22] Aditya Narayan, Yvain Thonnart, Pascal Vivet, Ayse Coskun, and Ajay Joshi. Architecting optically controlled phase change memory. ACM Transactions on Architecture and Code Optimization, 19(4): 48:1–48:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3533252.

Natarajan:2015:LTE

[NZ15] Ragavendra Natarajan and Antonia Zhai. Leveraging transactional execution for memory consistency model emulation. ACM Transactions on Architecture and Code Optimization, 12(3):29:1–29:??, October 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Orosa:2012:FIF

[OAB12] Lois Orosa, Elisardo Antelo, and Javier D. Bruguera. FlexSig: Implementing flexible hardware signatures. ACM Transactions on Architecture and Code Optimization, 8(4): 30:1–30:??, January 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Orosa:2019:AAF

[OAM19] Lois Orosa, Rodolfo Azevedo, and Onur Mutlu. AVPP: Address-first value-next predictor with value prefetching for improving the efficiency of load value prediction. ACM Transactions on Architecture and Code Optimization, 15(4): 49:1–49:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Olgun:2024:SDP

[OBdOJ^{+}24] Ataberk Olgun, F. Nisa Bostanci, Geraldo Francisco de Oliveira Junior, Yahya Can Tugrul, Rahul Bera, Abdullah Giray Yaglikci, Hasan Hassan, Oguz Ergin, and Onur Mutlu. Sectored DRAM: a practical energy-efficient and high-performance fine-grained DRAM architecture. ACM Transactions on Architecture and Code Optimization, 21(3): 60:1–60:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3673653.

Orozco:2012:THT

[OGK^{+}12] Daniel Orozco, Elkin Garcia, Rishi Khan, Kelly Livingston, and Guang R. Gao. Toward high-throughput algorithms on many-core architectures. ACM Transactions on Architecture and Code Optimization, 8(4): 49:1–49:??, January 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Omar:2021:PSH

[OK21] Hamza Omar and Omer Khan. PRISM: Strong hardware isolation-based soft-error resilient multicore architecture with high performance and availability at low hardware overheads. ACM Transactions on Architecture and Code Optimization, 18(3):31:1–31:25, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3450523.

Olson:2022:OAG

[OKJ^{+}22] M. Ben Olson, Brandon Kammerdiener, Michael R. Jantz, Kshitij A. Doshi, and Terry Jones. Online application guidance for heterogeneous memory systems. ACM Transactions on Architecture and Code Optimization, 19(3): 45:1–45:27, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3533855.

Olgun:2023:PHE

[OLK^{+}23] Ataberk Olgun, Juan Gómez Luna, Konstantinos Kanellopoulos, Behzad Salami, Hasan Hassan, Oguz Ergin, and Onur Mutlu. PiDRAM: a holistic end-to-end FPGAbased framework for processingin-DRAM. ACM Transactions on Architecture and Code Optimization, 20(1):8:1–8:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3563697.

Olson:2018:CLM

[OTR^{+}18] Matthew Benjamin Olson, Joseph T. Teague, Divyani Rao, Michael R. JANTZ, Kshitij A. Doshi, and Prasad A. Kulkarni. Cross-layer memory management to improve DRAM energy efficiency. ACM Transactions on Architecture and Code Optimization, 15(2): 20:1–20:??, June 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pananilath:2015:OCG

[PAVB15] Irshad Pananilath, Aravind Acharya, Vinay Vasista, and Uday Bondhugula. An optimizing code generator for a class of lattice-Boltzmann computations. ACM Transactions on Architecture and Code Optimization, 12(2): 14:1–14:??, July 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Panda:2015:CUD

[PB15] Biswabandan Panda and Shankar Balachandran. CAFFEINE: a utility-driven prefetcher aggressiveness engine for multicores. ACM Transactions on Architecture and Code Optimization, 12(3):30:1–30:??, October 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pompougnac:2022:WSR

[PBCB22] Hugo Pompougnac, Ulysse Beaugnon, Albert Cohen, and Dumitru Potop Butucaru. Weaving synchronous reactions into the fabric of SSA-form compilers. ACM Transactions on Architecture and Code Optimization, 19(2): 22:1–22:25, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3506706.

Pu:2017:PHS

[PBY^{+}17] Jing Pu, Steven Bell, Xuan Yang, Jeff Setter, Stephen Richardson, Jonathan Ragan-Kelley, and Mark Horowitz. Programming heterogeneous systems from an image processing DSL. ACM Transactions on Architecture and Code Optimization, 14(3): 26:1–26:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pop:2013:OED

[PC13] Antoniu Pop and Albert Cohen. OpenStream: Expressiveness and data-flow compilation of OpenMP streaming programs. ACM Transactions on Architecture and Code Optimization, 9(4):53:1–53:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Park:2016:CJP

[PCM16] Hyukwoo Park, Myungsu Cha, and Soo-Mook Moon. Concurrent JavaScript parsing for faster loading of Web apps. ACM Transactions on Architecture and Code Optimization, 13(4):41:1–41:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Patsilaras:2012:EEM

[PCT12] George Patsilaras, Niket K. Choudhary, and James Tuck. Efficiently exploiting memory level parallelism on asymmetric coupled cores in the dark silicon era. ACM Transactions on Architecture and Code Optimization, 8(4):28:1–28:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Peterson:2017:TCT

[PD17] Christina Peterson and Damian Dechev. A transactional correctness tool for abstract data types. ACM Transactions on Architecture and Code Optimization, 14(4):37:1–37:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pal:2024:CUA

[PDCS24] Asmita Pal, Keerthana Desai, Rahul Chatterjee, and Joshua San Miguel. Camouflage: Utility-aware obfuscation for accurate simulation of sensitive program traces. ACM Transactions on Architecture and Code Optimization, 21(2): 36:1–36:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3650110.

Peng:2023:FPS

[PDY^{+}23] Bo Peng, Yaozu Dong, Jianguo Yao, Fengguang Wu, and Haibing Guan. FlexHM: a practical system for heterogeneous memory with flexible and efficient performance optimizations. ACM Transactions on Architecture and Code Optimization, 20(1):13:1–13:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3565885.

Pericas:2018:EPA

[Per18] Miquel Pericàs. Elastic Places: an adaptive resource manager for scalable and portable performance. ACM Transactions on Architecture and Code Optimization, 15(2):19:1–19:??, June 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Patil:2017:HHA

[PG17] Adarsh Patil and Ramaswamy Govindarajan. HAShCache: Heterogeneity-aware shared DRAMCache for integrated heterogeneous systems. ACM Transactions on Architecture and Code Optimization, 14(4): 51:1–51:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pusukuri:2012:TTD

[PGB12] Kishore Kumar Pusukuri, Rajiv Gupta, and Laxmi N. Bhuyan. Thread tranquilizer: Dynamically reducing performance variation. ACM Transactions on Architecture and Code Optimization, 8(4):46:1– 46:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pusukuri:2013:AFC

[PGB13] Kishore Kumar Pusukuri, Rajiv Gupta, and Laxmi N.

Bhuyan. ADAPT: a framework for coscheduling multithreaded programs. ACM Transactions on Architecture and Code Optimization, 9(4): 45:1–45:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pusukuri:2016:TEL

[PGB16] Kishore Kumar Pusukuri, Rajiv Gupta, and Laxmi N. Bhuyan. Tumbler: an effective load-balancing technique for multi-CPU multicore systems. ACM Transactions on Architecture and Code Optimization, 12(4):36:1–36:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Proy:2017:CAL

[PHBC17] Julien Proy, Karine Heydemann, Alexandre Berzati, and Albert Cohen. Compilerassisted loop hardening against fault attacks. ACM Transactions on Architecture and Code Optimization, 14(4): 36:1–36:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Purnaprajna:2012:MWI

[PI12] Madhura Purnaprajna and Paolo Ienne. Making wideissue VLIW processors viable on FPGAs. ACM Transactions on Architecture and Code Optimization, 8(4):33:1–33:??, January 2012. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic).

Purini:2013:FGO

[PJ13] Suresh Purini and Lakshya Jain. Finding good optimization sequences covering program space. ACM Transactions on Architecture and Code Optimization, 9(4):56:1–56:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pradelle:2012:PPB

[PKC12] Benoit Pradelle, Alain Ketterlin, and Philippe Clauss. Polyhedral parallelization of binary code. ACM Transactions on Architecture and Code Optimization, 8(4):39:1–39:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Park:2024:HHS

[PKL^{+}24] Joongun Park, Seunghyo Kang, Sanghyeon Lee, Taehoon Kim, Jongse Park, Youngjin Kwon, and Jaehyuk Huh. Hardware-hardened sandbox enclaves for trusted serverless computing. ACM Transactions on Architecture and Code Optimization, 21(1): 13:1–13:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3632954.

Park:2019:ROC

[PKPM19] Hyukwoo Park, Sungkook Kim, Jung-Geun Park, and Soo-Mook Moon. Reusing the optimized code for JavaScript ahead-of-time compilation. ACM Transactions on Architecture and Code Optimization, 15(4):54:1–54:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Puthoor:2023:TBS

[PL23] Sooraj Puthoor and Mikko H.

Lipasti. Turn-based spatiotemporal coherence for GPUs. ACM Transactions on Architecture and Code Optimization, 20(3):33:1–33:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3593054.

Pereira:2019:SPS

[PLG19] Fernando Magno Quintão Pereira, Guilherme Vieira Leobas, and Abdoulaye Gamatié. Static prediction of silent stores. ACM Transactions on Architecture and Code Optimization, 15(4):44:1–44:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Park:2019:MCM

[PLK^{+}19] Jungwoo Park, Myoungjun Lee, Soontae Kim, Minho Ju, and Jeongkyu Hong. MH cache: a multi-retention STT-RAM-based low-power lastlevel cache for mobile hardware rendering systems. ACM Transactions on Architecture and Code Optimization, 16(3): 26:1–26:??, July 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pao:2010:MEP

[PLL10] Derek Pao, Wei Lin, and Bin Liu. A memory-efficient pipelined implementation of the Aho–Corasick stringmatching algorithm. ACM Transactions on Architecture and Code Optimization, 7(2): 10:1–10:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Peng:2024:LER

[PLLW24] Wangqi Peng, Yusen Li, Xiaoguang Liu, and Gang Wang. Lavender: an efficient resource partitioning framework for large-scale job colocation. ACM Transactions on Architecture and Code Optimization, 21(3):58:1–58:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674736.

Porter:2015:MMS

[PLT^{+}15] Leo Porter, Michael A. Laurenzano, Ananta Tiwari, Adam Jundt, William A. Ward, Jr., Roy Campbell, and Laura Carrington. Making the most of SMT in HPC: Systemand application-level perspectives. ACM Transactions on Architecture and Code Optimization, 11(4):59:1–59:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pricopi:2012:BPH

[PM12] Mihai Pricopi and Tulika Mitra. Bahurupi: a polymorphic heterogeneous multi-core architecture. ACM Transactions on Architecture and Code Optimization, 8(4):22:1–22:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Palangappa:2017:CCE

[PM17] Poovaiah M. Palangappa and Kartik Mohanram. CompEx++: Compression-expansion coding for energy, latency, and lifetime improvements in MLC/TLC NVMs. ACM Transactions on Architecture and Code Optimization, 14(1): 10:1–10:??, April 2017. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Prisacari:2013:FPS

[PRMH13] Bogdan Prisacari, German Rodriguez, Cyriel Minkenberg, and Torsten Hoefler. Fast pattern-specific routing for fat tree networks. ACM Transactions on Architecture and Code Optimization, 10(4): 36:1–36:??, December 2013.

CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Proficz:2021:AGA

[Pro21] Jerzy Proficz. All-gather algorithms resilient to imbalanced process arrival patterns. ACM Transactions on Architecture and Code Optimization, 18(4): 41:1–41:22, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3460122.

Premillieu:2012:SSR

[PS12] Nathanael Premillieu and Andre Seznec. SYRANT: SYmmetric Resource Allocation on Not-taken and Taken paths. ACM Transactions on Architecture and Code Optimization, 8(4):43:1–43:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Premillieu:2015:EOE

[PS15] Nathanael Prémillieu and André Seznec. Efficient outof-order execution of guarded ISAs. ACM Transactions on Architecture and Code Optimization, 11(4):41:1–41:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Perez:2023:UDO

[PSL^{+}23] Víctor Pérez, Lukas Sommer, Victor Lomüller, Kumudha Narasimhan, and Mehdi Goli.

User-driven online kernel fusion for SYCL. ACM Transactions on Architecture and Code Optimization, 20(2): 21:1–21:??, June 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3571284.

Patsilaras:2017:RRD

[PT17] George Patsilaras and James Tuck. ReDirect: Reconfigurable directories for multicore architectures. ACM Transactions on Architecture and Code Optimization, 14(4): 50:1–50:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Parasyris:2017:SAP

[PVA^{+}17] Konstantinos Parasyris, Vassilis Vassiliadis, Christos D. Antonopoulos, Spyros Lalis, and Nikolaos Bellas. Significanceaware program execution on unreliable hardware. ACM Transactions on Architecture and Code Optimization, 14(2): 12:1–12:??, July 2017. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pathania:2017:DTM

[PVS^{+}17] Anuj Pathania, Vanchinathan Venkataramani, Muhammad Shafique, Tulika Mitra, and Jörg Henkel. Defragmentation of tasks in many-core architecture. ACM Transactions on Architecture and Code Optimization, 14(1):2:1–2:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Peled:2020:NNP

[PWE20] Leeor Peled, Uri Weiser, and Yoav Etsion. A neural network prefetcher for arbitrary memory access patterns. ACM Transactions on Architecture and Code Optimization, 16(4): 37:1–37:27, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3345000.

Pirkelbauer:2019:BTF

[PWPD19] Peter Pirkelbauer, Amalee Wilson, Christina Peterson, and Damian Dechev. Blaze-Tasks: a framework for computing parallel reductions over tasks. ACM Transactions on Architecture and Code Optimization, 15(4):66:1–66:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Qararyah:2024:EHD

[QAT24] Fareed Qararyah, Muhammad Waqar Azhar, and Pedro Trancoso. An efficient hybrid deep learning accelerator for compact and heterogeneous CNNs. ACM Transactions on Architecture and Code Optimization, 21(2):25:1–25:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639823.

Qureshi:2021:GXM

[QSZ^{+}21] Yasir Mahmood Qureshi, William Andrew Simon, Marina Zapater, Katzalin Olcoz, and David Atienza. Gem5-X: a many-core heterogeneous simulation platform for architectural exploration and optimization. ACM Transactions on Architecture and Code Optimization, 18(4):44:1–44:27, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3461662.

Qi:2014:VVG

[QYZ^{+}14] Zhengwei Qi, Jianguo Yao, Chao Zhang, Miao Yu, Zhizhou Yang, and Haibing Guan. VGRIS: Virtualized GPU resource isolation and scheduling in cloud gaming. ACM Transactions on Architecture and Code Optimization, 11(2): 17:1–17:25, July 2014. ISSN 1544-3566 (print), 1544-3973 (electronic).

Rieber:2022:JPL

[RAF22] Dennis Rieber, Axel Acosta, and Holger Fröning. Joint program and layout transformations to enable convolutional operators on specialized hardware based on constraint programming. ACM Transactions on Architecture and Code Optimization, 19(1):7:1–7:26, March 2022. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3487922.

Ramashekar:2013:ADA

[RB13] Thejas Ramashekar and Uday Bondhugula. Automatic data allocation and buffer management for multi-GPU machines. ACM Transactions on Architecture and Code Optimization, 10(4):60:1–60:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rac:2024:CAS

[RB24] Samuel Rac and Mats Brorsson. Cost-aware service placement and scheduling in the edge-cloud continuum. ACM Transactions on Architecture and Code Optimization, 21(2): 29:1–29:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3640823.

Raghavan:2010:TTP

[RBM10] Arun Raghavan, Colin Blundell, and Milo M. K. Martin. Token tenure and PATCH: a predictive/adaptive tokencounting hybrid. ACM Transactions on Architecture and Code Optimization, 7(2):6:1– 6:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rakvic:2010:TMT

[RCG^{+}10a] R. Rakvic, Q. Cai, J. González, G. Magklis, P. Chaparro, and A. González. Threadmanagement techniques to maximize efficiency in multicore and simultaneous multithreaded microprocessors. ACM Transactions on Architecture and Code Optimization, 7(2):9:1–9:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Reddi:2010:EVE

[RCG^{+}10b] Vijay Janapa Reddi, Simone Campanoni, Meeta S. Gupta, Michael D. Smith, Gu-Yeon Wei, David Brooks, and Kim Hazelwood. Eliminating voltage emergencies via softwareguided code transformations. ACM Transactions on Architecture and Code Optimization, 7(2):12:1–12:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Reis:2005:SCF

[RCV^{+}05] George A. Reis, Jonathan Chang, Neil Vachharajani, Ram Rangan, David I. August, and Shubhendu S. Mukherjee. Software-controlled fault tolerance. ACM Transactions on Architecture and Code Optimization, 2(4):366–396, December 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rico:2012:SLS

[RCV^{+}12] Alejandro Rico, Felipe Cabarcas, Carlos Villavieja, Milan Pavlovic, Augusto Vega, Yoav Etsion, Alex Ramirez, and Mateo Valero. On the simulation of large-scale architectures using multiple application abstraction levels. ACM Transactions on Architecture and Code Optimization, 8(4): 36:1–36:??, January 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rolan:2013:VSC

[RFD13] Dyer Rolán, Basilio B. Fraguela, and Ramón Doallo. Virtually split cache: an efficient mechanism to distribute instructions and data 1. ACM Transactions on Architecture and Code Optimization, 10(4):27:1–27:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Radojkovic:2012:EIS

[RGG^{+}12] Petar Radojković, Sylvain Girbal, Arnaud Grasset, Eduardo Quiñones, Sami Yehia, and Francisco J. Cazorla. On the evaluation of the impact of shared resources in multithreaded COTS processors in time-critical environments. ACM Transactions on Architecture and Code Optimization, 8(4):34:1–34:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Reber:2023:CPS

[RGK^{+}23] Benjamin Reber, Matthew Gould, Alexander H. Kneipp, Fangzhou Liu, Ian Prechtl, Chen Ding, Linlin Chen, and Dorin Patru. Cache programming for scientific loops using leases. ACM Transactions on Architecture and Code Optimization, 20(3):39:1–39:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3600090.

Railing:2015:CEG

[RHC15] Brian P. Railing, Eric R. Hein, and Thomas M. Conte. Contech: Efficiently generating dynamic task graphs for arbitrary parallel programs. ACM Transactions on Architecture and Code Optimization, 12(2): 25:1–25:??, July 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ramachandran:2014:HFR

[RHLA14] Pradeep Ramachandran, Siva Kumar Sastry Hari, Manlap Li, and Sarita V. Adve. Hardware fault recovery for I/O intensive applications. ACM Transactions on Architecture and Code Optimization, 11(3): 33:1–33:??, October 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Raman:2024:NGD

[RJK24] Siddhartha Raman Sundara Raman, Lizy John, and Jaydeep P. Kulkarni. NEM-GNN: DAC/ADC-less, scalable, reconfigurable, graph and sparsity-aware near-memory accelerator for graph neural networks. ACM Transactions on Architecture and Code Optimization, 21(2):39:1–39:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3652607.

Rashidi:2018:IMP

[RJSA18] Saeed Rashidi, Majid Jalili, and Hamid Sarbazi-Azad. Improving MLC PCM performance through relaxed write and read for intermediate resistance levels. ACM Transactions on Architecture and Code Optimization, 15(1): 12:1–12:??, April 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Resch:2020:PBN

[RKC^{+}20] Salonik Resch, S. Karen Khatamifard, Zamshed Iqbal Chowdhury, Masoud Zabihi, Zhengyang Zhao, Jian-Ping Wang, Sachin S. Sapatnekar, and Ulya R. Karpuzcu. PIM-BALL: Binary neural networks in spintronic memory. ACM Transactions on Architecture and Code Optimization, 16(4): 41:1–41:26, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ravi:2023:TMA

[RKL23] Gokul Subramanian Ravi, Tushar Krishna, and Mikko Lipasti. TNT: a modular approach to traversing physically heterogeneous NOCs at bare-wire latency. ACM Transactions on Architecture and Code Optimization, 20(3): 35:1–35:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3597611.

Rubin:2015:MOM

[RLBBN15] Eri Rubin, Ely Levy, Amnon Barak, and Tal Ben-Nun. MAPS: Optimizing massively parallel applications using device-level memory abstraction. ACM Transactions on Architecture and Code Optimization, 11(4):44:1–44:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ruan:2013:BTB

[RLS13] Wenjia Ruan, Yujie Liu, and Michael Spear. Boosting timestamp-based transactional memory by exploiting hardware cycle counters. ACM Transactions on Architecture and Code Optimization, 10(4): 40:1–40:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ruan:2015:TRM

[RLS15] Wenjia Ruan, Yujie Liu, and Michael Spear. Transactional read-modify-write without aborts. ACM Transactions on Architecture and Code Optimization, 11(4):63:1–63:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ren:2014:POE

[RMA14] Bin Ren, Todd Mytkowicz, and Gagan Agrawal. A portable optimization engine for accelerating irregular datatraversal applications on SIMD architectures. ACM Transactions on Architecture and Code Optimization, 11(2): 16:1–16:??, June 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ryckbosch:2012:VSM

[RPE12] Frederick Ryckbosch, Stijn Polfliet, and Lieven Eeckhout. VSim: Simulating multi-server setups at near native hardware speed. ACM Transactions on Architecture and Code Optimization, 8(4):52:1–52:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rochecouste:2006:CCE

[RPS06] Olivier Rochecouste, Gilles Pokam, and André Seznec. A case for a complexity-effective, width-partitioned microarchitecture. ACM Transactions on Architecture and Code Optimization, 3(3):295–326, September 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

RodriguesCarvalho:2021:UCC

[RS21] Daniel Rodrigues Carvalho and André Seznec. Understanding cache compression. ACM Transactions on Architecture and Code Optimization, 18(3):36:1–36:27, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3457207.

Roy:2018:NCN

[RSK^{+}18] Probir Roy, Shuaiwen Leon Song, Sriram Krishnamoorthy, Abhinav Vishnu, Dipanjan Sengupta, and Xu Liu. NUMA-Caffe: NUMA-aware deep learning neural networks. ACM Transactions on Architecture and Code Optimization, 15(2):24:1–24:??, June 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rangan:2020:ZEZ

[RSU^{+}20] Ram Rangan, Mark W.

Stephenson, Aditya Ukarande, Shyam Murthy, Virat Agarwal, and Marc Blackstein. Zeroploit: Exploiting zero valued operands in interactive gaming applications. ACM Transactions on Architecture and Code Optimization, 17(3): 17:1–17:26, August 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3394284.

Rong:2007:SDS

[RTG^{+}07] Hongbo Rong, Zhizhong Tang, R. Govindarajan, Alban Douillet, and Guang R. Gao. Singledimension software pipelining for multidimensional loops. ACM Transactions on Architecture and Code Optimization, 4(1):7:1–7:44, March 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rodriguez:2015:VSR

[RTK15] Gabriel Rodríguez, Juan Touriño, and Mahmut T. Kandemir. Volatile STT–RAM scratchpad design and data allocation for low energy. ACM Transactions on Architecture and Code Optimization, 11(4): 38:1–38:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Riebler:2019:TAH

[RVKP19] Heinrich Riebler, Gavin Vaz, Tobias Kenter, and Christian Plessl. Transparent acceleration for heterogeneous platforms with compilation to OpenCL. ACM Transactions on Architecture and Code Optimization, 16(2):14:1–14:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rangan:2008:PSD

[RVOA08] Ram Rangan, Neil Vachharajani, Guilherme Ottoni, and David I. August. Performance scalability of decoupled software pipelining. ACM Transactions on Architecture and Code Optimization, 5(2):8:1– 8:??, August 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Roberts:2019:POS

[RWFJ19] Stephen I. Roberts, Steven A.

Wright, Suhaib A. Fahmy, and Stephen A. Jarvis. The poweroptimised software envelope. ACM Transactions on Architecture and Code Optimization, 16(3):21:1–21:??, July 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rohou:2013:VTI

[RWY13] Erven Rohou, Kevin Williams, and David Yuste. Vectorization technology to improve interpreter performance. ACM Transactions on Architecture and Code Optimization, 9(4): 26:1–26:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Silfa:2022:BEE

[SAG22] Franyell Silfa, Jose Maria Arnau, and Antonio González. E-BATCH: Energy-efficient and high-throughput RNN batching. ACM Transactions on Architecture and Code Optimization, 19(1):14:1–14:23, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3499757.

Song:2019:SAR

[SAL19] Yang Song, Olivier Alavoine, and Bill Lin. A self-aware resource management framework for heterogeneous multicore SoCs with diverse QoS targets. ACM Transactions on Architecture and Code Optimization, 16(2):16:1–16:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Susskind:2023:UNA

[SAM^{+}23] Zachary Susskind, Aman Arora, Igor D. S. Miranda, Alan T. L. Bacellar, Luis A. Q. Villon, Rafael F. Katopodis, Leandro S. de Araújo, Diego L. C. Dutra, Priscila M. V. Lima, Felipe M. G. França, Mauricio Breternitz Jr., and Lizy K. John. ULEEN: a novel architecture for ultralow-energy edge neural networks. ACM Transactions on Architecture and Code Optimization, 20(4):61:1–61:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3629522.

Siso:2020:EAV

[SAT20] Sergi Siso, Wes Armour, and Jeyarajan Thiyagalingam. Evaluating auto-vectorizing compilers through objective withdrawal of useful information. ACM Transactions on Architecture and Code Optimization, 16(4):40:1–40:23, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Strozek:2009:EAE

[SB09] Lukasz Strozek and David Brooks. Energy- and areaefficient architectures through application clustering and architectural heterogeneity. ACM Transactions on Architecture and Code Optimization, 6(1):4:1–4:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sharma:2005:SPE

[SBC05] Saurabh Sharma, Jesse G.

Beu, and Thomas M. Conte. Spectral prefetcher: An effective mechanism for L2 cache prefetching. ACM Transactions on Architecture and Code Optimization, 2(4):423– 450, December 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Saileshwar:2022:HLC

[SBC^{+}22] Gururaj Saileshwar, Rick Boivie, Tong Chen, Benjamin Segal, and Alper Buyuktosunoglu. HeapCheck: Lowcost hardware support for memory safety. ACM Transactions on Architecture and Code Optimization, 19(1): 10:1–10:24, March 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3495152.

Scolari:2016:SCP

[SBS16] Alberto Scolari, Davide Basilio Bartolini, and Marco Domenico Santambrogio. A software cache partitioning system for hash-based caches. ACM Transactions on Architecture and Code Optimization, 13(4): 57:1–57:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shi:2008:VMS

[SCEG08] Yunhe Shi, Kevin Casey, M. Anton Ertl, and David Gregg. Virtual machine showdown: Stack versus registers. ACM Transactions on Architecture and Code Optimization, 4(4):2:1–2:??, January 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Srikanthan:2022:MMA

[SCFD22] Sharanyan Srikanthan, Sayak Chakraborti, Princeton Ferro, and Sandhya Dwarkadas. MAPPER: Managing application performance via parallel efficiency regulation *. ACM Transactions on Architecture and Code Optimization, 19(2): 18:1–18:26, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3501767.

Seo:2021:SAI

[SCK^{+}21] Wonik Seo, Sanghoon Cha, Yeonjae Kim, Jaehyuk Huh, and Jongse Park. SLO-aware inference scheduler for heterogeneous processors in edge platforms. ACM Transactions on Architecture and Code Optimization, 18(4):43:1–43:26, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460352.

Sasongko:2022:RFY

[SCMU22] Muhammad Aditya Sasongko, Milind Chabbi, Mandana Bagheri Marzijarani, and Didem Unat. ReuseTracker: Fast yet accurate multicore reuse distance analyzer. ACM Transactions on Architecture and Code Optimization, 19(1):3:1–3:25, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3484199.

Stenstrom:2012:ISI

[SD12] Per Stenström and Koen De Bosschere. Introduction to the special issue on highperformance and embedded architectures and compilers. ACM Transactions on Architecture and Code Optimization, 8(4):18:1–18:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Streit:2015:GTP

[SDH^{+}15] Kevin Streit, Johannes Doerfert, Clemens Hammacher, Andreas Zeller, and Sebastian Hack. Generalized task parallelism. ACM Transactions on Architecture and Code Optimization, 12(1):8:1–8:??, April 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

So:2022:EES

[SDK^{+}22] Hwisoo So, Moslem Didehban, Yohan Ko, Aviral Shrivastava, and Kyoungwoo Lee. EXPER-TISE: an effective softwarelevel redundant multithreading scheme against hardware faults. ACM Transactions on Architecture and Code Optimization, 19(4):53:1–53:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3546073.

Song:2021:CRE

[SDS^{+}21] Weijia Song, Christina Delimitrou, Zhiming Shen, Robbert Van Renesse, Hakim Weatherspoon, Lotfi Benmohamed, Frederic De Vaulx, and Charif Mahmoudi. CacheInspector: Reverse engineering cache resources in public clouds. ACM Transactions on Architecture and Code Optimization, 18(3): 35:1–35:25, June 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3457373.

Singh:2021:SSM

[SDZ^{+}21] Abhishek Singh, Shail Dave, Pantea Zardoshti, Robert Brotzman, Chao Zhang, Xiaochen Guo, Aviral Shrivastava, Gang Tan, and Michael Spear. SPX64: a scratchpad memory for general-purpose microprocessors. ACM Transactions on Architecture and Code Optimization, 18(1): 14:1–14:26, January 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3436730.

Sadrosadati:2019:IIT

[SEF^{+}19] Mohammad Sadrosadati, Seyed Borna Ehsani, Hajar Falahati, Rachata Ausavarungnirun, Arash Tavakkol, Mojtaba Abaee, Lois Orosa, Yaohua Wang, Hamid Sarbazi-Azad, and Onur Mutlu. ITAP: Idle-time-aware power management for GPU execution units. ACM Transactions on Architecture and Code Optimization, 16(1):3:1–3:??, March 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Soteriou:2007:SDP

[SEP07] Vassos Soteriou, Noel Eisley, and Li-Shiuan Peh. Softwaredirected power-aware interconnection networks. ACM Transactions on Architecture and Code Optimization, 4(1):5:1– 5:40, March 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Stawinoga:2018:PTC

[SF18] Nicolai Stawinoga and Tony Field. Predictable thread coarsening. ACM Transactions on Architecture and Code Optimization, 15(2):23:1–23:??, June 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shobaki:2022:RPA

[SGM^{+}22] Ghassan Shobaki, Vahl Scott Gordon, Paul McHugh, Theodore Dubois, and Austin Kerbow. Register-pressure-aware instruction scheduling using ant colony optimization. ACM Transactions on Architecture and Code Optimization, 19(2): 23:1–23:23, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3505558.

Selva:2020:BPR

[SGS^{+}20] Manuel Selva, Fabian Gruber, Diogo Sampaio, Christophe Guillon, Louis-Noël Pouchet, and Fabrice Rastello. Building a polyhedral representation from an instrumented execution: Making dynamic analyses of nonaffine programs scalable. ACM Transactions on Architecture and Code Optimization, 16(4):45:1–45:26, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3363785.

She:2013:EEM

[SHC13] Dongrui She, Yifan He, and Henk Corporaal. An energyefficient method of supporting flexible special instructions in an embedded processor with compact ISA. ACM Transactions on Architecture and Code Optimization, 10(3): 15:1–15:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Suh:2015:DMR

[SHD15] Jinho Suh, Chieh-Ting Huang, and Michel Dubois. Dynamic MIPS rate stabilization for complex processors. ACM Transactions on Architecture and Code Optimization, 12(1): 4:1–4:??, April 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Samadi:2014:LGU

[SHLM14] Mehrzad Samadi, Amir Hormati, Janghaeng Lee, and Scott Mahlke. Leveraging GPUs using cooperative loop speculation. ACM Transactions on Architecture and Code Optimization, 11(1):3:1–3:26, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Stoltzfus:2020:TOS

[SHS^{+}20] Larisa Stoltzfus, Bastian Hagedorn, Michel Steuwer, Sergei Gorlatch, and Christophe Dubach. Tiling optimizations for stencil computations using rewrite rules in lift. ACM Transactions on Architecture and Code Optimization, 16(4): 52:1–52:25, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3368858.

Shen:2014:RSB

[SHY14] Bor-Yeh Shen, Wei-Chung Hsu, and Wuu Yang. A retargetable static binary translator for the ARM architecture. ACM Transactions on Architecture and Code Optimization, 11(2):18:1–18:??, June 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sharafeddine:2012:DOE

[SJA12] Mageda Sharafeddine, Komal Jothi, and Haitham Akkary. Disjoint out-of-order execution processor. ACM Transactions on Architecture and Code Optimization, 9(3):19:1–19:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Srikanth:2021:SIC

[SJC^{+}21] Sriseshan Srikanth, Anirudh Jain, Thomas M. Conte, Erik P. Debenedictis, and Jeanine Cook. SortCache: Intelligent cache management for accelerating sparse data workloads. ACM Transactions on Architecture and Code Optimization, 18(4):56:1–56:24, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473332.

Schlaak:2022:MAF

[SJD22] Christof Schlaak, Tzung-Han Juang, and Christophe Dubach. Memory-aware functional IR for higher-level synthesis of accelerators. ACM Transactions on Architecture and Code Optimization, 19(2): 16:1–16:26, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3501768.

Srikanth:2020:MAS

[SJL^{+}20] Sriseshan Srikanth, Anirudh Jain, Joseph M. Lennon, Thomas M. Conte, Erik Debenedictis, and Jeanine Cook. MetaStrider: Architectures for scalable memorycentric reduction of sparse data streams. ACM Transactions on Architecture and Code Optimization, 16(4): 35:1–35:26, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shahbahrami:2008:VES

[SJV08] Asadollah Shahbahrami, Ben Juurlink, and Stamatis Vassiliadis. Versatility of extended subwords and the matrix register file. ACM Transactions on Architecture and Code Optimization, 5(1):5:1–5:??, May 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Serres:2016:EPP

[SKAEG16] Olivier Serres, Abdullah Kayi, Ahmad Anbar, and Tarek El-Ghazawi. Enabling PGAS productivity with hardware support for shared address mapping: a UPC case study. ACM Transactions on Architecture and Code Optimization, 12(4): 52:1–52:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shi:2016:LLA

[SKH^{+}16] Qingchuan Shi, George Kurian, Farrukh Hijaz, Srinivas Devadas, and Omer Khan. LDAC: Locality-aware data access control for large-scale multicore cache hierarchies. ACM Transactions on Architecture and Code Optimization, 13(4):37:1–37:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sfakianakis:2018:QPB

[SKKB18] Yannis Sfakianakis, Christos Kozanitis, Christos Kozyrakis, and Angelos Bilas. QuMan: Profile-based improvement of cluster utilization. ACM Transactions on Architecture and Code Optimization, 15(3): 27:1–27:??, October 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Siddhu:2022:CII

[SKP^{+}22] Lokesh Siddhu, Rajesh Kedia, Shailja Pandey, Martin Rapp, Anuj Pathania, Jörg Henkel, and Preeti Ranjan Panda. CoMeT: an integrated interval thermal simulation toolchain for 2D, 2.5D, and 3D processor-memory systems. ACM Transactions on Architecture and Code Optimization, 19(3):44:1–44:25, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532185.

Shobaki:2019:EAC

[SKPD19] Ghassan Shobaki, Austin Kerbow, Christopher Pulido, and William Dobson. Exploring an alternative cost function for combinatorial registerpressure-aware instruction scheduling. ACM Transactions on Architecture and Code Optimization, 16(1):1:1–1:??, March 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sakalis:2023:DSS

[SKS23] Christos Sakalis, Stefanos Kaxiras, and Magnus Själander. Delay-on-squash: Stopping microarchitectural replay attacks in their tracks. ACM Transactions on Architecture and Code Optimization, 20(1): 9:1–9:??, March 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3563695.

Subramaniam:2009:DOS

[SL09] Samantika Subramaniam and Gabriel H. Loh. Design and optimization of the store vectors memory dependence predictor. ACM Transactions on Architecture and Code Optimization, 6(4):16:1–16:??, October 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Song:2020:IME

[SL20] Yang Song and Bill Lin. Improving memory efficiency in heterogeneous MPSoCs through row-buffer localityaware forwarding. ACM Transactions on Architecture and Code Optimization, 17(1): 6:1–6:26, March 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3377149.

Sasanka:2007:AES

[SLA^{+}07] Ruchira Sasanka, Man-Lap Li, Sarita V. Adve, Yen-Kuang Chen, and Eric Debes. ALP: Efficient support for all levels of parallelism for complex media applications. ACM Transactions on Architecture and Code Optimization, 4(1):3:1– 3:30, March 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shi:2020:OSB

[SLH^{+}20] Xuanhua Shi, Wei Liu, Ligang He, Hai Jin, Ming Li, and Yong Chen. Optimizing the SSD burst buffer by traffic detection. ACM Transactions on Architecture and Code Optimization, 17(1):8:1–8:26, March 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3377705.

Sangaiah:2018:SSA

[SLJ^{+}18] Karthik Sangaiah, Michael Lui, Radhika Jagtap, Stephan Diestelhorst, Siddharth Nilakantan, Ankit More, Baris Taskin, and Mark Hempstead. SynchroTrace: Synchronizationaware architecture-agnostic traces for lightweight multicore simulation of CMP and HPC workloads. ACM Transactions on Architecture and Code Optimization, 15(1):2:1– 2:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Su:2019:SSC

[SLJ^{+}19] Xing Su, Xiangke Liao, Hao Jiang, Canqun Yang, and Jingling Xue. SCP: Shared cache partitioning for highperformance GEMM. ACM Transactions on Architecture and Code Optimization, 15(4): 43:1–43:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Seghir:2012:IAT

[SLM12] Rachid Seghir, Vincent Loechner, and Benoît Meister. Integer affine transformations of parametric Z-polytopes and applications to loop nest optimization. ACM Transactions on Architecture and Code Optimization, 9(2):8:1–8:??, June 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sharkey:2008:RRP

[SLP08] Joseph J. Sharkey, Jason Loew, and Dmitry V. Ponomarev. Reducing register pressure in SMT processors through L2-miss-driven early register release. ACM Transactions on Architecture and Code Optimization, 5(3):13:1–13:??, November 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sha:2021:LIA

[SLS^{+}21] Zhibing Sha, Jun Li, Lihao Song, Jiewen Tang, Min Huang, Zhigang Cai, Lianju Qian, Jianwei Liao, and Zhiming Liu. Low I/O intensityaware partial GC scheduling to reduce long-tail latency in SSDs. ACM Transactions on Architecture and Code Optimization, 18(4):46:1–46:25, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3460433.

Sargaran:2019:SSA

[SM19] Sahar Sargaran and Naser Mohammadzadeh. SAQIP: a scalable architecture for quantum information processors. ACM Transactions on Architecture and Code Optimization, 16(2): 12:1–12:??, May 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sanchez:2010:ACI

[SMK10] Daniel Sanchez, George Michelogiannakis, and Christos Kozyrakis. An analysis of onchip interconnection networks for large-scale chip multiprocessors. ACM Transactions on Architecture and Code Optimization, 7(1):4:1–4:??, April 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Schaub:2015:ISW

[SMKH15] Thomas Schaub, Simon Moll, Ralf Karrenberg, and Sebastian Hack. The impact of the SIMD width on control-flow and memory divergence. ACM Transactions on Architecture and Code Optimization, 11(4): 54:1–54:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shah:2023:BSA

[SMM^{+}23] Nilesh Rajendra Shah, Ashitabh Misra, Antoine Miné, Rakesh Venkat, and Ramakrishna Upadrasta. BullsEye: Scalable and accurate approximation framework for cache miss calculation. ACM Transactions on Architecture and Code Optimization, 20(1):2:1–2:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3558003.

Soltaniyeh:2022:ASC

[SMN22] Mohammadreza Soltaniyeh, Richard P. Martin, and Santosh Nagarakatte. An accelerator for sparse convolutional neural networks leveraging systolic general matrix– matrix multiplication. ACM Transactions on Architecture and Code Optimization, 19(3): 42:1–42:26, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3532863.

Schuler:2023:XOT

[SMS23] Manuela Schuler, Richard Membarth, and Philipp Slusallek. XEngine: Optimal tensor rematerialization for neural networks in heterogeneous environments. ACM Transactions on Architecture and Code Optimization, 20(1):17:1–17:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3568956.

Shrivastava:2017:EEC

[SN17] Rahul Shrivastava and V. Krishna Nandivada. Energyefficient compilation of irregular task-parallel loops. ACM Transactions on Architecture and Code Optimization, 14(4): 35:1–35:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Suyeon:2023:FFF

[SNK^{+}23] Hur Suyeon, Seongmin Na, Dongup Kwon, Kim Joonsung, Andrew Boutros, Eriko Nurvitadhi, and Jangwoo Kim. A fast and flexible FPGA-based accelerator for natural language processing neural networks. ACM Transactions on Architecture and Code Optimization, 20(1):11:1–11:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564606.

Sankaralingam:2004:TPA

[SNL^{+}04] Karthikeyan Sankaralingam, Ramadass Nagarajan, Haiming Liu, Changkyu Kim, Jaehyuk Huh, Nitya Ranganathan, Doug Burger, Stephen W. Keckler, Robert G. McDonald, and Charles R. Moore. TRIPS: a polymorphous architecture for exploiting ILP, TLP, and DLP. ACM Transactions on Architecture and Code Optimization, 1(1):62–93, March 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shekofteh:2019:MSG

[SNN^{+}19] S.-Kazem Shekofteh, Hamid Noori, Mahmoud Naghibzadeh, Hadi Sadoghi Yazdi, and Holger Fröning. Metric selection for GPU kernel classification. ACM Transactions on Architecture and Code Optimization, 15(4):68:1–68:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sahni:2023:AAS

[SOAK23] Abdul Rasheed Sahni, Hamza Omar, Usman Ali, and Omer Khan. ASM: an adaptive secure multicore for colocated mutually distrusting processes. ACM Transactions on Architecture and Code Optimization, 20(3):32:1–32:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587480.

Sharkey:2006:IPT

[SPGE06] Joseph J. Sharkey, Dmitry V. Ponomarev, Kanad Ghose, and Oguz Ergin. Instruction packing: Toward fast and energy-efficient instruction scheduling. ACM Transactions on Architecture and Code Optimization, 3(2):156– 181, June 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Stanic:2017:IVS

[SPH^{+}17] Milan Stanic, Oscar Palomar, Timothy Hayes, Ivan Ratkovic, Adrian Cristal, Osman Unsal, and Mateo Valero.

An integrated vector-scalar design on an in-order ARM core. ACM Transactions on Architecture and Code Optimization, 14(2):17:1–17:??, July 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Swami:2017:EEC

[SPM17] Shivam Swami, Poovaiah M.

Palangappa, and Kartik Mohanram. ECS: Error-correcting strings for lifetime improvements in nonvolatile memories. ACM Transactions on Architecture and Code Optimization, 14(4):40:1–40:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Stock:2012:UML

[SPS12] Kevin Stock, Louis-Noël Pouchet, and P. Sadayappan. Using machine learning to improve automatic vectorization. ACM Transactions on Architecture and Code Optimization, 8(4): 50:1–50:??, January 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sridharan:2017:BPP

[SPS17] Aswinkumar Sridharan, Biswabandan Panda, and Andre Seznec. Band-pass prefetching: an effective prefetch management mechanism using prefetch-fraction metric in multi-core systems. ACM Transactions on Architecture and Code Optimization, 14(2):

19:1–19:??, July 2017. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Soni:2023:AC

[SPS23] Mitali Soni, Asmita Pal, and Joshua San Miguel. As-is approximate computing. ACM Transactions on Architecture and Code Optimization, 20(1): 3:1–3:??, March 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3559761.

Sabet:2020:RAU

[SQZK20] Amir Hossein Nodehi Sabet, Junqiao Qiu, Zhijia Zhao, and Sriram Krishnamoorthy. Reliability analysis for unreliable FSM computations. ACM Transactions on Architecture and Code Optimization, 17(2): 12:1–12:23, June 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3377456.

Sukumaran-Rajam:2016:PMN

[SRC16] Aravind Sukumaran-Rajam and Philippe Clauss. The polyhedral model of nonlinear loops. ACM Transactions on Architecture and Code Optimization, 12(4):48:1–48:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Santana:2004:LCF

[SRLPV04] Oliverio J. Santana, Alex Ramirez, Josep L. Larriba-Pey, and Mateo Valero. A low-complexity fetch architecture for high-performance superscalar processors. ACM Transactions on Architecture and Code Optimization, 1(2): 220–245, June 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sankaranarayanan:2004:PBA

[SS04] Karthik Sankaranarayanan and Kevin Skadron. Profilebased adaptation for cache decay. ACM Transactions on Architecture and Code Optimization, 1(3):305–322, September 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sioutas:2020:SSH

[SSB^{+}20] Savvas Sioutas, Sander Stuijk, Twan Basten, Henk Corporaal, and Lou Somers. Schedule synthesis for Halide pipelines on GPUs. ACM Transactions on Architecture and Code Optimization, 17(3): 23:1–23:25, August 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3406117.

Sanchez:2013:MIP

[SSC^{+}13] Daniel Sánchez, Yiannakis Sazeides, Juan M. Cebrián, José M. García, and Juan L. Aragón. Modeling the impact of permanent faults in caches. ACM Transactions on Architecture and Code Optimization, 10(4):29:1–29:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Subramaniam:2013:UFC

[SSH^{+}13] Samantika Subramaniam, Simon C. Steely, Will Hasenplaugh, Aamer Jaleel, Carl Beckmann, Tryggve Fossum, and Joel Emer. Using inflight chains to build a scalable cache coherence protocol. ACM Transactions on Architecture and Code Optimization, 10(4):28:1–28:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sadeghi:2019:TCN

[SSH19] Mohammad Sadegh Sadeghi, Siavash Bayat Sarmadi, and Shaahin Hessabi. Toward on-chip network security using runtime isolation mapping. ACM Transactions on Architecture and Code Optimization, 16(3):28:1–28:??, July 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Samih:2011:EPP

[SSK11] Ahmad Samih, Yan Solihin, and Anil Krishna. Evaluating placement policies for managing capacity sharing in CMP architectures with private caches. ACM Transactions on Architecture and Code Optimization, 8(3):15:1–15:??, October 2011. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic).

Singh:2023:HHP

[SSP^{+}23] Sarabjeet Singh, Neelam Surana, Kailash Prasad, Pranjali Jain, Joycee Mekie, and Manu Awasthi. HyGain: Highperformance, energy-efficient hybrid gain cell-based cache hierarchy. ACM Transactions on Architecture and Code Optimization, 20(2):24:1–24:??, June 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572839.

Strydis:2013:SAP

[SSPL^{+}13] Christos Strydis, Robert M.

Seepers, Pedro Peris-Lopez, Dimitrios Siskos, and Ioannis Sourdis. A system architecture, processor, and communication protocol for secure implants. ACM Transactions on Architecture and Code Optimization, 10(4):57:1–57:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shobaki:2013:PIS

[SSR13] Ghassan Shobaki, Maxim Shawabkeh, and Najm Eldeen Abu Rmaileh. Preallocation instruction scheduling with register pressure minimization using a combinatorial optimization approach. ACM Transactions on Architecture and Code Optimization, 10(3):

14:1–14:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Suresh:2015:IFM

[SSRS15] Arjun Suresh, Bharath Narasimha Swamy, Erven Rohou, and André Seznec. Intercepting functions for memoization: a case study using transcendental functions. ACM Transactions on Architecture and Code Optimization, 12(2): 18:1–18:??, July 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Skadron:2004:TAM

[SSS^{+}04] Kevin Skadron, Mircea R. Stan, Karthik Sankaranarayanan, Wei Huang, Sivakumar Velusamy, and David Tarjan. Temperatureaware microarchitecture: Modeling and implementation. ACM Transactions on Architecture and Code Optimization, 1(1):94–125, March 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shah:2023:TDS

[SSS^{+}23] Parth Shah, Ranjal Gautham Shenoy, Vaidyanathan Srinivasan, Pradip Bose, and Alper Buyuktosunoglu. TokenSmart: Distributed, scalable power management in the many-core era. ACM Transactions on Architecture and Code Optimization, 20(1):4:1–4:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3559762.

Singhal:2024:OHP

[SSS^{+}24] Vidush Singhal, Laith Sakka, Kirshanthan Sundararajah, Ryan Newton, and Milind Kulkarni. Orchard: Heterogeneous parallelism and finegrained fusion for complex tree traversals. ACM Transactions on Architecture and Code Optimization, 21(2):41:1–41:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3652605.

Stipic:2013:PGT

[SSU^{+}13] Srdan Stipić, Vesna Smiljković, Osman Unsal, Adrián Cristal, and Mateo Valero. Profileguided transaction coalescinglowering transactional overheads by merging transactions. ACM Transactions on Architecture and Code Optimization, 10(4):50:1–50:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sardashti:2016:YAC

[SSW16] Somayeh Sardashti, Andre Seznec, and David A. Wood. Yet another compressed cache: a low-cost yet effective compressed cache. ACM Transactions on Architecture and Code Optimization, 13(3): 27:1–27:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sioutas:2019:SSH

[SSW^{+}19] Savvas Sioutas, Sander Stuijk, Luc Waeijen, Twan Basten, Henk Corporaal, and Lou Somers. Schedule synthesis for Halide pipelines through reuse analysis. ACM Transactions on Architecture and Code Optimization, 16(2):10:1–10:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Saidi:2012:OED

[STLM12] Selma Saidi, Pranav Tendulkar, Thierry Lepley, and Oded Maler. Optimizing explicit data transfers for data parallel applications on the Cell architecture. ACM Transactions on Architecture and Code Optimization, 8(4):37:1– 37:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Salami:2005:DMI

[SV05] Esther Salamí and Mateo Valero. Dynamic memory interval test vs. interprocedural pointer analysis in multimedia applications. ACM Transactions on Architecture and Code Optimization, 2(2):199– 219, June 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shifer:2013:LLA

[SW13] Eran Shifer and Shlomo Weiss. Low-latency adaptive mode transitions and hierarchical power management in asymmetric clustered cores. ACM Transactions on Architecture and Code Optimization, 10(3): 10:1–10:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sardashti:2017:CCG

[SW17a] Somayeh Sardashti and David A.

Wood. Could compression be of general use? Evaluating memory compression across domains. ACM Transactions on Architecture and Code Optimization, 14(4):44:1–44:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sen:2017:PGE

[SW17b] Rathijit Sen and David A.

Wood. Pareto governors for energy-optimal computing. ACM Transactions on Architecture and Code Optimization, 14(1):6:1–6:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Spink:2016:HAC

[SWF16] Tom Spink, Harry Wagstaff, and Björn Franke. Hardwareaccelerated cross-architecture full-system virtualization. ACM Transactions on Architecture and Code Optimization, 13(4): 36:1–36:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shobaki:2009:OTS

[SWH09] Ghassan Shobaki, Kent Wilken, and Mark Heffernan. Optimal trace scheduling using enumeration. ACM Transactions on Architecture and Code Optimization, 5(4):19:1–19:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Stokes:2021:DMR

[SWO21] Michael Stokes, David Whalley, and Soner Onder. Decreasing the miss rate and eliminating the performance penalty of a data filter cache. ACM Transactions on Architecture and Code Optimization, 18(3): 28:1–28:22, June 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3449043.

Simon:2015:STH

[SWU^{+}15] Doug Simon, Christian Wimmer, Bernhard Urban, Gilles Duboscq, Lukas Stadler, and Thomas Würthinger. Snippets: Taking the high road to a low level. ACM Transactions on Architecture and Code Optimization, 12(2):20:1–20:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sato:2019:AFS

[SYE19] Yukinori Sato, Tomoya Yuki, and Toshio Endo. An autotuning framework for scalable execution of tiled code via iterative polyhedral compilation. ACM Transactions on Architecture and Code Optimization, 15(4):67:1–67:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Seshadri:2015:MPC

[SYX^{+}15] Vivek Seshadri, Samihan Yedkar, Hongyi Xin, Onur Mutlu, Phillip B. Gibbons, Michael A. Kozuch, and Todd C. Mowry. Mitigating prefetcher-caused pollution using informed caching policies for prefetched blocks. ACM Transactions on Architecture and Code Optimization, 11(4): 51:1–51:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Savrun-Yeniceri:2014:EHI

[SYZZ^{+}14] Gülfem Savrun-Yeniçeri, Wei Zhang, Huahan Zhang, Eric Seckler, Chen Li, Stefan Brunthaler, Per Larsen, and Michael Franz. Efficient hosted interpreters on the JVM. ACM Transactions on Architecture and Code Optimization, 11(1): 9:1–9:24, February 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/2532642.

Shihab:2018:RFD

[SZJK18] Mustafa M. Shihab, Jie Zhang, Myoungsoo Jung, and Mahmut Kandemir. ReveNAND: a fast-drift-aware resilient 3D NAND flash design. ACM Transactions on Architecture and Code Optimization, 15(2): 17:1–17:??, June 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tripathy:2021:PLG

[TAB^{+}21] Devashree Tripathy, Amirali Abdolrashidi, Laxmi Narayan Bhuyan, Liang Zhou, and Daniel Wong. PAVER: Locality graph-based thread block scheduling for GPUs. ACM Transactions on Architecture and Code Optimization, 18(3): 32:1–32:26, June 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3451164.

Therdsteerasukdi:2012:URI

[TBC^{+}12] Kanit Therdsteerasukdi, Gyungsu Byun, Jason Cong, M. Frank Chang, and Glenn Reinman. Utilizing RF-I and intelligent scheduling for better throughput/watt in a mobile GPU memory system. ACM Transactions on Architecture and Code Optimization, 8(4):51:1– 51:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tan:2006:BSS

[TBS06] Lin Tan, Brett Brotherton, and Timothy Sherwood. Bitsplit string-matching engines for intrusion detection and prevention. ACM Transactions on Architecture and Code Optimization, 3(1):3–34, March 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Terechko:2007:ICC

[TC07] A. S. Terechko and H. Corporaal. Inter-cluster communication in VLIW architectures. ACM Transactions on Architecture and Code Optimization, 4(2):11:1–11:??, June 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tirumalasetty:2022:RMP

[TCR^{+}22] Chandrahas Tirumalasetty, Chih Chieh Chou, Narasimha Reddy, Paul Gratz, and Ayman Abouelwafa. Reducing minor page fault overheads through enhanced page walker. ACM Transactions on Architecture and Code Optimization, 19(4):57:1–57:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547142.

Tan:2016:SEE

[TCS16] Li Tan, Zizhong Chen, and Shuaiwen Leon Song. Scalable energy efficiency with resilience for high performance computing systems: a quantitative methodology. ACM Transactions on Architecture and Code Optimization, 12(4): 35:1–35:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tino:2020:SXE

[TCS20] Anita Tino, Caroline Collange, and André Seznec. SIMT-X: Extending singleinstruction multi-threading to out-of-order cores. ACM Transactions on Architecture and Code Optimization, 17(2): 15:1–15:23, June 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3392032.

Theocharis:2016:BSC

[TD16] Panagiotis Theocharis and Bjorn De Sutter. A bimodal scheduler for coarse-grained reconfigurable arrays. ACM Transactions on Architecture and Code Optimization, 13(2): 15:1–15:??, June 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Totoni:2013:EFE

[TDG13] Ehsan Totoni, Mert Dikmen, and María Jesús Garzarán. Easy, fast, and energy-efficient object detection on heterogeneous on-chip architectures. ACM Transactions on Architecture and Code Optimization, 10(4):45:1–45:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tomusk:2016:FME

[TDO16a] Erik Tomusk, Christophe Dubach, and Michael O’Boyle. Four metrics to evaluate heterogeneous multicores. ACM Transactions on Architecture and Code Optimization, 12(4): 37:1–37:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tomusk:2016:SHC

[TDO16b] Erik Tomusk, Christophe Dubach, and Michael O’Boyle. Selecting heterogeneous cores for diversity. ACM Transactions on Architecture and Code Optimization, 13(4): 49:1–49:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Trinh:2015:EDE

[TDP15] Hong-Phuc Trinh, Marc Duranton, and Michel Paindavoine. Efficient data encoding for convolutional neural network application. ACM Transactions on Architecture and Code Optimization, 11(4): 49:1–49:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tomei:2021:BSC

[TDS^{+}21] Matthew Tomei, Shomit Das, Mohammad Seyedzadeh, Philip Bedoukian, Bradford Beckmann, Rakesh Kumar, and David Wood. Byte-select compression. ACM Transactions on Architecture and Code Optimization, 18(4):49:1–49:27, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3462209.

Tallam:2007:UCF

[TG07] Sriraman Tallam and Rajiv Gupta. Unified control flow and data dependence traces. ACM Transactions on Architecture and Code Optimization, 4(3):19:1–19:??, September 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Titos-Gil:2012:HTM

[TGAG^{+}12] Ruben Titos-Gil, Manuel E. Acacio, Jose M. Garcia, Tim Harris, Adrian Cristal, Osman Unsal, Ibrahim Hur, and Mateo Valero. Hardware transactional memory with softwaredefined conflicts. ACM Transactions on Architecture and Code Optimization, 8(4):31:1– 31:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tiwari:2021:PCP

[TGRK21] Sugandha Tiwari, Neel Gala, Chester Rebeiro, and V. Kamakoti. PERI: a configurable posit enabled RISC-V core. ACM Transactions on Architecture and Code Optimization, 18(3):25:1–25:26, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3446210.

Tavarageri:2021:PPO

[THA^{+}21] Sanket Tavarageri, Alexander Heinecke, Sasikanth Avancha, Bharat Kaul, Gagandeep Goyal, and Ramakrishna Upadrasta. PolyDL: Polyhedral optimizations for creation of high-performance DL primitives. ACM Transactions on Architecture and Code Optimization, 18(1):11:1–11:27, January 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3433103.

Tollenaere:2023:ACE

[TIP^{+}23] Nicolas Tollenaere, Guillaume Iooss, Stéphane Pouget, Hugo Brunie, Christophe Guillon, Albert Cohen, P. Sadayappan, and Fabrice Rastello. Autotuning convolutions is easier than you think. ACM Transactions on Architecture and Code Optimization, 20(2): 20:1–20:??, June 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3570641.

Tian:2013:TBM

[TKJ13] Yingying Tian, Samira M.

Khan, and Daniel A. Jiménez. Temporal-based multilevel correlating inclusive cache replacement. ACM Transactions on Architecture and Code Optimization, 10(4):33:1–33:??, December 2013. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic).

Tong:2015:OMT

[TKKM15] Xin Tong, Toshihiko Koju, Motohiro Kawahito, and Andreas Moshovos. Optimizing memory translation emulation in full system emulators. ACM Transactions on Architecture and Code Optimization, 11(4): 60:1–60:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tawa:2014:EEF

[TKM14] Venkata Kalyan Tawa, Ravi Kasha, and Madhu Mutyam. EFGR: an enhanced fine granularity refresh feature for highperformance DDR4 DRAM devices. ACM Transactions on Architecture and Code Optimization, 11(3):31:1–31:??, October 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tampouratzis:2016:AIH

[TMP16] Nikolaos Tampouratzis, Pavlos M. Mattheakis, and Ioannis Papaefstathiou. Accelerating intercommunication in highly parallel systems. ACM Transactions on Architecture and Code Optimization, 13(4): 40:1–40:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

TACO-294350005

[TMSR23] Syed Salauddin Mohammad Tariq, Lance Menard, Pengfei Su, and Probir Roy. Micro-Prof: Code-level attribution of unnecessary data transfer in microservice applications. ACM Transactions on Architecture and Code Optimization, 20(4):50:1–50:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3622787.

Thangamani:2020:ORC

[TN20] Arun Thangamani and V. Krishna Nandivada. Optimizing remote communication in X10. ACM Transactions on Architecture and Code Optimization, 16(4):34:1–34:26, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tayeb:2024:AAV

[TPB24] Hayfa Tayeb, Ludovic Paillat, and Bérenger Bramas. Autovesk: Automatic vectorized code generation from unstructured static kernels using graph transformations. ACM Transactions on Architecture and Code Optimization, 21(1): 4:1–4:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3631709.

Tampouratzis:2020:NHI

[TPN^{+}20] Nikolaos Tampouratzis, Ioannis Papaefstathiou, Antonios Nikitakis, Andreas Brokalakis, Stamatis Andrianakis, Apostolos Dollas, Marco Marcon, and Emanuele Plebani. A novel, highly integrated simulator for parallel and distributed systems. ACM Transactions on Architecture and Code Optimization, 17(1):2:1– 2:28, March 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3378934.

Tartara:2013:CLC

[TR13] Michele Tartara and Stefano Crespi Reghizzi. Continuous learning of compiler heuristics. ACM Transactions on Architecture and Code Optimization, 9(4):46:1–46:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tarjan:2005:MPG

[TS05] David Tarjan and Kevin Skadron. Merging path and gshare indexing in perceptron branch prediction. ACM Transactions on Architecture and Code Optimization, 2(3): 280–300, September 2005. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tabkhi:2015:JSH

[TS15] Hamed Tabkhi and Gunar Schirner. A joint SW/HW approach for reducing register file vulnerability. ACM Transactions on Architecture and Code Optimization, 12(2):

9:1–9:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tzilis:2019:EER

[TTS19] Stavros Tzilis, Pedro Trancoso, and Ioannis Sourdis. Energy-efficient runtime management of heterogeneous multicores using online projection. ACM Transactions on Architecture and Code Optimization, 15(4):63:1–63:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Thuerck:2021:FRA

[TWB21] Daniel Thuerck, Nicolas Weber, and Roberto Bifulco. Flynn’s reconciliation: Automating the register cache idiom for cross-accelerator programming. ACM Transactions on Architecture and Code Optimization, 18(3):37:1–37:26, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458357.

Tavana:2018:BCA

[TZK18] Mohammad Khavari Tavana, Amir Kavyan Ziabari, and David Kaeli. Block cooperation: Advancing lifetime of resistive memories by increasing utilization of error correcting codes. ACM Transactions on Architecture and Code Optimization, 15(3):36:1–36:??, October 2018. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic).

Uguen:2020:ASA

[UDLD20] Yohann Uguen, Florent De Dinechin, Victor Lezaud, and Steven Derrien. Applicationspecific arithmetic in highlevel synthesis tools. ACM Transactions on Architecture and Code Optimization, 17(1): 5:1–5:23, March 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3377403.

UlHuda:2015:UTM

[UJW15] Zia Ul Huda, Ali Jannesari, and Felix Wolf. Using template matching to infer parallel design patterns. ACM Transactions on Architecture and Code Optimization, 11(4): 64:1–64:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ukarande:2022:LAC

[UPR22] Aditya Ukarande, Suryakant Patidar, and Ram Rangan. Locality-aware CTA scheduling for gaming applications. ACM Transactions on Architecture and Code Optimization, 19(1):1:1–1:26, March 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3477497.

Usui:2016:DDA

[USCM16] Hiroyuki Usui, Lavanya Subramanian, Kevin Kai-Wei Chang, and Onur Mutlu. DASH: Deadline-aware highperformance memory scheduler for heterogeneous systems with hardware accelerators. ACM Transactions on Architecture and Code Optimization, 12(4):65:1–65:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

VanDenBraak:2016:RGR

[VC16] Gert-Jan Van Den Braak and Henk Corporaal. R-GPU: a reconfigurable GPU architecture. ACM Transactions on Architecture and Code Optimization, 13(1):12:1–12:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vocke:2017:EHI

[VCJ^{+}17] Sander Vocke, Henk Corporaal, Roel Jordans, Rosilde Corvino, and Rick Nas. Extending Halide to improve software development for imaging DSPs. ACM Transactions on Architecture and Code Optimization, 14(3):21:1–21:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Venkataramani:2009:MAM

[VDSP09] Guru Venkataramani, Ioannis Doudalis, Yan Solihin, and Milos Prvulovic. MemTracker:

An accelerator for memory debugging and monitoring. ACM Transactions on Architecture and Code Optimization, 6(2): 5:1–5:??, June 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

vanderVlag:2020:ECB

[vdVSAAS20] Michiel A. van der Vlag, Georgios Smaragdos, Zaid Al-Ars, and Christos Strydis. Exploring complex brain-simulation workloads on multi-GPU deployments. ACM Transactions on Architecture and Code Optimization, 16(4):53:1–53:25, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3371235.

VanCraeynest:2013:UFD

[VE13] Kenzo Van Craeynest and Lieven Eeckhout. Understanding fundamental design choices in single-ISA heterogeneous multicore architectures. ACM Transactions on Architecture and Code Optimization, 9(4):32:1–32:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Venstermans:2007:JOH

[VED07] Kris Venstermans, Lieven Eeckhout, and Koen De Bosschere. Java object header elimination for reduced memory consumption in 64-bit virtual machines. ACM Transactions on Architecture and Code Optimization, 4(3):17:1–17:??, September 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vermij:2017:AIN

[VFJ^{+}17] Erik Vermij, Leandro Fiorin, Rik Jongerius, Christoph Hagleitner, Jan Van Lunteren, and Koen Bertels. An architecture for integrated near-data processors. ACM Transactions on Architecture and Code Optimization, 14(3):30:1–30:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vaish:2016:OMT

[VFW16] Nilay Vaish, Michael C. Ferris, and David A. Wood. Optimization models for three onchip network problems. ACM Transactions on Architecture and Code Optimization, 13(3): 26:1–26:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vora:2016:SAE

[VGX16] Keval Vora, Rajiv Gupta, and Guoqing Xu. Synergistic analysis of evolving graphs. ACM Transactions on Architecture and Code Optimization, 13(4): 32:1–32:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Venkataramani:2011:DDS

[VHKP11] Guru Venkataramani, Christopher J. Hughes, Sanjeev Kumar, and Milos Prvulovic. DeFT: Design space exploration for on-the-fly detection of coherence misses. ACM Transactions on Architecture and Code Optimization, 8(2): 8:1–8:??, July 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Verdoolaege:2013:PPC

[VJC^{+}13] Sven Verdoolaege, Juan Carlos Juega, Albert Cohen, José Ignacio Gómez, Christian Tenllado, and Francky Catthoor. Polyhedral parallel code generation for CUDA. ACM Transactions on Architecture and Code Optimization, 9(4):54:1– 54:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Voss:2021:PRS

[VKM^{+}21] Nils Voss, Bastiaan Kwaadgras, Oskar Mencer, Wayne Luk, and Georgi Gaydadjiev. On predictable reconfigurable system design. ACM Transactions on Architecture and Code Optimization, 18(2): 17:1–17:28, March 2021. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3436995.

Vijaykumar:2022:MPO

[VOK^{+}22] Nandita Vijaykumar, Ataberk Olgun, Konstantinos Kanellopoulos, F. Nisa Bostanci, Hasan Hassan, Mehrshad Lotfi, Phillip B. Gibbons, and Onur Mutlu. MetaSys: a practical open-source metadata management system to implement and evaluate crosslayer optimizations. ACM Transactions on Architecture and Code Optimization, 19(2): 26:1–26:29, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3505250.

Vasilakis:2019:DFC

[VPTS19] Evangelos Vasilakis, Vassilis Papaefstathiou, Pedro Trancoso, and Ioannis Sourdis. Decoupled fused cache: Fusing a decoupled LLC with a DRAM cache. ACM Transactions on Architecture and Code Optimization, 15(4):65:1–65:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vandierendonck:2008:SRA

[VS08] Hans Vandierendonck and André Seznec. Speculative return address stack management revisited. ACM Transactions on Architecture and Code Optimization, 5(3):15:1–15:??, November 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vandierendonck:2011:MSR

[VS11] Hans Vandierendonck and André Seznec. Managing SMT resource usage through speculative instruction window weighting. ACM Transactions on Architecture and Code Optimization, 8(3):12:1–12:??, October 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vale:2016:PDT

[VSDL16] Tiago M. Vale, João A. Silva, Ricardo J. Dias, and João M. Lourenço. Pot: Deterministic transactional execution. ACM Transactions on Architecture and Code Optimization, 13(4):52:1–52:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Valero:2012:CRI

[VSP^{+}12] Alejandro Valero, Julio Sahuquillo, Salvador Petit, Pedro López, and José Duato. Combining recency of information with selective random and a victim cache in last-level caches. ACM Transactions on Architecture and Code Optimization, 9(3):16:1–16:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vandierendonck:2013:ADT

[VTN13] Hans Vandierendonck, George Tzenakis, and Dimitrios S. Nikolopoulos. Analysis of dependence tracking algorithms for task dataflow execution. ACM Transactions on Architecture and Code Optimization, 10(4):61:1–61:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vespa:2011:DFA

[VW11] Lucas Vespa and Ning Weng.

Deterministic finite automata characterization and optimization for scalable pattern matching. ACM Transactions on Architecture and Code Optimization, 8(1):4:1–4:??, April 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Voitsechov:2018:SDT

[VZS^{+}18] Dani Voitsechov, Arslan Zulfiqar, Mark Stephenson, Mark Gebhart, and Stephen W. Keckler. Software-directed techniques for improved GPU register file utilization. ACM Transactions on Architecture and Code Optimization, 15(3): 38:1–38:??, October 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vasilache:2020:NAL

[VZT^{+}20] Nicolas Vasilache, Oleksandr Zinenko, Theodoros Theodoridis, Priya Goyal, Zachary Devito, William S. Moses, Sven Verdoolaege, Andrew Adams, and Albert Cohen. The next 700 accelerated layers: From mathematical expressions of network computation graphs to accelerated GPU kernels, automatically. ACM Transactions on Architecture and Code Optimization, 16(4): 38:1–38:26, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Winter:2008:ATN

[WA08] Jonathan A. Winter and David H. Albonesi. Addressing thermal nonuniformity in SMT workloads. ACM Transactions on Architecture and Code Optimization, 5(1):4:1–4:??, May 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wibowo:2016:ACL

[WAST16] Bagus Wibowo, Abhinav Agrawal, Thomas Stanton, and James Tuck. An accurate cross-layer approach for online architectural vulnerability estimation. ACM Transactions on Architecture and Code Optimization, 13(3):30:1–30:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2016:HPC

[WCI^{+}16] Xueyang Wang, Sek Chai, Michael Isnardi, Sehoon Lim, and Ramesh Karri. Hardware performance counterbased malware identification and detection with adaptive compressive sensing. ACM Transactions on Architecture and Code Optimization, 13(1): 3:1–3:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2024:OSM

[WCW^{+}24] Yizhuo Wang, Fangli Chang, Bingxin Wei, Jianhua Gao, and Weixing Ji. Optimization of sparse matrix computation for algebraic multigrid on GPUs. ACM Transactions on Architecture and Code Optimization, 21(3):54:1–54:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3664924.

Wang:2024:EPM

[WDK^{+}24] Ziheng Wang, Xiaoshe Dong, Yan Kang, Heng Chen, and Qiang Wang. An example of parallel Merkle tree traversal: Post-quantum Leighton– Micali signature on the GPU. ACM Transactions on Architecture and Code Optimization, 21(3):44:1–44:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3659209.

Wang:2024:WWO

[WDS^{+}24] Zhonghua Wang, Chen Ding, Fengguang Song, Kai Lu, Jiguang Wan, Zhihu Tan, Changsheng Xie, and Guokuan Li. WIPE: a write-optimized learned index for persistent memory. ACM Transactions on Architecture and Code Optimization, 21(2):22:1–22:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3634915.

Wang:2022:OSS

[WDW^{+}22] Yufei Wang, Xiaoshe Dong, Longxiang Wang, Weiduo Chen, and Xingjun Zhang. Optimizing small-sample disk fault detection based on LSTM-GAN model. ACM Transactions on Architecture and Code Optimization, 19(1): 13:1–13:24, March 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3500917.

Wang:2014:PSR

[WDX14] Jue Wang, Xiangyu Dong, and Yuan Xie. Preventing STT-RAM last-level caches from port obstruction. ACM Transactions on Architecture and Code Optimization, 11(3): 23:1–23:??, October 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2015:BOM

[WDX15] Jue Wang, Xiangyu Dong, and Yuan Xie. Building and optimizing MRAM-based commodity memories. ACM Transactions on Architecture and Code Optimization, 11(4): 36:1–36:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wu:2024:CNI

[WDX24] Ke Wu, Dezun Dong, and Weixia Xu. COER: a network interface offloading architecture for RDMA and congestion control protocol codesign. ACM Transactions on Architecture and Code Optimization, 21(3):49:1–49:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660525.

Wang:2014:EAC

[WDXJ14] Jue Wang, Xiangyu Dong, Yuan Xie, and Norman P. Jouppi. Endurance-aware cache line management for non-volatile caches. ACM Transactions on Architecture and Code Optimization, 11(1): 4:1–4:24, February 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Woo:2010:CVI

[WFKL10] Dong Hyuk Woo, Joshua B.

Fryman, Allan D. Knies, and Hsien-Hsin S. Lee. Chameleon: Virtualizing idle acceleration cores of a heterogeneous multicore processor for caching and prefetching. ACM Transactions on Architecture and Code Optimization, 7(1):3:1– 3:??, April 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Weber:2017:MAL

[WG17] Nicolas Weber and Michael Goesele. MATOG: Array layout auto-tuning for CUDA. ACM Transactions on Architecture and Code Optimization, 14(3):28:1–28:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2024:RRR

[WGL^{+}24] Zhonghua Wang, Yixing Guo, Kai Lu, Jiguang Wan, Daohui Wang, Ting Yao, and Huatao Wu. Rcmp: Reconstructing RDMA-Based memory disaggregation via CXL. ACM Transactions on Architecture and Code Optimization, 21(1): 15:1–15:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3634916.

Wang:2015:APM

[WGO15] Zheng Wang, Dominik Grewe, and Michael F. P. O’Boyle. Automatic and portable mapping of data parallel programs to OpenCL for GPU-based heterogeneous systems. ACM Transactions on Architecture and Code Optimization, 11(4): 42:1–42:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wimmer:2013:MAV

[WHV^{+}13] Christian Wimmer, Michael Haupt, Michael L. Van De Vanter, Mick Jordan, Laurent Daynès, and Douglas Simon. Maxine: an approachable virtual machine for, and in, Java. ACM Transactions on Architecture and Code Optimization, 9(4):30:1–30:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wei:2017:HHM

[WJXC17] Wei Wei, Dejun Jiang, Jin Xiong, and Mingyu Chen. HAP: Hybrid-memory-aware partition in shared last-level cache. ACM Transactions on Architecture and Code Optimization, 14(3):24:1–24:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wegiel:2009:SRC

[WK09] Michal Wegiel and Chandra Krintz. The single-referent collector: Optimizing compaction for the common case. ACM Transactions on Architecture and Code Optimization, 6(4): 15:1–15:??, October 2009. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2012:TMA

[WKCS12] Qingping Wang, Sameer Kulkarni, John Cavazos, and Michael Spear. A transactional memory with automatic performance tuning. ACM Transactions on Architecture and Code Optimization, 8(4):54:1– 54:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2024:FCM

[WLJ^{+}24] Xueying Wang, Guangli Li, Zhen Jia, Xiaobing Feng, and Yida Wang. Fast convolution meets low precision: Exploring efficient quantized Winograd convolution on modern CPUs. ACM Transactions on Architecture and Code Optimization, 21(1):5:1–5:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632956.

Wang:2019:SSL

[WLL^{+}19] Xiaoyuan Wang, Haikun Liu, Xiaofei Liao, Ji Chen, Hai Jin, Yu Zhang, Long Zheng, Bingsheng He, and Song Jiang. Supporting superpages and lightweight page migration in hybrid memory systems. ACM Transactions on Architecture and Code Optimization, 16(2): 11:1–11:??, May 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wu:2020:MBS

[WLLW20] Hao Wu, Weizhi Liu, Huanxin Lin, and Cho-Li Wang. A model-based software solution for simultaneous multiple kernels on GPUs. ACM Transactions on Architecture and Code Optimization, 17(1):7:1–7:26, March 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3377138.

Wang:2019:PNW

[WLWB19] Yu Wang, Victor Lee, Gu-Yeon Wei, and David Brooks.

Predicting new workload or CPU performance by analyzing public datasets. ACM Transactions on Architecture and Code Optimization, 15(4): 53:1–53:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2024:INM

[WLX24] Zhuohao Wang, Lei Liu, and Limin Xiao. iSwap: a new memory page swap mechanism for reducing ineffective I/O operations in cloud environments. ACM Transactions on Architecture and Code Optimization, 21(3):47:1–47:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3653302.

Wu:2010:DEH

[WLZ^{+}10] Xiaoxia Wu, Jian Li, Lixin Zhang, Evan Speight, Ram Rajamony, and Yuan Xie. Design exploration of hybrid caches with disparate memory technologies. ACM Transactions on Architecture and Code Optimization, 7(3):15:1–15:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2013:MTD

[WLZ^{+}13] Chao Wang, Xi Li, Junneng Zhang, Xuehai Zhou, and Xiaoning Nie. MP-Tomasulo: a dependency-aware automatic parallel execution engine for sequential programs. ACM Transactions on Architecture and Code Optimization, 10(2): 9:1–9:??, May 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wimmer:2010:AFD

[WM10] Christian Wimmer and Hanspeter Mössenbösck. Automatic feedback-directed object fusing. ACM Transactions on Architecture and Code Optimization, 7(2):7:1–7:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wu:2011:ATR

[WM11] Carole-Jean Wu and Margaret Martonosi. Adaptive timekeeping replacement: Finegrained capacity management for shared CMP caches. ACM Transactions on Architecture and Code Optimization, 8(1): 3:1–3:??, April 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2019:GGC

[WMGS19] Ping Wang, Luke Mchale, Paul V. Gratz, and Alex Sprintson. GenMatcher: a generic clustering-based arbitrary matching framework. ACM Transactions on Architecture and Code Optimization, 15(4):51:1–51:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2013:UML

[WO13] Zheng Wang and Michael F. P.

O’Boyle. Using machine learning to partition streaming programs. ACM Transactions on Architecture and Code Optimization, 10(3):20:1–20:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2019:SSS

[WPJ19] Jiajun Wang, Reena Panda, and Lizy K. John. SelSMaP: a selective stride masking prefetching scheme. ACM Transactions on Architecture and Code Optimization, 15(4): 42:1–42:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2022:MOG

[WPR^{+}22] Qihan Wang, Zhen Peng, Bin Ren, Jie Chen, and Robert G. Edwards. MemHC: an optimized GPU memory management framework for accelerating many-body correlation. ACM Transactions on Architecture and Code Optimization, 19(2):24:1–24:26, June 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3506705.

Wang:2013:WWA

[WSC^{+}13] Zhe Wang, Shuchang Shan, Ting Cao, Junli Gu, Yi Xu, Shuai Mu, Yuan Xie, and Daniel A. Jiménez. WADE:

Writeback-aware dynamic cache management for NVM-based main memory system. ACM Transactions on Architecture and Code Optimization, 10(4): 51:1–51:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Walden:2021:MIN

[WSJ^{+}21] Candace Walden, Devesh Singh, Meenatchi Jagasivamani, Shang Li, Luyi Kang, Mehdi Asnaashari, Sylvain Dubois, Bruce Jacob, and Donald Yeung. Monolithically integrating non-volatile main memory over the lastlevel cache. ACM Transactions on Architecture and Code Optimization, 18(4):48:1–48:26, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462632.

Wang:2014:IPD

[WTFO14] Zheng Wang, Georgios Tournavitis, Björn Franke, and Michael F. P. O’Boyle. Integrating profile-driven parallelism detection and machinelearning-based mapping. ACM Transactions on Architecture and Code Optimization, 11(1): 2:1–2:26, February 2014. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2016:IIB

[WWC^{+}16] Yaohua Wang, Dong Wang, Shuming Chen, Zonglin Liu, Shenggang Chen, Xiaowen Chen, and Xu Zhou. Iteration interleaving–based SIMD lane partition. ACM Transactions on Architecture and Code Optimization, 12(4):58:1–58:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2022:SMS

[WWGS22] Ping Wang, Fei Wen, Paul V.

Gratz, and Alex Sprintson. SIMD-Matcher: a SIMD-based arbitrary matching framework. ACM Transactions on Architecture and Code Optimization, 19(3):30:1–30:20, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3514246.

Wang:2016:DMB

[WWH^{+}16] Zhigang Wang, Xiaolin Wang, Fang Hou, Yingwei Luo, and Zhenlin Wang. Dynamic memory balancing for virtualization. ACM Transactions on Architecture and Code Optimization, 13(1):2:1–2:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2021:GTU

[WWL^{+}21] Pengyu Wang, Jing Wang, Chao Li, Jianzong Wang, Haojin Zhu, and Minyi Guo. Grus: Toward unifiedmemory-efficient high-performance graph processing on GPU.

ACM Transactions on Architecture and Code Optimization, 18(2):22:1–22:25, March 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3444844.

Wang:2021:EPO

[WWW^{+}21] Yu Emma Wang, Carole-Jean Wu, Xiaodong Wang, Kim Hazelwood, and David Brooks. Exploiting parallelism opportunities with deep learning frameworks. ACM Transactions on Architecture and Code Optimization, 18(1):9:1–9:23, January 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3431388.

Wang:2013:RMM

[WWWL13] Xiaolin Wang, Lingmei Weng, Zhenlin Wang, and Yingwei Luo. Revisiting memory management on virtualized environments. ACM Transactions on Architecture and Code Optimization, 10(4):48:1–48:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2012:FSS

[WWY^{+}12] Zhenjiang Wang, Chenggang Wu, Pen-Chung Yew, Jianjun Li, and Di Xu. On-the-fly structure splitting for heap objects. ACM Transactions on Architecture and Code Optimization, 8(4):26:1–26:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wei:2024:RRI

[WWY^{+}24] Xiaohui Wei, Chenyang Wang, Hengshan Yue, Jingweijia Tan, Zeyu Guan, Nan Jiang, Xinyang Zheng, Jianpeng Zhao, and Meikang Qiu. ReIPE: Recycling idle PEs in CNN accelerator for vulnerable filters soft-error detection. ACM Transactions on Architecture and Code Optimization, 21(3):61:1–61:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674909.

Wei:2024:CCB

[WX24] Yan Wei and Zhang Xingjun.

A concise concurrent B+-tree for persistent memory. ACM Transactions on Architecture and Code Optimization, 21(2): 24:1–24:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3638717.

Wang:2011:PGS

[WYCC11] Po-Han Wang, Chia-Lin Yang, Yen-Ming Chen, and Yu-Jung Cheng. Power gating strategies on GPUs. ACM Transactions on Architecture and Code Optimization, 8(3):13:1–13:??, October 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2010:PAM

[WYJL10] Xiaohang Wang, Mei Yang, Yingtao Jiang, and Peng Liu. A power-aware mapping approach to map IP cores onto NoCs under bandwidth and latency constraints. ACM Transactions on Architecture and Code Optimization, 7(1): 1:1–1:??, April 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wu:2023:SDM

[WYZ^{+}23] Donglei Wu, Weihao Yang, Xiangyu Zou, Wen Xia, Shiyi Li, Zhenbo Hu, Weizhe Zhang, and Binxing Fang. Smart-DNN+: a memory-efficient neural networks compression framework for the model inference. ACM Transactions on Architecture and Code Optimization, 20(4):49:1–49:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617688.

Wu:2019:DPC

[WZG^{+}19] Song Wu, Fang Zhou, Xiang Gao, Hai Jin, and Jinglei Ren. Dual-page checkpointing: an architectural approach to efficient data persistence for inmemory applications. ACM Transactions on Architecture and Code Optimization, 15(4): 57:1–57:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2024:AMA

[WZW^{+}24] Luming Wang, Xu Zhang, Songyue Wang, Zhuolun Jiang, Tianyue Lu, Mingyu Chen, Siwei Luo, and Keji Huang. Asynchronous memory access unit: Exploiting massive parallelism for far memory access. ACM Transactions on Architecture and Code Optimization, 21(3):55:1–55:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663479.

Wei:2023:FOT

[WZWW23] Jia Wei, Xingjun Zhang, Longxiang Wang, and Zheng Wei. Fastensor: Optimise the Tensor I/O path from SSD to GPU for deep learning training. ACM Transactions on Architecture and Code Optimization, 20(4):62:1–62:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3630108.

Wu:2024:AAM

[WZX^{+}24] Lingxi Wu, Minxuan Zhou, Weihong Xu, Ashish Venkat, Tajana Rosing, and Kevin Skadron. Abakus: Accelerating k-mer counting with storage technology. ACM Transactions on Architecture and Code Optimization, 21(1): 10:1–10:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3632952.

Wu:2024:KAM

[WZY^{+}24a] Jiang Wu, Zhuo Zhang, Deheng Yang, Jianjun Xu, Jiayu He, and Xiaoguang Mao. Knowledge-augmented mutation-based bug localization for hardware design code. ACM Transactions on Architecture and Code Optimization, 21(3):45:1–45:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3660526.

Wu:2024:TAS

[WZY^{+}24b] Jiang Wu, Zhuo Zhang, Deheng Yang, Jianjun Xu, Jiayu He, and Xiaoguang Mao. Time-aware spectrum-based bug localization for hardware design code with data purification. ACM Transactions on Architecture and Code Optimization, 21(3):64:1–64:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3678009.

Wang:2020:CFS

[WZZ^{+}20] Qinggang Wang, Long Zheng, Jieshan Zhao, Xiaofei Liao, Hai Jin, and Jingling Xue. A conflict-free scheduler for highperformance graph processing on multi-pipeline FPGAs.

ACM Transactions on Architecture and Code Optimization, 17(2):14:1–14:26, June 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3390523.

Xue:2006:LOA

[XC06] Jingling Xue and Qiong Cai.

A lifetime optimal algorithm for speculative PRE. ACM Transactions on Architecture and Code Optimization, 3(2): 115–155, June 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xiao:2013:SAT

[XCC^{+}13] Chunhua Xiao, M-C. Frank Chang, Jason Cong, Michael Gill, Zhangqin Huang, Chunyue Liu, Glenn Reinman, and Hao Wu. Stream arbitration: Towards efficient bandwidth utilization for emerging on-chip interconnects. ACM Transactions on Architecture and Code Optimization, 9(4): 60:1–60:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xu:2023:SSC

[XDW^{+}23] Yemao Xu, Dezun Dong, Dongsheng Wang, Shi Xu, Enda Yu, Weixia Xu, and Xiangke Liao. SSD-SGD: Communication sparsification for distributed deep learning training. ACM Transactions on Architecture and Code Optimization, 20(1):7:1–7:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3563038.

Xu:2019:SSD

[XDXL19] Yemao Xu, Dezun Dong, Weixia Xu, and Xiangke Liao. SketchDLC: a sketch on distributed deep learning communication via trace capturing. ACM Transactions on Architecture and Code Optimization, 16(2):7:1–7:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xu:2019:PTA

[XFS^{+}19] Jingheng Xu, Haohuan Fu, Wen Shi, Lin Gan, Yuxuan Li, Wayne Luk, and Guangwen Yang. Performance tuning and analysis for stencil-based applications on POWER8 processor. ACM Transactions on Architecture and Code Optimization, 15(4): 41:1–41:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xie:2023:MMC

[XGD^{+}23] Xinfeng Xie, Peng Gu, Yufei Ding, Dimin Niu, Hongzhong Zheng, and Yuan Xie. MPU: Memory-centric SIMT processor via in-DRAM near-bank computing. ACM Transactions on Architecture and Code Optimization, 20(3): 40:1–40:??, September 2023.

CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3603113.

Xiong:2016:MAS

[XHJY16] Dongliang Xiong, Kai Huang, Xiaowen Jiang, and Xiaolang Yan. Memory access scheduling based on dynamic multilevel priority in shared DRAM systems. ACM Transactions on Architecture and Code Optimization, 13(4):42:1–42:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xiong:2017:PPP

[XHJY17] Dongliang Xiong, Kai Huang, Xiaowen Jiang, and Xiaolang Yan. Providing predictable performance via a slowdown estimation model. ACM Transactions on Architecture and Code Optimization, 14(3): 25:1–25:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xue:2024:TEG

[XHL^{+}24] Feng Xue, Chenji Han, Xinyu Li, Junliang Wu, Tingting Zhang, Tianyi Liu, Yifan Hao, Zidong Du, Qi Guo, and Fuxin Zhang. Tyche: an efficient and general prefetcher for indirect memory accesses. ACM Transactions on Architecture and Code Optimization, 21(2): 30:1–30:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3641853.

Xekalakis:2012:MSM

[XIC12] Polychronis Xekalakis, Nikolas Ioannou, and Marcelo Cintra. Mixed speculative multithreaded execution models. ACM Transactions on Architecture and Code Optimization, 9(3):18:1–18:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xiao:2007:VIS

[XL07] Shu Xiao and Edmund M.- K. Lai. VLIW instruction scheduling for minimal power variation. ACM Transactions on Architecture and Code Optimization, 4(3):18:1–18:??, September 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xu:2024:RAR

[XLD^{+}24] Jiahong Xu, Haikun Liu, Zhuohui Duan, Xiaofei Liao, Hai Jin, Xiaokang Yang, Huize Li, Cong Liu, Fubing Mao, and Yu Zhang. Re-Harvest: an ADC resourceharvesting crossbar architecture for ReRAM-based DNN accelerators. ACM Transactions on Architecture and Code Optimization, 21(3): 63:1–63:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3659208.

Xie:2024:WLT

[XLH^{+}24] Kunpeng Xie, Ye Lu, Xinyu He, Dezhi Yi, Huijuan Dong, and Yao Chen. Winols: a large-tiling sparse Winograd CNN accelerator on FPGAs. ACM Transactions on Architecture and Code Optimization, 21(2):31:1–31:??, June 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643682.

Xie:2004:IDV

[XMM04] Fen Xie, Margaret Martonosi, and Sharad Malik. Intraprogram dynamic voltage scaling: Bounding opportunities with analytic modeling. ACM Transactions on Architecture and Code Optimization, 1(3): 323–367, September 2004. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xu:2021:CMD

[XMW^{+}21] Rui Xu, Sheng Ma, Yaohua Wang, Xinhai Chen, and Yang Guo. Configurable multi-directional systolic array architecture for convolutional neural networks. ACM Transactions on Architecture and Code Optimization, 18(4): 42:1–42:24, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3460776.

Xu:2023:ACN

[XSF^{+}23] Weizhi Xu, Yintai Sun, Shengyu Fan, Hui Yu, and Xin Fu. Accelerating convolutional neural network by exploiting sparsity on GPUs. ACM Transactions on Architecture and Code Optimization, 20(3): 36:1–36:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3600092.

Xu:2009:TXP

[XT09] Weifeng Xu and Russell Tessier. Tetris-XL: a performancedriven spill reduction technique for embedded VLIW processors. ACM Transactions on Architecture and Code Optimization, 6(3):11:1–11:??, September 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xue:2020:NIA

[XVT20] Jiachen Xue, T. N. Vijaykumar, and Mithuna Thottethodi. Network interface architecture for remote indirect memory access (RIMA) in datacenters. ACM Transactions on Architecture and Code Optimization, 17(2):13:1–13:22, June 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3374215.

Xie:2024:IIA

[XYY^{+}24] Benyi Xie, Yue Yan, Chenghao Yan, Sicheng Tao, Zhuangzhuang Zhang, Xinyu Li, Yanzhi Lan, Xiang Wu, Tianyi Liu, Tingting Zhang, and Fuxin Zhang. An instruction inflation analyzing framework for dynamic binary translators. ACM Transactions on Architecture and Code Optimization, 21(2): 28:1–28:??, June 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3640813.

Xia:2020:DAB

[XZC^{+}20] Chunwei Xia, Jiacheng Zhao, Huimin Cui, Xiaobing Feng, and Jingling Xue. DNNTune: Automatic benchmarking DNN models for mobile-cloud computing. ACM Transactions on Architecture and Code Optimization, 16(4):49:1–49:26, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3368305.

Xu:2022:BFE

[XZW^{+}22] Peng Xu, Nannan Zhao, Jiguang Wan, Wei Liu, Shuning Chen, Yuanhui Zhou, Hadeel Albahar, Hanyang Liu, Liu Tang, and Zhihu Tan. Building a fast and efficient LSM-tree store by integrating local storage with cloud storage. ACM Transactions on Architecture and Code Optimization, 19(3):37:1–37:26, September 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3527452.

Yilmaz:2016:ARS

[YAG^{+}16] Buse Yilmaz, Baris Aktemur, MaríA J. Garzarán, Sam Kamin, and Furkan Kiraç. Autotuning runtime specialization for sparse matrix-vector multiplication. ACM Transactions on Architecture and Code Optimization, 13(1):5:1– 5:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yu:2019:ITL

[YBSY19] Chao Yu, Yuebin Bai, Qingxiao Sun, and Hailong Yang. Improving thread-level parallelism in GPUs through expanding register file to scratchpad memory. ACM Transactions on Architecture and Code Optimization, 15(4): 48:1–48:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yviquel:2018:CPU

[YCA18] Hervé Yviquel, Lauro Cruz, and Guido Araujo. Cluster programming using the OpenMP accelerator model. ACM Transactions on Architecture and Code Optimization, 15(3):35:1–35:??, October 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yeh:2011:MPP

[YCCY11] Chang-Ching Yeh, Kuei-Chung Chang, Tien-Fu Chen, and Chingwei Yeh. Maintaining performance on power gating of microprocessor functional units by using a predictive pre-wakeup strategy. ACM Transactions on Architecture and Code Optimization, 8(3):16:1–16:??, October 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ye:2017:CES

[YDL^{+}17] Chencheng Ye, Chen Ding, Hao Luo, Jacob Brock, Dong Chen, and Hai Jin. Cache exclusivity and sharing: Theory and optimization. ACM Transactions on Architecture and Code Optimization, 14(4): 34:1–34:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yuan:2019:RTL

[YDS^{+}19] Liang Yuan, Chen Ding, Wesley Smith, Peter Denning, and Yunquan Zhang. A relational theory of locality. ACM Transactions on Architecture and Code Optimization, 16(3):33:1–33:??, August 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3341109.

Yalcin:2014:EEC

[YEI^{+}14] Gulay Yalcin, Oguz Ergin, Emrah Islek, Osman Sabri Unsal, and Adrian Cristal. Exploiting existing comparators for fine-grained low-cost error detection. ACM Transactions on Architecture and Code Optimization, 11(3):32:1–32:??, October 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yadalam:2021:SXS

[YGB21] Sujay Yadalam, Vinod Ganapathy, and Arkaprava Basu. SG XL: Security and performance for enclaves using large pages. ACM Transactions on Architecture and Code Optimization, 18(1):12:1–12:25, January 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3433983.

Yasin:2020:MGM

[YHYBAM20] Ahmad Yasin, Jawad Haj-Yahya, Yosi Ben-Asher, and Avi Mendelson. A metricguided method for discovering impactful features and architectural insights for Skylakebased processors. ACM Transactions on Architecture and Code Optimization, 16(4): 46:1–46:25, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3369383.

Yan:2013:IPA

[YJTF13] Zhichao Yan, Hong Jiang, Yujuan Tan, and Dan Feng. An integrated pseudo-associativity and relaxed-order approach to hardware transactional memory. ACM Transactions on Architecture and Code Optimization, 9(4):42:1–42:??, January 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yang:2017:EJV

[YKM17] Byung-Sun Yang, Jae-Yun Kim, and Soo-Mook Moon. Exceptionization: a Java VM optimization for non-Java languages. ACM Transactions on Architecture and Code Optimization, 14(1):5:1–5:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yang:2004:TML

[YLTL04] Chia-Lin Yang, Alvin R.

Lebeck, Hung-Wei Tseng, and Chien-Hao Lee. Tolerating memory latency through push prefetching for pointerintensive applications. ACM Transactions on Architecture and Code Optimization, 1(4): 445–475, December 2004. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yu:2008:OCL

[YLW08] Zoe C. H. Yu, Francis C. M.

Lau, and Cho-Li Wang. Object co-location and memory reuse for Java programs. ACM Transactions on Architecture and Code Optimization, 4(4): 4:1–4:??, January 2008. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yoon:2015:EDM

[YMM^{+}15] Hanbin Yoon, Justin Meza, Naveen Muralimanohar, Norman P. Jouppi, and Onur Mutlu. Efficient data mapping and buffering techniques for multilevel cell phase-change memories. ACM Transactions on Architecture and Code Optimization, 11(4):40:1–40:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yazdanbakhsh:2016:RRF

[YPT^{+}16] Amir Yazdanbakhsh, Gennady Pekhimenko, Bradley Thwaites, Hadi Esmaeilzadeh, Onur Mutlu, and Todd C. Mowry. RFVP: Rollbackfree value prediction with safeto-approximate loads. ACM Transactions on Architecture and Code Optimization, 12(4): 62:1–62:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yebenes:2019:CSA

[YRGES^{+}19] Pedro Yebenes, Jose Rocher-Gonzalez, Jesus Escudero-Sahuquillo, Pedro Javier Garcia, Francisco J. Alfaro, Francisco J. Quiles, Crispín Gómez, and Jose Duato. Combining source-adaptive and oblivious routing with congestion control in high-performance interconnects using hybrid and direct topologies. ACM Transactions on Architecture and Code Optimization, 16(2): 17:1–17:??, May 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yiapanis:2013:OSR

[YRHBL13] Paraskevas Yiapanis, Demian Rosas-Ham, Gavin Brown, and Mikel Luján. Optimizing software runtime systems for speculative parallelization. ACM Transactions on Architecture and Code Optimization, 9(4): 39:1–39:??, January 2013. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yuzuguler:2023:SSA

[YSD^{+}23] Ahmet Caner Yüzügüler, Canberk Sönmez, Mario Drumond, Yunho Oh, Babak Falsafi, and Pascal Frossard. Scaleout systolic arrays. ACM Transactions on Architecture and Code Optimization, 20(2): 27:1–27:??, June 2023. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3572917.

Yousefzadeh:2022:EEM

[YSH^{+}22] Amirreza Yousefzadeh, Jan Stuijt, Martijn Hijdra, Hsiao-Hsuan Liu, Anteneh Gebregiorgis, Abhairaj Singh, Said Hamdioui, and Francky Catthoor. Energy-efficient inmemory address calculation. ACM Transactions on Architecture and Code Optimization, 19(4):52:1–52:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3546071.

Yang:2012:CGC

[YWXW12] Xuejun Yang, Li Wang, Jingling Xue, and Qingbo Wu. Comparability graph coloring for optimizing utilization of software-managed stream register files for stream processors. ACM Transactions on Architecture and Code Optimization, 9(1):5:1–5:??, March 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yang:2012:UOC

[YXK^{+}12] Yi Yang, Ping Xiang, Jingfei Kong, Mike Mantor, and Huiyang Zhou. A unified optimizing compiler framework for different GPGPU architectures. ACM Transactions on Architecture and Code Optimization, 9(2):9:1–9:??, June 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yu:2023:MFE

[YXMC23] Miao Yu, Tingting Xiang, Venkata Pavan Kumar Miriyala, and Trevor E. Carlson. Multiplyand-Fire: an event-driven sparse neural network accelerator. ACM Transactions on Architecture and Code Optimization, 20(4):59:1–59:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3630255.

Ye:2022:PAU

[YXS^{+}22] Chencheng Ye, Yuanchao Xu, Xipeng Shen, Hai Jin, Xiaofei Liao, and Yan Solihin. Preserving addressability upon GC-triggered data movements on non-volatile memory. ACM Transactions on Architecture and Code Optimization, 19(2): 28:1–28:26, June 2022. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3511706.

Yan:2008:EVR

[YZ08] Jun Yan and Wei Zhang. Exploiting virtual registers to reduce pressure on real registers. ACM Transactions on Architecture and Code Optimization, 4(4):3:1–3:??, January 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yang:2010:ERS

[YZL^{+}10] Xuejun Yang, Ying Zhang, Xicheng Lu, Jingling Xue, Ian Rogers, Gen Li, Guibin Wang, and Xudong Fang. Exploiting the reuse supplied by loopdependent stream references for stream processors. ACM Transactions on Architecture and Code Optimization, 7(2): 11:1–11:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yu:2023:RER

[YZZ^{+}23] Hui Yu, Yu Zhang, Jin Zhao, Yujian Liao, Zhiying Huang, Donghao He, Lin Gu, Hai Jin, Xiaofei Liao, Haikun Liu, Bingsheng He, and Jianhui Yue. RACE: an efficient redundancy-aware accelerator for dynamic graph neural network. ACM Transactions on Architecture and Code Optimization, 20(4):53:1–53:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617685.

Zhang:2022:AAS

[ZBC^{+}22] Chao Zhang, Maximilian Bremer, Cy Chan, John Shalf, and Xiaochen Guo. ASA: Accelerating sparse accumulation in column-wise SpGEMM. ACM Transactions on Architecture and Code Optimization, 19(4): 49:1–49:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3543068.

Zhao:2013:HPP

[ZBH^{+}13] Zhijia Zhao, Michael Bebenita, Dave Herman, Jianhua Sun, and Xipeng Shen. HPar: a practical parallel parser for HTML — taming HTML complexities for parallel parsing.

ACM Transactions on Architecture and Code Optimization, 10(4):44:1–44:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2020:FTF

[ZC20] Jie Zhao and Albert Cohen.

Flextended tiles: a flexible extension of overlapped tiles for polyhedral compilation. ACM Transactions on Architecture and Code Optimization, 16(4): 47:1–47:25, January 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3369382.

Zhao:2016:FMR

[ZCCD16] Lihang Zhao, Lizhong Chen, Woojin Choi, and Jeffrey Draper. A filtering mechanism to reduce network bandwidth utilization of transaction execution. ACM Transactions on Architecture and Code Optimization, 12(4):51:1–51:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhou:2023:IPS

[ZCDD23] Yufeng Zhou, Alan L. Cox, Sandhya Dwarkadas, and Xiaowan Dong. The impact of page size and microarchitecture on instruction address translation overhead. ACM Transactions on Architecture and Code Optimization, 20(3): 38:1–38:??, September 2023.

CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3600089.

Zoni:2018:DEP

[ZCF18] Davide Zoni, Luca Colombo, and William Fornaciari. Dark-Cache: Energy-performance optimization of tiled multicores by adaptively powergating LLC banks. ACM Transactions on Architecture and Code Optimization, 15(2): 21:1–21:??, June 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2019:BLA

[ZCQ^{+}19] Han Zhao, Quan Chen, Yuxian Qiu, Ming Wu, Yao Shen, Jingwen Leng, Chao Li, and Minyi Guo. Bandwidth and locality aware task-stealing for manycore architectures with bandwidth-asymmetric memory. ACM Transactions on Architecture and Code Optimization, 15(4):55:1–55:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2006:ATP

[ZCS06] Min Zhao, Bruce R. Childers, and Mary Lou Soffa. An approach toward profit-driven optimization. ACM Transactions on Architecture and Code Optimization, 3(3):231– 262, September 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2010:PPP

[ZCW10] Qin Zhao, Ioana Cutcutache, and Weng-Fai Wong. PiPA: Pipelined profiling and analysis on multicore systems. ACM Transactions on Architecture and Code Optimization, 7(3): 13:1–13:??, December 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhou:2012:WAP

[ZDC^{+}12] Miao Zhou, Yu Du, Bruce Childers, Rami Melhem, and Daniel Mossé. Writebackaware partitioning and replacement for last-level caches in phase change main memory systems. ACM Transactions on Architecture and Code Optimization, 8(4):53:1–53:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhou:2016:SAC

[ZDC^{+}16] Miao Zhou, Yu Du, Bruce Childers, Daniel Mosse, and Rami Melhem. Symmetryagnostic coordinated management of the memory hierarchy in multicore systems. ACM Transactions on Architecture and Code Optimization, 12(4): 61:1–61:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhou:2024:SSA

[ZF24] Hai Zhou and Dan Feng.

Stripe-schedule aware repair in erasure-coded clusters with heterogeneous star networks. ACM Transactions on Architecture and Code Optimization, 21(3):52:1–52:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3664926.

Zhao:2018:OCN

[ZFF^{+}18] Wenlai Zhao, Haohuan Fu, Jiarui Fang, Weijie Zheng, Lin Gan, and Guangwen Yang. Optimizing convolutional neural networks on the Sunway TaihuLight Supercomputer. ACM Transactions on Architecture and Code Optimization, 15(1):13:1–13:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zahedi:2018:MHD

[ZFL18] Seyed Majid Zahedi, Songchun Fan, and Benjamin C. Lee. Managing heterogeneous datacenters with tokens. ACM Transactions on Architecture and Code Optimization, 15(2): 18:1–18:??, June 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2018:CNC

[ZFT^{+}18] Yang Zhang, Dan Feng, Wei Tong, Yu Hua, Jingning Liu, Zhipeng Tan, Chengning Wang, Bing Wu, Zheng Li, and Gaoxiang Xu. CACF: a novel circuit architecture co-optimization framework for improving performance, reliability and energy of ReRAMbased main memory system. ACM Transactions on Architecture and Code Optimization, 15(2):22:1–22:??, June 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2005:WET

[ZG05] Xiangyu Zhang and Rajiv Gupta. Whole execution traces and their applications. ACM Transactions on Architecture and Code Optimization, 2(3): 301–334, September 2005. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2012:TPB

[ZGC^{+}12] Dongsong Zhang, Deke Guo, Fangyuan Chen, Fei Wu, Tong Wu, Ting Cao, and Shiyao Jin. TL-plane-based multicore energy-efficient real-time scheduling algorithm for sporadic tasks. ACM Transactions on Architecture and Code Optimization, 8(4):47:1–47:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhou:2015:OPS

[ZGP15] Xing Zhou, María J. Garzarán, and David A. Padua. Optimal parallelogram selection for hierarchical tiling. ACM Transactions on Architecture and Code Optimization, 11(4): 58:1–58:??, January 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zou:2022:PSB

[ZGX22] Changwei Zou, Yaoqing Gao, and Jingling Xue. Practical software-based shadow stacks on x86-64. ACM Transactions on Architecture and Code Optimization, 19(4):61:1–61:??, December 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3556977.

Zhang:2008:RCM

[Zha08]  Chuanjun Zhang.  Reduc-

ing cache misses through programmable decoders. ACM Transactions on Architecture and Code Optimization, 4(4): 5:1–5:??, January 2008. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zinenko:2018:VPM

[ZHB18] Oleksandr Zinenko, Stéphane Huot, and Cédric Bastoul. Visual program manipulation in the polyhedral model. ACM Transactions on Architecture and Code Optimization, 15(1): 16:1–16:??, April 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2004:RIC

[ZHD^{+}04] W. Zhang, J. S. Hu, V. Degalahal, M. Kandemir, N. Vijaykrishnan, and M. J. Irwin. Reducing instruction cache energy consumption using a compiler-based strategy. ACM Transactions on Architecture and Code Optimization, 1(1):3–33, March 2004. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2019:REU

[ZHS^{+}19] Jun Zhang, Rui Hou, Wei Song, Sally A. Mckee, Zhen Jia, Chen Zheng, Mingyu Chen, Lixin Zhang, and Dan Meng. RAGuard: an efficient and user-transparent hardware mechanism against ROP attacks. ACM Transactions on Architecture and Code Optimization, 15(4):50:1–50:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2015:BSS

[ZJJ^{+}15] Tao Zhang, Naifeng Jing, Kaiming Jiang, Wei Shu, Min-You Wu, and Xiaoyao Liang. Buddy SM: Sharing pipeline front-end for improved energy efficiency in GPGPUs. ACM Transactions on Architecture and Code Optimization, 12(2): 16:1–16:??, July 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2005:DIE

[ZK05] Lingli Zhang and Chandra Krintz. The design, implementation, and evaluation of adaptive code unloading for resource-constrained devices. ACM Transactions on Architecture and Code Optimization, 2(2):131–164, June 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zmily:2006:BAI

[ZK06] Ahmad Zmily and Christos Kozyrakis. Block-aware instruction set architecture. ACM Transactions on Architecture and Code Optimization, 3(3):327–357, September 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2015:BSB

[ZLC^{+}15] Jishen Zhao, Sheng Li, Jichuan Chang, John L. Byrne, Laura L. Ramirez, Kevin Lim, Yuan Xie, and Paolo Faraboschi. Buri: Scaling big-memory computing with hardware-based memory expansion. ACM Transactions on Architecture and Code Optimization, 12(3):31:1–31:??, October 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zheng:2018:ESG

[ZLJ18] Long Zheng, Xiaofei Liao, and Hai Jin. Efficient and scalable graph parallel processing with symbolic execution. ACM Transactions on Architecture and Code Optimization, 15(1): 3:1–3:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2023:MGA

[ZLM^{+}23] Yuwen Zhao, Fangfang Liu, Wenjing Ma, Huiyuan Li, Yuanchi Peng, and Cui Wang. MFFT: a GPU accelerated highly efficient mixed-precision large-scale FFT framework. ACM Transactions on Architecture and Code Optimization, 20(3):43:1–43:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3605148.

Zhang:2024:EVI

[ZLWS24] Dunbo Zhang, Qingjie Lang, Ruoxi Wang, and Li Shen. Extension VM: Interleaved data layout in vector memory. ACM Transactions on Architecture and Code Optimization, 21(1): 18:1–18:??, March 2024. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

URL https://dl.acm.org/
doi/10.1145/3631528.

Zhang:2018:SSM

[ZLYW18] Mingzhe Zhang, King Tin Lam, Xin Yao, and Cho-Li Wang. SIMPO: a scalable in-memory persistent object framework using NVRAM for reliable big data computing. ACM Transactions on Architecture and Code Optimization, 15(1):7:1–7:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2016:CPS

[ZLYZ16] Yunquan Zhang, Shigang Li, Shengen Yan, and Huiyang Zhou. A cross-platform SpMV framework on many-core architectures. ACM Transactions on Architecture and Code Optimization, 13(4):33:1–33:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zimmer:2015:NSM

[ZM15] Christopher Zimmer and Frank Mueller. NoCMsg: a scalable message-passing abstraction for network-on-chips. ACM Transactions on Architecture and Code Optimization, 12(1):1:1–1:??, April 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2024:SUE

[ZML^{+}24] Yunping Zhao, Sheng Ma, Heng Liu, Libo Huang, and Yi Dai. SAC: an ultraefficient spin-based architecture for compressed DNNs. ACM Transactions on Architecture and Code Optimization, 21(1):7:1–7:??, March 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3632957.

Zhao:2024:SOD

[ZMLL24] Yunping Zhao, Sheng Ma, Hengzhu Liu, and Dongsheng Li. SAL: Optimizing the dataflow of spin-based architectures for lightweight neural networks. ACM Transactions on Architecture and Code Optimization, 21(3):56:1–56:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673654.

Zhang:2023:CMC

[ZNTJE23] Shiqing Zhang, Mahmood Naderan-Tahan, Magnus Jahre, and Lieven Eeckhout. Characterizing multi-chip GPU data sharing. ACM Transactions on Architecture and Code Optimization, 20(4):56:1–56:??, December 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3629521.

Zhang:2024:ATE

[ZNX^{+}24] Feng Zhang, Fulin Nan, Binbin Xu, Zhirong Shen, Jiebin Zhai, Dmitrii Kalplun, and Jiwu Shu. Achieving tunable erasure coding with clusteraware redundancy transitioning. ACM Transactions on Architecture and Code Optimization, 21(3):59:1–59:??, September 2024. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672077.

Zhang:2006:EAR

[ZPC06] Lixin Zhang, Mike Parker, and John Carter. Efficient address remapping in distributed shared-memory systems. ACM Transactions on Architecture and Code Optimization, 3(2): 209–229, June 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2023:PSC

[ZPH^{+}23] Tuowen Zhao, Tobi Popoola, Mary Hall, Catherine Olschanowsky, and Michelle Strout. Polyhedral specification and code generation of sparse tensor contraction with co-iteration. ACM Transactions on Architecture and Code Optimization, 20(1):16:1–16:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3566054.

Zhang:2021:LED

[ZPL^{+}21] Yu Zhang, Da Peng, Xiaofei Liao, Hai Jin, Haikun Liu, Lin Gu, and Bingsheng He. LargeGraph: an efficient dependency-aware GPUaccelerated large-scale graph processing. ACM Transactions on Architecture and Code Optimization, 18(4):58:1–58:24, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3477603.

Zivanovic:2017:MMH

[ZPR^{+}17] Darko Zivanovic, Milan Pavlovic, Milan Radulovic, Hyunsung Shin, Jongpil Son, Sally A.

Mckee, Paul M. Carpenter, Petar Radojković, and Eduard Ayguadé. Main memory in HPC: Do we need more or could we live with less? ACM Transactions on Architecture and Code Optimization, 14(1): 3:1–3:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhai:2008:CHS

[ZSCM08] Antonia Zhai, J. Gregory Steffan, Christopher B. Colohan, and Todd C. Mowry. Compiler and hardware support for reducing the synchronization of speculative threads. ACM Transactions on Architecture and Code Optimization, 5(1): 3:1–3:??, May 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2013:OGE

[ZSLX13] Jishen Zhao, Guangyu Sun, Gabriel H. Loh, and Yuan Xie. Optimizing GPU energy efficiency with 3D die-stacking graphics memory and reconfigurable memory interface. ACM Transactions on Architecture and Code Optimization, 10(4): 24:1–24:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ziabari:2016:UHB

[ZSM^{+}16] Amir Kavyan Ziabari, Yifan Sun, Yenai Ma, Dana Schaa, José L. Abellán, Rafael Ubal, John Kim, Ajay Joshi, and David Kaeli. UMH: a hardware-based unified memory hierarchy for systems with multiple discrete GPUs. ACM Transactions on Architecture and Code Optimization, 13(4): 35:1–35:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2005:WHC

[ZVYN05] Chuanjun Zhang, Frank Vahid, Jun Yang, and Walid Najjar. A way-halting cache for lowenergy high-performance systems. ACM Transactions on Architecture and Code Optimization, 2(1):34–54, March 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2005:IWA

[ZWHM05] Wankang Zhao, David Whalley, Christopher Healy, and Frank Mueller. Improving WCET by applying a WC code-positioning optimization. ACM Transactions on Architecture and Code Optimization, 2(4):335–365, December 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhou:2019:SNS

[ZWL^{+}19] You Zhou, Fei Wu, Zhonghai Lu, Xubin He, Ping Huang, and Changsheng Xie. SCORE: a novel scheme to efficiently cache overlong ECCs in NAND flash memory. ACM Transactions on Architecture and Code Optimization, 15(4):

60:1–60:??, January 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhou:2016:ERI

[ZWS^{+}16] Mingzhou Zhou, Bo Wu, Xipeng Shen, Yaoqing Gao, and Graham Yiu. Examining and reducing the influence of sampling errors on feedbackdriven optimizations. ACM Transactions on Architecture and Code Optimization, 13(1): 6:1–6:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zheng:2017:WAD

[ZWY17] Wenguang Zheng, Hui Wu, and Qing Yang. WCET-aware dynamic I-cache locking for a single task. ACM Transactions on Architecture and Code Optimization, 14(1):4:1– 4:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhou:2016:CAE

[ZX16] Hao Zhou and Jingling Xue. A compiler approach for exploiting partial SIMD parallelism. ACM Transactions on Architecture and Code Optimization, 13(1):11:1–11:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2019:PPB

[ZX19] Feng Zhang and Jingling Xue.

Poker: Permutation-based SIMD execution of intensive tree search by path encoding. ACM Transactions on Architecture and Code Optimization, 15(4):46:1–46:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhang:2023:RRB

[ZXX23] Qiang Zhang, Lei Xu, and Baowen Xu. RegCPython: a register-based Python interpreter for better performance. ACM Transactions on Architecture and Code Optimization, 20(1):14:1–14:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3568973.

Zhou:2010:PAT

[ZYCZ10] Xiuyi Zhou, Jun Yang, Marek Chrobak, and Youtao Zhang. Performance-aware thermal management via task scheduling. ACM Transactions on Architecture and Code Optimization, 7(1):5:1–5:??, April 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zardoshti:2019:STM

[ZZB^{+}19] Pantea Zardoshti, Tingzhe Zhou, Pavithra Balaji, Michael L. Scott, and Michael Spear. Simplifying transactional memory support in C++. ACM Transactions on Architecture and Code Optimization, 16(3): 25:1–25:??, July 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhao:2023:GED

[ZZH^{+}23] Jin Zhao, Yu Zhang, Ligang He, Qikun Li, Xiang Zhang, Xinyu Jiang, Hui Yu, Xiaofei Liao, Hai Jin, Lin Gu, Haikun Liu, Bingsheng He, Ji Zhang, Xianzheng Song, Lin Wang, and Jun Zhou. GraphTune: an efficient dependency-aware substrate to alleviate irregularity in concurrent graph processing. ACM Transactions on Architecture and Code Optimization, 20(3):37:1–37:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3600091.

Zou:2021:SLE

[ZZL^{+}21] An Zou, Huifeng Zhu, Jingwen Leng, Xin He, Vijay Janapa Reddi, Christopher D. Gill, and Xuan Zhang. Systemlevel early-stage modeling and evaluation of IVR-assisted processor power delivery system. ACM Transactions on Architecture and Code Optimization, 18(4):52:1–52:27, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468145.

Zhou:2005:EFA

[ZZQ^{+}05] Yuanyuan Zhou, Pin Zhou, Feng Qin, Wei Liu, and Josep Torrellas. Efficient and flexible architectural support for dynamic monitoring. ACM Transactions on Architecture and Code Optimization, 2(1): 3–33, March 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).